### v1

# In-and-Out: Algorithmic Diffusion for Sampling Convex Bodies

Abstract
We present a new random walk for uniformly sampling high-dimensional convex bodies. It achieves state-of-the-art runtime complexity with stronger guarantees on the output than previously known, namely in Rényi divergence (which implies TV, $\mathcal{W}_{2}$, KL, $\chi^{2}$). The proof departs from known approaches for polytime algorithms for the problem — we utilize a stochastic diffusion perspective to show contraction to the target distribution with the rate of convergence determined by functional isoperimetric constants of the stationary density.
1 Introduction
Generating random samples from a high-dimensional convex body is a basic algorithmic problem with myriad connections and applications. The core of the celebrated result of [dyer1991random], giving a randomized polynomial-time algorithm for computing the volume of a convex body, was the first polynomial-time algorithm for uniformly sampling convex bodies. In the decades since, the study of sampling has led to a long series of improvements in its algorithmic complexity [lovasz1990mixing, lovasz1993random, kannan1997random, lovasz2006hit, cousins2018gaussian], often based on uncovering new mathematical/geometric structure, establishing connections to other fields (e.g., functional analysis, matrix concentration) and developing new tools for proving isoperimetric inequalities and analyzing Markov chains. With the proliferation of data and the increasing importance of machine learning, sampling has also become an essential algorithmic tool, with applications needing samplers in very high dimension, e.g., scientific computing [cousins2016practical, haraldsdottir2017chrr, kook2022sampling], systems biology [lewis2012constraining, thiele2013community], differential privacy [mcsherry2007mechanism, mironov2017renyi] and machine learning [bingham2019pyro, stan].
Samplers for convex bodies are based on Markov chains (see Appendix A for a summary). Their analysis is based on bounding the conductance of the associated Markov chain, which in turn bounds the mixing rate. Analyzing the conductance requires combining delicate geometric arguments with (Cheeger) isoperimetric inequalities for convex bodies. An archetypal example of the latter is the following: for any measurable partition $S_{1},S_{2},S_{3}$ of a convex body $\mathcal{K}\subset\mathbb{R}^{d}$, we have
$\operatorname{vol}(S_{3})\geq\frac{d(S_{1},S_{2})}{C_{\mathcal{K}}}\min\{\operatorname{vol}(S_{1}),\operatorname{vol}(S_{2})\}\,,$
where $d(\cdot,\cdot)$ is the (minimum) Euclidean distance, and $C_{\mathcal{K}}$ is an isoperimetric constant of the uniform distribution over $\mathcal{K}$. (The KLS conjecture posits that $C_{\mathcal{K}}=\mathcal{O}(1)$ for any convex body $\mathcal{K}$ in isotropic position, i.e., under the normalization that a random point from $\mathcal{K}$ has identity covariance). The coefficient $C_{\mathcal{K}}^{2}$ is bounded by the Poincaré constant of the uniform distribution over $\mathcal{K}$ (and they are in fact asymptotically equal). The classical proof of conductance uses geometric properties of the random walk at hand to reduce the analysis to a suitable isoperimetric inequality (see e.g., [lovasz1993random, vempala2005geometric]). The end result is a guarantee on the number of steps after which the total variation distance (TV distance) between the current distribution and the target is bounded by a desired error parameter. This framework has been widely used and effective in analyzing an array of candidate samplers, e.g., $\mathsf{Ball\ walk}$ [kannan1997random], $\mathsf{Hit\text{-}and\text{-}Run}$ [lovasz1999hit, lovasz2006hit], $\mathsf{Riemannian\ Hamiltonian\ Monte\ Carlo}$ [lee2018convergence] etc.
One successful approach, studied intensively over the past decade, is based on diffusion. The basic idea is to first analyze a continuous-time diffusion process, typically modeled by a stochastic differential equation (SDE), and then show that a suitable time-discretization of the process, sometimes together with a Metropolis filter, converges to the desired distribution efficiently. A major success along this line is the $\mathsf{Unadjusted\ Langevin\ Algorithm}$ and its variants [besag1995bayesian, dalalyan2012sparse, dalalyan2017further, durmus2019analysis, vempala2019rapid]. These algorithms have strong guarantees for sampling “nice” distributions, such as ones that are strongly log-concave, or more generally distributions satisfying isoperimetric inequalities, while also obeying some smoothness conditions. The analysis of these algorithms is markedly different from the conductance approach, and typically yields guarantees in stronger metrics such as the $\mathsf{KL}$-divergence.
Our starting point is the following question:
Can diffusion-based approaches be used for the problem of sampling convex bodies?
Despite remarkable progress, thus far, constrained sampling problems have evaded the diffusion approach, except as a high-level analogy (e.g., the $\mathsf{Ball\ walk}$ can be viewed as a discretization of Brownian motion, but this alone does not suggest a route for analysis) or with significantly worse convergence rates (e.g., [pmlr-v65-brosse17a, bubeck2018sampling]).
Our main finding is a simple diffusion-based algorithm that can be mapped to a stochastic process (and, importantly, to a pair of forward and backward processes), such that the rate of convergence is bounded directly by an appropriate functional inequality for the target distribution. As a consequence, for the first time, we obtain clean end-to-end guarantees in the Rényi divergence (which implies guarantees in other well known quantities such as $\mathcal{W}_{2},\mathsf{TV},\mathsf{KL},\chi^{2}$ etc.), while giving state-of-the-art runtime complexity for sampling convex bodies (e.g., $\mathsf{Ball\ walk}$ or $\mathsf{Speedy\ walk}$ [lovasz1993random, kannan1997random]). Besides being a stronger guarantee on the output, Rényi divergence is of particular interest for differential privacy [mironov2017renyi]. Perhaps most interesting is that our proof approach is completely different from prior work on convex body sampling. In summary,
  • The guarantees hold for the $q$-Rényi divergences while matching the rates of previous work (prior work only had guarantees in the TV distance).
  • The analysis is simple, modular, and easily extendable to several other settings.
Figure 1.1: Description of uniform samplers: (i) $\mathsf{Ball\ walk}$: proposes a uniform random point $z$ from $B_{\delta}(x_{1})$, but $z\notin\mathcal{K}$ so it stays at $x_{1}=x_{2}$. (ii) $\mathsf{Speedy\ walk}$: moves to $x_{2}$ drawn uniformly at random from $\mathcal{K}\cap B_{\delta}(x_{1})$. (iii) $\mathsf{In\text{-}and\text{-}Out}$: first moves to $y_{2}$ obtained by taking a Gaussian step from $x_{1}$, and then to $x_{2}$ obtained by sampling the truncated Gaussian $\mathcal{N}(y_{2},hI_{d})|_{\mathcal{K}}$.
1.1 Diffusion for uniform sampling
We propose the following $\mathsf{In\text{-}and\text{-}Out}$ [^1] This name reflects the “geometry” of how the iterates are moving. As we elaborate in Remark 1, the name ‘proximal sampler’ may be more familiar to those from an optimization background. sampler for uniformly sampling from $\mathcal{K}$. Each iteration consists of two steps, one that might leave the body and the second accepted only if it is (back) in the body.
Input: initial point $x_{0}\sim\pi_{0}$, convex body $\mathcal{K}\subset\mathbb{R}^{d}$, iterations $T$, threshold $N$, and $h>0$.
Output: $x_{T+1}$.
1:  for $i=0,\dotsc,T$ do
2:     Sample $y_{i+1}\sim\mathcal{N}(x_{i},hI_{d})$.
3:     Repeat: Sample $x_{i+1}\sim\mathcal{N}(y_{i+1},hI_{d})$ until $x_{i+1}\in\mathcal{K}$ or #attemptsi $\geq N$ (declare Failure).
4:  end for
Algorithm 1 $\mathsf{In\text{-}and\text{-}Out}$
It might be illuminating for the reader to compare this algorithm to the well-studied $\mathsf{Ball\ walk}$ (Algorithm 2); each proposed step is a uniform random point in a fixed-radius ball around the current point, and is accepted only if the proposed point is in the body $\mathcal{K}$. In contrast, each iteration of $\mathsf{In\text{-}and\text{-}Out}$ is a two-step process, where the first step (Line 2) ignores the boundary of the body, and the second step (Line 3) is accepted only if a proposal $x_{i+1}$ is a feasible point in $\mathcal{K}$. We will presently elaborate on the benefits of this variation.
Each successful iteration of the algorithm, i.e., one that is not declared “Failure”, can be called a proper step. We will see that the number of proper steps is directly bounded by isoperimetric constants (such as Poincaré and log-Sobolev) of the target distribution. In fact, this holds quite generally without assuming the convexity of $\mathcal{K}$. The implementation of an iteration is based on rejection sampling (Line 3), and our analysis of the efficiency of this step relies crucially on the convexity of $\mathcal{K}$. This is reminiscent of the $\mathsf{Speedy\ walk}$ in the literature on convex body sampling (Algorithm 3), which is used as a tool to analyze proper steps of the $\mathsf{Ball\ walk}$. We refer the reader to Appendix B for a brief survey on these and related walks.
This simple algorithm can be interpreted as a composition of “flows” in the space of measures. This view will allow us to use tools from stochastic analysis. In particular, we shall demonstrate how to interpret the two steps of one iteration of $\mathsf{In\text{-}and\text{-}Out}$ as alternating forward and backward heat flows.
We begin by defining an augmented probability measure on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ by
[…]
We denote by $\pi^{X},\pi^{X|Y}(\cdot|y)$ the marginal distribution of its first component (resp. conditional distribution given the second component), and similarly denote by $\pi^{Y},\pi^{Y|X}(\cdot|x)$ for the second component. In particular, the marginal in the first component $\pi^{X}$ is the uniform distribution over $\mathcal{K}$. Sampling from such a joint distribution to obtain the marginal on $X$ (say), can be more efficient than directly working only with $\pi^{X}$. This idea was utilized in Gaussian Cooling [cousins2018gaussian] and later as the restricted Gaussian Oracle (RGO) [lee2021structured, chen2022improved].
Under this notation, Algorithm 1 corresponds to a Gibbs sampling scheme from the two marginals of $\pi(x,y)$. To be precise, Line 2 and Line 3 correspond to sampling from
$y_{i+1}\sim\pi^{Y|X}(\cdot|x_{i})=\mathcal{N}(x_{i},hI_{d})\qquad\mbox{ and }\qquad x_{i+1}\sim\pi^{X|Y}(\cdot|y_{i+1})=\mathcal{N}(y_{i+1},hI_{d})|_{\mathcal{K}}\,.$
We implement the latter step through rejection sampling; if the number of trials in Line 3 hits the threshold $N$, then we halt and declare failure of the algorithm. It is well known that such a Gibbs sampling procedure will ensure the desired stationarity of $\pi(x,y)$.
Stochastic perspective: forward and backward heat flows.
Our algorithm can be viewed through the lens of stochastic analysis, due to an improved analysis for the proximal sampling [chen2022improved]. This view provides an interpolation in continuous-time, which is simple and powerful. To make this concrete, we borrow an exposition from [chewi2023log, Chapter 8.3]. Let us denote the successive laws of $x_{i}$ and $y_{i}$ by $\mu^{X}_{i}$ and $\mu^{Y}_{i}$, respectively. Recall that the first step of sampling from $\pi^{Y|X}(\cdot|x_{i})$ (Line 2) yields $\mu^{Y}_{i+1}=\int\pi^{Y|X=x}\,\mathrm{d}\mu^{X}_{i}(x)$. This is the result of evolving a probability measure under (forward) heat flow of $\mu_{i}^{X}$ for some time $h$, given precisely by the following stochastic differential equation: for $Z_{0}\sim\mu_{i}^{X}$,
$\mathrm{d}Z_{t}=\mathrm{d}B_{t}$ ($\mathsf{FH}$)
where $B_{t}$ is the standard Brownian process. We write $\mathsf{law}(Z_{t})=\mu_{i}^{X}P_{t}$. In particular, $Z_{h}=Z_{0}+\zeta\sim\mu_{i}^{X}*\mathcal{N}(0,hI_{d})=\mu_{i+1}^{Y}$ for $\zeta\sim\mathcal{N}(0,hI_{d})$. When $\mu_{i}^{X}=\pi^{X}$, the first step of Algorithm 1 gives
$\pi^{Y}(y)=[\pi^{X}*\mathcal{N}(0,hI_{d})](y)=\frac{1}{\operatorname{vol}(\mathcal{K})\,(2\pi h)^{d/2}}\int_{\mathcal{K}}\exp\bigl{(}-\frac{1}{2h}\|y-x\|^{2}\bigr{)}\,\mathrm{d}x\,.$ (1.1)
The second step of sampling from $\pi^{X|Y}(\cdot|y_{i+1})$ can be represented by $\mu_{i+1}^{X}=\int\pi^{X|Y=y}\,\mathrm{d}\mu_{i+1}^{Y}(y)$ (Line 3). The continuous-time process corresponding to this step might not be obvious. However, let us consider ($\mathsf{FH}$) with $Z_{0}\sim\pi^{X}$. Then, $Z_{h}\sim\pi^{Y}$, so the joint distribution of $(Z_{0},Z_{h})$ is simply $\pi$. This implies that $(Z_{0}|Z_{h}=y)\sim\pi^{X|Y=y}$. Imagine there is an SDE reversing the forward heat flow in a sense that if we are initialized deterministically at $Z_{h}\sim\delta_{y}$ at time 00, then the law of the SDE at time $h$ would be $\mathsf{law}(Z_{0}|Z_{h}=y)=\pi^{X|Y=y}$. Then, this SDE would serve as a continuous-time interpolation of the second step.
       Forward flow Backward flow
SDE $\mathrm{d}Z_{t}=\mathrm{d}B_{t}$ $\mathrm{d}Z_{t}^{\leftarrow}=\nabla\log(\pi^{X}P_{h-t})(Z_{t}^{\leftarrow})\,\mathrm{d}t+\mathrm{d}B_{t}$
Fokker-Planck $\partial_{t}\mu_{t}=\frac{1}{2}\Delta\mu_{t}$ $\partial_{t}\mu_{t}^{\leftarrow}=-\operatorname{div}\bigl{(}\mu_{t}^{\leftarrow}\nabla\log(\pi^{X}P_{h-t})\bigr{)}+\frac{1}{2}\Delta\mu_{t}^{\leftarrow}$
Table 1: The Fokker-Planck equations for the forward and backward heat flow describe how the laws of $Z_{t}$ and $Z_{t}^{\leftarrow}$ in ($\mathsf{FH}$) and ($\mathsf{BH}$) evolve over time. See Section 3.2 for details.
Such a time reversal SDE is indeed possible! The following SDE on $(Z_{t}^{\leftarrow})_{t\in[0,h]}$ initialized at $Z_{0}^{\leftarrow}\sim\pi^{Y}=\pi^{X}P_{h}$ ensures $Z_{h-t}\sim\mathsf{law}(Z_{t}^{\leftarrow})=\pi^{X}P_{h-t}$:
$\mathrm{d}Z_{t}^{\leftarrow}=\nabla\log(\pi^{X}P_{h-t})(Z_{t}^{\leftarrow})\,\mathrm{d}t+\mathrm{d}B_{t}\quad\text{for }t\in[0,h]\,.$ ($\mathsf{BH}$)
Although this is designed to reverse ($\mathsf{FH}$) initialized by $Z_{0}\sim\pi^{X}$ (so $Z_{h}=Z_{0}^{\leftarrow}\sim\pi^{Y})$, its construction also ensures that if $Z_{0}^{\leftarrow}\sim\delta_{y}$, a point mass, then $Z_{h}^{\leftarrow}\sim\mathsf{law}(Z_{0}|Z_{h}=y)=\pi^{X|Y=y}$. Thus, if we initialize ($\mathsf{BH}$) with $Z_{0}^{\leftarrow}\sim\mu_{i+1}^{Y}$, then the law of $Z_{h}^{\leftarrow}$ corresponds to $\int\pi^{X|Y=y}\,\mathrm{d}\mu_{i+1}^{Y}(y)=\mu_{i+1}^{X}$.
Remark 1.
We note that $\mathsf{In\text{-}and\text{-}Out}$ is exactly the proximal sampling scheme [lee2021structured, chen2022improved, fan2023improved] for uniform distributions. The proximal sampler with a target density proportional to $\exp(-V(x))$ considers an augmented distribution $\pi(x,y)\propto\exp(-V(x)-\frac{1}{2h}\|x-y\|^{2})$ and then repeats the following two steps: (1) $y_{i+1}\sim\pi^{Y|X=x_{i}}=\mathcal{N}(x_{i},hI_{d})$ and then (2) $x_{i+1}\sim\pi^{X|Y=y_{i+1}}$.
1.2 Results
Our model of computation is the classical general model for convex bodies [grotschel2012geometric]. We assume $\operatorname{vol}(\mathcal{K})>0$ throughout this paper. Below, $B_{r}(x)$ denotes the $d$-dimensional ball of radius $r$ centered at $x$.
Definition 1 (Convex body oracle).
A well-defined membership oracle for a convex body $\mathcal{K}\subset\mathbb{R}^{d}$ is given by a point $x_{0}\in\mathcal{K}$, a number $D>0$, with the guarantee that $B_{1}(x_{0})\subseteq\mathcal{K}\subseteq B_{D}(x_{0})$, and an oracle that correctly answers YES or NO to any query of the form “$x\in\mathcal{K}$?”
Definition 2 (Warmness).
A distribution $\mu$ is $M$-warm with respect to another distribution $\pi$, i.e., for every $x$ in the support of $\pi$, we have $\mathrm{d}\mu(x)\leq M\,\mathrm{d}\pi(x)$.
We now summarize our main result which is further elaborated in Section 3.4 (Theorem 27). Below, $\pi^{\mathcal{K}}$ is the uniform distribution over $\mathcal{K}$, and $\mathcal{R}_{q}$ is the Rényi-divergence of order $q$ (see Definition 6).
Theorem 3.
For any given $\eta,\varepsilon\in(0,1)$, $q\geq 1$, and any convex body $\mathcal{K}$ given by a well-defined membership oracle, there exist choices of parameters $h,N$ such that $\mathsf{In\text{-}and\text{-}Out}$, starting from an $M$-warm distribution, with probability at least $1-\eta$, returns $X\sim\mu$ such that $\mathcal{R}_{q}(\mu\mathbin{\|}\pi^{\mathcal{K}})\leq\varepsilon$. The number of proper steps is $\widetilde{\mathcal{O}}(qd^{2}\Lambda\log\frac{M}{\eta\varepsilon})$, and the expected total number of membership queries is $\widetilde{\mathcal{O}}(qMd^{2}\Lambda\log^{5}(1/\eta\varepsilon))$, where $\Lambda$ is the largest eigenvalue of the covariance of $\pi^{\mathcal{K}}$.
We note that for $X\sim\pi^{\mathcal{K}}$,
$\|\mathrm{Cov}(\pi^{\mathcal{K}})\|_{\mathsf{op}}\leq\operatorname{tr}(\mathrm{Cov}(\pi^{\mathcal{K}}))=\mathbb{E}\bigl{[}\|X-\mathbb{E}[X]\|^{2}\bigr{]}\leq D^{2}\,.$
The above guarantee in the Rényi divergence immediately provides $\mathcal{W}_{2},\mathsf{TV},\mathsf{KL}$, and $\chi^{2}$ guarantees as special cases. Previous guarantees for uniformly sampling convex bodies were only in the $\mathsf{TV}$-distance. For two distributions $\mu$ and $\pi$, we have
  1. 1.
    $\mathsf{KL}(\mu\mathbin{\|}\pi)=\lim_{q\to 1}\mathcal{R}_{q}(\mu\mathbin{\|}\pi)\leq\mathcal{R}_{q}(\mu\mathbin{\|}\pi)\leq\mathcal{R}_{q^{\prime}}(\mu\mathbin{\|}\pi)\leq\mathcal{R}_{\infty}(\mu\mathbin{\|}\pi)=\log\sup\frac{\mathrm{d}\mu}{\mathrm{d}\pi}$ for $1<q\leq q^{\prime}$.
  2. 2.
    $2\,\|\mu-\pi\|_{\mathsf{TV}}^{2}\leq\mathsf{KL}(\mu\mathbin{\|}\pi)\leq\log(\chi^{2}(\mu\mathbin{\|}\pi)+1)=\mathcal{R}_{2}(\mu\mathbin{\|}\pi)$.
  3. 3.
    $\mathcal{W}_{2}^{2}(\mu,\pi)\leq 2C_{\mathsf{LSI}}(\pi)\,\mathsf{KL}(\mu\mathbin{\|}\pi)$ (Talagrand’s $\mathsf{T}_{2}$-inequality) and $C_{\mathsf{LSI}}(\pi^{\mathcal{K}})\lesssim D^{2}$.
  4. 4.
    $\mathcal{W}_{2}^{2}(\mu,\pi)\leq 2C_{\mathsf{PI}}(\pi)\,\chi^{2}(\mu\mathbin{\|}\pi)$ [liu2020poincare] and $C_{\mathsf{PI}}(\pi^{\mathcal{K}})\lesssim\|\mathrm{Cov}(\pi^{\mathcal{K}})\|_{\mathsf{op}}\log d$.
The query complexity is better if the convex body is (near-)isotropic, i.e., the uniform distribution over the body has (near-)identity covariance. This relies on recent estimates of the worst-case Poincaré constant for isotropic log-concave distributions [kannan1995isoperimetric, klartag2023logarithmic].
Corollary 4.
Assume that $\pi^{\mathcal{K}}$ is isotropic. Under the same setting as above, $\mathsf{In\text{-}and\text{-}Out}$ succeeds with probability $1-\eta$, returning $X\sim\mu$ such that $\mathcal{R}_{q}(\mu\mathbin{\|}\pi^{\mathcal{K}})\leq\varepsilon$. The number of proper steps is $\widetilde{\mathcal{O}}(qd^{2}\log\frac{M}{\eta\varepsilon})$, and the expected total number of membership queries is $\widetilde{\mathcal{O}}(qMd^{2}\log^{5}(1/\eta\varepsilon))$.
Our analysis will in fact show that the bound on the number of proper steps holds for general non-convex bodies and any feasible start in $\mathcal{K}$. This is deduced under an $M$-warm start in Corollaries 28 and 29. We remark that such a bound for non-convex uniform sampling is not known for the $\mathsf{Ball\ walk}$ or the $\mathsf{Speedy\ walk}$.
Theorem 5.
For any given $\varepsilon\in(0,1)$ and set $\mathcal{K}\subset B_{D}(0)$ with $\operatorname{vol}(\mathcal{K})>0$, $\mathsf{In\text{-}and\text{-}Out}$ with variance $h$ and $M$-warm initial distribution achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after the following number of iterations:
$m=\min\begin{cases}\mathcal{O}\bigl{(}qh^{-1}C_{\mathsf{PI}}(\pi^{X})\log\frac{M}{\varepsilon}\bigr{)}&\text{for }q\geq 2\,,\\ \mathcal{O}\bigl{(}qh^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{\log M}{\varepsilon}\bigr{)}&\text{for }q\geq 1\,.\end{cases}$
We have two different convergence results above under (8) and (7). Under (8) we have a doubly-logarithmic dependence on the warmness parameter $M$. On the other hand, using (7), which is weaker than (8) (in general, $C_{\mathsf{PI}}\leq C_{\mathsf{LSI}}$), the dependence on $M$ is logarithmic. We discuss our results further in Section 1.3.
Outline of analysis.
We summarize our proof strategy below, which consists of two steps: (i) The current distribution should converge to the uniform distribution, (ii) within each iteration of the algorithm, the failure probability and the expected number of rejections should be small enough.
  • We need to demonstrate that the corresponding Markov chain is rapidly mixing. Here, we use the heat flow perspective to derive mixing rates under any suitable divergence measure (such as $\mathsf{KL}$, $\chi^{2}$, or $\mathcal{R}_{q}$), extending known results for the unconstrained setting [chen2022improved]. As a result, the mixing analysis reduces to a suitable functional inequality of the target distribution alone.
  • We show that the number of rejections in Line 3 over the entire execution of the algorithm is bounded with high probability. To do this, we apply a detailed argument involving local conductance and the convexity of $\mathcal{K}$, which relies on techniques from [belkin2006heat]. For this step, we show that with the appropriate choice of variance $h=\widetilde{\Theta}(d^{-2})$ and threshold $N=\widetilde{\Theta}(T\eta^{-1})$, the entire algorithm succeeds with probability $1-\eta$. The expected number of rejections is polylogarithmic.
While each individual component resembles pre-existing work in the literature, in their synthesis we will demonstrate how to interleave past developments in theoretical computer science, optimal transport, and functional analysis. The combination of these in this domain yields elegant and surprisingly simple proofs, as well as stronger results.
1.3 Discussion
Here we make a few remarks, contrasting our results with known ones.
No need to be lazy.
Previous uniform samplers like the $\mathsf{Ball\ walk}$ are made lazy (i.e., with probability $1/2$, the Markov chain does nothing), to ensure convergence to the target stationary distribution. However, our algorithm does not need this, since we directly show that our sampler contracts towards a uniform distribution.
Unified framework.
We remark that these two different bounds also place the previously known mixing guarantees for $\mathsf{Ball\ walk},\mathsf{Speedy\ walk}$ in a unified framework. Existing tight guarantees for $\mathsf{Speedy\ walk}$ are in TV distance and based on the log-Sobolev constant, assuming an oracle for implementing each step [lee2017eldan]. The known convergence guarantees of $\mathsf{Ball\ walk}$ (see Appendix B for details), namely the mixing time of $\widetilde{\mathcal{O}}(Md^{2}D^{2}\log\frac{1}{\varepsilon})$ for TV distance, are for the composite algorithm [$\mathsf{Speedy\ walk}+$rejection sampling]. Here $\mathsf{Speedy\ walk}$ records only the accepted steps of $\mathsf{Ball\ walk}$, so its stationary distribution differs slightly from the uniform distribution (and can be corrected with a post-processing step). On the other hand, $\mathsf{In\text{-}and\text{-}Out}$ actually converges to $\pi^{\mathcal{K}}$ without any adjustments and achieves stronger Rényi divergence bounds in the same asymptotic complexity. Our analysis shows that the mixing guarantee is determined by isoperimetric constants of the target (Poincaré or log-Sobolev).
Effective step size.
The $\mathsf{Ball\ walk}$’s largest possible step size is of order $1/\sqrt{d}$ (see Appendix B) to keep the rejection probability bounded by a constant. This bound could also be viewed as an “effective” step size of $\mathsf{In\text{-}and\text{-}Out}$. This follows from the fact that the $\ell_{2}$-norm of the Gaussian $\mathcal{N}(0,hI)$ is concentrated around $\sqrt{hd}$, and we will set the variance $h$ of $\mathsf{In\text{-}and\text{-}Out}$ to $\widetilde{\mathcal{O}}(1/d^{2})$, so we have $\sqrt{hd}\approx 1/\sqrt{d}$.
What has really changed?
$\mathsf{In\text{-}and\text{-}Out}$ has clear similarities to both $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$. What then are the changes that allow us to use continuous-time interpolation? One step of $\mathsf{Ball\ walk}$ is [random step ($y\in B_{\delta}(x)$) $+$ Metropolis-filter (accept if $y\in\mathcal{K}$)]. This filtering is an abrupt discrete step, and it is unclear how to control contraction. It could be replaced by a step of $\mathsf{Speedy\ walk}$ ($x\sim\text{Unif}(B_{\delta}(y)\cap\mathcal{K})$). Then, each iteration of $\mathsf{In\text{-}and\text{-}Out}$ can be viewed as a Gaussian version of a $\mathsf{Ball\ walk^{\prime}s\ proposal}+\mathsf{Speedy\ walk}$ algorithm.
How can we compare $\mathsf{In\text{-}and\text{-}Out}$ with $\mathsf{Speedy\ walk}?$ Iterating speedy steps leads to a biased distribution (one that is proportional to the local conductance). As clarified in Remark 2, one step of (a Gaussian version of) $\mathsf{Speedy\ walk}$ can be understood as a step of backward heat flow. Therefore, if one can control the isoperimetric constants of the biased distribution along the trajectory of the backward flow, then contraction of $\mathsf{Speedy\ walk}$ toward the biased distribution will follow from the simultaneous backward analysis.
2 Preliminaries
[…]
To quantify the convergence rate, we introduce some common divergences between distributions.
Definition 6 (Distance and divergence).
For two measures $\mu,\nu$ on $\mathbb{R}^{d}$, the total variation distance between them is defined by
$\|\mu-\nu\|_{\mathsf{TV}}:=\sup_{B\in\mathcal{F}}\lvert\mu(B)-\nu(B)\rvert\,,$
where $\mathcal{F}$ is the collection of all measurable subsets of $\mathbb{R}^{d}\,$. The $2$-Wasserstein distance is given by
$\mathcal{W}_{2}^{2}(\mu,\nu):=\inf_{\gamma\in\Gamma(\mu,\nu)}\mathbb{E}_{(X,Y)\sim\gamma}[\|X-Y\|^{2}]\,,$
where $\Gamma$ is the set of all couplings between $\mu,\nu$. Next, we define the $f$-divergence of $\mu$ towards $\nu$ with $\mu\ll\nu$ (i.e., $\mu$ is absolutely continuous with respect to $\nu$) as, for some convex function $f:\mathbb{R}_{+}\to\mathbb{R}$ with $f(1)=0$ and $f^{\prime}(\infty)=\infty$,
$D_{f}(\mu\mathbin{\|}\nu):=\int f\Bigl{(}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\Bigr{)}\,\mathrm{d}\nu\,.$
The $\mathsf{KL}$-divergence arises when taking $f(x)=x\log x$, the $\chi^{q}$-divergence when taking $f(x)=x^{q}-1$, and the $q$-Rényi divergence is given by
$\mathcal{R}_{q}(\mu\mathbin{\|}\nu):=\frac{1}{q-1}\log\bigl{(}\chi^{q}(\mu\mathbin{\|}\nu)+1\bigr{)}\,.$
We recall two important functional inequalities of a distribution. We use $\nu$ to denote a probability measure over $\mathbb{R}^{d}$.
Definition 7.
We say that $\nu$ satisfies a Poincaré inequality (PI) with parameter $C_{\mathsf{PI}}(\nu)$ if for all smooth functions $f:\mathbb{R}^{d}\to\mathbb{R}$,
$\mathrm{Var}_{\nu}f\leq C_{\mathsf{PI}}(\nu)\,\mathbb{E}_{\nu}[\|\nabla f\|^{2}]\,,$
where $\mathrm{Var}_{\nu}f\coloneqq\mathbb{E}_{\nu}\lvert f-\mathbb{E}_{\nu}f\rvert^{2}$.
The Poincaré inequality is implied by the log-Sobolev inequality.
Definition 8.
We say that $\nu$ satisfies a log-Sobolev inequality (LSI) with parameter $C_{\mathsf{LSI}}(\nu)$ if for all smooth functions $f:\mathbb{R}^{d}\to\mathbb{R}$,
$\mathsf{Ent}_{\nu}(f^{2})\leq 2C_{\mathsf{LSI}}(\nu)\,\mathbb{E}_{\nu}[\|\nabla f\|^{2}]\,,$
where $\mathsf{Ent}_{\nu}(f^{2}):=\mathbb{E}_{\nu}[f^{2}\log f^{2}]-\mathbb{E}_{\nu}[f^{2}]\log(\mathbb{E}_{\nu}[f^{2}])$. Equivalently, for any probability measure $\mu$ over $\mathbb{R}^{d}$ with $\mu\ll\nu$,
$\mathsf{KL}(\mu\mathbin{\|}\nu)\leq\frac{C_{\mathsf{LSI}}(\nu)}{2}\,\mathsf{FI}(\mu\mathbin{\|}\nu)\,,$
where $\mathsf{FI}(\mu\mathbin{\|}\nu):=\mathbb{E}_{\mu}[\|\nabla\log\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\|^{2}]$ is the Fisher information of $\mu$ with respect to $\nu$.
We recall the data-processing inequality for Rényi divergence and $f$-divergence.
Lemma 9 (Data-processing inequality).
For measures $\mu,\nu$, Markov kernel $P$, $f$-divergence $D_{f}$, and $q\geq 1$, it holds that
$D_{f}(\mu P\mathbin{\|}\nu P)\leq D_{f}(\mu\mathbin{\|}\nu)\,,\quad\text{and}\quad\mathcal{R}_{q}(\mu P\mathbin{\|}\nu P)\leq\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\,.$
The aforementioned functional inequalities allow us to show exponential contraction of various divergences, through the following helpful inequality.
Lemma 10 (Grönwall).
Suppose that $u,g:[0,T]\to\mathbb{R}$ are two continuous functions, with $u$ being differentiable on $[0,T]$ and satisfying
$u^{\prime}(t)\leq g(t)\,u(t)\qquad\text{for all }t\in[0,T]\,.$
Then,
$u(t)\leq\exp\Bigl{(}\int_{0}^{t}g(s)\,\mathrm{d}s\Bigr{)}\,u(0)\qquad\text{forall }t\in[0,T]\,.$
3 Analysis
We begin this section by proving some introductory lemmas, which will streamline our later exposition.
We first record two fundamental lemmas, which introduce the mathematical formalism for our analysis. The first is the existence of forward and backward heat flows (Lemma 22), which will interpolate each line in Algorithm 1. These flow equations describe how the laws of $Z_{t}$ and $Z_{t}^{\leftarrow}$ in ($\mathsf{FH}$) and ($\mathsf{BH}$) evolve respectively over time.
Lemma 11.
The forward heat flow equation with initial distribution $\mu_{0}$ is given by
$\partial_{t}\mu_{t}=\frac{1}{2}\Delta\mu_{t}\,,$
and its backward heat flow equation is given by
$\partial_{t}\mu_{t}^{\leftarrow}=-\operatorname{div}\bigl{(}\mu_{t}^{\leftarrow}\nabla\log(\pi^{X}P_{h-t})\bigr{)}+\frac{1}{2}\Delta\mu_{t}^{\leftarrow}\quad\text{with }\mu_{0}^{\leftarrow}=\mu_{h}\,.$
These admit (weak) solutions on $[0,h]$ for any initial distribution $\mu_{0}$ with $\frac{\mathrm{d}\mu_{0}}{\mathrm{d}\pi^{X}}\leq M<\infty$.
One successful iteration of $\mathsf{In\text{-}and\text{-}Out}$ is exactly the same as the composition of running the forward heat flow and then backward heat flow, both for time $h$. Its validity is justified in Section 1.1, and we give its proof below.
Lemma 12.
Let $\mu_{k}^{X}$ be the law of the $k$-th iterate $x_{k}$ of $\mathsf{In\text{-}and\text{-}Out}$. If ($\mathsf{FH}$) is initialized with $\mathsf{law}(Z_{0})=\mu_{k}^{X}$, then $\mathsf{law}(Z_{h})=\mu_{k+1}^{Y}$. If ($\mathsf{BH}$) is initialized with $\mathsf{law}(Z_{0}^{\leftarrow})=\mu_{k+1}^{Y}$, then $\mathsf{law}(Z_{h}^{\leftarrow})=\mu_{k+1}^{X}$.
Proof.
We explicitly show that the forward and backward heat flows indeed interpolate the two discrete steps given in Algorithm 1. For the forward part, we have $Z_{h}=Z_{0}+\zeta$ for $\zeta\sim\mathcal{N}(0,hI_{d})$, so
$\mathsf{law}(Z_{h})=\mathsf{law}(Z_{0})*\mathcal{N}(0,hI_{d})=\mu_{k}^{X}*\mathcal{N}(0,hI_{d})=\mu_{k+1}^{Y}\,.$
Regarding the backward part, it is known from [chen2022improved, Lemma 14] that the construction of the time-reversal SDE ensures that $(Z_{h}^{\leftarrow},Z_{0}^{\leftarrow})$ and $(Z_{0},Z_{h})$ have the same joint distribution, when $Z_{0}\sim\pi^{X}$ (and so $Z_{h}\sim\pi^{Y}$). Hence, $\mathsf{law}(Z_{h}^{\leftarrow}|Z_{0}^{\leftarrow}=y)=\mathsf{law}(Z_{0}|Z_{h}=y)=\pi^{X|Y=y}$, where the last equality follows from $(Z_{0},Z_{h})\sim\pi$. Since we initialize ($\mathsf{BH}$) with $Z_{0}^{\leftarrow}=y\sim\mu_{k+1}^{Y}$, we have
$\mathsf{law}(Z_{h}^{\leftarrow})=\int\mathsf{law}(Z_{h}^{\leftarrow}|Z_{0}^{\leftarrow}=y)\,\mu_{k+1}^{Y}(\mathrm{d}y)=\int\pi^{X|Y}(\cdot|y)\,\mu_{k+1}^{Y}(\mathrm{d}y)=\mu_{k+1}^{X}\,,$
where the last follows from the definition of Line 3. ∎
Our analysis of these objects consists of two parts: (1) demonstrating the mixing of $\mathsf{In\text{-}and\text{-}Out}$, i.e., how many outer iterations are needed to be sufficiently close to the uniform distribution, and (2) quantifying the failure probability and wasted steps in Line 3.
For (1), we collect in Section 3.1 some important implications of functional inequalities, e.g. the Poincaré and log-Sobolev inequalities, for the uniform distribution. Then in Section 3.2, we exploit the flow perspective of the algorithm to obtain the mixing guarantees. To this end, we revisit the proofs for the contraction results of forward and backward heat flows, and then extend them to our constrained setting:
Theorem 13 (Adapted for uniform distributions, [chen2022improved]).
Let $\mu^{X}_{k}$ be the law of the $k$-th output of $\mathsf{In\text{-}and\text{-}Out}$ with initial distribution $\mu_{0}^{X}$. Let $C_{\mathsf{LSI}}$ be the (8) constant of the uniform distribution $\pi^{X}$ over $\mathcal{K}$. Then, for any $q\geq 1$,
$\mathcal{R}_{q}(\mu_{k}^{X}\mathbin{\|}\pi^{X})\leq\frac{\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})}{(1+h/C_{\mathsf{LSI}})^{2k/q}}\,.$
For $C_{\mathsf{PI}}$ the (7) constant of $\pi^{X}$ and any $q\geq 2$,
$\mathcal{R}_{q}(\mu_{k}^{X}\mathbin{\|}\pi^{X})\leq\begin{cases}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-\frac{2k\log(1+h/C_{\mathsf{PI}})}{q}&\text{if }k\leq\frac{q}{2\log(1+h/C_{\mathsf{PI}})}\,\bigl{(}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-1\bigr{)}\,,\\ (1+h/C_{\mathsf{PI}})^{-2(k-k_{0})/q}&\text{if }k\geq k_{0}:=\lceil\frac{q}{2\log(1+h/C_{\mathsf{PI}})}\,\bigl{(}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-1\bigr{)}\rceil\,.\end{cases}$
This reduces the problem of obtaining a mixing guarantee to that of demonstrating a functional inequality on the target distribution. For this, it is not strictly necessary that $\mathcal{K}$ be convex.
On the other hand, convexity of $\mathcal{K}$ is crucial for the proof of (2). We show in Section 3.3 that the failure probability remains under control by taking a suitable variance $h$ and threshold $N$. We then show that the expected number of trials per iteration is of order $\log N$, not $N$:
Lemma 14 (Per-iteration guarantees).
Let $\mathcal{K}$ be any convex body in $\mathbb{R}^{d}$ presented by a well-defined membership oracle, $\pi^{X}$ the uniform distribution over $\mathcal{K}$, and $\mu$ an $M$-warm initial distribution with respect to $\pi^{X}$. For any given $m\in\mathbb{N}$ and $\eta\in(0,1)$, set $Z=\frac{9mM}{\eta}(\geq 9)$, $h=\frac{\log\log Z}{2d^{2}\log Z}$ and $N=Z(\log Z)^{4}=\widetilde{\mathcal{O}}(\frac{mM}{\eta})$. Then, the failure probability of one iteration of $\mathsf{In\text{-}and\text{-}Out}$ is at most $\eta/m$. Moreover, the expected number of membership queries needed per iteration is $\mathcal{O}\bigl{(}M(\log\frac{mM}{\eta})^{4}\bigr{)}$.
3.1 Functional inequalities
The contraction of an outer loop of our algorithm is controlled by isoperimetry of the uniform distribution $\pi^{X}$, which is described precisely by a functional inequality. The most natural ones to consider in this setting are the Poincaré inequality (7) and log-Sobolev inequality (8). In Appendix C, we provide a more detailed discussion of how these are related to other important notions of isoperimetry, such as the Cheeger and log-Cheeger inequalities.
Below, we use $\mu,\nu$ to denote two arbitrary probability measures over $\mathbb{R}^{d}$. The relationship between a Poincaré inequality and the $\chi^{2}$-divergence is derived by substituting $f=\frac{\mathrm{d}\nu}{\mathrm{d}\mu}$ into (7).
Lemma 15.
Assume that $\nu$ satisfies (7) with parameter $C_{\mathsf{PI}}(\nu)$. For any probability measure $\mu$ over $\mathbb{R}^{d}$ with $\mu\ll\nu$, it holds that
$\chi^{2}(\mu\mathbin{\|}\nu)\leq\frac{C_{\mathsf{PI}}(\nu)}{2}\,\mathbb{E}_{\nu}\bigl{[}\bigl{\|}\nabla\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\bigr{\|}^{2}\bigr{]}\,.$
The Poincaré inequality implies functional inequalities for the Rényi divergence.
Lemma 16 ([vempala2019rapid, Lemma 9]).
Assume that $\nu$ satisfies (7) with parameter $C_{\mathsf{PI}}(\nu)$. For any $q\geq 2$ and probability measure $\mu$ over $\mathbb{R}^{d}$, it holds that
$1-\exp(-\mathcal{R}_{q}(\mu\mathbin{\|}\nu))\leq\frac{q\,C_{\mathsf{PI}}(\nu)}{4}\,\mathsf{RF}_{q}(\mu\mathbin{\|}\nu)\,,$
where $\mathsf{RF}_{q}(\mu\mathbin{\|}\nu):=q\,\mathbb{E}_{\nu}\bigl{[}\bigl{(}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\bigr{)}^{q}\|\nabla\log\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\|^{2}\bigr{]}\,/\,\mathbb{E}_{\nu}\bigl{[}\bigl{(}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\bigr{)}^{q}\bigr{]}$ is the Rényi Fisher information of order $q$ of $\mu$ with respect to $\nu$.
The log-Sobolev inequality paired with the KL-divergence (8) can be understood as a special case of the following inequality [^2] Such inequalities are often called Polyak-Łojasiewicz inequalities, which say for $f:\mathbb{R}^{d}\to\mathbb{R}$, and all $y\in\mathbb{R}^{d}$ that $f(y)\leq c\,\|\nabla f(y)\|^{2}$ for some constant $c$, if $\min f(x)=0$. paired with the $q$-Rényi divergence for $q\geq 1$.
Lemma 17 ([vempala2019rapid, Lemma 5]).
Assume that $\nu$ satisfies (8) with parameter $C_{\mathsf{LSI}}(\nu)$. For any $q\geq 1$ and probability measure $\mu$ over $\mathbb{R}^{d}$, it holds that
$\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq\frac{q\,C_{\mathsf{LSI}}(\nu)}{2}\,\mathsf{RF}_{q}(\mu\mathbin{\|}\nu)\,.$
Note that $\lim_{q\to 1}\mathcal{R}_{q}=\mathsf{KL}$ and $\mathsf{RF}_{1}=\mathsf{FI}$.
We have collected below the functional inequalities used to establish the mixing of our algorithm (see Appendix C for a detailed presentation).
Lemma 18.
Let $\mathcal{K}\subset\mathbb{R}^{d}$ be a convex body with diameter $D$, and $\pi$ be the uniform distribution over $\mathcal{K}$. Then, $C_{\mathsf{PI}}(\pi)\lesssim\|\mathrm{Cov}(\pi)\|_{\mathsf{op}}\,\log d$ and $C_{\mathsf{LSI}}(\pi)\lesssim D^{2}$. If $\pi$ is isotropic, then $C_{\mathsf{PI}}(\pi)\lesssim\log d$ and $C_{\mathsf{LSI}}(\pi)\lesssim D$.
3.2 Contraction and mixing
We start by analyzing how many outer iterations of $\mathsf{In\text{-}and\text{-}Out}$ are required to be $\varepsilon$-close to $\pi^{X}$, the uniform distribution over $\mathcal{K}$. The contraction of Algorithm 1 comes from analyzing Lines 2 and 3 through the perspective of heat flows (see Section 1.1). To exploit this view, we first revisit the previous contraction analysis in [chen2022improved], which is carried out for distributions with smooth densities. Although the uniform distribution is not even continuous, we prove a technical lemma (Lemma 22) that enables us to extend previously known results to the uniform distribution. Lastly, combining the previous results with our technical lemma, we obtain clean contraction results of Algorithm 1 toward the uniform distribution $\pi^{X}$ in Theorem 23.
Part I: Contraction analysis for smooth distributions.
In this part, we review the contraction results for heat flow and its time-reversal [chen2022improved], which are intimately connected with our algorithm. We also provide key technical ingredients needed for its proof, such as the computations for measures evolving under simultaneous forward/backward heat flows. We refer interested readers to Appendix D for additional details. Only in Part I, we assume that $\nu$ denotes a probability measure with smooth density.
Forward heat flow.
We begin by introducing the “heat flow” equation (or also known as the Fokker-Planck equation), which describes the evolution of the law of $Z_{t}$ under ($\mathsf{FH}$),
$\partial_{t}\mu_{t}=\frac{1}{2}\,\Delta\mu_{t}=\frac{1}{2}\operatorname{div}(\mu_{t}\nabla\log\mu_{t})\,.$ ($\mathsf{FP\text{-}FH}$)
It is well known that one can realize this equation in discrete time through a Gaussian transition density, in the sense that, for $\mu_{h}$ (the solution at time $h>0$ to ($\mathsf{FP\text{-}FH}$) with initial condition $\mu_{0}$), and for any smooth function $f:\mathbb{R}^{d}\to\mathbb{R}$,
$\mathbb{E}_{\mu_{h}}[f(x)]=\mathbb{E}_{\mu_{0}}[P_{h}f(x)]\,,$
where $P_{h}f(x)=\mathbb{E}_{\mathcal{N}(x,hI_{d})}[f]$.333$\{P_{h}\}_{h\geq 0}$ is often called the heat semigroup. By this we can formally identify $\mu_{h}=\mu_{0}P_{h}$, and also write $\mu_{h}$ for the law of $Z_{h}$, where $\{Z_{h}\}_{h\geq 0}$ solves ($\mathsf{FH}$).
Backward heat flow.
Although there are many ways to define a “reversal” of $P_{h}$, we will use the notion of adjoint introduced by [klartag2021spectral], which is the most immediately useful.
[…]
$\partial_{t}\mu_{t}^{\leftarrow}=-\operatorname{div}\bigl{(}\mu_{t}^{\leftarrow}\nabla\log(\nu P_{h-t})\bigr{)}+\frac{1}{2}\Delta\mu_{t}^{\leftarrow}\quad\text{for }t\in[0,h]\,.$ ($\mathsf{FP\text{-}BH}$)
Write $\mu_{t}^{\leftarrow}=\mu_{0}^{\leftarrow}Q_{t}^{\nu,h}$, where $\{Q_{t}^{\nu,h}\}_{t\in[0,h]}$ is a family of transition densities. Write $\mathbf{P}_{0,h}$ for the joint distribution of the $(Z_{0},Z_{h})$-marginals of ($\mathsf{FH}$), when $Z_{0}\sim\nu$, and $\mathbf{P}_{0|h}$ for the conditional. Note that $\mathbf{P}_{h|0}(\cdot|x)=\mathcal{N}(x,hI_{d})$. It is also known that ($\mathsf{FP\text{-}BH}$) gives a time-reversal of the heat equation at the SDE level, in the sense that we can interpret $\delta_{x}Q_{h}^{\nu,h}=\mathbf{P}_{0|h}(\cdot|Z_{h}=x)$. Thus $\mu_{0}^{\leftarrow}Q_{h}^{\nu,h}=\int\mathbf{P}_{0|h}(\cdot|Z_{h}=x)\,\mu_{0}^{\leftarrow}(\mathrm{d}x)$, and $\nu P_{h}Q_{t}^{\nu,h}=\nu P_{h-t}$ for all $t\in[0,h]$.
The ultimate purpose of this machinery is to affirm our earlier description of the Gibbs sampling procedure as alternating forward and backward heat flows. Indeed, notice that, if $\mu_{i}^{X}$ is the law of the iterate at some iteration $i$, then $\mu_{i}^{X}P_{h}$ is precisely $\mu_{i+1}^{Y}$ under our scheme, while $(\mu_{i}^{X}P_{h})Q_{h}^{\pi^{X},h}$ is $\mu_{i+1}^{X}$, assuming $Q_{h}^{\pi^{X},h}$ is well defined for non-smooth measures $\pi^{X}$. Thus, while Algorithm 1 is implemented via discrete steps, it can be exactly analyzed through arguments in continuous time. We shall see the benefits of this shortly.
Instead of considering the change in metrics along the evolution of $\mu P_{t}$ with respect to “fixed” $\nu$, it will be useful to consider the simultaneous evolution of $\mu P_{t},\nu P_{t}$ (and similarly $\mu Q_{t},(\nu P_{h})Q_{t}$). This type of computation was carried out for specific metrics in earlier work [vempala2019rapid, chen2022improved]. The following is a more generalized form of one appearing in [yuan2023class, Lemma 2]. In the lemma below, we consider an arbitrary diffusion equation with corresponding Fokker-Planck equation:
$\mathrm{d}X_{t}=b_{t}(X_{t})\,\mathrm{d}t+\,\mathrm{d}B_{t}\,\quad\text{and}\quad\partial_{t}\mu_{t}=-\nabla\cdot(b_{t}\mu_{t})+\frac{1}{2}\Delta\mu_{t}\,$ (3.1)
where $b_{t}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is smooth, $X_{t}\in\mathbb{R}^{d}$, and $\mu_{t}=\mathsf{Law}(X_{t})$ if $X_{0}\sim\mu_{0}$.
Lemma 19 (Decay along forward/backward heat flows).
Let $(\mu_{t})_{t\geq 0},(\nu_{t})_{t\geq 0}$ denote the laws of the solutions to (3.1) starting at $\mu_{0},\nu_{0}$ respectively. Then, for any differentiable function $g$,
$\partial_{t}g\bigl{(}D_{f}(\mu_{t}\mathbin{\|}\nu_{t})\bigr{)}=-\frac{1}{2}\,g^{\prime}\bigl{(}D_{f}(\mu_{t}\mathbin{\|}\nu_{t})\bigr{)}\times\mathbb{E}_{\mu_{t}}\Bigl{\langle}\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)},\nabla\log\frac{\mu_{t}}{\nu_{t}}\Bigr{\rangle}\,.$
Proof.
The case where $g\neq\mathsf{id}$ is an application of the chain rule, so it suffices to take $g=\mathsf{id}$ and simply differentiate an $f$-divergence.
For brevity, we drop the variable $x$ of functions involved, and proceed as follows:
$\partial_{t}D_{f}(\mu_{t}\mathbin{\|}\nu_{t})$ =$=\int\Bigl{\{}\bigl{(}f\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\partial_{t}\nu_{t}+\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\bigl{(}\frac{\mu_{t}}{\nu_{t}}\bigr{)}^{\prime}\nu_{t}\Bigr{\}}\,\mathrm{d}x$
=$=\int\Bigl{\{}\partial_{t}\nu_{t}\Bigl{(}\bigl{(}f\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}-\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\Bigr{)}+\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\partial_{t}\mu_{t}\Bigr{\}}\,\mathrm{d}x$
=$\underset{(i)}{=}\int\bigl{[}-\nabla\cdot(b_{t}\nu_{t})+\frac{1}{2}\Delta\nu_{t}\bigr{]}\Bigl{(}\bigl{(}f\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}-\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\Bigr{)}\,\mathrm{d}x$
$\qquad+\int\bigl{[}-\nabla\cdot(b_{t}\mu_{t})+\frac{1}{2}\Delta\mu_{t}\bigr{]}\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\,\mathrm{d}x\,,$
where in $(i)$ we substitute the F-P equation from (3.1). Integrating by parts (i.e., $\int f\operatorname{div}(\mathbf{G})=-\int\langle\nabla f,\mathbf{G}\rangle$ for a real-valued function $f$ and vector-valued function $\mathbf{G}$), we have that
$\int\bigl{[}-\nabla\cdot(b_{t}\nu_{t})\bigr{]}\bigl{(}f\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\,\mathrm{d}x=\int\Bigl{\langle}b_{t}\nu_{t},\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\nabla\frac{\mu_{t}}{\nu_{t}}\Bigr{\rangle}\,\mathrm{d}x\,.$ (3.2)
On the other hand, we have that
$-\int\bigl{[}-\nabla\cdot(b_{t}\nu_{t})\bigr{]}\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\,\mathrm{d}x=-\int\Bigl{\langle}b_{t}\nu_{t},\frac{\mu_{t}}{\nu_{t}}\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}+\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\nabla\frac{\mu_{t}}{\nu_{t}}\Bigr{\rangle}\,\mathrm{d}x\,.$
The second term cancels with the RHS of (3.2). We have a similar cancellation for the $\frac{1}{2}\Delta\nu_{t}$ term:
$\int\frac{1}{2}\Delta\nu_{t}\,\bigl{(}f\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\,\mathrm{d}x=-\int\frac{1}{2}\Bigl{\langle}\nabla\nu_{t},\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\nabla\frac{\mu_{t}}{\nu_{t}}\Bigr{\rangle}\,\mathrm{d}x\,,$
and
$-\int\frac{1}{2}\Delta\nu_{t}\,\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\,\mathrm{d}x=\int\frac{1}{2}\Bigl{\langle}\nabla\nu_{t},\frac{\mu_{t}}{\nu_{t}}\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}+\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\nabla\frac{\mu_{t}}{\nu_{t}}\Bigr{\rangle}\,\mathrm{d}x\,.$
Combining these, we are left with
$\int\bigl{[}-\nabla\cdot(b_{t}\nu_{t})+\frac{1}{2}\Delta\nu_{t}\bigr{]}\Bigl{(}\bigl{(}f\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}-\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\Bigr{)}\,\mathrm{d}x$ $=-\int\Bigl{\langle}b_{t}\nu_{t}-\frac{1}{2}\nabla\nu_{t},\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\Bigr{\rangle}\,\mathrm{d}x$
$=-\int\Bigl{\langle}b_{t}\mu_{t}-\frac{1}{2}\mu_{t}\nabla\log\nu_{t},\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\Bigr{\rangle}\,\mathrm{d}x\,.$
Finally, we note that
$\int\bigl{[}-\nabla\cdot(b_{t}\mu_{t})+\frac{1}{2}\Delta\mu_{t}\bigr{]}\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\,\mathrm{d}x$ $=\int\Bigl{\langle}b_{t}\mu_{t}-\frac{1}{2}\nabla\mu_{t},\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\Bigr{\rangle}\,\mathrm{d}x$
$=\int\Bigl{\langle}b_{t}\mu_{t}-\frac{1}{2}\mu_{t}\nabla\log\mu_{t},\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\Bigr{\rangle}\,\mathrm{d}x\,.$
Putting it all together, noticing that the drift terms cancel, we are left with
$\partial_{t}D_{f}(\mu_{t}\mathbin{\|}\nu_{t})=-\int\frac{1}{2}\Bigl{\langle}\mu_{t}\nabla\log\frac{\mu_{t}}{\nu_{t}},\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\Bigr{\rangle}\,\mathrm{d}x=-\frac{1}{2}\mathbb{E}_{\mu_{t}}\Bigl{\langle}\nabla\log\frac{\mu_{t}}{\nu_{t}},\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\Bigr{\rangle}\,,$
which completes the proof. ∎
To recover the decay result for the $q$-Rényi divergence, one can substitute $g(x)=\frac{1}{q-1}\log x$ and $f(x)=x^{q}-1$. For the $\chi^{2}$-divergence, instead substitute $g(x)=x$ and $f(x)=x^{2}-1$. From this, we can obtain a single step of decay for the Rényi and $\chi^{2}$-divergences under different functional inequalities.
Before proceeding, we need a standard lemma on functional inequalities under ($\mathsf{FH}$).
Lemma 20 (Functional inequalities under Gaussian convolutions, [chafai2004entropies, Corollary 13]).
The following inequality holds for any $\pi$ with finite log-Sobolev and Poincaré constants,
$C_{\mathsf{PI}}(\pi P_{t})\leq C_{\mathsf{PI}}(\pi)+t\,,\qquad\text{and}\qquad C_{\mathsf{LSI}}(\pi P_{t})\leq C_{\mathsf{LSI}}(\pi)+t\,.$
Combining the previous two lemmas, we can establish contraction between $\mu P_{h}Q_{h}$ and $\nu$ after one forward/backward iteration.
Theorem 21 ([chen2022improved, Theorem 3 and 4]).
Assume $\nu$, a measure with smooth density, satisfies (8) with constant $C_{\mathsf{LSI}}$. For any $q\geq 1$ and initial distribution $\mu$ with a smooth density, denoting again $Q_{h}:=Q_{h}^{\nu,h}$,
$\mathcal{R}_{q}(\mu P_{h}Q_{h}\mathbin{\|}\nu)\leq\frac{\mathcal{R}_{q}(\mu\mathbin{\|}\nu)}{(1+h/C_{\mathsf{LSI}})^{2/q}}\,.$
If $\nu$ satisfies (7) with constant $C_{\mathsf{PI}}$, then it follows that
$\chi^{2}(\mu P_{h}Q_{h}\mathbin{\|}\nu)\leq\frac{\chi^{2}(\mu\mathbin{\|}\nu)}{(1+h/C_{\mathsf{PI}})^{2}}\,.$
Moreover, for all $q\geq 2$,
$\mathcal{R}_{q}(\mu P_{h}Q_{h}\mathbin{\|}\nu)\leq\begin{cases}\mathcal{R}_{q}(\mu\mathbin{\|}\nu)-\frac{2\log(1+h/C_{\mathsf{PI}})}{q}&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\geq 1\,,\\ \frac{\mathcal{R}_{q}(\mu\mathbin{\|}\nu)}{(1+h/C_{\mathsf{PI}})^{2}}&\text{if}\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq 1\,.\end{cases}$
Proof.
Since the SDE in (3.1) captures the forward heat flow ($\mathsf{FH}$), we set $\mu_{0}$ and $\nu_{0}$ in Lemma 19 to $\mu$ and $\nu$, respectively, obtaining contraction along the forward heat flow as follows: Substituting the $q$-Rényi into Lemma 19, we have, from the definition of the Rényi divergence as $\mathcal{R}_{q}(\mu\mathbin{\|}\nu):=\frac{1}{q-1}\log\mathopen{}\mathclose{{}\left(D_{f}(\mu\mathbin{\|}\nu)+1}\right),$ with $f(x)=x^{q}-1$ and $g(x)=\frac{1}{q-1}\log(x+1)$,
$\partial_{t}\mathcal{R}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})$ $=-\frac{q}{2}\,\frac{\mathbb{E}_{\mu P_{t}}\Bigl{[}\Bigl{\langle}\nabla\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q-1},\nabla\log\frac{\mu P_{t}}{\nu P_{t}}\Bigr{\rangle}\Bigr{]}}{(q-1)\,\mathbb{E}_{\nu P_{t}}\Bigl{[}\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q}\Bigr{]}}\underset{(i)}{=}-\frac{q}{2}\,\frac{\mathbb{E}_{\mu P_{t}}\Bigl{[}\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q-2}\langle\nabla\frac{\mu P_{t}}{\nu P_{t}},\nabla\log\frac{\mu P_{t}}{\nu P_{t}}\rangle\Bigr{]}}{\mathbb{E}_{\nu P_{t}}\Bigl{[}\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q}\Bigr{]}}$
=$\underset{(ii)}{=}-\frac{q}{2}\,\frac{\mathbb{E}_{\nu P_{t}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{\mu P_{t}}{\nu P_{t}}}\right)^{q}\|\nabla\log\frac{\mu P_{t}}{\nu P_{t}}\|^{2}}\right]}{\mathbb{E}_{\nu P_{t}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{\mu P_{t}}{\nu P_{t}}}\right)^{q}}\right]}=-\frac{1}{2}\,\mathsf{RF}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})\,,$
where in $(i)$, we use again that $\nabla\bigl{[}f^{\prime}\bigl{(}\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\bigr{]}=\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\cdot\frac{\mu_{t}}{\nu_{t}}+f^{\prime}\bigl{(}\frac{\mu_{t}}{\nu_{t}}\bigr{)}\nabla\frac{\mu_{t}}{\nu_{t}}$, and $(ii)$ uses that $\nabla\frac{\mu P_{t}}{\nu P_{t}}=\frac{\mu P_{t}}{\nu P_{t}}\nabla\log\frac{\mu P_{t}}{\nu P_{t}}$, and the last equality recalls the definition of the Rényi Fisher information. This yields
$\partial_{t}\mathcal{R}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})=-\frac{1}{2}\mathsf{RF}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})\underset{(i)}{\leq}-\frac{1}{q}\frac{\mathcal{R}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})}{C_{\mathsf{LSI}}(\nu P_{t})}\underset{(ii)}{\leq}-\frac{1}{q}\frac{\mathcal{R}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})}{C_{\mathsf{LSI}}+t}\,,$
where we used Lemma 17 in $(i)$ and Lemma 20 in $(ii)$. Applying Grönwall’s inequality (Lemma 10),
$\mathcal{R}_{q}(\mu P_{h}\mathbin{\|}\nu P_{h})\leq\exp\Bigl{(}-\frac{1}{q}\int_{0}^{h}\frac{1}{C_{\mathsf{LSI}}+t}\,\mathrm{d}t\Bigr{)}\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq\frac{\mathcal{R}_{q}(\mu\mathbin{\|}\nu)}{(1+h/C_{\mathsf{LSI}})^{1/q}}\,.$
Since the SDE (3.1) also captures the backward equation ($\mathsf{BH}$), we set $\mu_{0}$ and $\nu_{0}$ in Lemma 19 to $\mu P_{h}$ and $\tilde{\nu}:=\nu P_{h}$ respectively, obtaining contraction along the backward heat flow:
$\partial_{t}\mathcal{R}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})=-\frac{1}{2}\,\mathsf{RF}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})\leq-\frac{1}{q}\,\frac{\mathcal{R}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})}{C_{\mathsf{LSI}}(\tilde{\nu}Q_{t})}\underset{(i)}{\leq}-\frac{1}{q}\,\frac{\mathcal{R}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})}{C_{\mathsf{LSI}}+{h-t}}\,,$
where $(i)$ follows from that $\tilde{\nu}Q_{t}=\nu P_{h}Q_{t}=\nu P_{h-t}$ and $C_{\mathsf{LSI}}(\tilde{\nu}Q_{t})\leq C_{\mathsf{LSI}}+h-t$ due to Lemma 20. Applying Lemma 10 again yields
$\mathcal{R}_{q}(\mu P_{h}Q_{h}\mathbin{\|}\nu)\leq\frac{\mathcal{R}_{q}(\mu P_{h}\mathbin{\|}\tilde{\nu})}{(1+h/C_{\mathsf{LSI}})^{1/q}}\,.$
Composing these two inequalities leads to the decay rate claimed in the theorem.
The result in the $\chi^{2}$-divergence can be derived entirely analogously. For instance, the decay from the forward part can be shown as follows:
$\partial_{t}\chi^{2}(\mu P_{t}\mathbin{\|}\nu P_{t})=-\frac{1}{2}\,\mathbb{E}_{\nu P_{t}}\Bigl{[}\Bigl{\|}\nabla\frac{\mu P_{t}}{\nu P_{t}}\Bigr{\|}^{2}\Bigr{]}\underset{(i)}{\leq}-\frac{\chi^{2}(\mu P_{t}\mathbin{\|}\nu P_{t})}{C_{\mathsf{PI}}(\nu P_{t})}\leq-\frac{\chi^{2}(\mu P_{t}\mathbin{\|}\nu P_{t})}{C_{\mathsf{PI}}+t}\,,$
where $(i)$ follows from Lemma 15. Applying Grönwall’s inequality then gives
$\chi^{2}(\mu P_{h}\mathbin{\|}\nu P_{h})\leq\exp\Bigl{(}-\int_{0}^{h}\frac{1}{C_{\mathsf{PI}}+t}\,\mathrm{d}t\Bigr{)}\,\chi^{2}(\mu\mathbin{\|}\nu)\leq\frac{\chi^{2}(\mu\mathbin{\|}\nu)}{1+h/C_{\mathsf{PI}}}\,.$
The decay along the backward heat flow in $\chi^{2}$ is entirely analogous to the Rényi case. Then we combine two contraction results from the forward and backward flows, completing the proof.
The result in the $\mathcal{R}_{q}$ under (7) can be shown in a similar manner. Only difference is that in forward and backward computations, one should use the functional inequality in Lemma 16 and the following standard inequalities:
$1-\exp\bigl{(}-\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\bigr{)}\geq\begin{cases}\frac{1}{2}&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\geq 1\,,\\ \frac{1}{2}\mathcal{R}_{q}(\mu\mathbin{\|}\nu)&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq 1\,.\end{cases}$
Part II: Extension to constrained distributions.
We now prove a technical lemma that extends the contraction results to constrained distributions. This lemma guarantees the existence of weak solutions to two stochastic processes that describe the evolution of distributions involved in Line 2 and 3 in $\mathsf{In\text{-}and\text{-}Out}$, in addition to lower-semicontinuity of $f$-divergence. We shall prove it for any measure that is absolutely continuous with respect to $\pi^{X}$, since this imposes no additional technical hurdles.
Lemma 22.
Let $\nu$ be a measure, absolutely continuous with respect to the uniform measure $\pi^{X}$. The forward and backward heat flow equations given by
$\partial_{t}\mu_{t}$ $=\frac{1}{2}\Delta\mu_{t}\,,$
$\partial_{t}\mu_{t}^{\leftarrow}$ $=-\operatorname{div}\bigl{(}\mu_{t}^{\leftarrow}\nabla\log(\nu P_{h-t})\bigr{)}+\frac{1}{2}\Delta\mu_{t}^{\leftarrow}\quad\text{with }\mu_{0}^{\leftarrow}=\mu_{h}\,,$
admit solutions on $(0,h]$, and the weak limit $\lim_{t\to h}\mu_{t}^{\leftarrow}=\mu_{h}^{\leftarrow}$ exists for any initial measure $\mu_{0}$ with bounded support. Moreover, for any $f$-divergence with $f$ lower semi-continuous,
$D_{f}(\mu_{h}^{\leftarrow}\mathbin{\|}\nu)\leq\lim_{t\downarrow 0}D_{f}(\mu_{h-t}^{\leftarrow}\mathbin{\|}\nu_{t})\,.$
Proof.
The existence of weak solutions for the forward equation is well-known, since $\mu_{0}$ can be weakly approximated by measures with continuous density, for which the heat equation admits a unique solution for all time. In particular, the weak solution is $C^{\infty}$ for $t>0$.
The reverse SDE is more subtle, since $\nabla\log\nu P_{t}$ will in general cease to be Lipschitz as $t\to 0$. On the other hand, for any $h>0$, we can write explicitly
$\mu_{h}(x)=\frac{1}{(2\pi h)^{d/2}}\int\exp\bigl{(}-\frac{\|x-y\|^{2}}{2h}\bigr{)}\,\mathrm{d}\mu_{0}(y)\,.$
If one considers the system started at $\tilde{\mu}_{0}=\mu_{\epsilon}=\nu P_{\epsilon}$ and solve the forward-backward Fokker-Planck equations on times $[0,h-\epsilon]$, then $\tilde{\mu}_{h-\epsilon}=\mu_{h}=\mu_{0}^{\leftarrow}=\tilde{\mu}_{0}^{\leftarrow}$ and
$\mu_{h-\epsilon}^{\leftarrow}(x)=\tilde{\mu}_{h-\epsilon}^{\leftarrow}(x)=\int\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(x)}{\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)\,.$
This follows from that if we consider system started at time $\epsilon>0$, with initial distribution $\mu_{\epsilon}$, then we obtain the above through the Bayesian perspective on the forward and reverse heat semigroups, elaborated in Appendix D.
We now show that the following integral is indeed integrable, so $\tilde{\mu}_{h}^{\leftarrow}$ is well-defined:
$\tilde{\mu}_{h}^{\leftarrow}(x):=\int\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2h}\bigr{)}\nu(x)}{\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)\,.$
For fixed $x$ and $\epsilon<h/2$,
$\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2(h-\epsilon)}\bigr{)}\nu(z)\,\mathrm{d}z\gtrsim\exp\bigl{(}-\frac{(\|y-x_{0}\|+D)^{2}}{2(h-\epsilon)}\bigr{)}\,,$
as the support of $\nu$ is constrained to $\mathcal{K}\subset B_{D}(x_{0})$. Since $\mu_{0}$ has bounded support, $\mu_{h}(y)\lesssim\exp(-\frac{\|y\|^{2}}{a})$ for some constant $a>0$. Thus,
$\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2(h-\epsilon)}\bigr{)}\,\mu_{h}(y)}{\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z}\lesssim\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2(h-\epsilon)}\bigr{)}\,\mu_{h}(y)}{\exp(-\frac{(\|y-x_{0}\|+D)^{2}}{2(h-\epsilon)})}\lesssim\exp\Bigl{(}\frac{\langle 2(x-x_{0}),y\rangle+2D\|y-x_{0}\|}{h}-\frac{\|y\|^{2}}{a}\Bigr{)}\,,$
and the last bound is integrable in $y$.
We then show the pointwise convergence of $\tilde{\mu}_{h-\epsilon}^{\leftarrow}$ to $\tilde{\mu}_{h}^{\leftarrow}$ as $\epsilon\to 0$. Note that $\nu P_{\epsilon}\to\nu$, as $\nu$ has a bounded support. Also, the denominator is independent of $\epsilon$ due to
$\frac{1}{(2\pi(h-\epsilon))^{d/2}}\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z=\mathcal{N}(0,(h-\epsilon)I)*\nu P_{\epsilon}=\nu*\mathcal{N}(0,hI)\,.$
Hence, for $\epsilon\leq d^{-1}$,
$\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2(h-\epsilon)}\bigr{)}}{\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z}\leq\Bigl{(}\frac{h}{h-\epsilon}\Bigr{)}^{d/2}\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2h}\bigr{)}}{\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}\lesssim\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2h}\bigr{)}}{\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}\,.$
As shown above, the last bound is integrable with respect to $\mu_{h}$, so the dominated convergence theorem implies
$\lim_{\epsilon\to 0}\int\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2(h-\epsilon)}\bigr{)}}{\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)=\int\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2h}\bigr{)}}{\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)\,,$
Thus, the pointwise convergence follows. Note that if we take $\nu(x)=\pi^{X}(x)=\frac{\mathds{1}_{\mathcal{K}}(x)}{\operatorname{vol}(\mathcal{K})}$, then $\tilde{\mu}_{h}^{\leftarrow}$ is the distribution of the backwards step of our algorithm. In particular, this corresponds to first sampling $x\sim\mu_{h}$, then $y\sim Q_{h}^{\nu,h}(\cdot|x)$, which is precisely the law of $\mu_{h}^{\leftarrow}$ given by ($\mathsf{FP\text{-}BH}$).
As for the second statement, it follows from Scheffé’s lemma [billingsley2017probability, Theorem 16.12] that the pointwise convergence of $\mu_{h-\varepsilon}^{\leftarrow}\to\mu_{h}^{\leftarrow}$ leads to its $\mathsf{TV}$-convergence, which in turn implies the weak convergence. It follows from lower semicontinuity of $D_{f}$ [ambrosio2000functions, Theorem 2.34] that the weak convergence ensures $D_{f}(\mu_{h}^{\leftarrow}\mathbin{\|}\nu)\leq\lim_{t\downarrow 0}D_{f}(\mu_{h-t}^{\leftarrow}\mathbin{\|}\nu^{\leftarrow}_{h-t})$. ∎
In the sequel, we will only consider $\nu=\pi^{X}$. Since the Rényi divergence is a continuous function of the $\chi^{q}$ divergence (see Definition 6), which itself is an $f$-divergence, it enjoys the same lower-semicontinuity properties. Using this lower-semicontinuity together with the decay results in Theorem 21, we can easily derive the contraction results of $\mathsf{In\text{-}and\text{-}Out}$ in $\mathcal{R}_{q}$ and $\chi^{q}$ for any $q\geq 1$. We remark that this result does not require convexity of $\mathcal{K}$.
Theorem 23.
Let $\mu^{X}_{k}$ be the law of the $k$-th output of $\mathsf{In\text{-}and\text{-}Out}$ with initial distribution $\mu_{0}^{X}$. Let $C_{\mathsf{LSI}}$ be the (8) constant of the uniform distribution $\pi^{X}$ over $\mathcal{K}$. Then, for any $q\geq 1$,
$\mathcal{R}_{q}(\mu_{k}^{X}\mathbin{\|}\pi^{X})\leq\frac{\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})}{(1+h/C_{\mathsf{LSI}})^{2k/q}}\,.$
For $C_{\mathsf{PI}}$ the (7) constant of $\pi^{X}$,
$\chi^{2}(\mu^{X}_{k}\mathbin{\|}\pi^{X})\leq\frac{\chi^{2}(\mu^{X}_{0}\mathbin{\|}\pi^{X})}{(1+h/C_{\mathsf{PI}})^{2k}}\,.$
Furthermore, for any $q\geq 2$,
$\mathcal{R}_{q}(\mu_{k}^{X}\mathbin{\|}\pi^{X})\leq\begin{cases}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-\frac{2k\log(1+h/C_{\mathsf{PI}})}{q}&\text{if }k\leq\frac{q}{2\log(1+h/C_{\mathsf{PI}})}\,\bigl{(}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-1\bigr{)}\,,\\ (1+h/C_{\mathsf{PI}})^{-2(k-k_{0})/q}&\text{if }k\geq k_{0}:=\lceil\frac{q}{2\log(1+h/C_{\mathsf{PI}})}\,\bigl{(}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-1\bigr{)}\rceil\,.\end{cases}$
Proof.
Let us set $\mu_{0}=\mu_{0}^{X}$ and $\pi_{0}=\pi^{X}$. Then, $\mu_{h}=\mu_{0}^{\leftarrow}=\mu_{1}^{Y}$, $\pi_{h}=\pi_{0}^{\leftarrow}=\pi^{Y}$, and $\mu_{h}^{\leftarrow}=\mu_{1}^{X}$, $\pi_{h}^{\leftarrow}=\pi^{X}$. For small $\epsilon>0$, as $\mu_{\epsilon}=(\mu_{0}^{X})_{\epsilon}=\mu_{0}^{X}*\mathcal{N}(0,\epsilon I_{d})$ is $C^{\infty}$-smooth, we can now invoke the decay results with step size $h-\epsilon$ in Theorem 21. Thus, for contraction constants $C_{\epsilon}=(1+\frac{h-\epsilon}{C_{\mathsf{LSI}}+\epsilon})^{-2/q}$ and $C_{\epsilon}=(1+\frac{h-\epsilon}{C_{\mathsf{PI}}+\epsilon})^{-2}$ respectively when $\Phi=\mathcal{R}_{q}$ and $\Phi=\chi^{2}$,
$\Phi(\mu_{h-\epsilon}^{\leftarrow}\mathbin{\|}\pi_{\epsilon})\leq C_{\epsilon}\cdot\Phi(\mu_{\epsilon}\mathbin{\|}\pi_{\epsilon})\leq C_{\epsilon}\cdot\Phi(\mu_{0}\mathbin{\|}\pi_{0})\,,$
where we used the data-processing inequality for the last inequality. By the second result of Lemma 22, sending $\epsilon\to 0$ leads to
$\Phi(\mu_{1}^{X}\mathbin{\|}\pi^{X})=\Phi(\mu_{h}^{\leftarrow}\mathbin{\|}\pi_{0})\leq C\cdot\Phi(\mu_{0}\mathbin{\|}\pi_{0})=C\cdot\Phi(\mu_{0}^{X}\mathbin{\|}\pi^{X})\,.$
Repeating this argument $k$ times completes the proof. ∎
3.3 Failure probability and wasted steps
We begin by defining a suitable version of local conductance [kannan1997random].
Definition 24 (Local conductance).
The local conductance $\ell$ on $\mathbb{R}^{d}$ is defined by
$\ell(x)\stackrel{{\scriptstyle\mathrm{{\scriptscriptstyle def}}}}{{=}}\frac{\int_{\mathcal{K}}\exp(-\frac{1}{2h}\|x-y\|^{2})\,\mathrm{d}y}{\int_{\mathbb{R}^{d}}\exp(-\frac{1}{2h}\|x-y\|^{2})\,\mathrm{d}y}=\frac{\int_{\mathcal{K}}\exp(-\frac{1}{2h}\|x-y\|^{2})\,\mathrm{d}y}{(2\pi h)^{d/2}}\,.$
The local conductance at $y$ quantifies the success probability of the proposal at $y$ in Line 3. Then the expected number of trials until the first success of Line 3 is $1/\ell(y)$. Revisiting (1.1), we can notice $\pi^{Y}(y)=\ell(y)/\operatorname{vol}(\mathcal{K})$.
Naïve analysis for expected number of trials.
Starting from $\pi^{X}$, when we just naïvely sample from $\pi^{Y|X}(\cdot|x)$ for all $x$ without imposing any failure condition, the expected number of trials for one iteration is that for the probability density $p_{x}$ of $\mathcal{N}(x,hI_{d})$,
$\int_{\mathcal{K}}\int_{\mathbb{R}^{d}}\frac{1}{\ell(y)}\,p_{x}(\mathrm{d}y)\pi^{X}(\mathrm{d}x)=\int_{\mathbb{R}^{d}}\frac{1}{\ell(y)}\,\pi^{Y}(\mathrm{d}y)=\int_{\mathbb{R}^{d}}\frac{1}{\ell(y)}\,\frac{\ell(y)}{\operatorname{vol}(\mathcal{K})}\,\mathrm{d}y=\infty\,.$
This suggests that one should consider the algorithm as having “failed” if the number of trials exceeds some threshold.
Refined analysis under a failure condition.
Going forward, we assume an $M$-warm start as in previous work for uniform sampling algorithms. By induction we have $\frac{\mathrm{d}\mu^{X}_{i}}{\mathrm{d}\pi^{X}}\leq M$ for all $i$.
Lemma 25 (Propagation of warm-start).
From an $M$-warm start, we have $\nicefrac{{\mathrm{d}\mu_{i}^{X}}}{{\mathrm{d}\pi^{X}}}\leq M$ for all $i$.
Proof.
Assume that $\mu_{i}^{X}$ satisfies the $M$-warm start. Then, for any measurable $S$ and the transition kernel $T_{x}$ of Algorithm 1 at $x$,
$\mu_{i+1}^{X}(S)=\int_{\mathcal{K}}T_{x}(S)\,\mathrm{d}\mu_{i}^{X}(x)\leq M\int_{\mathcal{K}}T_{x}(S)\,\mathrm{d}\pi^{X}(x)=M\pi^{X}(S)\,,$
where the last equality follows from the stationarity of $\pi$. Hence, $\mathrm{d}\mu_{i+1}^{X}/\mathrm{d}\pi^{X}\leq M$. ∎
We now establish a lemma that comes in handy when analyzing the failure probability of the algorithm. In essence, this lemma bounds the probability that taking a Gaussian step from $\pi^{X}$ in Line 2 gets $\delta$-distance away from $\mathcal{K}$. Let us denote the $\delta$-blowup of $\mathcal{K}$ by $\mathcal{K}_{\delta}:=\{x\in\mathbb{R}^{d}:d(x,\mathcal{K})\leq\delta\}$.
Lemma 26.
$\pi^{Y}(\mathcal{K}_{\delta}^{c})\leq e^{c_{1}^{2}}\exp(-\nicefrac{{t^{2}}}{{8c_{1}^{2}}})$ for $\delta=t/d$ and variance $h=c_{1}^{2}/d^{2}$, where $c_{1}$ is some constant and $t\geq 2c_{1}(c_{1}+1)$.
Proof.
For $y\in\partial\mathcal{K}_{\delta}$, we can take the supporting half-space $H$ at $\mathsf{proj}_{\mathcal{K}}(y)$ containing $\mathcal{K}$, due to convexity of $\mathcal{K}$. Then,
$\pi^{Y}(\mathcal{K}_{\delta}^{c})$ $=\frac{1}{\operatorname{vol}(\mathcal{K})}\int_{\mathcal{K}_{\delta}^{c}}\int_{\mathcal{K}}\frac{\exp\bigl{(}-\frac{1}{2h}\|y-x\|^{2}\bigr{)}}{(2\pi h)^{d/2}}\,\mathrm{d}x\,\mathrm{d}y\leq\frac{1}{\operatorname{vol}(\mathcal{K})}\int_{\mathcal{K}_{\delta}^{c}}\int_{H}\frac{\exp\bigl{(}-\frac{1}{2h}\|y-x\|^{2}\bigr{)}}{(2\pi h)^{d/2}}\,\mathrm{d}x\,\mathrm{d}y$
$=\frac{1}{\operatorname{vol}(\mathcal{K})}\int_{\mathcal{K}_{\delta}^{c}}\int_{d(y,\mathcal{K})}^{\infty}\frac{\exp\bigl{(}-\frac{1}{2h}z^{2}\bigr{)}}{\sqrt{2\pi h}}\,\mathrm{d}z\,\mathrm{d}y\,.$ (3.3)
Let us denote the tail probability of the $1$-dimensional Gaussian with variance $h$ by
$\mathsf{T}(s):=\mathbb{P}_{Z\sim\mathcal{N}(0,h)}(Z\geq s)=\frac{1}{\sqrt{2\pi h}}\int_{s}^{\infty}\exp\bigl{(}-\frac{1}{2h}z^{2}\bigr{)}\,\mathrm{d}z\,.$
By the co-area formula and integration by parts,
$\int_{\mathcal{K}_{\delta}^{c}}\int_{d(y,\mathcal{K})}^{\infty}\frac{\exp\bigl{(}-\frac{1}{2h}z^{2}\bigr{)}}{\sqrt{2\pi h}}\,\mathrm{d}z\mathrm{d}y=\int_{\delta}^{\infty}\mathsf{T}(s)\operatorname{vol}(\partial\mathcal{K}_{s})\,\mathrm{d}s$
$=\Bigl{[}\underbrace{\mathsf{T}(s)\int_{0}^{s}\operatorname{vol}(\partial\mathcal{K}_{z})\,\mathrm{d}z}_{\eqqcolon\mathsf{F}}\Bigr{]}_{\delta}^{\infty}+\int_{\delta}^{\infty}\frac{1}{\sqrt{2\pi h}}\exp\bigl{(}-\frac{s^{2}}{2h}\bigr{)}\int_{0}^{s}\operatorname{vol}(\partial\mathcal{K}_{z})\,\mathrm{d}z\,\mathrm{d}s\,.$ (3.4)
Recall that $\mathsf{T}(s)\leq\frac{1}{2}\exp(-\frac{1}{2}(sh^{-1/2})^{2})$ for $s\geq\delta=t/d$ due to a standard tail bound on a Gaussian distribution. This tail bound, combined with
$\int_{0}^{s}\operatorname{vol}(\partial\mathcal{K}_{z})\,\mathrm{d}z=\operatorname{vol}(\mathcal{K}_{s})-\operatorname{vol}(\mathcal{K})\leq\operatorname{vol}\bigl{(}(1+s)\,\mathcal{K}\bigr{)}-\operatorname{vol}(\mathcal{K})=\bigl{(}(1+s)^{d}-1\bigr{)}\operatorname{vol}(\mathcal{K})\,,$
ensures that $\mathsf{F}$ vanishes at $s=\infty$. Hence, bounding the first term in (3.4) by 00 results in
$\int_{\mathcal{K}_{\delta}^{c}}\int_{d(y,\mathcal{K})}^{\infty}\frac{\exp\bigl{(}-\frac{1}{2h}z^{2}\bigr{)}}{\sqrt{2\pi h}}\,\mathrm{d}z\,\mathrm{d}y$ $\leq\frac{1}{\sqrt{2\pi h}}\int_{\delta}^{\infty}\exp\bigl{(}-\frac{s^{2}}{2h}\bigr{)}\bigl{(}\underbrace{(1+s)^{d}}_{\leq\exp(sd)}-1\bigr{)}\operatorname{vol}(\mathcal{K})\,\mathrm{d}s$
$\leq\frac{\operatorname{vol}(\mathcal{K})}{\sqrt{2\pi h}}\exp(hd^{2}/2)\int_{\delta}^{\infty}\exp\Bigl{(}-\frac{1}{2h}\,(s-hd)^{2}\Bigr{)}\,\mathrm{d}s$
$\underset{(i)}{\leq}\operatorname{vol}(\mathcal{K})\exp(hd^{2}/2)\exp\bigl{(}-\frac{1}{2}\,(\delta h^{-1/2}-dh^{1/2})^{2}\bigr{)}$
$\underset{(ii)}{\leq}\operatorname{vol}(\mathcal{K})\exp(c_{1}^{2}/2)\exp\bigl{(}-\frac{t^{2}}{8c_{1}^{2}}\bigr{)}\,,$
where in $(i)$ we used the tail bound of a Gaussian, and $(ii)$ follows from that in the regime of $h=c_{1}^{2}d^{-2}$, we have $\delta h^{-1/2}-dh^{1/2}-1\geq t/c_{1}-(c_{1}+1)\geq\nicefrac{{t}}{{2c_{1}}}$ for $t\geq 2c_{1}(c_{1}+1)$. Putting the last bound into (3.3) completes the proof of the claim. ∎
Now we choose a suitable threshold $N$ for bounding the failure probability. Following (3.3) in the proof, one can notice that for $y\in\mathcal{K}^{c}_{\delta}$, $\delta=\Omega(1/d)$, and $h=\Theta(d^{-2})$,
$\ell(y)\leq\int_{d(y,\mathcal{K})}^{\infty}\frac{\exp\bigl{(}-\frac{1}{2h}z^{2}\bigr{)}}{\sqrt{2\pi h}}\,\mathrm{d}z=\mathbb{P}_{Z\sim\mathcal{N}(0,h)}(Z\geq\delta)\leq\exp(-\Omega(t^{2}))\,.$
Thus, the expected number of trials from $\mathcal{K}^{c}_{\delta}$ for the rejection sampling in Line 3 is $\ell(y)^{-1}\geq\exp(\Omega(t^{2}))$. Intuitively, one can ignore whatever happens in $\mathcal{K}^{c}_{\delta}$, since $\mathcal{K}_{\delta}$ takes up most of measure of $\pi^{Y}$. As the number of trials from $\mathcal{K}_{\delta}^{c}$ is at least $\exp(\Omega(t^{2}))$ in expectation, the most straightforward way to ignore algorithmic behaviors from $\mathcal{K}_{\delta}^{c}$ is simply to set the threshold to $N=\widetilde{\mathcal{O}}(\exp(t^{2}))$. Even though the threshold is $N$, the expected number of trials is much lower.
Lemma 14 bounds the failure probability and expected number of trials per iteration.
Proof of Lemma 14.
For $\mu_{h}:=\mu*\mathcal{N}(0,hI_{d})$, the failure probability is $\mathbb{E}_{\mu_{h}}[(1-\ell)^{N}]$. Since $\mathrm{d}\mu/\mathrm{d}\pi^{X}\leq M$ implies $\mathrm{d}\mu_{h}/\mathrm{d}(\pi^{X})_{h}=\mathrm{d}\mu_{h}/\mathrm{d}\pi^{Y}\leq M$, it follows that
$\mathbb{E}_{\mu_{h}}[(1-\ell)^{N}]\leq M\,\mathbb{E}_{\pi^{Y}}[(1-\ell)^{N}]\,.$
Then,
$\int_{\mathbb{R}^{d}}\underbrace{(1-\ell)^{N}\,\mathrm{d}\pi^{Y}}_{\eqqcolon\mathsf{A}}$ $=\int_{\mathcal{K}_{\delta}^{c}}\mathsf{A}+\int_{\mathcal{K}_{\delta}\cap[\ell\geq N^{-1}\log(3mM/\eta)]}\mathsf{A}+\int_{\mathcal{K}_{\delta}\cap[\ell<N^{-1}\log(3mM/\eta)]}\mathsf{A}$
$\leq\pi^{Y}(\mathcal{K}_{\delta}^{c})+\int_{[\ell\geq N^{-1}\log(3mM/\eta)]}\exp(-\ell N)\,\mathrm{d}\pi^{Y}+\int_{\mathcal{K}_{\delta}\cap[\ell<N^{-1}\log(3mM/\eta)]}\frac{\ell(y)}{\operatorname{vol}(\mathcal{K})}\,\mathrm{d}y$
$\leq e^{c_{1}^{2}}\exp(-\frac{t^{2}}{8c_{1}^{2}})+\frac{\eta}{3mM}+\frac{\log(3mM/\eta)}{N}\,\frac{\operatorname{vol}(\mathcal{K}_{\delta})}{\operatorname{vol}(\mathcal{K})}$
$\leq e^{c_{1}^{2}}\exp(-\frac{t^{2}}{8c_{1}^{2}})+\frac{\eta}{3mM}+\frac{e^{t}}{N}\,\log\frac{3mM}{\eta}\,,$
where we used $\operatorname{vol}(\mathcal{K}_{\delta})\subset\operatorname{vol}\bigl{(}(1+\delta)\mathcal{K}\bigr{)}=(1+\delta)^{d}\operatorname{vol}(\mathcal{K})\leq e^{t}\operatorname{vol}(\mathcal{K})$. Taking $c_{1}^{2}=\frac{\log\log Z}{2\log Z}$, $t=\sqrt{8}\log\log Z$, and $N=Z(\log Z)^{4}$, we can bound the last line by $\frac{\eta}{mM}$. Therefore,
$\mathbb{E}_{\mu_{h}}[(1-\ell(\cdot))^{N}]\leq M\,\mathbb{E}_{\pi^{Y}}[(1-\ell(\cdot))^{N}]\leq\frac{\eta}{m}\,.$
We now bound the expected number of trials per iteration. Let $X$ be the minimum of the threshold $N$ and the number of trials until the first success. Then the expected number of trials per step is bounded by $M\mathbb{E}_{\pi^{Y}}[X]$ due to $\mathrm{d}\mu_{h}/\mathrm{d}\pi^{Y}\leq M$. Thus,
$\int_{\mathbb{R}^{d}}\Bigl{(}\frac{1}{\ell}\wedge N\Bigr{)}\,\mathrm{d}\pi^{Y}$ $\leq\int_{\mathcal{K}_{\delta}}\frac{1}{\ell}\,\mathrm{d}\pi^{Y}+N\pi^{Y}(\mathcal{K}_{\delta}^{c})=\frac{\operatorname{vol}(\mathcal{K}_{\delta})}{\operatorname{vol}(\mathcal{K})}+N\pi^{Y}(\mathcal{K}_{\delta}^{c})$
$\leq e^{t}+N\exp(-\Omega(t^{2}))\leq(\log Z)^{3}+3(\log Z)^{4}=\mathcal{O}\Bigl{(}\bigl{(}\log\frac{mM}{\eta}\bigr{)}^{4}\Bigr{)}\,.$
Therefore, the expected number of trials per step is $\mathcal{O}\bigl{(}M(\log\frac{mM}{\eta})^{4}\bigr{)}$, and the claim follows since each trial uses one query to the membership oracle of $\mathcal{K}$. ∎
3.4 Putting it together
We can now show that $\mathsf{In\text{-}and\text{-}Out}$ subsumes previous results on uniform sampling from convex bodies (such as $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$), providing detailed versions of the main results in Section 1.2.
We first establish that the query complexity of $\mathsf{In\text{-}and\text{-}Out}$ matches that of the $\mathsf{Ball\ walk}$ under stronger divergences. Recall that $2\|\cdot\|_{\mathsf{TV}}^{2}\leq\mathsf{KL}\leq\log(1+\chi^{2})\leq\chi^{2}$.
Theorem 27.
For any given $\eta,\varepsilon\in(0,1)$, $q\geq 1$, $m\in\mathbb{N}$ defined below and any convex body $\mathcal{K}$ given by a well-defined membership oracle, consider $\mathsf{In\text{-}and\text{-}Out}$ (Algorithm 1) with an $M$-warm initial distribution $\mu_{0}^{X}$, $h=(2d^{2}\log\frac{9mM}{\eta})^{-1}$, and $N=\widetilde{\mathcal{O}}(\frac{mM}{\eta})$. For $\pi^{X}$ the uniform distribution over $\mathcal{K}$,
  • It achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{\mathcal{O}}(qd^{2}\|\mathrm{Cov}(\pi^{X})\|_{\mathsf{op}}\log\frac{M}{\eta\varepsilon})$ iterations. With probability $1-\eta$, the algorithm iterates this many times without failure, using $\widetilde{\mathcal{O}}(qMd^{2}\|\mathrm{Cov}(\pi^{X})\|_{\mathsf{op}}(\log\frac{1}{\eta\varepsilon})^{5})$ expected number of membership queries in total.
  • For isotropic $\pi^{X}$, with probability $1-\eta$, the algorithm achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ with $m=\widetilde{\mathcal{O}}(qd^{2}\log\frac{M}{\eta\varepsilon})$ iterations, using $\widetilde{\mathcal{O}}(qMd^{2}(\log\frac{1}{\eta\varepsilon})^{5})$ membership queries in expectation.
Proof.
We just put together Lemma 14 and Theorem 23. For target accuracy $\varepsilon>0$, we use the $\mathcal{R}_{q}$-decay under (7) for $q\geq 2$ in Theorem 23. The $M$-warm start assumption guarantees $\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})\lesssim\log M$. Due to $C_{\mathsf{PI}}(\pi^{X})=\mathcal{O}(\|\mathrm{Cov}(\pi^{X})\|_{\mathsf{op}}\,\log d)$ (Lemma 18), $\mathsf{In\text{-}and\text{-}Out}$ can achieve $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{\mathcal{O}}(qd^{2}\|\mathrm{Cov}(\pi^{X})\|_{\mathsf{op}}\log\frac{M}{\eta\varepsilon})$ iterations. Since each iteration has $\eta/m$-failure probability by Lemma 14, the union bound ensures that the total failure probability is at most $\eta$ throughout $m$ iterations. Lastly, each iteration requires $\widetilde{\mathcal{O}}(M(\log\frac{1}{\eta\varepsilon})^{4})$ membership queries in expectation by Lemma 14. Therefore, $\mathsf{In\text{-}and\text{-}Out}$ uses $\widetilde{\mathcal{O}}(qMd^{2}\min(D^{2},\|\mathrm{Cov}(\pi^{X})\|_{\mathsf{op}})(\log\frac{1}{\eta\varepsilon})^{5})$ expected number of membership queries over $m$ iterations. Since $\mathcal{R}_{q}$ is non-decreasing in $q$, we can obtain the desired bound on $\mathcal{R}_{q}$ for $q\in[1,2)$.
For isotropic $\pi^{X}$, we have $\mathrm{Cov}(\pi^{X})=I_{d}$, so the claim immediately follows from $C_{\mathsf{PI}}(\pi^{X})=\mathcal{O}(\log d)$ (see Lemma 18). ∎
We now show that the number of proper steps is bounded as claimed for general non-convex bodies and any feasible start in $\mathcal{K}$. We first establish this result under an $M$-warm start (Theorem 5).
Proof of Theorem 5..
By the Rényi-decay under (8) in Theorem 23, $\mathsf{In\text{-}and\text{-}Out}$ can achieve $\varepsilon$-distance to $\pi^{X}$ after $\mathcal{O}\bigl{(}qh^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{\mathcal{R}_{q}(\mu^{X}_{1}\mathbin{\|}\pi^{X})}{\varepsilon}\bigr{)}$ iterations for $q\geq 1$.
For $q\geq 2$, we use the decay result under (7). In this case, $\mathsf{In\text{-}and\text{-}Out}$ decays under two different rates depending on the value of $\mathcal{R}_{q}(\cdot\mathbin{\|}\pi^{X})$. It first needs $\mathcal{O}(qh^{-1}C_{\mathsf{PI}}(\pi^{X})\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X}))$ iterations until $\mathcal{R}_{q}(\cdot\mathbin{\|}\pi^{X})$ reaches $1$. Then, $\mathsf{In\text{-}and\text{-}Out}$ additionally needs $\mathcal{O}(qh^{-1}C_{\mathsf{PI}}(\pi^{X})\log\frac{1}{\varepsilon})$ iterations, and thus it needs $\mathcal{O}(qh^{-1}C_{\mathsf{PI}}(\pi^{X})\bigl{(}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})+\log\frac{1}{\varepsilon}\bigr{)})$ iterations in total. By substituting $\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})\lesssim\log M$, we complete the proof. ∎
Next, we show that $\mathsf{In\text{-}and\text{-}Out}$ mixes from any start.
Corollary 28.
For any given $\varepsilon\in(0,1)$ and set $\mathcal{K}\subset B_{D}(0)$, $\mathsf{In\text{-}and\text{-}Out}$ with variance $h$ and any feasible start $x_{0}\in\mathcal{K}$ achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{\mathcal{O}}(qh^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{d+D^{2}/h}{\varepsilon})$ iterations.
Proof.
We first bound the warmness of $\mu^{X}_{1}$ w.r.t. $\pi^{X}$ when $\mu^{X}_{0}=\delta_{x_{0}}$. One can readily check that
$\mu_{1}^{X}(x)=\mathds{1}_{\mathcal{K}}(x)\cdot\int\frac{\exp\bigl{(}-\frac{1}{2h}\|y-x\|^{2}\bigr{)}\exp\bigl{(}-\frac{1}{2h}\|y-x_{0}\|^{2}\bigr{)}}{(2\pi h)^{d/2}\int_{\mathcal{K}}\exp\bigl{(}-\frac{1}{2h}\|y-x\|^{2}\bigr{)}\,\mathrm{d}x}\,\mathrm{d}y\,.$
By Young’s inequality, $\|y-x\|^{2}\leq(\|y\|+D)^{2}\leq\frac{3}{2}\|y\|^{2}+3D^{2}$ for $x\in\mathcal{K}$. Hence,
$\int\frac{\exp\bigl{(}-\frac{1}{2h}\|y-x\|^{2}\bigr{)}\exp\bigl{(}-\frac{1}{2h}\|y-x_{0}\|^{2}\bigr{)}}{\int_{\mathcal{K}}\exp\bigl{(}-\frac{1}{2h}\|y-x\|^{2}\bigr{)}\,\mathrm{d}x}\,\mathrm{d}y$
$\leq$ $\frac{\exp(2h^{-1}D^{2})}{\operatorname{vol}(\mathcal{K})}\int\exp\Bigl{(}-\frac{1}{2h}\bigl{(}\|y-x\|^{2}+\|y-x_{0}\|^{2}-\frac{3}{2}\|y\|^{2}\bigr{)}\Bigr{)}\,\mathrm{d}y$
$=$ $\frac{\exp(2h^{-1}D^{2})}{\operatorname{vol}(\mathcal{K})}\int\exp\Bigl{(}-\frac{1}{2h}\bigl{(}\frac{1}{2}\|y-2(x+x_{0})\|^{2}+(\|x\|^{2}+\|x_{0}\|^{2}-2\|x+x_{0}\|^{2})\bigr{)}\Bigr{)}\,\mathrm{d}y$
$\leq$ $\frac{\exp(5h^{-1}D^{2})}{\operatorname{vol}(\mathcal{K})}\int\exp\Bigl{(}-\frac{1}{4h}\|y-2(x+x_{0})\|^{2}\Bigr{)}\,\mathrm{d}y$
$=$ $\frac{\exp(5h^{-1}D^{2})}{\operatorname{vol}(\mathcal{K})}(4\pi h)^{d/2}\,.$
Therefore, $M=\operatorname{ess\,sup}\frac{\mu_{1}^{X}}{\pi^{X}}\leq 2^{d/2}\exp(5h^{-1}D^{2})$. By Theorem 5 under (8), $\mathsf{In\text{-}and\text{-}Out}$ needs $\widetilde{\mathcal{O}}(qh^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{d+D^{2}/h}{\varepsilon})$ iterations. ∎
We then obtain the following corollary for a convex body $\mathcal{K}$.
Corollary 29.
For any given $\varepsilon\in(0,1)$ and convex body $\mathcal{K}\subset B_{D}(0)$, $\mathsf{In\text{-}and\text{-}Out}$ with variance $h$ and an $M$-warm initial distribution achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{\mathcal{O}}(qh^{-1}D^{2}\log\frac{1}{\varepsilon})$ iterations. If $\pi^{X}$ is isotropic, then $\mathsf{In\text{-}and\text{-}Out}$ only needs $\widetilde{\mathcal{O}}(qh^{-1}D\log\frac{d+d^{2}/h}{\varepsilon})$ iterations.
Proof.
For convex $\mathcal{K}$, it follows from Lemma 18 that $C_{\mathsf{LSI}}(\pi^{X})=\mathcal{O}(D^{2})$ and $C_{\mathsf{LSI}}(\pi^{X})=\mathcal{O}(D)$ for isotropic $\mathcal{K}$. The rest of the proof can be completed in a similar way. ∎
For $h=\tilde{\Theta}(d^{-2})$, $\mathsf{In\text{-}and\text{-}Out}$ requires $\widetilde{\mathcal{O}}(qd^{2}D^{2})$ iterations and in particular $\widetilde{\mathcal{O}}(qd^{2}D)$ iteration for isotropic uniform distributions. These results match those of $\mathsf{Speedy\ walk}$ [kannan2006blocking, lee2017eldan] (see Theorem 31).
Acknowledgements.
We are deeply grateful to Andre Wibisono and Sinho Chewi for helpful comments and pointers to the literature for Lemma 19. This work was supported in part by NSF award 210644, NSERC through the CGS-D award, and a Simons Investigator award.
[…]
Appendix A Related work
Sampling from constrained log-concave distributions is a fundamental task arising in many fields. Uniform sampling with convex constraints is its simplest manifestation, which was first studied as a core subroutine for a randomized volume-computation algorithm [dyer1991random]. Since then, this fundamental problem has been studied for over three decades [lovasz1990mixing, lovasz1993random, kannan1997random, lovasz2003hit, lovasz2006hit, bubeck2018sampling, pmlr-v65-brosse17a]. We review these algorithms, grouping them under three categories — geometric random walks, structured samplers, and diffusion-type samplers. Below, $\mathcal{K}$ is convex.
Geometric random walk.
We discuss two geometric random walks – $\mathsf{Ball\ walk}$ [lovasz1993random, kannan1997random] and $\mathsf{Hit\text{-}and\text{-}Run}$ [smith1984efficient, lovasz1999hit].
$\mathsf{Ball\ walk}$ is a simple metropolized random walk; it draws $y$ uniformly at random from a ball of radius $\delta$ centered at a current point $x$, and moves to $y$ if $y\in\mathcal{K}$ and stays at $x$ otherwise. In the literature, $\mathsf{Ball\ walk}$ actually refers to a composite algorithm consisting of [$\mathsf{Speedy\ walk}+$ rejection sampling], where $\mathsf{Speedy\ walk}$ records only the accepted steps of $\mathsf{Ball\ walk}$ (see Appendix B for details). The step size $\delta$ should be set to $\mathcal{O}(d^{-1/2})$ to avoid stepping outside of $\mathcal{K}$. [kannan1997random] showed that $\mathsf{Ball\ walk}$ needs $\widetilde{\mathcal{O}}(Md^{2}D^{2}\log\frac{1}{\varepsilon})$ membership queries to be $\varepsilon$-close to $\pi^{\mathcal{K}}$ in $\mathsf{TV}$, where $D$ is the diameter of $\mathcal{K}$, and the warmness parameter $M$ measures the closeness of the initial distribution to the target uniform distribution $\pi^{\mathcal{K}}$.
$\mathsf{Hit\text{-}and\text{-}Run}$ is another zeroth-order algorithm that needs no step size; it picks a uniform random line $\ell$ passing a current point, and move to a uniform random point on $\ell\cap\mathcal{K}$. [lovasz2006hit] shows that, if we define the second moment as $R^{2}:=\mathbb{E}_{X\sim\pi^{\mathcal{K}}}[\|X-\mathbb{E}[X]\|^{2}]$, then $\mathsf{Hit\text{-}and\text{-}Run}$ requires $\mathcal{O}(d^{2}R^{2}\log\frac{M}{\varepsilon})$ queries. Notably, this algorithm has a poly-logarithmic dependence on $M$ as opposed to $\mathsf{Ball\ walk}$.
Both algorithm are affected by skewed shape of $\mathcal{K}$ (i.e., large $D$ or $R$), so these samplers are combined with pre-processing step called rounding. This procedure finds a linear transformation that makes the geometry of $\mathcal{K}$ less skewed and so more amenable to sampling. In literature, there exists a randomized algorithm [jia2021reducing] that rounds $\mathcal{K}$ and generates a good warm start (i.e., $M=\mathcal{O}(1)$), with $\mathsf{Ball\ walk}$ used as a core subroutine. This algorithm takes up $\widetilde{\mathcal{O}}(d^{3})$ queries in total, and in such position with the good warm start, $\mathsf{Ball\ walk}$ only needs $\widetilde{\mathcal{O}}(d^{2}\log\frac{1}{\varepsilon})$ queries to sample from $\pi^{\mathcal{K}}$.
Structured samplers.
The aforementioned samplers based on geometric random walks require only access to the membership oracle of the convex body without any additional structural assumptions. The alternate paradigm of geometry-aware sampling attempts to exploit the structure of convex constraints, with the aim of expediting the convergence of the resultant sampling schemes. One common assumption is to make available a self-concordant barrier function $\phi$ which has regularity on its high-order derivatives and blows up when approaching the boundary $\partial\mathcal{K}$. The Hessian of $\phi$ encodes the local geometry of the constraint, and the samplers often work directly with $\nabla^{2}\phi$.
The first canonical example of such a zeroth-order sampler is $\mathsf{Dikin\ walk}$ used when $\mathcal{K}$ is given by $m$ linear constraints [kannan2012random]; it draws a uniform sample from an ellipsoid (characterized by $\nabla^{2}\phi$) of fixed radius around a current point, and is often combined with a Metropolis adjustment. [kannan2012random] shows that $\mathsf{Dikin\ walk}$ mixes in $\mathcal{O}(md\log\frac{M}{\varepsilon})$ steps, although each iteration is slightly more expensive than one membership query. This algorithm requires no rounding, but still needs a good warm-start, which can be achieved by an annealing-type algorithm using $\widetilde{\mathcal{O}}(md)$ iterations of $\mathsf{Dikin\ walk}$ [kook2024gaussian].
$\mathsf{Riemannian\ Hamiltonian\ Monte\ Carlo}$ is a structured sampler that exploits the first-order information of the potential (i.e., $\nabla\log(1/\pi)$) [girolami2011riemann]; its proposal is given as the solution to the Hamilton’s ODE equation, followed by the Metropolis-filter. In the linear-constraint setting above, this sampler requires $\mathcal{O}(md^{2/3}\log\frac{M}{\varepsilon})$ many iterations to achieve $\varepsilon$-close distance to $\pi^{\mathcal{K}}$ [lee2018convergence]. This sampler is further analyzed for practical ODE solvers [kook2022condition] and for more sophisticated self-concordant barriers [gatmiry2023sampling].
Similarly, $\mathsf{Mirror\ Langevin}$ [zhang2020wasserstein, jiang2021mirror, ahn2021efficient, li2022mirror] is a class of algorithms which converts the constrained problem into an unconstrained one obtained by considering the pushforward of the constrained space by $\nabla\phi$. The algorithm can also be metropolized [srinivasan2023fast]. The best known rates for this algorithm are $\widetilde{\mathcal{O}}(d\log\frac{1}{\varepsilon})$ under some strong assumptions on $\phi$.
Diffusion-type samplers.
Samplers based on discretizations of Itô diffusions, stochastic processes which rapidly mix to $\pi$ in continuous time, have long been used for sampling without constraints [besag1995bayesian, dalalyan2012sparse, dalalyan2017further, chewi2023log]. While the underlying stochastic processes generalize easily to constrained settings, the discretization analysis relies crucially on the smoothness of the target distribution. This is clearly impossible to achieve in the constrained setting, and so some techniques are required to circumvent this difficulty. These algorithms, however, generalize easily to the more general problem of sampling from distributions of the form $\tilde{\pi}^{X}\propto e^{-f}\mathds{1}_{\mathcal{K}}$, by naturally incorporating first order information from $f$.
The first approach for adapting diffusion-based samplers [bubeck2015finite, bubeck2018sampling, lehec2023langevin] iterates a two-step procedure. First, a random step is taken, with $x_{k+1/2}\sim\mathcal{N}(x_{k},2hI_{d})$ for some appropriately chosen step $h$, [^4] A gradient step can be added in the more general case, for sampling from $\tilde{\pi}^{X}$ and then project it to $\mathcal{K}$, i.e., $x_{k+1}=\mathsf{proj}_{\mathcal{K}}(x_{k+1/2})$. The complexity is given in terms of queries to a projection oracle, each call to which can be implemented with a polynomial number of membership oracle queries; a total of $\tilde{\mathcal{O}}(\frac{d^{2}D^{3}}{\varepsilon^{4}})$ queries are needed to be $\varepsilon$-close in $\mathcal{W}_{2}$ to $\pi^{X}$. Another approach, which uses an algorithmically designed “soft” penalty instead of a projection, was proposed in [gurbuzbalaban2022penalized], and achieves a rate estimate of $\tilde{\mathcal{O}}(d/\varepsilon^{10})$.
A second approach, suggested by [pmlr-v65-brosse17a], considers a different proximal scheme, which performs a “soft projection” onto $\mathcal{K}$, by taking steps like $\mathcal{N}((1-h\lambda^{-1})x_{k}+h\mathsf{proj}_{\mathcal{K}}(x_{k}),2hI_{d})$. It is called Moreau-Yosida regularized Langevin, named after an analogous regularization scheme for constrained optimization. This scheme also relies on access to a projection oracle for $\mathcal{K}$, and quantifies their query complexity accordingly. Their final rate estimate is $\tilde{\mathcal{O}}(\frac{d^{5}}{\varepsilon^{6}})$ to be $\varepsilon$-close in $\mathsf{TV}$ distance to $\pi^{X}$.
Observing the prior work integrating diffusion-based sampling with convex constraints, the dependence on the key parameters $d,\varepsilon$, while polynomial, are many orders worse than the rates for zeroth-order samplers such as $\mathsf{Ball\ walk},\mathsf{Hit\text{-}and\text{-}Run}$. In contrast, our analysis not only recovers but in some sense surpasses the known rates for $\mathsf{Ball\ walk},\mathsf{Hit\text{-}and\text{-}Run}$, while harmonizing well with the continuous-time perspective of diffusions.
Proximal schemes for sampling.
The Gibbs sampling scheme used in this paper was inspired by the restricted Gaussian oracle introduced in [lee2021structured] (in turn inspired by Gaussian Cooling [cousins2018gaussian]), which alternately iterates between a pure Gaussian step, and a “proximal” step (which we elaborate in our exposition). This scheme was given novel interpretations by [chen2022improved], which showed that it interpolates the forward and backward heat flows, in the sense defined by [klartag2021spectral]. The backward heat flow itself is intimately related to stochastic localization schemes, invented and popularized in [eldan2013thin, chen2021almost].
This formulation proved surprisingly powerful, allowing many existing rates in unconstrained sampling to be recovered from a relatively simple analysis. This was further extended by [fan2023improved] to achieve the current state-of-the-art rate in unconstrained sampling. Finally, [gopi2023algorithmic] suggest that this could be applied to tackle some constrained problems. However, the assumptions in this final mentioned work are not compatible with the uniform sampling problem on general convex bodies.
Appendix B $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$
We restate the previously known guarantees for uniform sampling by $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$. Below, let $B_{r}(x)$ denote the $d$-dimensional ball of radius $r$ centered at $x$.
Input: initial distribution $\pi_{0}$, convex body $\mathcal{K}\subset\mathbb{R}^{d}$, iterations $T$, step size $\delta>0$.
1:  Sample $x_{0}\sim\pi_{0}$.
2:  for $i=1,\dotsc,T$ do
3:     Sample $y\sim\mathcal{\textnormal{Unif}}(B_{\delta}(x_{i-1}))$.
4:     If $y\in\mathcal{K}$, then $x_{i}\leftarrow y$. Else, $x_{i}\leftarrow x_{i-1}$.
5:  end for
Algorithm 2 $\mathsf{Ball\ walk}$
$\mathsf{Ball\ walk}$ is particularly simple; draw a uniform random point from $B_{\delta}$ around the current point, and go there if the drawn point is inside of $\mathcal{K}$ and stay at the current point otherwise. Its stationary distribution can be easily seen to be $\pi\propto\mathds{1}_{\mathcal{K}}$, the uniform distribution over $\mathcal{K}$.
In the literature, there are two approaches to analyzing the convergence rate of this sampler: (i) a direct analysis via the $s$-conductance of $\mathsf{Ball\ walk}$ and (ii) an indirect approach which first passes through $\mathsf{Speedy\ walk}$.
Direct analysis.
The following $\mathsf{TV}$-guarantee is obtained by lower bounding the $s$-conductance of $\mathsf{Ball\ walk}$, which requires a one-step coupling argument and the Cheeger inequality for $\pi$. We refer interested readers to [vempala2005geometric, Section 5].
Theorem 30 (Convergence of $\mathsf{Ball\ walk}$).
For any $\varepsilon\in(0,1)$ and convex body $\mathcal{K}\subset\mathbb{R}^{d}$ presented by a well-defined membership oracle, let $\pi_{t}$ be the distribution after $t$ steps of $\mathsf{Ball\ walk}$ with an $M$-warm initial distribution $\pi_{0}$. Then, $\mathsf{Ball\ walk}$ with step size $\delta=\Theta(\frac{\varepsilon}{M\sqrt{d}})$ achieves $\|\pi_{t}-\pi\|_{\mathsf{TV}}\leq\varepsilon$ for $t\gtrsim d^{2}D^{2}\frac{M^{2}}{\varepsilon^{2}}\log\frac{M}{\varepsilon}$. If $\pi$ is isotropic, then $\mathsf{Ball\ walk}$ only needs $\mathcal{O}(d^{2}\log d\frac{M^{2}}{\varepsilon^{2}}\log\frac{M}{\varepsilon})$ iterations.
The mixing time of $\mathsf{Ball\ walk}$ under this approach has a polynomial dependence on $1/\varepsilon$, rather than a polylogarithmic dependence.
Indirect analysis through $\mathsf{Speedy\ walk}$.
[kannan1997random] introduced $\mathsf{Speedy\ walk}$, which could be viewed as a version of $\mathsf{Ball\ walk}$ and converges to a speedy distribution (see Proposition 1), which is slightly biased from $\pi$. Then, $\mathsf{Speedy\ walk}$ is used together with another algorithmic component (rejection sampling) [kannan1997random, Algorithm 4.15] that converts the speedy distribution to the uniform distribution. In the literature, $\mathsf{Ball\ walk}$ often refers to ‘$\mathsf{Speedy\ walk}$ combined with the conversion step’, rather than a direct implementation of Algorithm 2. Strictly speaking, a mixing guarantee of this combined algorithm should not be referred to as a provable guarantee of $\mathsf{Ball\ walk}$.
Input: initial distribution $\pi_{0}$, convex body $\mathcal{K}\subset\mathbb{R}^{d}$, iterations $T$, step size $\delta>0$.
1:  Sample $x_{0}\sim\pi_{0}$.
2:  for $i=1,\dotsc,T$ do
3:     Sample $x_{i}\sim\textnormal{Unif}\mathopen{}\mathclose{{}\left(\mathcal{K}\cap B_{\delta}(x_{i-1})}\right)$.
4:  end for
Algorithm 3 $\mathsf{Speedy\ walk}$
As opposed to $\mathsf{Ball\ walk}$, $\mathsf{Speedy\ walk}$ always takes some step at each iteration. However, the problem of sampling from $x_{i}\sim\textnormal{Unif}\mathopen{}\mathclose{{}\left(\mathcal{K}\cap B_{\delta}(x_{i-1})}\right)$ in Line 3 is not straightforward. This step admits the following implementation based on rejection sampling, via a procedure denoted by ()($\ast$)( ∗ ):
  • Propose $y\sim\textnormal{Unif}(B_{\delta}(x_{i-1}))$.
  • Set $x_{i+1}\leftarrow y$ if $y\in\mathcal{K}$. Otherwise, repeat the proposal.
Each actual step (indexed by $i$) in $\mathsf{Speedy\ walk}$ is called a proper step, and rejected steps during ()($\ast$)( ∗ ) are called improper steps. For example, if $x_{1},x_{1},x_{2},x_{3},x_{3},x_{3},x_{4},\dotsc$ are the positions produced by $\mathsf{Ball\ walk}$, then only proper steps $x_{1},x_{2},x_{3},x_{4},\dotsc$ are recorded by $\mathsf{Speedy\ walk}$.
To describe the theoretical guarantees of $\mathsf{Speedy\ walk}$, we define the local conductance $\ell(x)$ at $x\in\mathcal{K}$, which measures the success probability of the rejection sampling scheme in ()($\ast$)( ∗ ):
$\ell(x):=\frac{\operatorname{vol}(\mathcal{K}\cap B_{\delta}(x))}{\operatorname{vol}(B_{\delta}(x))}\,,$
and define the average conductance:
$\lambda:=\mathbb{E}_{\pi}[\ell]=\frac{1}{\operatorname{vol}(\mathcal{K})}\int_{\mathcal{K}}\ell(x)\,\mathrm{d}x\,.$
Proposition 1 ([kannan1997random]).
The stationary distribution $\nu$ of $\mathsf{Speedy\ walk}$ has density
[…]
The speedy distribution $\nu$ is indeed different from the uniform distribution $\pi$, and this discrepancy is quantified in terms of the average conductance.
Proposition 2 ([kannan1997random, Page 22]).
$\|\nu-\pi\|_{\mathsf{TV}}\leq\frac{1-\lambda}{\lambda}$.
One can relate the step size $\delta$ to the average conductance.
Proposition 3 (Bound on average conductance, [kannan1997random, Corollary 4.5]).
$\lambda\geq 1-\frac{\delta\sqrt{d}}{2}$.
The best known result for $\mathsf{Speedy\ walk}$’s mixing is due to [kannan2006blocking] devising the blocking conductance and using the log-Cheeger inequality. When $\nu$ is isotropic (i.e., it has covariance proportional to the identity matrix), [lee2017eldan] improves the mixing bound via the log-Cheeger constant.
Theorem 31 (Mixing of $\mathsf{Speedy\ walk}$).
For any $\varepsilon\in(0,1)$ and convex body $\mathcal{K}\subset\mathbb{R}^{d}$ presented by a well-defined membership oracle, let $\nu_{t}$ be the distribution after $t$ proper steps of $\mathsf{Speedy\ walk}$ started at any feasible point $x_{0}\in\mathcal{K}$. Then, $\mathsf{Speedy\ walk}$ with step size $\delta=\Theta(d^{-1/2})$ achieves $\|\nu_{t}-\nu\|_{\mathsf{TV}}\leq\varepsilon$ for $t\gtrsim(D^{2}+\log(D\sqrt{d}))\,d^{2}\log\frac{1}{\varepsilon}$. From an $M$-warm start, the expected number of improper steps during $t$ iterations is $\widetilde{\mathcal{O}}(tM)$. When $\nu$ is isotropic, $\mathsf{Speedy\ walk}$ needs $\mathcal{O}(d^{2}D\log\frac{1}{\varepsilon}\log\log D)$ proper steps to achieve $\varepsilon$-$\mathsf{TV}$ distance to $\nu$.
Then, [kannan1997random] uses the following post-processing step to obtain an approximately uniform distribution on $\mathcal{K}$, with a provable guarantee.
$\mathcal{A}$: Call $\mathsf{Speedy\ walk}$ to obtain a sample $X\sim\nu_{t}$ until $\frac{2d}{2d-1}\,X\in\mathcal{K}$. If so, return $\bar{X}=\frac{2d}{2d-1}\,X$.
Proposition 4 ([kannan1997random, Theorem 4.16]).
Under the same setting above, assume $\|\nu_{t}-\nu\|_{\mathsf{TV}}\leq\varepsilon$ for step size $\delta\leq(8d\log\frac{d}{\varepsilon})^{-1/2}$ and fixed $t\in\mathbb{N}$. For $\bar{\nu}=\mathsf{law}(\bar{X})$ given by $\mathcal{A}$, it holds that $\|\bar{\nu}-\pi\|_{\mathsf{TV}}\leq\varepsilon$, and the expected number of calls on the conversion algorithm is at most $2$.
Combining the previous two results, we conclude that the total expected number of membership queries to obtain a sample $\varepsilon$-close to $\pi$ in $\mathsf{TV}$ is $\widetilde{\mathcal{O}}(Md^{2}D^{2}\log\frac{1}{\varepsilon})$, which now has a poly-logarithmic dependence on $1/\varepsilon$.
Remark 2 (Backward heat flow analysis of $\mathsf{Speedy\ walk}$).
Consider a Gaussian version of $\mathsf{Speedy\ walk}$, whose one-step corresponds to $x_{i+1}\sim\mathcal{N}(x_{i},hI_{d})|_{\mathcal{K}}$, and this transition kernel exactly matches integrating ($\mathsf{BH}$) for time $h$. Thus, $\nu Q^{\pi^{X},h}_{h}=\nu$ due to the stationarity of $\nu$ under $\mathsf{Speedy\ walk}$, where $Q^{\pi^{X},h}_{h}$ is the transition kernel defined by the backward heat flow for time $h$ that reverses $\pi^{X}*\mathcal{N}(0,hI_{d})$ to $\pi^{X}$. Hence, if we can control the LSI/PI constants of $\nu$ along the backward heat-flow’s trajectory, then we could directly analyze $\mathsf{Speedy\ walk}$ by emulating computations in Lemma 21.
Appendix C Functional inequalities
We provide full details on functional inequalities omitted in Section 3.1. We use $\mu$ and $\mu_{\mathsf{LC}}$ to denote a probability measure and log-concave probability measure over $\mathbb{R}^{d}$, respectively.
Cheeger and PI constants.
The Cheeger isoperimetric constant $C_{\mathsf{Ch}}(\mu)$ measures how large surface area a measurable subset with larger volume has, defined by
$C_{\mathsf{Ch}}(\mu):=\inf_{S\subset\mathbb{R}^{d}}\frac{\mu^{+}(S)}{\min(\mu(S),\mu(S^{c}))}\,,$
where the infimum is taken over all measurable subsets $S$, and $\mu^{+}(S)$ is the Minkowski content of $S$ under $\mu$ defined as, for $S^{\varepsilon}:=\{x\in X:d(x,S)<\varepsilon\}$,
$\mu^{+}(S):=\liminf_{\varepsilon\to 0}\frac{\mu(S^{\varepsilon})-\mu(S)}{\varepsilon}\,.$
[cheeger1970lower] established $C_{\mathsf{PI}}(\mu)\lesssim C_{\mathsf{Ch}}^{-2}(\mu)$ [^5] The opposite direction $C_{\mathsf{PI}}(\mu_{\mathsf{LC}})\gtrsim C_{\mathsf{Ch}}^{-2}(\mu_{\mathsf{LC}})$ also holds for log-concave distributions due to [milman2009role], while $C_{\mathsf{PI}}(\mu)\gtrsim C_{\mathsf{Ch}}^{-2}(\mu)/d$ for general distributions due to [ledoux1994simple]., and then [kannan1995isoperimetric] showed that for covariance matrix $\Sigma_{\mu}:=\mathbb{E}_{\mu}[(\cdot-\mathbb{E}_{\mu}X)(\cdot-\mathbb{E}_{\mu}X)^{\mathsf{T}}]$,
$C_{\mathsf{Ch}}(\mu_{\mathsf{LC}})\gtrsim\frac{1}{(\mathbb{E}_{\mu_{\mathsf{LC}}}[\|X-\mathbb{E}_{\mu_{\mathsf{LC}}}X\|^{2}])^{1/2}}=\frac{1}{(\operatorname{tr}(\Sigma_{\mu_{\mathsf{LC}}}))^{1/2}}\,.$ (C.1)
This immediately leads to $C_{\mathsf{PI}}(\pi)\lesssim(\mathbb{E}_{\pi}[\|X-\mathbb{E}_{\pi}X\|^{2}])^{1/2}\leq D^{2}$ for the uniform distribution $\pi$ over a convex body $\mathcal{K}$ with diameter $D>0$.
Kannan et al. proposed the KLS conjecture in the same paper, which says that for the spectral norm $\|\cdot\|_{2}$,
$C_{\mathsf{Ch}}(\mu_{\mathsf{LS}})\gtrsim\frac{1}{\|\Sigma_{\mu_{\mathsf{LS}}}\|_{2}^{1/2}}\,.$
While the original result in [kannan1995isoperimetric] ensures $C_{\mathsf{Ch}}\gtrsim d^{-1/2}$ for an isotropic log-concave distribution (due to $\Sigma=I_{d}$), this conjecture indeed claims $C_{\mathsf{Ch}}\gtrsim 1$ for such case. Following a line of work [lee2017eldan, chen2021almost, klartag2022bourgain, klartag2023logarithmic], the current bound is
$C_{\mathsf{Ch}}(\mu_{\mathsf{LS}})\gtrsim\frac{(\log d)^{-1/2}}{\|\Sigma_{\mu_{\mathsf{LS}}}\|_{2}^{1/2}}\,,$
which implies that $C_{\mathsf{PI}}(\pi)\lesssim\log d$ when $\pi$ is isotropic for convex $\mathcal{K}$.
Log-Cheeger and LSI constants.
Just as the Cheeger and PI constants are related above, there are known connections between LSI and log-Cheeger constants. The log-Cheeger constant $C_{\mathsf{logCh}}(\mu)$ of a distribution $\mu\in\mathcal{P}(\mathbb{R}^{d})$ is defined as
$C_{\mathsf{logCh}}(\mu):=\inf_{S\subset\mathbb{R}^{d}:\mu(S)\leq\frac{1}{2}}\frac{\mu^{+}(S)}{\mu(S)\sqrt{\log\frac{1}{\mu(S)}}}\,.$
[ledoux1994simple] established that $C_{\mathsf{LSI}}(\mu)\lesssim C_{\mathsf{logCh}}^{-2}(\mu)$ [^6] The opposite direction holds under dimension-scaling due to [ledoux1994simple]: $C_{\mathsf{LSI}}(\mu)\gtrsim C_{\mathsf{logCh}}^{-2}(\mu)/d$., and [kannan2006blocking] showed that any log-concave distributions with support of diameter $D>0$ satisfy $C_{\mathsf{logCh}}(\mu_{\mathsf{LS}})\gtrsim D^{-1}$. Later in 2016, [lee2017eldan] improved this to $C_{\mathsf{logCh}}(\mu_{\mathsf{LS}})\gtrsim D^{-1/2}$ under isotropy. Therefore, for convex $\mathcal{K}$, it follows that $C_{\mathsf{LSI}}(\pi)\lesssim D^{2}$ and that $C_{\mathsf{LSI}}(\pi)\lesssim D$ if $\pi$ is isotropic.
Appendix D The Wasserstein geometry
We present additional technical background on the Wasserstein geometry and Markov semigroup theory. Interested readers can refer to [villani2009optimal, ambrosio2005gradient, chewi2023log] for standard references on Wasserstein spaces and applications to sampling.
Wasserstein gradient.
Let $\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$ be the space of probability measures admitting densities on $\mathbb{R}^{d}$ with finite second moment. Although there are many ways to metrize $\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$, the geometry induced by the Wasserstein-$2$ distance $\mathcal{W}_{2}$ is a particularly useful structure for analysis.
Under the $\mathcal{W}_{2}$-geometry, one can define a “gradient” of a functional defined over $\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$. Specifically, for a functional $\mathcal{F}:\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})\to\mathbb{R}\cup\{\infty\}$, the Wasserstein gradient of $\mathcal{F}$ at $\mu\in\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$ is defined as $\nabla_{\mathcal{W}_{2}}\mathcal{F}(\mu)=\nabla(\delta\mathcal{F})(\mu)\in L^{2}(\mu)$, where $\nabla$ is the standard gradient and $\delta\mathcal{F}$ is the first variation of $\mathcal{F}$ [^7] The first variation can be defined, for any measures $\nu_{0},\nu_{1}\in\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$, as $\lim_{t\downarrow 0}\frac{\mathcal{F}((1-t)\nu_{0}+t\nu_{1})-\mathcal{F}(\nu_{0})}{t}=\langle\delta\mathcal{F}(\nu_{0}),\nu_{1}-\nu_{0}\rangle$. This definition is unique up to an additive constant, which is irrelevant as we are only concerned with its gradient.. Equipped with this $\mathcal{W}_{2}$-gradient, one can define the Wasserstein gradient flow of $\mathcal{F}$ that describes the evolution of a measure $\{\mu_{t}\}_{t\geq 0}$, from some initial measure $\mu_{0}$, as follows:
$\partial_{t}\mu_{t}=\operatorname{div}\bigl{(}\mu_{t}\nabla_{\mathcal{W}_{2}}\mathcal{F}(\mu_{t})\bigr{)}\,.$
More generally, we can identify the Wasserstein “velocity” for some measure $\mu_{t}$ as $v_{t}$ if the time derivative of $\mu_{t}$ can be written in the form
$\partial_{t}\mu_{t}=-\operatorname{div}(\mu_{t}v_{t})\,.$
Under this identification, the time derivative of a functional $\mathcal{F}$ on $\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$ with smooth Wasserstein gradient under these dynamics can be written as
$\partial_{t}\mathcal{F}(\mu_{t})=\mathbb{E}_{\mu_{t}}\langle\nabla_{\mathcal{W}_{2}}\mathcal{F}(\mu_{t}),v_{t}\rangle\,,$
when $v_{t}\in\overline{\{\nabla\psi:\psi\in C_{c}^{\infty}(\mathbb{R}^{d})\}}^{L^{2}(\mu_{t})}$, where {}¯$\overline{\{\cdot\}}^{L^{2}(\mu_{t})}$ denotes the closure of a set with respect to $L^{2}(\mu_{t})$. This is the appropriate notion of tangent space in this geometry.
For instance, when we take the functional to be the entropy of the measure, $\mathcal{H}(\mu)\coloneqq\frac{1}{2}\int\mu\log\mu$, one can verify $\nabla_{\mathcal{W}_{2}}\mathcal{H}(\mu)=\frac{1}{2}\nabla\log\mu$. The heat flow equation can be written as $\partial_{t}\mu_{t}=\frac{1}{2}\Delta\mu_{t}=\frac{1}{2}\operatorname{div}(\nabla\mu_{t})=\frac{1}{2}\operatorname{div}(\mu_{t}\nabla\log\mu_{t})$, which indicates that the velocity of measures $\mu_{t}$ under the heat flow is $v_{t}=-\frac{1}{2}\nabla\log\mu_{t}$. Hence, we can notice that $\nabla_{\mathcal{W}_{2}}\mathcal{H}(\mu_{t})=-v_{t}$, and thus recover the heat flow as the Wasserstein gradient flow of the entropy of the measure.
Fokker-Planck equation and time-reversal of SDE.
Consider a stochastic differential equation $(X_{t})$ given by
$\mathrm{d}X_{t}=-a_{t}(X_{t})\,\mathrm{d}t+\,\mathrm{d}B_{t}\qquad\text{with }X_{0}\sim\mu_{0}\,.$ (D.1)
It is well known that measures $\mu_{t}$ described by
$\partial_{t}\mu_{t}=\mathsf{div}(\mu_{t}a_{t})+\frac{1}{2}\Delta\mu_{t}\,,$ (D.2)
correspond to $\mathsf{law}(X_{t})$. In this context, (D.2) is referred to as the Fokker-Planck equation corresponding to (D.1).
From this equation, one can deduce the Fokker-Planck equation of the time reversal $\mu_{t}^{\leftarrow}:=\mu_{T-t}$:
$\partial_{t}\mu_{t}^{\leftarrow}=-\operatorname{div}(\mu_{t}^{\leftarrow}a_{T-t})-\frac{1}{2}\Delta\mu_{t}^{\leftarrow}=-\operatorname{div}\bigl{(}\mu_{t}^{\leftarrow}(a_{T-t}+\nabla\log\mu_{T-t})\bigr{)}+\frac{1}{2}\Delta\mu_{t}^{\leftarrow}$
In particular, this describes the evolution of $\mathsf{law}(X_{t})$ of the stochastic differential equation:
$\mathrm{d}X_{t}=\bigl{(}a_{T-t}(X_{t})+\nabla\log\mu_{T-t}(X_{t})\bigr{)}\,\mathrm{d}t+\,\mathrm{d}B_{t}\qquad\text{with }X_{0}\sim\mu_{0}^{\leftarrow}=\mu_{T}\,.$ (D.3)
While the law of this process will give $\mu_{T}^{\leftarrow}=\mu_{0}$ at time $T$, it is also true that it will give $\mu_{0|T}(\cdot|z)$ if one starts (D.3) at $X_{0}=z$. This is a subtle fact, whose justification requires the introduction of a tool called Doob’s $h$-transform. The presentation of this subject is beyond the scope of this paper, and we refer interested readers to [klartag2021spectral] as a reference to its application in this context.

### v2

@@
 Generating random samples from a high-dimensional convex body is a basic algorithmic problem with myriad connections and applications.
-The core of the celebrated result of [dyer1991random], giving a randomized polynomial-time algorithm for computing the volume of a convex body, was the first polynomial-time algorithm for uniformly sampling convex bodies.
-In the decades since, the study of sampling has led to a long series of improvements in its algorithmic complexity [lovasz1990mixing, lovasz1993random, kannan1997random, lovasz2006hit, cousins2018gaussian], often based on uncovering new mathematical/geometric structure, establishing connections to other fields (e.g., functional analysis, matrix concentration) and developing new tools for proving isoperimetric inequalities and analyzing Markov chains.
-With the proliferation of data and the increasing importance of machine learning, sampling has also become an essential algorithmic tool, with applications needing samplers in very high dimension, e.g., scientific computing [cousins2016practical, haraldsdottir2017chrr, kook2022sampling], systems biology [lewis2012constraining, thiele2013community], differential privacy [mcsherry2007mechanism, mironov2017renyi] and machine learning [bingham2019pyro, stan].
-Samplers for convex bodies are based on Markov chains (see Appendix A for a summary).
+The core of the celebrated result of [DFK91], giving a randomized polynomial-time algorithm for computing the volume of a convex body, was the first polynomial-time algorithm for uniformly sampling convex bodies.
+In the decades since, the study of sampling has led to a long series of improvements in its algorithmic complexity [LS90, LS93, KLS97, LV06, CV18], often based on uncovering new mathematical/geometric structure, establishing connections to other fields (e.g., functional analysis, matrix concentration) and developing new tools for proving isoperimetric inequalities and analyzing Markov chains.
+With the proliferation of data and the increasing importance of machine learning, sampling has also become an essential algorithmic tool, with applications needing samplers in very high dimension, e.g., scientific computing [CV16, HCT+17, KLSV22], systems biology [LNP12, TSF+13], differential privacy [MT07, Mir17] and machine learning [BCJ+19, Sta20].
+Samplers for convex bodies are based on Markov chains (see §A for a summary).
 Their analysis is based on bounding the conductance of the associated Markov chain, which in turn bounds the mixing rate.
@@
 An archetypal example of the latter is the following: for any measurable partition $S_{1},S_{2},S_{3}$ of a convex body $\mathcal{K}\subset\mathbb{R}^{d}$, we have
-$\operatorname{vol}(S_{3})\geq\frac{d(S_{1},S_{2})}{C_{\mathcal{K}}}\min\{\operatorname{vol}(S_{1}),\operatorname{vol}(S_{2})\}\,,$
+$\operatorname{vol}(S_{3})\geq\tfrac{d(S_{1},S_{2})}{C_{\mathcal{K}}}\min\{\operatorname{vol}(S_{1}),\operatorname{vol}(S_{2})\}\,,$
 where $d(\cdot,\cdot)$ is the (minimum) Euclidean distance, and $C_{\mathcal{K}}$ is an isoperimetric constant of the uniform distribution over $\mathcal{K}$.
@@
 The coefficient $C_{\mathcal{K}}^{2}$ is bounded by the Poincaré constant of the uniform distribution over $\mathcal{K}$ (and they are in fact asymptotically equal).
-The classical proof of conductance uses geometric properties of the random walk at hand to reduce the analysis to a suitable isoperimetric inequality (see e.g., [lovasz1993random, vempala2005geometric]).
+The classical proof of conductance uses geometric properties of the random walk at hand to reduce the analysis to a suitable isoperimetric inequality (see e.g., [LS93, Vem05]).
 The end result is a guarantee on the number of steps after which the total variation distance (TV distance) between the current distribution and the target is bounded by a desired error parameter.
-This framework has been widely used and effective in analyzing an array of candidate samplers, e.g., $\mathsf{Ball\ walk}$ [kannan1997random], $\mathsf{Hit\text{-}and\text{-}Run}$ [lovasz1999hit, lovasz2006hit], $\mathsf{Riemannian\ Hamiltonian\ Monte\ Carlo}$ [lee2018convergence] etc.
+This framework has been widely used and effective in analyzing an array of candidate samplers, e.g., $\mathsf{Ball\ walk}$ [KLS97], $\mathsf{Hit\text{-}and\text{-}Run}$ [Lov99, LV06], $\mathsf{Riemannian\ Hamiltonian\ Monte\ Carlo}$ [LV18] etc.
 One successful approach, studied intensively over the past decade, is based on diffusion.
 The basic idea is to first analyze a continuous-time diffusion process, typically modeled by a stochastic differential equation (SDE), and then show that a suitable time-discretization of the process, sometimes together with a Metropolis filter, converges to the desired distribution efficiently.
-A major success along this line is the $\mathsf{Unadjusted\ Langevin\ Algorithm}$ and its variants [besag1995bayesian, dalalyan2012sparse, dalalyan2017further, durmus2019analysis, vempala2019rapid].
-These algorithms have strong guarantees for sampling “nice” distributions, such as ones that are strongly log-concave, or more generally distributions satisfying isoperimetric inequalities, while also obeying some smoothness conditions.
+A major success along this line is the $\mathsf{Unadjusted\ Langevin\ Algorithm}$ and its variants, studied first in [BGHM95].
+These algorithms have strong guarantees for sampling “nice” distributions [DT12, Dal17, DMM19, VW19], such as ones that are strongly log-concave, or more generally distributions satisfying isoperimetric inequalities, while also obeying some smoothness conditions.
 The analysis of these algorithms is markedly different from the conductance approach, and typically yields guarantees in stronger metrics such as the $\mathsf{KL}$-divergence.
@@
 Can diffusion-based approaches be used for the problem of sampling convex bodies?
-Despite remarkable progress, thus far, constrained sampling problems have evaded the diffusion approach, except as a high-level analogy (e.g., the $\mathsf{Ball\ walk}$ can be viewed as a discretization of Brownian motion, but this alone does not suggest a route for analysis) or with significantly worse convergence rates (e.g., [pmlr-v65-brosse17a, bubeck2018sampling]).
+Despite remarkable progress, thus far, constrained sampling problems have evaded the diffusion approach, except as a high-level analogy (e.g., the $\mathsf{Ball\ walk}$ can be viewed as a discretization of Brownian motion, but this alone does not suggest a route for analysis) or with significantly worse convergence rates (e.g., [BDMP17, BEL18]).
 Our main finding is a simple diffusion-based algorithm that can be mapped to a stochastic process (and, importantly, to a pair of forward and backward processes), such that the rate of convergence is bounded directly by an appropriate functional inequality for the target distribution.
-As a consequence, for the first time, we obtain clean end-to-end guarantees in the Rényi divergence (which implies guarantees in other well known quantities such as $\mathcal{W}_{2},\mathsf{TV},\mathsf{KL},\chi^{2}$ etc.), while giving state-of-the-art runtime complexity for sampling convex bodies (e.g., $\mathsf{Ball\ walk}$ or $\mathsf{Speedy\ walk}$ [lovasz1993random, kannan1997random]).
-Besides being a stronger guarantee on the output, Rényi divergence is of particular interest for differential privacy [mironov2017renyi].
+As a consequence, for the first time, we obtain clean end-to-end guarantees in the Rényi divergence (which implies guarantees in other well-known quantities such as $\mathcal{W}_{2},\mathsf{TV},\mathsf{KL},\chi^{2}$ etc.), while giving state-of-the-art runtime complexity for sampling convex bodies (e.g., $\mathsf{Ball\ walk}$ or $\mathsf{Speedy\ walk}$ [LS93, KLS97]).
+Besides being a stronger guarantee on the output, Rényi divergence is of particular interest for differential privacy [Mir17].
 Perhaps most interesting is that our proof approach is completely different from prior work on convex body sampling.
@@
 sampler for uniformly sampling from $\mathcal{K}$.
-Each iteration consists of two steps, one that might leave the body and the second accepted only if it is (back) in the body.
+Each iteration consists of two steps, one that might leave the body and the second accepted only if it is (back) in $\mathcal{K}$.
 Input: initial point $x_{0}\sim\pi_{0}$, convex body $\mathcal{K}\subset\mathbb{R}^{d}$, iterations $T$, threshold $N$, and $h>0$.
@@
 This is reminiscent of the $\mathsf{Speedy\ walk}$ in the literature on convex body sampling (Algorithm 3), which is used as a tool to analyze proper steps of the $\mathsf{Ball\ walk}$.
-We refer the reader to Appendix B for a brief survey on these and related walks.
+We refer the reader to §B for a brief survey on these and related walks.
 This simple algorithm can be interpreted as a composition of “flows” in the space of measures.
@@
 Sampling from such a joint distribution to obtain the marginal on $X$ (say), can be more efficient than directly working only with $\pi^{X}$.
-This idea was utilized in Gaussian Cooling [cousins2018gaussian] and later as the restricted Gaussian Oracle (RGO) [lee2021structured, chen2022improved].
+This idea was utilized in Gaussian Cooling [CV18] and later as the restricted Gaussian Oracle (RGO) [LST21, CCSW22].
 Under this notation, Algorithm 1 corresponds to a Gibbs sampling scheme from the two marginals of $\pi(x,y)$. To be precise, Line 2 and Line 3 correspond to sampling from
@@
 It is well known that such a Gibbs sampling procedure will ensure the desired stationarity of $\pi(x,y)$.
+Note that, conditioned on the event that the algorithm does not fail, the resulting iterate will be an unbiased sample from the correct distribution.
 Stochastic perspective: forward and backward heat flows.
-Our algorithm can be viewed through the lens of stochastic analysis, due to an improved analysis for the proximal sampling [chen2022improved].
+Our algorithm can be viewed through the lens of stochastic analysis, due to an improved analysis for the proximal sampling [CCSW22].
 This view provides an interpolation in continuous-time, which is simple and powerful.
-To make this concrete, we borrow an exposition from [chewi2023log, Chapter 8.3].
-Let us denote the successive laws of $x_{i}$ and $y_{i}$ by $\mu^{X}_{i}$ and $\mu^{Y}_{i}$, respectively.
+To make this concrete, we borrow an exposition from [Che24, §8.3].
+We denote the successive laws of $x_{i}$ and $y_{i}$ by $\mu^{X}_{i}$ and $\mu^{Y}_{i}$, respectively.
 Recall that the first step of sampling from $\pi^{Y|X}(\cdot|x_{i})$ (Line 2) yields $\mu^{Y}_{i+1}=\int\pi^{Y|X=x}\,\mathrm{d}\mu^{X}_{i}(x)$.
-This is the result of evolving a probability measure under (forward) heat flow of $\mu_{i}^{X}$ for some time $h$, given precisely by the following stochastic differential equation: for $Z_{0}\sim\mu_{i}^{X}$,
+This is the result of evolving a probability measure under (forward) heat flow of $\mu_{i}^{X}$ for some time $h$, given by the following stochastic differential equation: for $Z_{0}\sim\mu_{i}^{X}$,
 $\mathrm{d}Z_{t}=\mathrm{d}B_{t}$ ($\mathsf{FH}$)
@@
 When $\mu_{i}^{X}=\pi^{X}$, the first step of Algorithm 1 gives
-$\pi^{Y}(y)=[\pi^{X}*\mathcal{N}(0,hI_{d})](y)=\frac{1}{\operatorname{vol}(\mathcal{K})\,(2\pi h)^{d/2}}\int_{\mathcal{K}}\exp\bigl{(}-\frac{1}{2h}\|y-x\|^{2}\bigr{)}\,\mathrm{d}x\,.$ (1.1)
+$\pi^{Y}(y)=[\pi^{X}*\mathcal{N}(0,hI_{d})](y)=\frac{1}{\operatorname{vol}(\mathcal{K})\,(2\pi h)^{d/2}}\int_{\mathcal{K}}\exp\bigl{(}-\frac{1}{2h}\lVert y-x\rVert^{2}\bigr{)}\,\mathrm{d}x\,.$ (1.1)
 The second step of sampling from $\pi^{X|Y}(\cdot|y_{i+1})$ can be represented by $\mu_{i+1}^{X}=\int\pi^{X|Y=y}\,\mathrm{d}\mu_{i+1}^{Y}(y)$ (Line 3).
@@
 Table 1: The Fokker-Planck equations for the forward and backward heat flow describe how the laws of $Z_{t}$ and $Z_{t}^{\leftarrow}$ in ($\mathsf{FH}$) and ($\mathsf{BH}$) evolve over time.
-See Section 3.2 for details.
+See §3.2 for details.
 Such a time reversal SDE is indeed possible!
@@
 Remark 1.
-We note that $\mathsf{In\text{-}and\text{-}Out}$ is exactly the proximal sampling scheme [lee2021structured, chen2022improved, fan2023improved] for uniform distributions.
-The proximal sampler with a target density proportional to $\exp(-V(x))$ considers an augmented distribution $\pi(x,y)\propto\exp(-V(x)-\frac{1}{2h}\|x-y\|^{2})$ and then repeats the following two steps: (1) $y_{i+1}\sim\pi^{Y|X=x_{i}}=\mathcal{N}(x_{i},hI_{d})$ and then (2) $x_{i+1}\sim\pi^{X|Y=y_{i+1}}$.
+We note that $\mathsf{In\text{-}and\text{-}Out}$ is exactly the proximal sampling scheme [LST21, CCSW22, FYC23] for uniform distributions.
+The proximal sampler with a target density proportional to $\exp(-V(x))$ considers an augmented distribution $\pi(x,y)\propto\exp(-V(x)-\frac{1}{2h}\lVert x-y\rVert^{2})$ and then repeats the following two steps: (1) $y_{i+1}\sim\pi^{Y|X=x_{i}}=\mathcal{N}(x_{i},hI_{d})$ and then (2) $x_{i+1}\sim\pi^{X|Y=y_{i+1}}$.
+Naïvely, the proximal sampler is implemented by performing rejection sampling, with the Gaussian centered at the minimizer of $\log\pi^{\cdot|Y=y_{i+1}}$ as the proposal.
+Realizing this would require a projection oracle (to $\mathcal{K}$), which is only known to be implementable with $\mathcal{O}(d^{2})$ membership queries.
+$\mathsf{In\text{-}and\text{-}Out}$ completely avoids the need for a projection oracle.
 1.2 Results
-Our model of computation is the classical general model for convex bodies [grotschel2012geometric].
-We assume $\operatorname{vol}(\mathcal{K})>0$ throughout this paper.
+Our computational model is the classical general model for convex bodies [GLS88].
+We assume $\operatorname{vol}\mathcal{K}>0$ in this paper.
 Below, $B_{r}(x)$ denotes the $d$-dimensional ball of radius $r$ centered at $x$.
@@
 Definition 2 (Warmness).
-A distribution $\mu$ is $M$-warm with respect to another distribution $\pi$, i.e., for every $x$ in the support of $\pi$, we have $\mathrm{d}\mu(x)\leq M\,\mathrm{d}\pi(x)$.
-We now summarize our main result which is further elaborated in Section 3.4 (Theorem 27).
+A distribution $\mu$ is $M$-warm with respect to another distribution $\pi$ if for every $x$ in the support of $\pi$, we have $\mathrm{d}\mu(x)\leq M\,\mathrm{d}\pi(x)$.
+We now summarize our main result, which is further elaborated in §3.4.
 Below, $\pi^{\mathcal{K}}$ is the uniform distribution over $\mathcal{K}$, and $\mathcal{R}_{q}$ is the Rényi-divergence of order $q$ (see Definition 6).
-Theorem 3.
+Theorem 3 (Informal version of Theorem 27).
 For any given $\eta,\varepsilon\in(0,1)$, $q\geq 1$, and any convex body $\mathcal{K}$ given by a well-defined membership oracle, there exist choices of parameters $h,N$ such that $\mathsf{In\text{-}and\text{-}Out}$, starting from an $M$-warm distribution, with probability at least $1-\eta$, returns $X\sim\mu$ such that $\mathcal{R}_{q}(\mu\mathbin{\|}\pi^{\mathcal{K}})\leq\varepsilon$.
-The number of proper steps is $\widetilde{\mathcal{O}}(qd^{2}\Lambda\log\frac{M}{\eta\varepsilon})$, and the expected total number of membership queries is $\widetilde{\mathcal{O}}(qMd^{2}\Lambda\log^{5}(1/\eta\varepsilon))$, where $\Lambda$ is the largest eigenvalue of the covariance of $\pi^{\mathcal{K}}$.
+The number of proper steps is $\widetilde{\mathcal{O}}(qd^{2}\Lambda\log^{2}\nicefrac{{M}}{{\eta\varepsilon}})$, and the expected total number of membership queries is $\widetilde{\mathcal{O}}(qMd^{2}\Lambda\log^{6}\nicefrac{{1}}{{\eta\varepsilon}})$, where $\Lambda$ is the largest eigenvalue of the covariance of $\pi^{\mathcal{K}}$.
+Remark 2.
+Despite our guarantee being in the much stronger “metric” of $\mathcal{R}_{q}$ compared to the $\mathsf{TV}$ guarantees of $\mathsf{Ball\ walk}$, we do not have to incur any additional asymptotic complexity.
+To obtain this result, one should choose the following values for the parameters: $h^{-1}=\widetilde{\Theta}(d^{2}\log\frac{qM\Lambda}{\eta}\log\log\frac{1}{\varepsilon})$, $N=\widetilde{\Theta}(\frac{qMd^{2}\Lambda\log^{5}(1/\varepsilon)}{\eta})$.
+See Lemma 14 for more details.
+Finally, while the assumption of warmness for the initialization may seem strong at the outset, for well-rounded convex bodies ($\mathbb{E}_{\pi}[\lVert X\rVert^{2}]\leq C^{2}d$ for some constant $C$), it is possible to generate an $\mathcal{O}(1)$ warm-start with complexity $\widetilde{\mathcal{O}}(d^{3})$.
+See [CV18, KZ25] for details.
 We note that for $X\sim\pi^{\mathcal{K}}$,
-$\|\mathrm{Cov}(\pi^{\mathcal{K}})\|_{\mathsf{op}}\leq\operatorname{tr}(\mathrm{Cov}(\pi^{\mathcal{K}}))=\mathbb{E}\bigl{[}\|X-\mathbb{E}[X]\|^{2}\bigr{]}\leq D^{2}\,.$
-The above guarantee in the Rényi divergence immediately provides $\mathcal{W}_{2},\mathsf{TV},\mathsf{KL}$, and $\chi^{2}$ guarantees as special cases.
-Previous guarantees for uniformly sampling convex bodies were only in the $\mathsf{TV}$-distance.
-For two distributions $\mu$ and $\pi$, we have
+$\lVert\operatorname{cov}\pi^{\mathcal{K}}\rVert_{\mathsf{op}}\leq\operatorname{tr}(\operatorname{cov}\pi^{\mathcal{K}})=\mathbb{E}[\lVert X-\mathbb{E}X\rVert^{2}]\leq D^{2}\,.$
+The above guarantee in the Rényi divergence immediately provides $\mathcal{W}_{2},\mathsf{TV},\mathsf{KL}$, and $\chi^{2}$ guarantees as special cases. For two distributions $\mu$ and $\pi$, we have
   1. 1.
-    $\mathsf{KL}(\mu\mathbin{\|}\pi)=\lim_{q\to 1}\mathcal{R}_{q}(\mu\mathbin{\|}\pi)\leq\mathcal{R}_{q}(\mu\mathbin{\|}\pi)\leq\mathcal{R}_{q^{\prime}}(\mu\mathbin{\|}\pi)\leq\mathcal{R}_{\infty}(\mu\mathbin{\|}\pi)=\log\sup\frac{\mathrm{d}\mu}{\mathrm{d}\pi}$ for $1<q\leq q^{\prime}$.
+    $\mathsf{KL}(\mu\mathbin{\|}\pi)=\lim_{q\downarrow 1}\mathcal{R}_{q}(\mu\mathbin{\|}\pi)\leq\mathcal{R}_{q}(\mu\mathbin{\|}\pi)\leq\mathcal{R}_{q^{\prime}}(\mu\mathbin{\|}\pi)\leq\mathcal{R}_{\infty}(\mu\mathbin{\|}\pi)=\log\sup\frac{\mathrm{d}\mu}{\mathrm{d}\pi}$, $1<q\leq q^{\prime}$.
   2. 2.
-    $2\,\|\mu-\pi\|_{\mathsf{TV}}^{2}\leq\mathsf{KL}(\mu\mathbin{\|}\pi)\leq\log(\chi^{2}(\mu\mathbin{\|}\pi)+1)=\mathcal{R}_{2}(\mu\mathbin{\|}\pi)$.
+    $2\,\lVert\mu-\pi\rVert_{\mathsf{TV}}^{2}\leq\mathsf{KL}(\mu\mathbin{\|}\pi)\leq\log(\chi^{2}(\mu\mathbin{\|}\pi)+1)=\mathcal{R}_{2}(\mu\mathbin{\|}\pi)$.
   3. 3.
@@
   4. 4.
-    $\mathcal{W}_{2}^{2}(\mu,\pi)\leq 2C_{\mathsf{PI}}(\pi)\,\chi^{2}(\mu\mathbin{\|}\pi)$ [liu2020poincare] and $C_{\mathsf{PI}}(\pi^{\mathcal{K}})\lesssim\|\mathrm{Cov}(\pi^{\mathcal{K}})\|_{\mathsf{op}}\log d$.
+    $\mathcal{W}_{2}^{2}(\mu,\pi)\leq 2C_{\mathsf{PI}}(\pi)\,\chi^{2}(\mu\mathbin{\|}\pi)$ [Liu20] and $C_{\mathsf{PI}}(\pi^{\mathcal{K}})\lesssim\lVert\operatorname{cov}(\pi^{\mathcal{K}})\rVert_{\mathsf{op}}\log d$.
 The query complexity is better if the convex body is (near-)isotropic, i.e., the uniform distribution over the body has (near-)identity covariance.
-This relies on recent estimates of the worst-case Poincaré constant for isotropic log-concave distributions [kannan1995isoperimetric, klartag2023logarithmic].
+This relies on recent estimates of the worst-case Poincaré constant for isotropic log-concave distributions [KLS95, Kla23].
+The condition that the convex body is isotropic can be achieved in practice through a rounding procedure [JLLV21, KZ24].
+See §A for more details.
 Corollary 4.
-Assume that $\pi^{\mathcal{K}}$ is isotropic.
+Assume that $\pi^{\mathcal{K}}$ is near-isotropic, i.e., the operator norm of its covariance is $\mathcal{O}(1)$.
 Under the same setting as above, $\mathsf{In\text{-}and\text{-}Out}$ succeeds with probability $1-\eta$, returning $X\sim\mu$ such that $\mathcal{R}_{q}(\mu\mathbin{\|}\pi^{\mathcal{K}})\leq\varepsilon$.
-The number of proper steps is $\widetilde{\mathcal{O}}(qd^{2}\log\frac{M}{\eta\varepsilon})$, and the expected total number of membership queries is $\widetilde{\mathcal{O}}(qMd^{2}\log^{5}(1/\eta\varepsilon))$.
+The number of proper steps is $\widetilde{\mathcal{O}}(qd^{2}\log^{2}\frac{M}{\eta\varepsilon})$, and the expected total number of membership queries is $\widetilde{\mathcal{O}}(qMd^{2}\log^{6}\tfrac{1}{\eta\varepsilon})$.
 Our analysis will in fact show that the bound on the number of proper steps holds for general non-convex bodies and any feasible start in $\mathcal{K}$.
@@
 On the other hand, using (7), which is weaker than (8) (in general, $C_{\mathsf{PI}}\leq C_{\mathsf{LSI}}$), the dependence on $M$ is logarithmic.
-We discuss our results further in Section 1.3.
+We discuss implications of our results further in §1.3.
 Outline of analysis.
-We summarize our proof strategy below, which consists of two steps: (i) The current distribution should converge to the uniform distribution, (ii) within each iteration of the algorithm, the failure probability and the expected number of rejections should be small enough.
+We summarize our proof strategy below, which requires us to demonstrate two facts: (i) The current distribution should converge to the uniform distribution, (ii) within each iteration of the algorithm, the failure probability and the expected number of rejections should be small enough.
   • We need to demonstrate that the corresponding Markov chain is rapidly mixing.
-Here, we use the heat flow perspective to derive mixing rates under any suitable divergence measure (such as $\mathsf{KL}$, $\chi^{2}$, or $\mathcal{R}_{q}$), extending known results for the unconstrained setting [chen2022improved].
-As a result, the mixing analysis reduces to a suitable functional inequality of the target distribution alone.
+Here, we use the heat flow perspective to derive mixing rates under any suitable divergence measure (such as $\mathsf{KL}$, $\chi^{2}$, or $\mathcal{R}_{q}$).
+This extends known results for the unconstrained setting [CCSW22].
+To summarize the proof, by considering instead the solutions after small time $t$, we invoke known contraction results from [CCSW22] and then use a continuity argument to conclude the proof.
   • We show that the number of rejections in Line 3 over the entire execution of the algorithm is bounded with high probability.
-To do this, we apply a detailed argument involving local conductance and the convexity of $\mathcal{K}$, which relies on techniques from [belkin2006heat].
+To do this, we apply a detailed argument involving local conductance and the convexity of $\mathcal{K}$, which relies on techniques from [BNN06].
 For this step, we show that with the appropriate choice of variance $h=\widetilde{\Theta}(d^{-2})$ and threshold $N=\widetilde{\Theta}(T\eta^{-1})$, the entire algorithm succeeds with probability $1-\eta$.
@@
 1.3 Discussion
-Here we make a few remarks, contrasting our results with known ones.
 No need to be lazy.
-Previous uniform samplers like the $\mathsf{Ball\ walk}$ are made lazy (i.e., with probability $1/2$, the Markov chain does nothing), to ensure convergence to the target stationary distribution.
-However, our algorithm does not need this, since we directly show that our sampler contracts towards a uniform distribution.
+Previous uniform samplers like $\mathsf{Ball\ walk}$ are made lazy (i.e., with probability $1/2$, it does nothing), to ensure convergence to the target stationary distribution.
+However, our algorithm does not need this, as our sampler is shown to directly contract towards the target.
 Unified framework.
-We remark that these two different bounds also place the previously known mixing guarantees for $\mathsf{Ball\ walk},\mathsf{Speedy\ walk}$ in a unified framework.
-Existing tight guarantees for $\mathsf{Speedy\ walk}$ are in TV distance and based on the log-Sobolev constant, assuming an oracle for implementing each step [lee2017eldan].
-The known convergence guarantees of $\mathsf{Ball\ walk}$ (see Appendix B for details), namely the mixing time of $\widetilde{\mathcal{O}}(Md^{2}D^{2}\log\frac{1}{\varepsilon})$ for TV distance, are for the composite algorithm [$\mathsf{Speedy\ walk}+$rejection sampling].
+We remark that Theorem 5 places the previously known mixing guarantees for $\mathsf{Ball\ walk},\mathsf{Speedy\ walk}$ in a unified framework.
+Existing tight guarantees for $\mathsf{Speedy\ walk}$ are in TV distance and based on the log-Sobolev constant, assuming an oracle for implementing each step [LV17].
+The known convergence guarantees of $\mathsf{Ball\ walk}$ (see §B for details), namely the mixing time of $\widetilde{\mathcal{O}}(Md^{2}D^{2}\log\frac{1}{\varepsilon})$ for TV distance, are for the composite algorithm [$\mathsf{Speedy\ walk}+$rejection sampling].
 Here $\mathsf{Speedy\ walk}$ records only the accepted steps of $\mathsf{Ball\ walk}$, so its stationary distribution differs slightly from the uniform distribution (and can be corrected with a post-processing step).
@@
 Effective step size.
-The $\mathsf{Ball\ walk}$’s largest possible step size is of order $1/\sqrt{d}$ (see Appendix B) to keep the rejection probability bounded by a constant.
-This bound could also be viewed as an “effective” step size of $\mathsf{In\text{-}and\text{-}Out}$.
-This follows from the fact that the $\ell_{2}$-norm of the Gaussian $\mathcal{N}(0,hI)$ is concentrated around $\sqrt{hd}$, and we will set the variance $h$ of $\mathsf{In\text{-}and\text{-}Out}$ to $\widetilde{\mathcal{O}}(1/d^{2})$, so we have $\sqrt{hd}\approx 1/\sqrt{d}$.
+The $\mathsf{Ball\ walk}$’s largest possible step size is of order $1/\sqrt{d}$ (see §B) to keep the rejection probability bounded by a constant.
+This bound could also be viewed as an “effective” step size of $\mathsf{In\text{-}and\text{-}Out}$, since the $\ell_{2}$-norm of the Gaussian $\mathcal{N}(0,hI)$ is concentrated around $\sqrt{hd}$ and we will set the variance $h$ of $\mathsf{In\text{-}and\text{-}Out}$ to $\widetilde{\mathcal{O}}(1/d^{2})$, so we have $\sqrt{hd}\approx 1/\sqrt{d}$.
 What has really changed?
@@
 Then, each iteration of $\mathsf{In\text{-}and\text{-}Out}$ can be viewed as a Gaussian version of a $\mathsf{Ball\ walk^{\prime}s\ proposal}+\mathsf{Speedy\ walk}$ algorithm.
-How can we compare $\mathsf{In\text{-}and\text{-}Out}$ with $\mathsf{Speedy\ walk}?$ Iterating speedy steps leads to a biased distribution (one that is proportional to the local conductance).
+How can we compare $\mathsf{In\text{-}and\text{-}Out}$ with $\mathsf{Speedy\ walk}?$ Iterating speedy steps leads to a biased distribution.
 As clarified in Remark 3, one step of (a Gaussian version of) $\mathsf{Speedy\ walk}$ can be understood as a step of backward heat flow.
@@
 For two measures $\mu,\nu$ on $\mathbb{R}^{d}$, the total variation distance between them is defined by
-$\|\mu-\nu\|_{\mathsf{TV}}:=\sup_{B\in\mathcal{F}}\lvert\mu(B)-\nu(B)\rvert\,,$
+$\lVert\mu-\nu\rVert_{\mathsf{TV}}:=\sup_{B\in\mathcal{F}}\lvert\mu(B)-\nu(B)\rvert\,,$
 where $\mathcal{F}$ is the collection of all measurable subsets of $\mathbb{R}^{d}\,$. The $2$-Wasserstein distance is given by
-$\mathcal{W}_{2}^{2}(\mu,\nu):=\inf_{\gamma\in\Gamma(\mu,\nu)}\mathbb{E}_{(X,Y)\sim\gamma}[\|X-Y\|^{2}]\,,$
+$\mathcal{W}_{2}^{2}(\mu,\nu):=\inf_{\gamma\in\Gamma(\mu,\nu)}\mathbb{E}_{(X,Y)\sim\gamma}[\lVert X-Y\rVert^{2}]\,,$
 where $\Gamma$ is the set of all couplings between $\mu,\nu$.
@@
 $\mathcal{R}_{q}(\mu\mathbin{\|}\nu):=\frac{1}{q-1}\log\bigl{(}\chi^{q}(\mu\mathbin{\|}\nu)+1\bigr{)}\,.$
-We recall two important functional inequalities of a distribution. We use $\nu$ to denote a probability measure over $\mathbb{R}^{d}$.
+We recall two important functional inequalities of a distribution.
 Definition 7.
-We say that $\nu$ satisfies a Poincaré inequality (PI) with parameter $C_{\mathsf{PI}}(\nu)$ if for all smooth functions $f:\mathbb{R}^{d}\to\mathbb{R}$,
-$\mathrm{Var}_{\nu}f\leq C_{\mathsf{PI}}(\nu)\,\mathbb{E}_{\nu}[\|\nabla f\|^{2}]\,,$
-where $\mathrm{Var}_{\nu}f\coloneqq\mathbb{E}_{\nu}\lvert f-\mathbb{E}_{\nu}f\rvert^{2}$.
+We say that a probability measure $\nu$ on $\mathbb{R}^{d}$ satisfies a Poincaré inequality (PI) with parameter $C_{\mathsf{PI}}(\nu)$ if for all smooth functions $f:\mathbb{R}^{d}\to\mathbb{R}$,
+$\operatorname{var}_{\nu}f\leq C_{\mathsf{PI}}(\nu)\,\mathbb{E}_{\nu}[\lVert\nabla f\rVert^{2}]\,,$
+where $\operatorname{var}_{\nu}f\coloneqq\mathbb{E}_{\nu}[\lvert f-\mathbb{E}_{\nu}f\rvert^{2}]$.
 The Poincaré inequality is implied by the log-Sobolev inequality.
 Definition 8.
-We say that $\nu$ satisfies a log-Sobolev inequality (LSI) with parameter $C_{\mathsf{LSI}}(\nu)$ if for all smooth functions $f:\mathbb{R}^{d}\to\mathbb{R}$,
-$\mathsf{Ent}_{\nu}(f^{2})\leq 2C_{\mathsf{LSI}}(\nu)\,\mathbb{E}_{\nu}[\|\nabla f\|^{2}]\,,$
+We say that a probability measure $\nu$ on $\mathbb{R}^{d}$ satisfies a log-Sobolev inequality (LSI) with parameter $C_{\mathsf{LSI}}(\nu)$ if for all smooth functions $f:\mathbb{R}^{d}\to\mathbb{R}$,
+$\mathsf{Ent}_{\nu}(f^{2})\leq 2C_{\mathsf{LSI}}(\nu)\,\mathbb{E}_{\nu}[\lVert\nabla f\rVert^{2}]\,,$
 where $\mathsf{Ent}_{\nu}(f^{2}):=\mathbb{E}_{\nu}[f^{2}\log f^{2}]-\mathbb{E}_{\nu}[f^{2}]\log(\mathbb{E}_{\nu}[f^{2}])$.
@@
 $\mathsf{KL}(\mu\mathbin{\|}\nu)\leq\frac{C_{\mathsf{LSI}}(\nu)}{2}\,\mathsf{FI}(\mu\mathbin{\|}\nu)\,,$
-where $\mathsf{FI}(\mu\mathbin{\|}\nu):=\mathbb{E}_{\mu}[\|\nabla\log\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\|^{2}]$ is the Fisher information of $\mu$ with respect to $\nu$.
-We recall the data-processing inequality for Rényi divergence and $f$-divergence.
+where $\mathsf{FI}(\mu\mathbin{\|}\nu):=\mathbb{E}_{\mu}[\lVert\nabla\log\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\rVert^{2}]$ is the Fisher information of $\mu$ with respect to $\nu$.
+We state two important lemmas which are needed for our proofs. The first is the data-processing inequality for Rényi divergence and $f$-divergence, given below.
 Lemma 9 (Data-processing inequality).
@@
 $D_{f}(\mu P\mathbin{\|}\nu P)\leq D_{f}(\mu\mathbin{\|}\nu)\,,\quad\text{and}\quad\mathcal{R}_{q}(\mu P\mathbin{\|}\nu P)\leq\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\,.$
-The aforementioned functional inequalities allow us to show exponential contraction of various divergences, through the following helpful inequality.
+Functional inequalities allow us to show exponential contraction of various divergences, through the following helpful inequality.
 Lemma 10 (Grönwall).
@@
 One successful iteration of $\mathsf{In\text{-}and\text{-}Out}$ is exactly the same as the composition of running the forward heat flow and then backward heat flow, both for time $h$.
-Its validity is justified in Section 1.1, and we give its proof below.
 Lemma 12.
@@
 $\mathsf{law}(Z_{h})=\mathsf{law}(Z_{0})*\mathcal{N}(0,hI_{d})=\mu_{k}^{X}*\mathcal{N}(0,hI_{d})=\mu_{k+1}^{Y}\,.$
-Regarding the backward part, it is known from [chen2022improved, Lemma 14] that the construction of the time-reversal SDE ensures that $(Z_{h}^{\leftarrow},Z_{0}^{\leftarrow})$ and $(Z_{0},Z_{h})$ have the same joint distribution, when $Z_{0}\sim\pi^{X}$ (and so $Z_{h}\sim\pi^{Y}$).
+Regarding the backward part, it is known from [CCSW22, Lemma 14] that the construction of the time-reversal SDE ensures that $(Z_{h}^{\leftarrow},Z_{0}^{\leftarrow})$ and $(Z_{0},Z_{h})$ have the same joint distribution, when $Z_{0}\sim\pi^{X}$ (and so $Z_{h}\sim\pi^{Y}$).
 Hence, $\mathsf{law}(Z_{h}^{\leftarrow}|Z_{0}^{\leftarrow}=y)=\mathsf{law}(Z_{0}|Z_{h}=y)=\pi^{X|Y=y}$, where the last equality follows from $(Z_{0},Z_{h})\sim\pi$.
@@
 Our analysis of these objects consists of two parts: (1) demonstrating the mixing of $\mathsf{In\text{-}and\text{-}Out}$, i.e., how many outer iterations are needed to be sufficiently close to the uniform distribution, and (2) quantifying the failure probability and wasted steps in Line 3.
-For (1), we collect in Section 3.1 some important implications of functional inequalities, e.g.
+For (1), we collect in §3.1 some important implications of functional inequalities, e.g.
 the Poincaré and log-Sobolev inequalities, for the uniform distribution.
-Then in Section 3.2, we exploit the flow perspective of the algorithm to obtain the mixing guarantees.
+Then in §3.2, we exploit the flow perspective of the algorithm to obtain the mixing guarantees.
 To this end, we revisit the proofs for the contraction results of forward and backward heat flows, and then extend them to our constrained setting:
-Theorem 13 (Adapted for uniform distributions, [chen2022improved]).
+Theorem 13 (Adapted for uniform distributions, [CCSW22]).
 Let $\mu^{X}_{k}$ be the law of the $k$-th output of $\mathsf{In\text{-}and\text{-}Out}$ with initial distribution $\mu_{0}^{X}$.
@@
 $\mathcal{R}_{q}(\mu_{k}^{X}\mathbin{\|}\pi^{X})\leq\begin{cases}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-\frac{2k\log(1+h/C_{\mathsf{PI}})}{q}&\text{if }k\leq\frac{q}{2\log(1+h/C_{\mathsf{PI}})}\,\bigl{(}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-1\bigr{)}\,,\\ (1+h/C_{\mathsf{PI}})^{-2(k-k_{0})/q}&\text{if }k\geq k_{0}:=\lceil\frac{q}{2\log(1+h/C_{\mathsf{PI}})}\,\bigl{(}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-1\bigr{)}\rceil\,.\end{cases}$
-This reduces the problem of obtaining a mixing guarantee to that of demonstrating a functional inequality on the target distribution.
+The result reduces the problem of obtaining a mixing guarantee to that of demonstrating a functional inequality on the target distribution.
 For this, it is not strictly necessary that $\mathcal{K}$ be convex.
-On the other hand, convexity of $\mathcal{K}$ is crucial for the proof of (2).
-We show in Section 3.3 that the failure probability remains under control by taking a suitable variance $h$ and threshold $N$.
-We then show that the expected number of trials per iteration is of order $\log N$, not $N$:
+Convexity of $\mathcal{K}$ is crucial this time unlike (1).
+We show in §3.3 that the failure probability remains under control by taking a suitable variance $h$ and threshold $N$, and that the expected number of trials per iteration is of order $\log N$, not $N$.
 Lemma 14 (Per-iteration guarantees).
 Let $\mathcal{K}$ be any convex body in $\mathbb{R}^{d}$ presented by a well-defined membership oracle, $\pi^{X}$ the uniform distribution over $\mathcal{K}$, and $\mu$ an $M$-warm initial distribution with respect to $\pi^{X}$.
-For any given $m\in\mathbb{N}$ and $\eta\in(0,1)$, set $Z=\frac{9mM}{\eta}(\geq 9)$, $h=\frac{\log\log Z}{2d^{2}\log Z}$ and $N=Z(\log Z)^{4}=\widetilde{\mathcal{O}}(\frac{mM}{\eta})$.
-Then, the failure probability of one iteration of $\mathsf{In\text{-}and\text{-}Out}$ is at most $\eta/m$.
-Moreover, the expected number of membership queries needed per iteration is $\mathcal{O}\bigl{(}M(\log\frac{mM}{\eta})^{4}\bigr{)}$.
+For any given $m\in\mathbb{N}$ and $\eta\in(0,1)$, set $Z=\frac{9mM}{\eta}(\geq 9)$, $h=\frac{\log\log Z}{2d^{2}\log Z}$ and $N=Z\log^{4}Z=\widetilde{\mathcal{O}}(\nicefrac{{mM}}{{\eta}})$.
+Then, the failure probability of one iteration of $\mathsf{In\text{-}and\text{-}Out}$ is at most $\eta/m$, and the expected membership queries per iteration is $\mathcal{O}\bigl{(}M\log^{4}\nicefrac{{mM}}{{\eta}}\bigr{)}$.
 3.1 Functional inequalities
@@
 The most natural ones to consider in this setting are the Poincaré inequality (7) and log-Sobolev inequality (8).
-In Appendix C, we provide a more detailed discussion of how these are related to other important notions of isoperimetry, such as the Cheeger and log-Cheeger inequalities.
+In §C, we provide a more detailed discussion of how these are related to other important notions of isoperimetry, such as the Cheeger and log-Cheeger inequalities.
 Below, we use $\mu,\nu$ to denote two arbitrary probability measures over $\mathbb{R}^{d}$.
@@
 The Poincaré inequality implies functional inequalities for the Rényi divergence.
-Lemma 16 ([vempala2019rapid, Lemma 9]).
+Lemma 16 ([VW19, Lemma 9]).
 Assume that $\nu$ satisfies (7) with parameter $C_{\mathsf{PI}}(\nu)$. For any $q\geq 2$ and probability measure $\mu$ over $\mathbb{R}^{d}$, it holds that
 $1-\exp(-\mathcal{R}_{q}(\mu\mathbin{\|}\nu))\leq\frac{q\,C_{\mathsf{PI}}(\nu)}{4}\,\mathsf{RF}_{q}(\mu\mathbin{\|}\nu)\,,$
-where $\mathsf{RF}_{q}(\mu\mathbin{\|}\nu):=q\,\mathbb{E}_{\nu}\bigl{[}\bigl{(}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\bigr{)}^{q}\|\nabla\log\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\|^{2}\bigr{]}\,/\,\mathbb{E}_{\nu}\bigl{[}\bigl{(}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\bigr{)}^{q}\bigr{]}$ is the Rényi Fisher information of order $q$ of $\mu$ with respect to $\nu$.
-The log-Sobolev inequality paired with the KL-divergence (8) can be understood as a special case of the following inequality [^2] Such inequalities are often called Polyak-Łojasiewicz inequalities, which say for $f:\mathbb{R}^{d}\to\mathbb{R}$, and all $y\in\mathbb{R}^{d}$ that $f(y)\leq c\,\|\nabla f(y)\|^{2}$ for some constant $c$, if $\min f(x)=0$.
+where $\mathsf{RF}_{q}(\mu\mathbin{\|}\nu):=q\,\mathbb{E}_{\nu}\bigl{[}\bigl{(}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\bigr{)}^{q}\lVert\nabla\log\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\rVert^{2}\bigr{]}\,/\,\mathbb{E}_{\nu}\bigl{[}\bigl{(}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\bigr{)}^{q}\bigr{]}$ is the Rényi Fisher information of order $q$ of $\mu$ with respect to $\nu$.
+The log-Sobolev inequality paired with the KL-divergence (8) can be understood as a special case of the following inequality [^2] Such inequalities are often called Polyak-Łojasiewicz inequalities, which say for $f:\mathbb{R}^{d}\to\mathbb{R}$, and all $y\in\mathbb{R}^{d}$ that $f(y)\leq c\,\lVert\nabla f(y)\rVert^{2}$ for some constant $c$, if $\min f(x)=0$.
 paired with the $q$-Rényi divergence for $q\geq 1$.
-Lemma 17 ([vempala2019rapid, Lemma 5]).
+Lemma 17 ([VW19, Lemma 5]).
 Assume that $\nu$ satisfies (8) with parameter $C_{\mathsf{LSI}}(\nu)$. For any $q\geq 1$ and probability measure $\mu$ over $\mathbb{R}^{d}$, it holds that
@@
 Note that $\lim_{q\to 1}\mathcal{R}_{q}=\mathsf{KL}$ and $\mathsf{RF}_{1}=\mathsf{FI}$.
-We have collected below the functional inequalities used to establish the mixing of our algorithm (see Appendix C for a detailed presentation).
+We have collected below the functional inequalities used to establish the mixing of our algorithm (see §C for a detailed presentation).
 Lemma 18.
 Let $\mathcal{K}\subset\mathbb{R}^{d}$ be a convex body with diameter $D$, and $\pi$ be the uniform distribution over $\mathcal{K}$.
-Then, $C_{\mathsf{PI}}(\pi)\lesssim\|\mathrm{Cov}(\pi)\|_{\mathsf{op}}\,\log d$ and $C_{\mathsf{LSI}}(\pi)\lesssim D^{2}$.
+Then, $C_{\mathsf{PI}}(\pi)\lesssim\lVert\operatorname{cov}\pi\rVert_{\mathsf{op}}\,\log d$ and $C_{\mathsf{LSI}}(\pi)\lesssim D^{2}$.
 If $\pi$ is isotropic, then $C_{\mathsf{PI}}(\pi)\lesssim\log d$ and $C_{\mathsf{LSI}}(\pi)\lesssim D$.
@@
 We start by analyzing how many outer iterations of $\mathsf{In\text{-}and\text{-}Out}$ are required to be $\varepsilon$-close to $\pi^{X}$, the uniform distribution over $\mathcal{K}$.
-The contraction of Algorithm 1 comes from analyzing Lines 2 and 3 through the perspective of heat flows (see Section 1.1).
-To exploit this view, we first revisit the previous contraction analysis in [chen2022improved], which is carried out for distributions with smooth densities.
+The contraction of Algorithm 1 comes from analyzing Lines 2 and 3 through the perspective of heat flows (see §1.1).
+To exploit this view, we first revisit the previous contraction analysis in [CCSW22], which is carried out for distributions with smooth densities.
 Although the uniform distribution is not even continuous, we prove a technical lemma (Lemma 22) that enables us to extend previously known results to the uniform distribution.
@@
 Part I: Contraction analysis for smooth distributions.
-In this part, we review the contraction results for heat flow and its time-reversal [chen2022improved], which are intimately connected with our algorithm.
+In this part, we review the contraction results for heat flow and its time-reversal [CCSW22], which are intimately connected with our algorithm.
 We also provide key technical ingredients needed for its proof, such as the computations for measures evolving under simultaneous forward/backward heat flows.
-We refer interested readers to Appendix D for additional details.
+We refer interested readers to §D for additional details.
 Only in Part I, we assume that $\nu$ denotes a probability measure with smooth density.
@@
 Backward heat flow.
-Although there are many ways to define a “reversal” of $P_{h}$, we will use the notion of adjoint introduced by [klartag2021spectral], which is the most immediately useful.
+Although there are many ways to define a “reversal” of $P_{h}$, we will use the notion of adjoint introduced by [KP21], which is the most immediately useful.
 […]
@@
 Instead of considering the change in metrics along the evolution of $\mu P_{t}$ with respect to “fixed” $\nu$, it will be useful to consider the simultaneous evolution of $\mu P_{t},\nu P_{t}$ (and similarly $\mu Q_{t},(\nu P_{h})Q_{t}$).
-This type of computation was carried out for specific metrics in earlier work [vempala2019rapid, chen2022improved].
-The following is a more generalized form of one appearing in [yuan2023class, Lemma 2].
+This type of computation was carried out for specific metrics in earlier work [VW19, CCSW22].
+The following is a more generalized form of one appearing in [YFL+23, Lemma 2].
 In the lemma below, we consider an arbitrary diffusion equation with corresponding Fokker-Planck equation:
@@
 Combining these, we are left with
-$\int\bigl{[}-\nabla\cdot(b_{t}\nu_{t})+\frac{1}{2}\Delta\nu_{t}\bigr{]}\Bigl{(}\bigl{(}f\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}-\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\Bigr{)}\,\mathrm{d}x$ $=-\int\Bigl{\langle}b_{t}\nu_{t}-\frac{1}{2}\nabla\nu_{t},\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\Bigr{\rangle}\,\mathrm{d}x$
+$\int\bigl{[}-\nabla\cdot(b_{t}\nu_{t})+$ $\frac{1}{2}\Delta\nu_{t}\bigr{]}\Bigl{(}\bigl{(}f\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}-\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\Bigr{)}\,\mathrm{d}x$
+$=-\int\Bigl{\langle}b_{t}\nu_{t}-\frac{1}{2}\nabla\nu_{t},\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\Bigr{\rangle}\,\mathrm{d}x$
 $=-\int\Bigl{\langle}b_{t}\mu_{t}-\frac{1}{2}\mu_{t}\nabla\log\nu_{t},\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\Bigr{\rangle}\,\mathrm{d}x\,.$
@@
 Before proceeding, we need a standard lemma on functional inequalities under ($\mathsf{FH}$).
-Lemma 20 (Functional inequalities under Gaussian convolutions, [chafai2004entropies, Corollary 13]).
+Lemma 20 (Functional inequalities under Gaussian convolutions, [Cha04, Corollary 13]).
 The following inequality holds for any $\pi$ with finite log-Sobolev and Poincaré constants,
@@
 Combining the previous two lemmas, we can establish contraction between $\mu P_{h}Q_{h}$ and $\nu$ after one forward/backward iteration.
-Theorem 21 ([chen2022improved, Theorem 3 and 4]).
+Theorem 21 ([CCSW22, Theorem 3 and 4]).
 Assume $\nu$, a measure with smooth density, satisfies (8) with constant $C_{\mathsf{LSI}}$.
@@
 Moreover, for all $q\geq 2$,
-$\mathcal{R}_{q}(\mu P_{h}Q_{h}\mathbin{\|}\nu)\leq\begin{cases}\mathcal{R}_{q}(\mu\mathbin{\|}\nu)-\frac{2\log(1+h/C_{\mathsf{PI}})}{q}&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\geq 1\,,\\ \frac{\mathcal{R}_{q}(\mu\mathbin{\|}\nu)}{(1+h/C_{\mathsf{PI}})^{2}}&\text{if}\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq 1\,.\end{cases}$
+$\mathcal{R}_{q}(\mu P_{h}Q_{h}\mathbin{\|}\nu)\leq\begin{cases}\mathcal{R}_{q}(\mu\mathbin{\|}\nu)-\frac{2\log(1+h/C_{\mathsf{PI}})}{q}&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\geq 1\,,\\ \frac{\mathcal{R}_{q}(\mu\mathbin{\|}\nu)}{(1+h/C_{\mathsf{PI}})^{2/q}}&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq 1\,.\end{cases}$
 Proof.
 Since the SDE in (3.1) captures the forward heat flow ($\mathsf{FH}$), we set $\mu_{0}$ and $\nu_{0}$ in Lemma 19 to $\mu$ and $\nu$, respectively, obtaining contraction along the forward heat flow as follows: Substituting the $q$-Rényi into Lemma 19, we have, from the definition of the Rényi divergence as $\mathcal{R}_{q}(\mu\mathbin{\|}\nu):=\frac{1}{q-1}\log\mathopen{}\mathclose{{}\left(D_{f}(\mu\mathbin{\|}\nu)+1}\right),$ with $f(x)=x^{q}-1$ and $g(x)=\frac{1}{q-1}\log(x+1)$,
-$\partial_{t}\mathcal{R}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})$ $=-\frac{q}{2}\,\frac{\mathbb{E}_{\mu P_{t}}\Bigl{[}\Bigl{\langle}\nabla\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q-1},\nabla\log\frac{\mu P_{t}}{\nu P_{t}}\Bigr{\rangle}\Bigr{]}}{(q-1)\,\mathbb{E}_{\nu P_{t}}\Bigl{[}\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q}\Bigr{]}}\underset{(i)}{=}-\frac{q}{2}\,\frac{\mathbb{E}_{\mu P_{t}}\Bigl{[}\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q-2}\langle\nabla\frac{\mu P_{t}}{\nu P_{t}},\nabla\log\frac{\mu P_{t}}{\nu P_{t}}\rangle\Bigr{]}}{\mathbb{E}_{\nu P_{t}}\Bigl{[}\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q}\Bigr{]}}$
-=$\underset{(ii)}{=}-\frac{q}{2}\,\frac{\mathbb{E}_{\nu P_{t}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{\mu P_{t}}{\nu P_{t}}}\right)^{q}\|\nabla\log\frac{\mu P_{t}}{\nu P_{t}}\|^{2}}\right]}{\mathbb{E}_{\nu P_{t}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{\mu P_{t}}{\nu P_{t}}}\right)^{q}}\right]}=-\frac{1}{2}\,\mathsf{RF}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})\,,$
+$\partial_{t}\mathcal{R}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})$ $=-\frac{q}{2}\,\frac{\mathbb{E}_{\mu P_{t}}\Bigl{[}\Bigl{\langle}\nabla\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q-1},\nabla\log\frac{\mu P_{t}}{\nu P_{t}}\Bigr{\rangle}\Bigr{]}}{(q-1)\,\mathbb{E}_{\nu P_{t}}\Bigl{[}\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q}\Bigr{]}}$
+=$\underset{(i)}{=}-\frac{q}{2}\,\frac{\mathbb{E}_{\mu P_{t}}\Bigl{[}\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q-2}\langle\nabla\frac{\mu P_{t}}{\nu P_{t}},\nabla\log\frac{\mu P_{t}}{\nu P_{t}}\rangle\Bigr{]}}{\mathbb{E}_{\nu P_{t}}\Bigl{[}\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q}\Bigr{]}}$
+=$\underset{(ii)}{=}-\frac{q}{2}\,\frac{\mathbb{E}_{\nu P_{t}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{\mu P_{t}}{\nu P_{t}}}\right)^{q}\lVert\nabla\log\frac{\mu P_{t}}{\nu P_{t}}\rVert^{2}}\right]}{\mathbb{E}_{\nu P_{t}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{\mu P_{t}}{\nu P_{t}}}\right)^{q}}\right]}=-\frac{1}{2}\,\mathsf{RF}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})\,,$
 where in $(i)$, we use again that $\nabla\bigl{[}f^{\prime}\bigl{(}\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\bigr{]}=\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\cdot\frac{\mu_{t}}{\nu_{t}}+f^{\prime}\bigl{(}\frac{\mu_{t}}{\nu_{t}}\bigr{)}\nabla\frac{\mu_{t}}{\nu_{t}}$, and $(ii)$ uses that $\nabla\frac{\mu P_{t}}{\nu P_{t}}=\frac{\mu P_{t}}{\nu P_{t}}\nabla\log\frac{\mu P_{t}}{\nu P_{t}}$, and the last equality recalls the definition of the Rényi Fisher information.
@@
 Since the SDE (3.1) also captures the backward equation ($\mathsf{BH}$), we set $\mu_{0}$ and $\nu_{0}$ in Lemma 19 to $\mu P_{h}$ and $\tilde{\nu}:=\nu P_{h}$ respectively, obtaining contraction along the backward heat flow:
-$\partial_{t}\mathcal{R}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})=-\frac{1}{2}\,\mathsf{RF}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})\leq-\frac{1}{q}\,\frac{\mathcal{R}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})}{C_{\mathsf{LSI}}(\tilde{\nu}Q_{t})}\underset{(i)}{\leq}-\frac{1}{q}\,\frac{\mathcal{R}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})}{C_{\mathsf{LSI}}+{h-t}}\,,$
+$\partial_{t}\mathcal{R}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})$ $=-\frac{1}{2}\,\mathsf{RF}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})$
+$\leq-\frac{1}{q}\,\frac{\mathcal{R}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})}{C_{\mathsf{LSI}}(\tilde{\nu}Q_{t})}\underset{(i)}{\leq}-\frac{1}{q}\,\frac{\mathcal{R}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})}{C_{\mathsf{LSI}}+{h-t}}\,,$
 where $(i)$ follows from that $\tilde{\nu}Q_{t}=\nu P_{h}Q_{t}=\nu P_{h-t}$ and $C_{\mathsf{LSI}}(\tilde{\nu}Q_{t})\leq C_{\mathsf{LSI}}+h-t$ due to Lemma 20. Applying Lemma 10 again yields
@@
 The result in the $\mathcal{R}_{q}$ under (7) can be shown in a similar manner.
-Only difference is that in forward and backward computations, one should use the functional inequality in Lemma 16 and the following standard inequalities:
-$1-\exp\bigl{(}-\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\bigr{)}\geq\begin{cases}\frac{1}{2}&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\geq 1\,,\\ \frac{1}{2}\mathcal{R}_{q}(\mu\mathbin{\|}\nu)&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq 1\,.\end{cases}$
+The only difference is that in forward and backward computations, one should use the functional inequality in Lemma 16 and the following standard inequalities:
+$1-\exp\bigl{(}-\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\bigr{)}\geq\begin{cases}\frac{1}{2}&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\geq 1\,,\\ \frac{1}{2}\mathcal{R}_{q}(\mu\mathbin{\|}\nu)&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq 1\,.\end{cases}\qed$
 Part II: Extension to constrained distributions.
@@
 The reverse SDE is more subtle, since $\nabla\log\nu P_{t}$ will in general cease to be Lipschitz as $t\to 0$. On the other hand, for any $h>0$, we can write explicitly
-$\mu_{h}(x)=\frac{1}{(2\pi h)^{d/2}}\int\exp\bigl{(}-\frac{\|x-y\|^{2}}{2h}\bigr{)}\,\mathrm{d}\mu_{0}(y)\,.$
+$\mu_{h}(x)=\frac{1}{(2\pi h)^{d/2}}\int\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2h}\bigr{)}\,\mathrm{d}\mu_{0}(y)\,.$
 If one considers the system started at $\tilde{\mu}_{0}=\mu_{\epsilon}=\nu P_{\epsilon}$ and solve the forward-backward Fokker-Planck equations on times $[0,h-\epsilon]$, then $\tilde{\mu}_{h-\epsilon}=\mu_{h}=\mu_{0}^{\leftarrow}=\tilde{\mu}_{0}^{\leftarrow}$ and
-$\mu_{h-\epsilon}^{\leftarrow}(x)=\tilde{\mu}_{h-\epsilon}^{\leftarrow}(x)=\int\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(x)}{\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)\,.$
-This follows from that if we consider system started at time $\epsilon>0$, with initial distribution $\mu_{\epsilon}$, then we obtain the above through the Bayesian perspective on the forward and reverse heat semigroups, elaborated in Appendix D.
-We now show that the following integral is indeed integrable, so $\tilde{\mu}_{h}^{\leftarrow}$ is well-defined:
-$\tilde{\mu}_{h}^{\leftarrow}(x):=\int\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2h}\bigr{)}\nu(x)}{\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)\,.$
-For fixed $x$ and $\epsilon<h/2$,
-$\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2(h-\epsilon)}\bigr{)}\nu(z)\,\mathrm{d}z\gtrsim\exp\bigl{(}-\frac{(\|y-x_{0}\|+D)^{2}}{2(h-\epsilon)}\bigr{)}\,,$
-as the support of $\nu$ is constrained to $\mathcal{K}\subset B_{D}(x_{0})$. Since $\mu_{0}$ has bounded support, $\mu_{h}(y)\lesssim\exp(-\frac{\|y\|^{2}}{a})$ for some constant $a>0$. Thus,
-$\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2(h-\epsilon)}\bigr{)}\,\mu_{h}(y)}{\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z}\lesssim\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2(h-\epsilon)}\bigr{)}\,\mu_{h}(y)}{\exp(-\frac{(\|y-x_{0}\|+D)^{2}}{2(h-\epsilon)})}\lesssim\exp\Bigl{(}\frac{\langle 2(x-x_{0}),y\rangle+2D\|y-x_{0}\|}{h}-\frac{\|y\|^{2}}{a}\Bigr{)}\,,$
-and the last bound is integrable in $y$.
-We then show the pointwise convergence of $\tilde{\mu}_{h-\epsilon}^{\leftarrow}$ to $\tilde{\mu}_{h}^{\leftarrow}$ as $\epsilon\to 0$.
+$\mu_{h-\epsilon}^{\leftarrow}(x)=\tilde{\mu}_{h-\epsilon}^{\leftarrow}(x)=\int\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(x)}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)\,.$
+This follows from that if we consider system started at time $\epsilon>0$, with initial distribution $\mu_{\epsilon}$, then we obtain the above through the Bayesian perspective on the forward and reverse heat semigroups, elaborated in §D.
+We note that $\tilde{\mu}_{h}^{\leftarrow}$ defined below is obviously integrable with respect to $x$, since this is a probability density function of $\mu_{h}Q$, where $Q$ is the backward kernel:
+$\tilde{\mu}_{h}^{\leftarrow}(x):=\int\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2h}\bigr{)}\nu(x)}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)\,.$
+Hence, $\tilde{\mu}_{h}^{\leftarrow}$ is well-defined over the support of $\nu$ almost surely.
+We now show the pointwise convergence of $\tilde{\mu}_{h-\epsilon}^{\leftarrow}$ to $\tilde{\mu}_{h}^{\leftarrow}$ as $\epsilon\to 0$.
 Note that $\nu P_{\epsilon}\to\nu$, as $\nu$ has a bounded support.
 Also, the denominator is independent of $\epsilon$ due to
-$\frac{1}{(2\pi(h-\epsilon))^{d/2}}\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z=\mathcal{N}(0,(h-\epsilon)I)*\nu P_{\epsilon}=\nu*\mathcal{N}(0,hI)\,.$
+$\frac{1}{(2\pi(h-\epsilon))^{d/2}}\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z=\mathcal{N}\bigl{(}0,(h-\epsilon)I\bigr{)}*\nu P_{\epsilon}=\nu*\mathcal{N}(0,hI)\,.$
 Hence, for $\epsilon\leq d^{-1}$,
-$\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2(h-\epsilon)}\bigr{)}}{\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z}\leq\Bigl{(}\frac{h}{h-\epsilon}\Bigr{)}^{d/2}\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2h}\bigr{)}}{\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}\lesssim\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2h}\bigr{)}}{\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}\,.$
-As shown above, the last bound is integrable with respect to $\mu_{h}$, so the dominated convergence theorem implies
-$\lim_{\epsilon\to 0}\int\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2(h-\epsilon)}\bigr{)}}{\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)=\int\frac{\exp\bigl{(}-\frac{\|x-y\|^{2}}{2h}\bigr{)}}{\int\exp\bigl{(}-\frac{\|z-y\|^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)\,,$
+$\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z}$ $\leq\Bigl{(}\frac{h}{h-\epsilon}\Bigr{)}^{d/2}\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2h}\bigr{)}}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}$
+$\lesssim\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2h}\bigr{)}}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}\,.$
+Since the last bound is integrable with respect to $\mu_{h}$, the dominated convergence theorem implies
+$\lim_{\epsilon\to 0}\int\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)=\int\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2h}\bigr{)}}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)\,,$
 Thus, the pointwise convergence follows.
-Note that if we take $\nu(x)=\pi^{X}(x)=\frac{\mathds{1}_{\mathcal{K}}(x)}{\operatorname{vol}(\mathcal{K})}$, then $\tilde{\mu}_{h}^{\leftarrow}$ is the distribution of the backwards step of our algorithm.
+Note that if we take $\nu(x)=\pi^{X}(x)=\frac{\mathds{1}_{\mathcal{K}}(x)}{\operatorname{vol}(\mathcal{K})}$, then $\tilde{\mu}_{h}^{\leftarrow}$ is the distribution of the backward step of our algorithm.
 In particular, this corresponds to first sampling $x\sim\mu_{h}$, then $y\sim Q_{h}^{\nu,h}(\cdot|x)$, which is precisely the law of $\mu_{h}^{\leftarrow}$ given by ($\mathsf{FP\text{-}BH}$).
-As for the second statement, it follows from Scheffé’s lemma [billingsley2017probability, Theorem 16.12] that the pointwise convergence of $\mu_{h-\varepsilon}^{\leftarrow}\to\mu_{h}^{\leftarrow}$ leads to its $\mathsf{TV}$-convergence, which in turn implies the weak convergence.
-It follows from lower semicontinuity of $D_{f}$ [ambrosio2000functions, Theorem 2.34] that the weak convergence ensures $D_{f}(\mu_{h}^{\leftarrow}\mathbin{\|}\nu)\leq\lim_{t\downarrow 0}D_{f}(\mu_{h-t}^{\leftarrow}\mathbin{\|}\nu^{\leftarrow}_{h-t})$.
+As for the second statement, it follows from Scheffé’s lemma [Bil95, Theorem 16.12] that the pointwise convergence of $\mu_{h-\varepsilon}^{\leftarrow}\to\mu_{h}^{\leftarrow}$ leads to its $\mathsf{TV}$-convergence, which in turn implies the weak convergence.
+It follows from lower semicontinuity of $D_{f}$ [AFP00, Theorem 2.34] that the weak convergence ensures $D_{f}(\mu_{h}^{\leftarrow}\mathbin{\|}\nu)\leq\lim_{t\downarrow 0}D_{f}(\mu_{h-t}^{\leftarrow}\mathbin{\|}\nu^{\leftarrow}_{h-t})$.
 ∎
@@
 3.3 Failure probability and wasted steps
-We begin by defining a suitable version of local conductance [kannan1997random].
+We begin by defining a suitable version of local conductance [KLS97].
 Definition 24 (Local conductance).
 The local conductance $\ell$ on $\mathbb{R}^{d}$ is defined by
-$\ell(x)\stackrel{{\scriptstyle\mathrm{{\scriptscriptstyle def}}}}{{=}}\frac{\int_{\mathcal{K}}\exp(-\frac{1}{2h}\|x-y\|^{2})\,\mathrm{d}y}{\int_{\mathbb{R}^{d}}\exp(-\frac{1}{2h}\|x-y\|^{2})\,\mathrm{d}y}=\frac{\int_{\mathcal{K}}\exp(-\frac{1}{2h}\|x-y\|^{2})\,\mathrm{d}y}{(2\pi h)^{d/2}}\,.$
+$\ell(x)\stackrel{{\scriptstyle\mathrm{{\scriptscriptstyle def}}}}{{=}}\frac{\int_{\mathcal{K}}\exp(-\frac{1}{2h}\lVert x-y\rVert^{2})\,\mathrm{d}y}{\int_{\mathbb{R}^{d}}\exp(-\frac{1}{2h}\lVert x-y\rVert^{2})\,\mathrm{d}y}=\frac{\int_{\mathcal{K}}\exp(-\frac{1}{2h}\lVert x-y\rVert^{2})\,\mathrm{d}y}{(2\pi h)^{d/2}}\,.$
 The local conductance at $y$ quantifies the success probability of the proposal at $y$ in Line 3.
@@
 Lemma 26.
-$\pi^{Y}(\mathcal{K}_{\delta}^{c})\leq e^{c_{1}^{2}}\exp(-\nicefrac{{t^{2}}}{{8c_{1}^{2}}})$ for $\delta=t/d$ and variance $h=c_{1}^{2}/d^{2}$, where $c_{1}$ is some constant and $t\geq 2c_{1}(c_{1}+1)$.
+For a convex body $\mathcal{K}\subset\mathbb{R}^{d}$ containing a unit ball $B_{1}(0)$,
+$\pi^{Y}(\mathcal{K}_{\delta}^{c})\leq\exp\bigl{(}-\frac{\delta^{2}}{2h}+\delta d\bigr{)}\,.$
 Proof.
-For $y\in\partial\mathcal{K}_{\delta}$, we can take the supporting half-space $H$ at $\mathsf{proj}_{\mathcal{K}}(y)$ containing $\mathcal{K}$, due to convexity of $\mathcal{K}$. Then,
-$\pi^{Y}(\mathcal{K}_{\delta}^{c})$ $=\frac{1}{\operatorname{vol}(\mathcal{K})}\int_{\mathcal{K}_{\delta}^{c}}\int_{\mathcal{K}}\frac{\exp\bigl{(}-\frac{1}{2h}\|y-x\|^{2}\bigr{)}}{(2\pi h)^{d/2}}\,\mathrm{d}x\,\mathrm{d}y\leq\frac{1}{\operatorname{vol}(\mathcal{K})}\int_{\mathcal{K}_{\delta}^{c}}\int_{H}\frac{\exp\bigl{(}-\frac{1}{2h}\|y-x\|^{2}\bigr{)}}{(2\pi h)^{d/2}}\,\mathrm{d}x\,\mathrm{d}y$
-$=\frac{1}{\operatorname{vol}(\mathcal{K})}\int_{\mathcal{K}_{\delta}^{c}}\int_{d(y,\mathcal{K})}^{\infty}\frac{\exp\bigl{(}-\frac{1}{2h}z^{2}\bigr{)}}{\sqrt{2\pi h}}\,\mathrm{d}z\,\mathrm{d}y\,.$ (3.3)
+For $y\in\partial\mathcal{K}_{\delta}$, we can take the supporting half-space $H(y)$ at $\mathsf{proj}_{\mathcal{K}}(y)$ containing $\mathcal{K}$, due to convexity of $\mathcal{K}$. Then,
+$\pi^{Y}(\mathcal{K}_{\delta}^{c})$ $=\frac{1}{\operatorname{vol}(\mathcal{K})}\int_{\mathcal{K}_{\delta}^{c}}\int_{\mathcal{K}}\frac{\exp(-\frac{1}{2h}\lVert y-x\rVert^{2})}{(2\pi h)^{d/2}}\,\mathrm{d}x\,\mathrm{d}y$
+$\leq\frac{1}{\operatorname{vol}(\mathcal{K})}\int_{\mathcal{K}_{\delta}^{c}}\int_{H(y)}\frac{\exp(-\frac{1}{2h}\lVert y-x\rVert^{2})}{(2\pi h)^{d/2}}\,\mathrm{d}x\,\mathrm{d}y$
+$=\frac{1}{\operatorname{vol}(\mathcal{K})}\int_{\mathcal{K}_{\delta}^{c}}\int_{d(y,\mathcal{K})}^{\infty}\frac{\exp(-\frac{z^{2}}{2h})}{\sqrt{2\pi h}}\,\mathrm{d}z\,\mathrm{d}y\,.$ (3.3)
 Let us denote the tail probability of the $1$-dimensional Gaussian with variance $h$ by
-$\mathsf{T}(s):=\mathbb{P}_{Z\sim\mathcal{N}(0,h)}(Z\geq s)=\frac{1}{\sqrt{2\pi h}}\int_{s}^{\infty}\exp\bigl{(}-\frac{1}{2h}z^{2}\bigr{)}\,\mathrm{d}z\,.$
-By the co-area formula and integration by parts,
-$\int_{\mathcal{K}_{\delta}^{c}}\int_{d(y,\mathcal{K})}^{\infty}\frac{\exp\bigl{(}-\frac{1}{2h}z^{2}\bigr{)}}{\sqrt{2\pi h}}\,\mathrm{d}z\mathrm{d}y=\int_{\delta}^{\infty}\mathsf{T}(s)\operatorname{vol}(\partial\mathcal{K}_{s})\,\mathrm{d}s$
+$\mathsf{T}(s):=\mathbb{P}_{\mathcal{N}(0,h)}(Z\geq s)=1-\Phi(h^{-1/2}s)\,,$
+where $\Phi$ is the CDF of the standard Gaussian. By the co-area formula and integration by parts,
+$\int_{\mathcal{K}_{\delta}^{c}}\int_{d(y,\mathcal{K})}^{\infty}\frac{\exp(-\frac{1}{2h}z^{2})}{\sqrt{2\pi h}}\,\mathrm{d}z\mathrm{d}y=\int_{\delta}^{\infty}\mathsf{T}(s)\operatorname{vol}(\partial\mathcal{K}_{s})\,\mathrm{d}s$
 $=\Bigl{[}\underbrace{\mathsf{T}(s)\int_{0}^{s}\operatorname{vol}(\partial\mathcal{K}_{z})\,\mathrm{d}z}_{\eqqcolon\mathsf{F}}\Bigr{]}_{\delta}^{\infty}+\int_{\delta}^{\infty}\frac{1}{\sqrt{2\pi h}}\exp\bigl{(}-\frac{s^{2}}{2h}\bigr{)}\int_{0}^{s}\operatorname{vol}(\partial\mathcal{K}_{z})\,\mathrm{d}z\,\mathrm{d}s\,.$ (3.4)
-Recall that $\mathsf{T}(s)\leq\frac{1}{2}\exp(-\frac{1}{2}(sh^{-1/2})^{2})$ for $s\geq\delta=t/d$ due to a standard tail bound on a Gaussian distribution. This tail bound, combined with
+Recall that $\mathsf{T}(s)\leq\frac{1}{2}\exp(-\frac{1}{2}(h^{-1/2}s)^{2})$ for $h^{-1/2}s\geq 0$ due to a standard tail bound on a Gaussian distribution. This tail bound, combined with
 $\int_{0}^{s}\operatorname{vol}(\partial\mathcal{K}_{z})\,\mathrm{d}z=\operatorname{vol}(\mathcal{K}_{s})-\operatorname{vol}(\mathcal{K})\leq\operatorname{vol}\bigl{(}(1+s)\,\mathcal{K}\bigr{)}-\operatorname{vol}(\mathcal{K})=\bigl{(}(1+s)^{d}-1\bigr{)}\operatorname{vol}(\mathcal{K})\,,$
 ensures that $\mathsf{F}$ vanishes at $s=\infty$. Hence, bounding the first term in (3.4) by 00 results in
-$\int_{\mathcal{K}_{\delta}^{c}}\int_{d(y,\mathcal{K})}^{\infty}\frac{\exp\bigl{(}-\frac{1}{2h}z^{2}\bigr{)}}{\sqrt{2\pi h}}\,\mathrm{d}z\,\mathrm{d}y$ $\leq\frac{1}{\sqrt{2\pi h}}\int_{\delta}^{\infty}\exp\bigl{(}-\frac{s^{2}}{2h}\bigr{)}\bigl{(}\underbrace{(1+s)^{d}}_{\leq\exp(sd)}-1\bigr{)}\operatorname{vol}(\mathcal{K})\,\mathrm{d}s$
+$\int_{\mathcal{K}_{\delta}^{c}}\int_{d(y,\mathcal{K})}^{\infty}\frac{\exp(-\frac{1}{2h}z^{2})}{\sqrt{2\pi h}}\,\mathrm{d}z\,\mathrm{d}y$ $\leq\frac{1}{\sqrt{2\pi h}}\int_{\delta}^{\infty}\exp\bigl{(}-\frac{s^{2}}{2h}\bigr{)}\bigl{(}\underbrace{(1+s)^{d}}_{\leq\exp(sd)}-1\bigr{)}\operatorname{vol}(\mathcal{K})\,\mathrm{d}s$
 $\leq\frac{\operatorname{vol}(\mathcal{K})}{\sqrt{2\pi h}}\exp(hd^{2}/2)\int_{\delta}^{\infty}\exp\bigl{(}-\frac{1}{2h}\,(s-hd)^{2}\bigr{)}\,\mathrm{d}s$
-$\underset{(i)}{\leq}\operatorname{vol}(\mathcal{K})\exp(hd^{2}/2)\exp\bigl{(}-\frac{1}{2}\,(\delta h^{-1/2}-dh^{1/2})^{2}\bigr{)}$
-$\underset{(ii)}{\leq}\operatorname{vol}(\mathcal{K})\exp(c_{1}^{2}/2)\exp\bigl{(}-\frac{t^{2}}{8c_{1}^{2}}\bigr{)}\,,$
-where in $(i)$ we used the tail bound of a Gaussian, and $(ii)$ follows from that in the regime of $h=c_{1}^{2}d^{-2}$, we have $\delta h^{-1/2}-dh^{1/2}-1\geq t/c_{1}-(c_{1}+1)\geq\nicefrac{{t}}{{2c_{1}}}$ for $t\geq 2c_{1}(c_{1}+1)$.
-Putting the last bound into (3.3) completes the proof of the claim.
-∎
+$\underset{(i)}{\leq}\operatorname{vol}(\mathcal{K})\exp(hd^{2}/2)\exp\bigl{(}-\frac{(\delta-hd)^{2}}{2h}\bigr{)}$
+$=\operatorname{vol}(\mathcal{K})\exp\bigl{(}-\frac{\delta^{2}}{2h}+\delta d\bigr{)}\,,$
+where in $(i)$ we used the tail bound for a Gaussian. ∎
+This core lemma suggests taking $\delta=\nicefrac{{t}}{{d}}$ and $h=\nicefrac{{c}}{{d^{2}}}$ for some $t,c>0$, under which we have
+$\pi^{Y}(\mathcal{K}_{\delta}^{c})\leq\exp\bigl{(}-\frac{t^{2}}{2c}+t\bigr{)}\,.$
 Now we choose a suitable threshold $N$ for bounding the failure probability.
 Following (3.3) in the proof, one can notice that for $y\in\mathcal{K}^{c}_{\delta}$, $\delta=\Omega(1/d)$, and $h=\Theta(d^{-2})$,
-$\ell(y)\leq\int_{d(y,\mathcal{K})}^{\infty}\frac{\exp\bigl{(}-\frac{1}{2h}z^{2}\bigr{)}}{\sqrt{2\pi h}}\,\mathrm{d}z=\mathbb{P}_{Z\sim\mathcal{N}(0,h)}(Z\geq\delta)\leq\exp(-\Omega(t^{2}))\,.$
+$\ell(y)\leq\int_{d(y,\mathcal{K})}^{\infty}\frac{\exp(-\frac{1}{2h}z^{2})}{\sqrt{2\pi h}}\,\mathrm{d}z=\mathbb{P}_{\mathcal{N}(0,h)}(Z\geq\delta)\leq\exp(-\Omega(t^{2}))\,.$
 Thus, the expected number of trials from $\mathcal{K}^{c}_{\delta}$ for the rejection sampling in Line 3 is $\ell(y)^{-1}\geq\exp(\Omega(t^{2}))$.
@@
 $\int_{\mathbb{R}^{d}}\underbrace{(1-\ell)^{N}\,\mathrm{d}\pi^{Y}}_{\eqqcolon\mathsf{A}}$ $=\int_{\mathcal{K}_{\delta}^{c}}\mathsf{A}+\int_{\mathcal{K}_{\delta}\cap[\ell\geq N^{-1}\log(3mM/\eta)]}\mathsf{A}+\int_{\mathcal{K}_{\delta}\cap[\ell<N^{-1}\log(3mM/\eta)]}\mathsf{A}$
-$\leq\pi^{Y}(\mathcal{K}_{\delta}^{c})+\int_{[\ell\geq N^{-1}\log(3mM/\eta)]}\exp(-\ell N)\,\mathrm{d}\pi^{Y}+\int_{\mathcal{K}_{\delta}\cap[\ell<N^{-1}\log(3mM/\eta)]}\frac{\ell(y)}{\operatorname{vol}(\mathcal{K})}\,\mathrm{d}y$
-$\leq e^{c_{1}^{2}}\exp(-\frac{t^{2}}{8c_{1}^{2}})+\frac{\eta}{3mM}+\frac{\log(3mM/\eta)}{N}\,\frac{\operatorname{vol}(\mathcal{K}_{\delta})}{\operatorname{vol}(\mathcal{K})}$
-$\leq e^{c_{1}^{2}}\exp(-\frac{t^{2}}{8c_{1}^{2}})+\frac{\eta}{3mM}+\frac{e^{t}}{N}\,\log\frac{3mM}{\eta}\,,$
+$\leq\pi^{Y}(\mathcal{K}_{\delta}^{c})+\int_{[\ell\geq N^{-1}\log(3mM/\eta)]}\exp(-\ell N)\,\mathrm{d}\pi^{Y}$
+$\qquad\qquad+\int_{\mathcal{K}_{\delta}\cap[\ell<N^{-1}\log(3mM/\eta)]}\frac{\ell(y)}{\operatorname{vol}(\mathcal{K})}\,\mathrm{d}y$
+$\leq\exp\bigl{(}-\frac{t^{2}}{2c}+t\bigr{)}+\frac{\eta}{3mM}+\frac{\log(3mM/\eta)}{N}\,\frac{\operatorname{vol}(\mathcal{K}_{\delta})}{\operatorname{vol}(\mathcal{K})}$
+$\leq\exp\bigl{(}-\frac{t^{2}}{2c}+t\bigr{)}+\frac{\eta}{3mM}+\frac{e^{t}}{N}\,\log\frac{3mM}{\eta}\,,$
 where we used $\operatorname{vol}(\mathcal{K}_{\delta})\subset\operatorname{vol}\bigl{(}(1+\delta)\mathcal{K}\bigr{)}=(1+\delta)^{d}\operatorname{vol}(\mathcal{K})\leq e^{t}\operatorname{vol}(\mathcal{K})$.
-Taking $c_{1}^{2}=\frac{\log\log Z}{2\log Z}$, $t=\sqrt{8}\log\log Z$, and $N=Z(\log Z)^{4}$, we can bound the last line by $\frac{\eta}{mM}$.
+Taking $c=\frac{\log\log Z}{2\log Z}$, $t=\sqrt{8}\log\log Z$, and $N=Z(\log Z)^{4}$, we can bound the last line by $\frac{\eta}{mM}$.
 Therefore,
@@
 Let $X$ be the minimum of the threshold $N$ and the number of trials until the first success.
-Then the expected number of trials per step is bounded by $M\mathbb{E}_{\pi^{Y}}[X]$ due to $\mathrm{d}\mu_{h}/\mathrm{d}\pi^{Y}\leq M$.
+Then the expected number of trials per step is bounded by $M\mathbb{E}_{\pi^{Y}}X$ since $\mathrm{d}\mu_{h}/\mathrm{d}\pi^{Y}\leq M$.
 Thus,
 $\int_{\mathbb{R}^{d}}\bigl{(}\frac{1}{\ell}\wedge N\bigr{)}\,\mathrm{d}\pi^{Y}$ $\leq\int_{\mathcal{K}_{\delta}}\frac{1}{\ell}\,\mathrm{d}\pi^{Y}+N\pi^{Y}(\mathcal{K}_{\delta}^{c})=\frac{\operatorname{vol}(\mathcal{K}_{\delta})}{\operatorname{vol}(\mathcal{K})}+N\pi^{Y}(\mathcal{K}_{\delta}^{c})$
-$\leq e^{t}+N\exp(-\Omega(t^{2}))\leq(\log Z)^{3}+3(\log Z)^{4}=\mathcal{O}\Bigl{(}\bigl{(}\log\frac{mM}{\eta}\bigr{)}^{4}\Bigr{)}\,.$
-Therefore, the expected number of trials per step is $\mathcal{O}\bigl{(}M(\log\frac{mM}{\eta})^{4}\bigr{)}$, and the claim follows since each trial uses one query to the membership oracle of $\mathcal{K}$.
-∎
+$\leq e^{t}+N\exp\bigl{(}-\frac{t^{2}}{2c}+t\bigr{)}\leq(\log Z)^{3}+3(\log Z)^{4}=\mathcal{O}\bigl{(}\log^{4}\frac{mM}{\eta}\bigr{)}\,.$
+Therefore, the expected number of trials per step is $\mathcal{O}(M\log^{4}\frac{mM}{\eta})$, and the claim follows since each trial uses one query to the membership oracle of $\mathcal{K}$. ∎
 3.4 Putting it together
-We can now show that $\mathsf{In\text{-}and\text{-}Out}$ subsumes previous results on uniform sampling from convex bodies (such as $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$), providing detailed versions of the main results in Section 1.2.
+We can now show that $\mathsf{In\text{-}and\text{-}Out}$ subsumes previous results on uniform sampling from convex bodies (such as $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$), providing detailed versions of the main results in §1.2.
 We first establish that the query complexity of $\mathsf{In\text{-}and\text{-}Out}$ matches that of the $\mathsf{Ball\ walk}$ under stronger divergences.
-Recall that $2\|\cdot\|_{\mathsf{TV}}^{2}\leq\mathsf{KL}\leq\log(1+\chi^{2})\leq\chi^{2}$.
+Recall that $2\,\lVert\cdot\rVert_{\mathsf{TV}}^{2}\leq\mathsf{KL}\leq\log(1+\chi^{2})\leq\chi^{2}$.
 Theorem 27.
@@
 For $\pi^{X}$ the uniform distribution over $\mathcal{K}$,
-  • It achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{\mathcal{O}}(qd^{2}\|\mathrm{Cov}(\pi^{X})\|_{\mathsf{op}}\log\frac{M}{\eta\varepsilon})$ iterations.
-With probability $1-\eta$, the algorithm iterates this many times without failure, using $\widetilde{\mathcal{O}}(qMd^{2}\|\mathrm{Cov}(\pi^{X})\|_{\mathsf{op}}(\log\frac{1}{\eta\varepsilon})^{5})$ expected number of membership queries in total.
-  • For isotropic $\pi^{X}$, with probability $1-\eta$, the algorithm achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ with $m=\widetilde{\mathcal{O}}(qd^{2}\log\frac{M}{\eta\varepsilon})$ iterations, using $\widetilde{\mathcal{O}}(qMd^{2}(\log\frac{1}{\eta\varepsilon})^{5})$ membership queries in expectation.
+  • It achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{\mathcal{O}}(qd^{2}\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{2}\frac{M}{\eta\varepsilon})$ iterations.
+With probability $1-\eta$, the algorithm iterates this many times without failure, using $\widetilde{\mathcal{O}}(qMd^{2}\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{6}\frac{1}{\eta\varepsilon})$ expected number of membership queries in total.
+  • For isotropic $\pi^{X}$, with probability $1-\eta$, the algorithm achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ with $m=\widetilde{\mathcal{O}}(qd^{2}\log^{2}\frac{M}{\eta\varepsilon})$ iterations, using $\widetilde{\mathcal{O}}(qMd^{2}\log^{6}\frac{1}{\eta\varepsilon})$ membership queries in expectation.
 Proof.
@@
 The $M$-warm start assumption guarantees $\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})\lesssim\log M$.
-Due to $C_{\mathsf{PI}}(\pi^{X})=\mathcal{O}(\|\mathrm{Cov}(\pi^{X})\|_{\mathsf{op}}\,\log d)$ (Lemma 18), $\mathsf{In\text{-}and\text{-}Out}$ can achieve $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{\mathcal{O}}(qd^{2}\|\mathrm{Cov}(\pi^{X})\|_{\mathsf{op}}\log\frac{M}{\eta\varepsilon})$ iterations.
+Due to $C_{\mathsf{PI}}(\pi^{X})=\mathcal{O}(\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\,\log d)$ (Lemma 18), $\mathsf{In\text{-}and\text{-}Out}$ can achieve $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{\mathcal{O}}(qd^{2}\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{2}\nicefrac{{M}}{{\eta\varepsilon}})$ iterations.
 Since each iteration has $\eta/m$-failure probability by Lemma 14, the union bound ensures that the total failure probability is at most $\eta$ throughout $m$ iterations.
-Lastly, each iteration requires $\widetilde{\mathcal{O}}(M(\log\frac{1}{\eta\varepsilon})^{4})$ membership queries in expectation by Lemma 14.
-Therefore, $\mathsf{In\text{-}and\text{-}Out}$ uses $\widetilde{\mathcal{O}}(qMd^{2}\min(D^{2},\|\mathrm{Cov}(\pi^{X})\|_{\mathsf{op}})(\log\frac{1}{\eta\varepsilon})^{5})$ expected number of membership queries over $m$ iterations.
+Lastly, each iteration requires $\widetilde{\mathcal{O}}(M\log^{4}\nicefrac{{1}}{{\eta\varepsilon}})$ membership queries in expectation by Lemma 14.
+Therefore, $\mathsf{In\text{-}and\text{-}Out}$ uses $\widetilde{\mathcal{O}}(qMd^{2}\min(D^{2},\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}})\log^{6}\nicefrac{{1}}{{\eta\varepsilon}})$ expected number of membership queries over $m$ iterations.
 Since $\mathcal{R}_{q}$ is non-decreasing in $q$, we can obtain the desired bound on $\mathcal{R}_{q}$ for $q\in[1,2)$.
-For isotropic $\pi^{X}$, we have $\mathrm{Cov}(\pi^{X})=I_{d}$, so the claim immediately follows from $C_{\mathsf{PI}}(\pi^{X})=\mathcal{O}(\log d)$ (see Lemma 18). ∎
+For isotropic $\pi^{X}$, we have $\operatorname{cov}\pi^{X}=I_{d}$, so the claim immediately follows from $C_{\mathsf{PI}}(\pi^{X})=\mathcal{O}(\log d)$ (see Lemma 18). ∎
 We now show that the number of proper steps is bounded as claimed for general non-convex bodies and any feasible start in $\mathcal{K}$.
@@
 We first bound the warmness of $\mu^{X}_{1}$ w.r.t. $\pi^{X}$ when $\mu^{X}_{0}=\delta_{x_{0}}$. One can readily check that
-$\mu_{1}^{X}(x)=\mathds{1}_{\mathcal{K}}(x)\cdot\int\frac{\exp\bigl{(}-\frac{1}{2h}\|y-x\|^{2}\bigr{)}\exp\bigl{(}-\frac{1}{2h}\|y-x_{0}\|^{2}\bigr{)}}{(2\pi h)^{d/2}\int_{\mathcal{K}}\exp\bigl{(}-\frac{1}{2h}\|y-x\|^{2}\bigr{)}\,\mathrm{d}x}\,\mathrm{d}y\,.$
-By Young’s inequality, $\|y-x\|^{2}\leq(\|y\|+D)^{2}\leq\frac{3}{2}\|y\|^{2}+3D^{2}$ for $x\in\mathcal{K}$. Hence,
-$\int\frac{\exp\bigl{(}-\frac{1}{2h}\|y-x\|^{2}\bigr{)}\exp\bigl{(}-\frac{1}{2h}\|y-x_{0}\|^{2}\bigr{)}}{\int_{\mathcal{K}}\exp\bigl{(}-\frac{1}{2h}\|y-x\|^{2}\bigr{)}\,\mathrm{d}x}\,\mathrm{d}y$
-$\leq$ $\frac{\exp(2h^{-1}D^{2})}{\operatorname{vol}(\mathcal{K})}\int\exp\Bigl{(}-\frac{1}{2h}\bigl{(}\|y-x\|^{2}+\|y-x_{0}\|^{2}-\frac{3}{2}\|y\|^{2}\bigr{)}\Bigr{)}\,\mathrm{d}y$
-$=$ $\frac{\exp(2h^{-1}D^{2})}{\operatorname{vol}(\mathcal{K})}\int\exp\Bigl{(}-\frac{1}{2h}\bigl{(}\frac{1}{2}\|y-2(x+x_{0})\|^{2}+(\|x\|^{2}+\|x_{0}\|^{2}-2\|x+x_{0}\|^{2})\bigr{)}\Bigr{)}\,\mathrm{d}y$
-$\leq$ $\frac{\exp(5h^{-1}D^{2})}{\operatorname{vol}(\mathcal{K})}\int\exp\Bigl{(}-\frac{1}{4h}\|y-2(x+x_{0})\|^{2}\Bigr{)}\,\mathrm{d}y$
+$\mu_{1}^{X}(x)=\mathds{1}_{\mathcal{K}}(x)\cdot\int\frac{\exp(-\frac{1}{2h}\,\lVert y-x\rVert^{2})\exp(-\frac{1}{2h}\,\lVert y-x_{0}\rVert^{2})}{(2\pi h)^{d/2}\int_{\mathcal{K}}\exp(-\frac{1}{2h}\,\lVert y-x\rVert^{2})\,\mathrm{d}x}\,\mathrm{d}y\,.$
+By Young’s inequality, $\lVert y-x\rVert^{2}\leq(\lVert y\rVert+D)^{2}\leq\frac{3}{2}\,\lVert y\rVert^{2}+3D^{2}$ for $x\in\mathcal{K}$. Hence,
+$\int\frac{\exp(-\frac{1}{2h}\,\lVert y-x\rVert^{2})\exp(-\frac{1}{2h}\lVert y-x_{0}\rVert^{2})}{\int_{\mathcal{K}}\exp(-\frac{1}{2h}\,\lVert y-x\rVert^{2})\,\mathrm{d}x}\,\mathrm{d}y$
+$\leq$ $\frac{\exp(2h^{-1}D^{2})}{\operatorname{vol}(\mathcal{K})}\int\exp\Bigl{(}-\frac{1}{2h}\,(\lVert y-x\rVert^{2}+\lVert y-x_{0}\rVert^{2}-\frac{3}{2}\lVert y\rVert^{2})\Bigr{)}\,\mathrm{d}y$
+$=$ $\frac{\exp(2h^{-1}D^{2})}{\operatorname{vol}(\mathcal{K})}\int\exp\Bigl{(}-\frac{1}{2h}\bigl{(}\frac{1}{2}\lVert y-2(x+x_{0})\rVert^{2}+(\lVert x\rVert^{2}+\lVert x_{0}\rVert^{2}-2\lVert x+x_{0}\rVert^{2})\bigr{)}\Bigr{)}\,\mathrm{d}y$
+$\leq$ $\frac{\exp(5h^{-1}D^{2})}{\operatorname{vol}(\mathcal{K})}\int\exp\Bigl{(}-\frac{1}{4h}\lVert y-2(x+x_{0})\rVert^{2}\Bigr{)}\,\mathrm{d}y$
 $=$ $\frac{\exp(5h^{-1}D^{2})}{\operatorname{vol}(\mathcal{K})}(4\pi h)^{d/2}\,.$
@@
 For $h=\tilde{\Theta}(d^{-2})$, $\mathsf{In\text{-}and\text{-}Out}$ requires $\widetilde{\mathcal{O}}(qd^{2}D^{2})$ iterations and in particular $\widetilde{\mathcal{O}}(qd^{2}D)$ iteration for isotropic uniform distributions.
-These results match those of $\mathsf{Speedy\ walk}$ [kannan2006blocking, lee2017eldan] (see Theorem 31).
+These results match those of $\mathsf{Speedy\ walk}$ [KLM06, LV17] (see Theorem 31).
 Acknowledgements.
 We are deeply grateful to Andre Wibisono and Sinho Chewi for helpful comments and pointers to the literature for Lemma 19.
-This work was supported in part by NSF award 210644, NSERC through the CGS-D award, and a Simons Investigator award.
-[…]
+This work was supported in part by NSF award 2106444, NSERC through the CGS-D award, and a Simons Investigator award.
 Appendix A Related work
 Sampling from constrained log-concave distributions is a fundamental task arising in many fields.
-Uniform sampling with convex constraints is its simplest manifestation, which was first studied as a core subroutine for a randomized volume-computation algorithm [dyer1991random].
-Since then, this fundamental problem has been studied for over three decades [lovasz1990mixing, lovasz1993random, kannan1997random, lovasz2003hit, lovasz2006hit, bubeck2018sampling, pmlr-v65-brosse17a].
+Uniform sampling with convex constraints is its simplest manifestation, which was first studied as a core subroutine for a randomized volume-computation algorithm [DFK91].
+Since then, this fundamental problem has been studied for over three decades [LS90, LS93, KLS97, LV03, LV06, BEL18, BDMP17].
 We review these algorithms, grouping them under three categories — geometric random walks, structured samplers, and diffusion-type samplers.
@@
 Geometric random walk.
-We discuss two geometric random walks – $\mathsf{Ball\ walk}$ [lovasz1993random, kannan1997random] and $\mathsf{Hit\text{-}and\text{-}Run}$ [smith1984efficient, lovasz1999hit].
+We discuss two geometric random walks – $\mathsf{Ball\ walk}$ [LS93, KLS97] and $\mathsf{Hit\text{-}and\text{-}Run}$ [Smi84, Lov99].
 $\mathsf{Ball\ walk}$ is a simple metropolized random walk; it draws $y$ uniformly at random from a ball of radius $\delta$ centered at a current point $x$, and moves to $y$ if $y\in\mathcal{K}$ and stays at $x$ otherwise.
-In the literature, $\mathsf{Ball\ walk}$ actually refers to a composite algorithm consisting of [$\mathsf{Speedy\ walk}+$ rejection sampling], where $\mathsf{Speedy\ walk}$ records only the accepted steps of $\mathsf{Ball\ walk}$ (see Appendix B for details).
+In the literature, $\mathsf{Ball\ walk}$ actually refers to a composite algorithm consisting of [$\mathsf{Speedy\ walk}+$ rejection sampling], where $\mathsf{Speedy\ walk}$ records only the accepted steps of $\mathsf{Ball\ walk}$ (see §B for details).
 The step size $\delta$ should be set to $\mathcal{O}(d^{-1/2})$ to avoid stepping outside of $\mathcal{K}$.
-[kannan1997random] showed that $\mathsf{Ball\ walk}$ needs $\widetilde{\mathcal{O}}(Md^{2}D^{2}\log\frac{1}{\varepsilon})$ membership queries to be $\varepsilon$-close to $\pi^{\mathcal{K}}$ in $\mathsf{TV}$, where $D$ is the diameter of $\mathcal{K}$, and the warmness parameter $M$ measures the closeness of the initial distribution to the target uniform distribution $\pi^{\mathcal{K}}$.
+[KLS97] showed that $\mathsf{Ball\ walk}$ needs $\widetilde{\mathcal{O}}(Md^{2}D^{2}\log\frac{1}{\varepsilon})$ membership queries to be $\varepsilon$-close to $\pi^{\mathcal{K}}$ in $\mathsf{TV}$, where $D$ is the diameter of $\mathcal{K}$, and the warmness parameter $M$ measures the closeness of the initial distribution to the target uniform distribution $\pi^{\mathcal{K}}$.
 $\mathsf{Hit\text{-}and\text{-}Run}$ is another zeroth-order algorithm that needs no step size; it picks a uniform random line $\ell$ passing a current point, and move to a uniform random point on $\ell\cap\mathcal{K}$.
-[lovasz2006hit] shows that, if we define the second moment as $R^{2}:=\mathbb{E}_{X\sim\pi^{\mathcal{K}}}[\|X-\mathbb{E}[X]\|^{2}]$, then $\mathsf{Hit\text{-}and\text{-}Run}$ requires $\mathcal{O}(d^{2}R^{2}\log\frac{M}{\varepsilon})$ queries.
+[LV06] shows that, if we define the second moment as $R^{2}:=\mathbb{E}_{\pi^{\mathcal{K}}}[\lVert X-\mathbb{E}X\rVert^{2}]$, then $\mathsf{Hit\text{-}and\text{-}Run}$ requires $\mathcal{O}(d^{2}R^{2}\log\frac{M}{\varepsilon})$ queries.
 Notably, this algorithm has a poly-logarithmic dependence on $M$ as opposed to $\mathsf{Ball\ walk}$.
@@
 This procedure finds a linear transformation that makes the geometry of $\mathcal{K}$ less skewed and so more amenable to sampling.
-In literature, there exists a randomized algorithm [jia2021reducing] that rounds $\mathcal{K}$ and generates a good warm start (i.e., $M=\mathcal{O}(1)$), with $\mathsf{Ball\ walk}$ used as a core subroutine.
-This algorithm takes up $\widetilde{\mathcal{O}}(d^{3})$ queries in total, and in such position with the good warm start, $\mathsf{Ball\ walk}$ only needs $\widetilde{\mathcal{O}}(d^{2}\log\frac{1}{\varepsilon})$ queries to sample from $\pi^{\mathcal{K}}$.
+In literature, there exists a randomized algorithm [JLLV21] that rounds $\mathcal{K}$ and generates a good warm start (i.e., $M=\mathcal{O}(1)$), with $\mathsf{Ball\ walk}$ used as a core subroutine.
+We refer readers to [KZ24] for a streamlined proof, with $\mathsf{In\text{-}and\text{-}Out}$ used as a main sampler.
+This algorithm takes up $\widetilde{\mathcal{O}}(d^{3.5})$ queries in total, and in such position with the good warm start, $\mathsf{Ball\ walk}$ only needs $\widetilde{\mathcal{O}}(d^{2}\log\frac{1}{\varepsilon})$ queries to sample from $\pi^{\mathcal{K}}$.
 Structured samplers.
@@
 The Hessian of $\phi$ encodes the local geometry of the constraint, and the samplers often work directly with $\nabla^{2}\phi$.
-The first canonical example of such a zeroth-order sampler is $\mathsf{Dikin\ walk}$ used when $\mathcal{K}$ is given by $m$ linear constraints [kannan2012random]; it draws a uniform sample from an ellipsoid (characterized by $\nabla^{2}\phi$) of fixed radius around a current point, and is often combined with a Metropolis adjustment.
-[kannan2012random] shows that $\mathsf{Dikin\ walk}$ mixes in $\mathcal{O}(md\log\frac{M}{\varepsilon})$ steps, although each iteration is slightly more expensive than one membership query.
-This algorithm requires no rounding, but still needs a good warm-start, which can be achieved by an annealing-type algorithm using $\widetilde{\mathcal{O}}(md)$ iterations of $\mathsf{Dikin\ walk}$ [kook2024gaussian].
-$\mathsf{Riemannian\ Hamiltonian\ Monte\ Carlo}$ is a structured sampler that exploits the first-order information of the potential (i.e., $\nabla\log(1/\pi)$) [girolami2011riemann]; its proposal is given as the solution to the Hamilton’s ODE equation, followed by the Metropolis-filter.
-In the linear-constraint setting above, this sampler requires $\mathcal{O}(md^{2/3}\log\frac{M}{\varepsilon})$ many iterations to achieve $\varepsilon$-close distance to $\pi^{\mathcal{K}}$ [lee2018convergence].
-This sampler is further analyzed for practical ODE solvers [kook2022condition] and for more sophisticated self-concordant barriers [gatmiry2023sampling].
-Similarly, $\mathsf{Mirror\ Langevin}$ [zhang2020wasserstein, jiang2021mirror, ahn2021efficient, li2022mirror] is a class of algorithms which converts the constrained problem into an unconstrained one obtained by considering the pushforward of the constrained space by $\nabla\phi$.
-The algorithm can also be metropolized [srinivasan2023fast].
-The best known rates for this algorithm are $\widetilde{\mathcal{O}}(d\log\frac{1}{\varepsilon})$ under some strong assumptions on $\phi$.
+The first canonical example of such a zeroth-order sampler is $\mathsf{Dikin\ walk}$ used when $\mathcal{K}$ is given by $m$ linear constraints [KN12]; it draws a uniform sample from an ellipsoid (characterized by $\nabla^{2}\phi$) of fixed radius around a current point, and is often combined with a Metropolis adjustment.
+[KN12] shows that $\mathsf{Dikin\ walk}$ mixes in $\mathcal{O}(md\log\frac{M}{\varepsilon})$ steps, although each iteration is slightly more expensive than one membership query.
+This algorithm requires no rounding, but still needs a good warm-start, which can be achieved by an annealing-type algorithm using $\widetilde{\mathcal{O}}(md)$ iterations of $\mathsf{Dikin\ walk}$ [KV24].
+$\mathsf{Riemannian\ Hamiltonian\ Monte\ Carlo}$ is a structured sampler that exploits the first-order information of the potential (i.e., $\nabla\log(1/\pi)$) [GC11]; its proposal is given as the solution to the Hamilton’s ODE equation, followed by the Metropolis-filter.
+In the linear-constraint setting above, it requires $\mathcal{O}(md^{2/3}\log\frac{M}{\varepsilon})$ many iterations to achieve $\varepsilon$-close distance to $\pi^{\mathcal{K}}$ [LV18].
+This sampler is further analyzed for practical ODE solvers [KLSV23] and for more sophisticated self-concordant barriers [GKV23].
+Similarly, $\mathsf{Mirror\ Langevin}$ [ZPFP20, Jia21, AC21, LTVW22] is a class of algorithms which converts the constrained problem into an unconstrained one obtained by considering the pushforward of the constrained space by $\nabla\phi$.
+The algorithm can also be metropolized [SWW24].
+The best known rate for this algorithm is $\widetilde{\mathcal{O}}(d\log\frac{1}{\varepsilon})$ under some strong assumptions on $\phi$.
 Diffusion-type samplers.
-Samplers based on discretizations of Itô diffusions, stochastic processes which rapidly mix to $\pi$ in continuous time, have long been used for sampling without constraints [besag1995bayesian, dalalyan2012sparse, dalalyan2017further, chewi2023log].
+Samplers based on discretizations of Itô diffusions, stochastic processes which rapidly mix to $\pi$ in continuous time, have long been used for sampling without constraints [BGHM95, DT12, Dal17, Che24].
 While the underlying stochastic processes generalize easily to constrained settings, the discretization analysis relies crucially on the smoothness of the target distribution.
-This is clearly impossible to achieve in the constrained setting, and so some techniques are required to circumvent this difficulty.
-These algorithms, however, generalize easily to the more general problem of sampling from distributions of the form $\tilde{\pi}^{X}\propto e^{-f}\mathds{1}_{\mathcal{K}}$, by naturally incorporating first order information from $f$.
-The first approach for adapting diffusion-based samplers [bubeck2015finite, bubeck2018sampling, lehec2023langevin] iterates a two-step procedure.
-First, a random step is taken, with $x_{k+1/2}\sim\mathcal{N}(x_{k},2hI_{d})$ for some appropriately chosen step $h$, [^4] A gradient step can be added in the more general case, for sampling from $\tilde{\pi}^{X}$ and then project it to $\mathcal{K}$, i.e., $x_{k+1}=\mathsf{proj}_{\mathcal{K}}(x_{k+1/2})$.
-The complexity is given in terms of queries to a projection oracle, each call to which can be implemented with a polynomial number of membership oracle queries; a total of $\tilde{\mathcal{O}}(\frac{d^{2}D^{3}}{\varepsilon^{4}})$ queries are needed to be $\varepsilon$-close in $\mathcal{W}_{2}$ to $\pi^{X}$.
-Another approach, which uses an algorithmically designed “soft” penalty instead of a projection, was proposed in [gurbuzbalaban2022penalized], and achieves a rate estimate of $\tilde{\mathcal{O}}(d/\varepsilon^{10})$.
-A second approach, suggested by [pmlr-v65-brosse17a], considers a different proximal scheme, which performs a “soft projection” onto $\mathcal{K}$, by taking steps like $\mathcal{N}((1-h\lambda^{-1})x_{k}+h\mathsf{proj}_{\mathcal{K}}(x_{k}),2hI_{d})$.
+This is clearly impossible to achieve in the constrained setting, so some techniques are required to circumvent this difficulty.
+These algorithms, however, generalize easily to the more general problem of sampling from distributions of the form $\tilde{\pi}^{X}\propto e^{-f}\mathds{1}_{\mathcal{K}}$, by incorporating first-order information from $f$.
+The first approach for adapting diffusion-based samplers [BEL15, BEL18, Leh23] iterates a two-step procedure.
+First, a random step is taken, with $x_{k+1/2}\sim\mathcal{N}(x_{k},2hI_{d})$ for some appropriately chosen step $h$, [^4] A gradient step can be added in the more general case, for sampling from $\tilde{\pi}^{X}$.
+and then project it to $\mathcal{K}$, i.e., $x_{k+1}=\mathsf{proj}_{\mathcal{K}}(x_{k+1/2})$.
+The complexity is given in terms of queries to a projection oracle, each call to which can be implemented with a polynomial number of membership oracle queries; a total of $\tilde{\mathcal{O}}(\nicefrac{{d^{2}D^{3}}}{{\varepsilon^{4}}})$ queries are needed to be $\varepsilon$-close in $\mathcal{W}_{2}$ to $\pi^{X}$.
+Another approach, which uses an algorithmically designed “soft” penalty instead of a projection, was proposed in [GHZ22], and achieves a rate estimate of $\tilde{\mathcal{O}}(\nicefrac{{d}}{{\varepsilon^{10}}})$.
+A second approach, suggested by [BDMP17], considers a different proximal scheme, which performs a “soft projection” onto $\mathcal{K}$, by taking steps like $\mathcal{N}((1-h\lambda^{-1})x_{k}+h\mathsf{proj}_{\mathcal{K}}(x_{k}),2hI_{d})$.
 It is called Moreau-Yosida regularized Langevin, named after an analogous regularization scheme for constrained optimization.
 This scheme also relies on access to a projection oracle for $\mathcal{K}$, and quantifies their query complexity accordingly.
-Their final rate estimate is $\tilde{\mathcal{O}}(\frac{d^{5}}{\varepsilon^{6}})$ to be $\varepsilon$-close in $\mathsf{TV}$ distance to $\pi^{X}$.
+Their final rate estimate is $\tilde{\mathcal{O}}(\nicefrac{{d^{5}}}{{\varepsilon^{6}}})$ to be $\varepsilon$-close in $\mathsf{TV}$ distance to $\pi^{X}$.
 Observing the prior work integrating diffusion-based sampling with convex constraints, the dependence on the key parameters $d,\varepsilon$, while polynomial, are many orders worse than the rates for zeroth-order samplers such as $\mathsf{Ball\ walk},\mathsf{Hit\text{-}and\text{-}Run}$.
@@
 Proximal schemes for sampling.
-The Gibbs sampling scheme used in this paper was inspired by the restricted Gaussian oracle introduced in [lee2021structured] (in turn inspired by Gaussian Cooling [cousins2018gaussian]), which alternately iterates between a pure Gaussian step, and a “proximal” step (which we elaborate in our exposition).
-This scheme was given novel interpretations by [chen2022improved], which showed that it interpolates the forward and backward heat flows, in the sense defined by [klartag2021spectral].
-The backward heat flow itself is intimately related to stochastic localization schemes, invented and popularized in [eldan2013thin, chen2021almost].
+The Gibbs sampling scheme used in this paper was inspired by the restricted Gaussian oracle introduced in [LST21] (in turn inspired by Gaussian Cooling [CV18]), which alternately iterates between a pure Gaussian step, and a “proximal” step (which we elaborate in our exposition).
+This scheme was given novel interpretations by [CCSW22], which showed that it interpolates the forward and backward heat flows, in the sense defined by [KP21].
+The backward heat flow itself is intimately related to stochastic localization schemes, invented and popularized in [Eld13, Che21].
 This formulation proved surprisingly powerful, allowing many existing rates in unconstrained sampling to be recovered from a relatively simple analysis.
-This was further extended by [fan2023improved] to achieve the current state-of-the-art rate in unconstrained sampling.
-Finally, [gopi2023algorithmic] suggest that this could be applied to tackle some constrained problems.
+This was further extended by [FYC23] to achieve the current state-of-the-art rate in unconstrained sampling.
+Finally, [GLL+23] suggests that this could be applied to tackle some constrained problems.
 However, the assumptions in this final mentioned work are not compatible with the uniform sampling problem on general convex bodies.
@@
 The following $\mathsf{TV}$-guarantee is obtained by lower bounding the $s$-conductance of $\mathsf{Ball\ walk}$, which requires a one-step coupling argument and the Cheeger inequality for $\pi$.
-We refer interested readers to [vempala2005geometric, Section 5].
+We refer interested readers to [Vem05, §5].
 Theorem 30 (Convergence of $\mathsf{Ball\ walk}$).
 For any $\varepsilon\in(0,1)$ and convex body $\mathcal{K}\subset\mathbb{R}^{d}$ presented by a well-defined membership oracle, let $\pi_{t}$ be the distribution after $t$ steps of $\mathsf{Ball\ walk}$ with an $M$-warm initial distribution $\pi_{0}$.
-Then, $\mathsf{Ball\ walk}$ with step size $\delta=\Theta(\frac{\varepsilon}{M\sqrt{d}})$ achieves $\|\pi_{t}-\pi\|_{\mathsf{TV}}\leq\varepsilon$ for $t\gtrsim d^{2}D^{2}\frac{M^{2}}{\varepsilon^{2}}\log\frac{M}{\varepsilon}$.
-If $\pi$ is isotropic, then $\mathsf{Ball\ walk}$ only needs $\mathcal{O}(d^{2}\log d\frac{M^{2}}{\varepsilon^{2}}\log\frac{M}{\varepsilon})$ iterations.
+Then, $\mathsf{Ball\ walk}$ with step size $\delta=\Theta(\frac{\varepsilon}{M\sqrt{d}})$ achieves $\lVert\pi_{t}-\pi\rVert_{\mathsf{TV}}\leq\varepsilon$ for $t\gtrsim d^{2}D^{2}\frac{M^{2}}{\varepsilon^{2}}\log\frac{M}{\varepsilon}$.
+If $\pi$ is isotropic, then $\mathsf{Ball\ walk}$ needs $\mathcal{O}(d^{2}\frac{M^{2}}{\varepsilon^{2}}\log d\log\frac{M}{\varepsilon})$ iterations.
 The mixing time of $\mathsf{Ball\ walk}$ under this approach has a polynomial dependence on $1/\varepsilon$, rather than a polylogarithmic dependence.
 Indirect analysis through $\mathsf{Speedy\ walk}$.
-[kannan1997random] introduced $\mathsf{Speedy\ walk}$, which could be viewed as a version of $\mathsf{Ball\ walk}$ and converges to a speedy distribution (see Proposition 1), which is slightly biased from $\pi$.
-Then, $\mathsf{Speedy\ walk}$ is used together with another algorithmic component (rejection sampling) [kannan1997random, Algorithm 4.15] that converts the speedy distribution to the uniform distribution.
+[KLS97] introduced $\mathsf{Speedy\ walk}$, which could be viewed as a version of $\mathsf{Ball\ walk}$ and converges to a speedy distribution (see Proposition 1), which is slightly biased from $\pi$.
+Then, $\mathsf{Speedy\ walk}$ is used together with another algorithmic component (rejection sampling) [KLS97, Algorithm 4.15] that converts the speedy distribution to the uniform distribution.
 In the literature, $\mathsf{Ball\ walk}$ often refers to ‘$\mathsf{Speedy\ walk}$ combined with the conversion step’, rather than a direct implementation of Algorithm 2.
@@
 and define the average conductance:
-$\lambda:=\mathbb{E}_{\pi}[\ell]=\frac{1}{\operatorname{vol}(\mathcal{K})}\int_{\mathcal{K}}\ell(x)\,\mathrm{d}x\,.$
-Proposition 1 ([kannan1997random]).
+$\lambda:=\mathbb{E}_{\pi}\ell=\frac{1}{\operatorname{vol}(\mathcal{K})}\int_{\mathcal{K}}\ell(x)\,\mathrm{d}x\,.$
+Proposition 1 ([KLS97]).
 The stationary distribution $\nu$ of $\mathsf{Speedy\ walk}$ has density
@@
 The speedy distribution $\nu$ is indeed different from the uniform distribution $\pi$, and this discrepancy is quantified in terms of the average conductance.
-Proposition 2 ([kannan1997random, Page 22]).
-$\|\nu-\pi\|_{\mathsf{TV}}\leq\frac{1-\lambda}{\lambda}$.
+Proposition 2 ([KLS97, Page 22]).
+$\lVert\nu-\pi\rVert_{\mathsf{TV}}\leq\frac{1-\lambda}{\lambda}$.
 One can relate the step size $\delta$ to the average conductance.
-Proposition 3 (Bound on average conductance, [kannan1997random, Corollary 4.5]).
+Proposition 3 (Bound on average conductance, [KLS97, Corollary 4.5]).
 $\lambda\geq 1-\frac{\delta\sqrt{d}}{2}$.
-The best known result for $\mathsf{Speedy\ walk}$’s mixing is due to [kannan2006blocking] devising the blocking conductance and using the log-Cheeger inequality.
-When $\nu$ is isotropic (i.e., it has covariance proportional to the identity matrix), [lee2017eldan] improves the mixing bound via the log-Cheeger constant.
+The best known result for $\mathsf{Speedy\ walk}$’s mixing is due to [KLM06] devising the blocking conductance and using the log-Cheeger inequality.
+When $\nu$ is isotropic (i.e., it has covariance proportional to the identity matrix), [LV17] improves the mixing bound via the log-Cheeger constant.
 Theorem 31 (Mixing of $\mathsf{Speedy\ walk}$).
 For any $\varepsilon\in(0,1)$ and convex body $\mathcal{K}\subset\mathbb{R}^{d}$ presented by a well-defined membership oracle, let $\nu_{t}$ be the distribution after $t$ proper steps of $\mathsf{Speedy\ walk}$ started at any feasible point $x_{0}\in\mathcal{K}$.
-Then, $\mathsf{Speedy\ walk}$ with step size $\delta=\Theta(d^{-1/2})$ achieves $\|\nu_{t}-\nu\|_{\mathsf{TV}}\leq\varepsilon$ for $t\gtrsim(D^{2}+\log(D\sqrt{d}))\,d^{2}\log\frac{1}{\varepsilon}$.
+Then, $\mathsf{Speedy\ walk}$ with step size $\delta=\Theta(d^{-1/2})$ achieves $\lVert\nu_{t}-\nu\rVert_{\mathsf{TV}}\leq\varepsilon$ for $t\gtrsim(D^{2}+\log(D\sqrt{d}))\,d^{2}\log\frac{1}{\varepsilon}$.
 From an $M$-warm start, the expected number of improper steps during $t$ iterations is $\widetilde{\mathcal{O}}(tM)$.
 When $\nu$ is isotropic, $\mathsf{Speedy\ walk}$ needs $\mathcal{O}(d^{2}D\log\frac{1}{\varepsilon}\log\log D)$ proper steps to achieve $\varepsilon$-$\mathsf{TV}$ distance to $\nu$.
-Then, [kannan1997random] uses the following post-processing step to obtain an approximately uniform distribution on $\mathcal{K}$, with a provable guarantee.
+Then, [KLS97] uses the following post-processing step to obtain an approximately uniform distribution on $\mathcal{K}$, with a provable guarantee.
 $\mathcal{A}$: Call $\mathsf{Speedy\ walk}$ to obtain a sample $X\sim\nu_{t}$ until $\frac{2d}{2d-1}\,X\in\mathcal{K}$. If so, return $\bar{X}=\frac{2d}{2d-1}\,X$.
-Proposition 4 ([kannan1997random, Theorem 4.16]).
-Under the same setting above, assume $\|\nu_{t}-\nu\|_{\mathsf{TV}}\leq\varepsilon$ for step size $\delta\leq(8d\log\frac{d}{\varepsilon})^{-1/2}$ and fixed $t\in\mathbb{N}$.
-For $\bar{\nu}=\mathsf{law}(\bar{X})$ given by $\mathcal{A}$, it holds that $\|\bar{\nu}-\pi\|_{\mathsf{TV}}\leq\varepsilon$, and the expected number of calls on the conversion algorithm is at most $2$.
+Proposition 4 ([KLS97, Theorem 4.16]).
+Under the same setting above, assume $\lVert\nu_{t}-\nu\rVert_{\mathsf{TV}}\leq\varepsilon$ for step size $\delta\leq(8d\log\frac{d}{\varepsilon})^{-1/2}$ and fixed $t\in\mathbb{N}$.
+For $\bar{\nu}=\mathsf{law}(\bar{X})$ given by $\mathcal{A}$, it holds that $\lVert\bar{\nu}-\pi\rVert_{\mathsf{TV}}\leq\varepsilon$, and the expected number of calls on the conversion algorithm is at most $2$.
 Combining the previous two results, we conclude that the total expected number of membership queries to obtain a sample $\varepsilon$-close to $\pi$ in $\mathsf{TV}$ is $\widetilde{\mathcal{O}}(Md^{2}D^{2}\log\frac{1}{\varepsilon})$, which now has a poly-logarithmic dependence on $1/\varepsilon$.
@@
 Appendix C Functional inequalities
-We provide full details on functional inequalities omitted in Section 3.1.
+We provide full details on functional inequalities omitted in §3.1.
 We use $\mu$ and $\mu_{\mathsf{LC}}$ to denote a probability measure and log-concave probability measure over $\mathbb{R}^{d}$, respectively.
@@
 $\mu^{+}(S):=\liminf_{\varepsilon\to 0}\frac{\mu(S^{\varepsilon})-\mu(S)}{\varepsilon}\,.$
-[cheeger1970lower] established $C_{\mathsf{PI}}(\mu)\lesssim C_{\mathsf{Ch}}^{-2}(\mu)$ [^5] The opposite direction $C_{\mathsf{PI}}(\mu_{\mathsf{LC}})\gtrsim C_{\mathsf{Ch}}^{-2}(\mu_{\mathsf{LC}})$ also holds for log-concave distributions due to [milman2009role], while $C_{\mathsf{PI}}(\mu)\gtrsim C_{\mathsf{Ch}}^{-2}(\mu)/d$ for general distributions due to [ledoux1994simple]., and then [kannan1995isoperimetric] showed that for covariance matrix $\Sigma_{\mu}:=\mathbb{E}_{\mu}[(\cdot-\mathbb{E}_{\mu}X)(\cdot-\mathbb{E}_{\mu}X)^{\mathsf{T}}]$,
-$C_{\mathsf{Ch}}(\mu_{\mathsf{LC}})\gtrsim\frac{1}{(\mathbb{E}_{\mu_{\mathsf{LC}}}[\|X-\mathbb{E}_{\mu_{\mathsf{LC}}}X\|^{2}])^{1/2}}=\frac{1}{(\operatorname{tr}(\Sigma_{\mu_{\mathsf{LC}}}))^{1/2}}\,.$ (C.1)
-This immediately leads to $C_{\mathsf{PI}}(\pi)\lesssim(\mathbb{E}_{\pi}[\|X-\mathbb{E}_{\pi}X\|^{2}])^{1/2}\leq D^{2}$ for the uniform distribution $\pi$ over a convex body $\mathcal{K}$ with diameter $D>0$.
-Kannan et al. proposed the KLS conjecture in the same paper, which says that for the spectral norm $\|\cdot\|_{2}$,
-$C_{\mathsf{Ch}}(\mu_{\mathsf{LS}})\gtrsim\frac{1}{\|\Sigma_{\mu_{\mathsf{LS}}}\|_{2}^{1/2}}\,.$
-While the original result in [kannan1995isoperimetric] ensures $C_{\mathsf{Ch}}\gtrsim d^{-1/2}$ for an isotropic log-concave distribution (due to $\Sigma=I_{d}$), this conjecture indeed claims $C_{\mathsf{Ch}}\gtrsim 1$ for such case.
-Following a line of work [lee2017eldan, chen2021almost, klartag2022bourgain, klartag2023logarithmic], the current bound is
-$C_{\mathsf{Ch}}(\mu_{\mathsf{LS}})\gtrsim\frac{(\log d)^{-1/2}}{\|\Sigma_{\mu_{\mathsf{LS}}}\|_{2}^{1/2}}\,,$
+[Che70] established $C_{\mathsf{PI}}(\mu)\lesssim C_{\mathsf{Ch}}^{-2}(\mu)$ [^5] The opposite direction $C_{\mathsf{PI}}(\mu_{\mathsf{LC}})\gtrsim C_{\mathsf{Ch}}^{-2}(\mu_{\mathsf{LC}})$ also holds for log-concave distributions due to [Mil09], while $C_{\mathsf{PI}}(\mu)\gtrsim C_{\mathsf{Ch}}^{-2}(\mu)/d$ for general distributions due to [Led94]., and then [KLS95] showed that for covariance matrix $\Sigma_{\mu}:=\mathbb{E}_{\mu}[(\cdot-\mathbb{E}_{\mu}X)(\cdot-\mathbb{E}_{\mu}X)^{\mathsf{T}}]$,
+$C_{\mathsf{Ch}}(\mu_{\mathsf{LC}})\gtrsim\frac{1}{(\mathbb{E}_{\mu_{\mathsf{LC}}}[\lVert X-\mathbb{E}_{\mu_{\mathsf{LC}}}X\rVert^{2}])^{1/2}}=\frac{1}{(\operatorname{tr}\Sigma_{\mu_{\mathsf{LC}}})^{1/2}}\,.$ (C.1)
+This immediately leads to $C_{\mathsf{PI}}(\pi)\lesssim(\mathbb{E}_{\pi}[\lVert X-\mathbb{E}_{\pi}X\rVert^{2}])^{1/2}\leq D^{2}$ for the uniform distribution $\pi$ over a convex body $\mathcal{K}$ with diameter $D>0$.
+Kannan et al. proposed the KLS conjecture in the same paper, which says that for the spectral norm $\lVert\cdot\rVert_{2}$,
+$C_{\mathsf{Ch}}(\mu_{\mathsf{LS}})\gtrsim\frac{1}{\lVert\Sigma_{\mu_{\mathsf{LS}}}\rVert_{2}^{1/2}}\,.$
+While the original result in [KLS95] ensures $C_{\mathsf{Ch}}\gtrsim d^{-1/2}$ for an isotropic log-concave distribution (due to $\Sigma=I_{d}$), this conjecture indeed claims $C_{\mathsf{Ch}}\gtrsim 1$ for such case.
+Following a line of work [LV17, Che21, KL22, Kla23], the current bound is
+$C_{\mathsf{Ch}}(\mu_{\mathsf{LS}})\gtrsim\frac{(\log d)^{-1/2}}{\lVert\Sigma_{\mu_{\mathsf{LS}}}\rVert_{2}^{1/2}}\,,$
 which implies that $C_{\mathsf{PI}}(\pi)\lesssim\log d$ when $\pi$ is isotropic for convex $\mathcal{K}$.
@@
 $C_{\mathsf{logCh}}(\mu):=\inf_{S\subset\mathbb{R}^{d}:\mu(S)\leq\frac{1}{2}}\frac{\mu^{+}(S)}{\mu(S)\sqrt{\log\frac{1}{\mu(S)}}}\,.$
-[ledoux1994simple] established that $C_{\mathsf{LSI}}(\mu)\lesssim C_{\mathsf{logCh}}^{-2}(\mu)$ [^6] The opposite direction holds under dimension-scaling due to [ledoux1994simple]: $C_{\mathsf{LSI}}(\mu)\gtrsim C_{\mathsf{logCh}}^{-2}(\mu)/d$., and [kannan2006blocking] showed that any log-concave distributions with support of diameter $D>0$ satisfy $C_{\mathsf{logCh}}(\mu_{\mathsf{LS}})\gtrsim D^{-1}$.
-Later in 2016, [lee2017eldan] improved this to $C_{\mathsf{logCh}}(\mu_{\mathsf{LS}})\gtrsim D^{-1/2}$ under isotropy.
+[Led94] established that $C_{\mathsf{LSI}}(\mu)\lesssim C_{\mathsf{logCh}}^{-2}(\mu)$ and $C_{\mathsf{LSI}}(\mu_{\mathsf{LS}})\gtrsim C_{\mathsf{logCh}}^{-2}(\mu_{\mathsf{LS}})$.
+[KLM06] showed that any log-concave distributions with support of diameter $D>0$ satisfy $C_{\mathsf{logCh}}(\mu_{\mathsf{LS}})\gtrsim D^{-1}$.
+Later in 2016, [LV17] improved this to $C_{\mathsf{logCh}}(\mu_{\mathsf{LS}})\gtrsim D^{-1/2}$ under isotropy.
 Therefore, for convex $\mathcal{K}$, it follows that $C_{\mathsf{LSI}}(\pi)\lesssim D^{2}$ and that $C_{\mathsf{LSI}}(\pi)\lesssim D$ if $\pi$ is isotropic.
@@
 We present additional technical background on the Wasserstein geometry and Markov semigroup theory.
-Interested readers can refer to [villani2009optimal, ambrosio2005gradient, chewi2023log] for standard references on Wasserstein spaces and applications to sampling.
+Interested readers can refer to [Vil09, AGS05, Che24] for standard references on Wasserstein spaces and applications to sampling.
 Wasserstein gradient.
@@
 This is a subtle fact, whose justification requires the introduction of a tool called Doob’s $h$-transform.
-The presentation of this subject is beyond the scope of this paper, and we refer interested readers to [klartag2021spectral] as a reference to its application in this context.
+The presentation of this subject is beyond the scope of this paper, and we refer interested readers to [KP21] as a reference to its application in this context.

### v3

@@
 It achieves state-of-the-art runtime complexity with stronger guarantees on the output than previously known, namely in Rényi divergence (which implies TV, $\mathcal{W}_{2}$, KL, $\chi^{2}$).
-The proof departs from known approaches for polytime algorithms for the problem — we utilize a stochastic diffusion perspective to show contraction to the target distribution with the rate of convergence determined by functional isoperimetric constants of the stationary density.
+The proof departs from known approaches for polytime algorithms for the problem — we utilize a stochastic diffusion perspective to show contraction to the target distribution with the rate of convergence determined by functional isoperimetric constants of the target distribution.
 1 Introduction
@@
 With the proliferation of data and the increasing importance of machine learning, sampling has also become an essential algorithmic tool, with applications needing samplers in very high dimension, e.g., scientific computing [CV16, HCT+17, KLSV22], systems biology [LNP12, TSF+13], differential privacy [MT07, Mir17] and machine learning [BCJ+19, Sta20].
-Samplers for convex bodies are based on Markov chains (see §A for a summary).
+Samplers for convex bodies are based on Markov chains (see §1.4 for a summary).
 Their analysis is based on bounding the conductance of the associated Markov chain, which in turn bounds the mixing rate.
@@
 An archetypal example of the latter is the following: for any measurable partition $S_{1},S_{2},S_{3}$ of a convex body $\mathcal{K}\subset\mathbb{R}^{d}$, we have
-$\operatorname{vol}(S_{3})\geq\tfrac{d(S_{1},S_{2})}{C_{\mathcal{K}}}\min\{\operatorname{vol}(S_{1}),\operatorname{vol}(S_{2})\}\,,$
+$\operatorname{vol}(S_{3})\geq\frac{d(S_{1},S_{2})}{C_{\mathcal{K}}}\min\{\operatorname{vol}(S_{1}),\operatorname{vol}(S_{2})\}\,$,
 where $d(\cdot,\cdot)$ is the (minimum) Euclidean distance, and $C_{\mathcal{K}}$ is an isoperimetric constant of the uniform distribution over $\mathcal{K}$.
-(The KLS conjecture posits that $C_{\mathcal{K}}=\mathcal{O}(1)$ for any convex body $\mathcal{K}$ in isotropic position, i.e., under the normalization that a random point from $\mathcal{K}$ has identity covariance).
+(The KLS conjecture posits that $C_{\mathcal{K}}=O(1)$ for any convex body $\mathcal{K}$ in isotropic position, i.e., under the normalization that a random point from $\mathcal{K}$ has identity covariance).
 The coefficient $C_{\mathcal{K}}^{2}$ is bounded by the Poincaré constant of the uniform distribution over $\mathcal{K}$ (and they are in fact asymptotically equal).
@@
 This framework has been widely used and effective in analyzing an array of candidate samplers, e.g., $\mathsf{Ball\ walk}$ [KLS97], $\mathsf{Hit\text{-}and\text{-}Run}$ [Lov99, LV06], $\mathsf{Riemannian\ Hamiltonian\ Monte\ Carlo}$ [LV18] etc.
-One successful approach, studied intensively over the past decade, is based on diffusion.
+A different approach, studied intensively over the past decade, is based on diffusion.
 The basic idea is to first analyze a continuous-time diffusion process, typically modeled by a stochastic differential equation (SDE), and then show that a suitable time-discretization of the process, sometimes together with a Metropolis filter, converges to the desired distribution efficiently.
 A major success along this line is the $\mathsf{Unadjusted\ Langevin\ Algorithm}$ and its variants, studied first in [BGHM95].
-These algorithms have strong guarantees for sampling “nice” distributions [DT12, Dal17, DMM19, VW19], such as ones that are strongly log-concave, or more generally distributions satisfying isoperimetric inequalities, while also obeying some smoothness conditions.
+These algorithms have strong guarantees for sampling “nice” distributions [DT12, Dal17, DMM19, VW23], such as ones that are strongly logconcave, or more generally distributions satisfying isoperimetric inequalities, while also obeying some smoothness conditions.
 The analysis of these algorithms is markedly different from the conductance approach, and typically yields guarantees in stronger metrics such as the $\mathsf{KL}$-divergence.
@@
 Despite remarkable progress, thus far, constrained sampling problems have evaded the diffusion approach, except as a high-level analogy (e.g., the $\mathsf{Ball\ walk}$ can be viewed as a discretization of Brownian motion, but this alone does not suggest a route for analysis) or with significantly worse convergence rates (e.g., [BDMP17, BEL18]).
-Our main finding is a simple diffusion-based algorithm that can be mapped to a stochastic process (and, importantly, to a pair of forward and backward processes), such that the rate of convergence is bounded directly by an appropriate functional inequality for the target distribution.
+Our main finding is a simple diffusion-based algorithm that can be mapped to a stochastic process such that the rate of convergence is bounded directly by an appropriate functional inequality for the target distribution.
 As a consequence, for the first time, we obtain clean end-to-end guarantees in the Rényi divergence (which implies guarantees in other well-known quantities such as $\mathcal{W}_{2},\mathsf{TV},\mathsf{KL},\chi^{2}$ etc.), while giving state-of-the-art runtime complexity for sampling convex bodies (e.g., $\mathsf{Ball\ walk}$ or $\mathsf{Speedy\ walk}$ [LS93, KLS97]).
 Besides being a stronger guarantee on the output, Rényi divergence is of particular interest for differential privacy [Mir17].
-Perhaps most interesting is that our proof approach is completely different from prior work on convex body sampling.
+Perhaps most interesting is that our proof approach is quite different from prior work on convex body sampling.
 In summary,
@@
 This is reminiscent of the $\mathsf{Speedy\ walk}$ in the literature on convex body sampling (Algorithm 3), which is used as a tool to analyze proper steps of the $\mathsf{Ball\ walk}$.
-We refer the reader to §B for a brief survey on these and related walks.
+We refer the reader to §A for a brief survey on these and related walks.
 This simple algorithm can be interpreted as a composition of “flows” in the space of measures.
@@
 We begin by defining an augmented probability measure on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ by
-[…]
-We denote by $\pi^{X},\pi^{X|Y}(\cdot|y)$ the marginal distribution of its first component (resp.
-conditional distribution given the second component), and similarly denote by $\pi^{Y},\pi^{Y|X}(\cdot|x)$ for the second component.
+$\pi(x,y)\propto\exp\bigl(-\frac{1}{2h}\,\lvert x-y\rvert^{2}\bigr)\,\mathds{1}_{\mathcal{K}}(x)\,.$
+We denote by $\pi^{X},\pi^{X|Y}(\cdot\,|\,y)$ the marginal distribution of its first component (i.e., conditional distribution given the second component), and similarly denote by $\pi^{Y},\pi^{Y|X}(\cdot\,|\,x)$ for the second component.
 In particular, the marginal in the first component $\pi^{X}$ is the uniform distribution over $\mathcal{K}$.
@@
 This idea was utilized in Gaussian cooling [CV18] and later as the restricted Gaussian Oracle (RGO) [LST21, CCSW22].
-Under this notation, Algorithm 1 corresponds to a Gibbs sampling scheme from the two marginals of $\pi(x,y)$. To be precise, Line 2 and Line 3 correspond to sampling from
+Under this notation, Algorithm 1 corresponds to a Gibbs sampling scheme from the two marginals of $\pi(x,y)$. To be precise, Line 2 and Line 3 correspond to sampling
 $y_{i+1}\sim\pi^{Y|X}(\cdot\,|\,x_{i})=\mathcal{N}(x_{i},hI_{d})\qquad\mbox{ and }\qquad x_{i+1}\sim\pi^{X|Y}(\cdot\,|\,y_{i+1})=\mathcal{N}(y_{i+1},hI_{d})|_{\mathcal{K}}\,.$
@@
 Note that, conditioned on the event that the algorithm does not fail, the resulting iterate will be an unbiased sample from the correct distribution.
-Stochastic perspective: forward and backward heat flows.
-Our algorithm can be viewed through the lens of stochastic analysis, due to an improved analysis for the proximal sampling [CCSW22].
+Stochastic perspective: heat flows.
+Our algorithm can be viewed through the lens of stochastic analysis of proximal sampling [CCSW22].
 This view provides an interpolation in continuous-time, which is simple and powerful.
-To make this concrete, we borrow an exposition from [Che24, §8.3].
+To make this concrete, we follow the exposition from [Che25, §8.3].
 We denote the successive laws of $x_{i}$ and $y_{i}$ by $\mu^{X}_{i}$ and $\mu^{Y}_{i}$, respectively.
@@
 The second step of sampling from $\pi^{X|Y}(\cdot\,|\,y_{i+1})$ can be represented by $\mu_{i+1}^{X}=\int\pi^{X|Y=y}\,\mathrm{d}\mu_{i+1}^{Y}(y)$ (Line 3).
-The continuous-time process corresponding to this step might not be obvious.
-However, let us consider ($\mathsf{FH}$) with $Z_{0}\sim\pi^{X}$.
+Although we do not require it in the analysis, it is interesting to note that the backward step can also be viewed as a continuous-time process.
+Consider ($\mathsf{FH}$) with $Z_{0}\sim\pi^{X}$.
 Then, $Z_{h}\sim\pi^{Y}$, so the joint distribution of $(Z_{0},Z_{h})$ is simply $\pi$.
 This implies that $(Z_{0}|Z_{h}=y)\sim\pi^{X|Y=y}$.
-Imagine there is an SDE reversing the forward heat flow in a sense that if we are initialized deterministically at $Z_{h}\sim\delta_{y}$ at time 00, then the law of the SDE at time $h$ would be $\mathsf{law}(Z_{0}|Z_{h}=y)=\pi^{X|Y=y}$.
+Imagine there is an SDE reversing the forward heat flow in a sense that if we initialize deterministically at $Z_{h}\sim\delta_{y}$ at time 0, then the law of the SDE at time $h$ would be $\mathsf{law}(Z_{0}|Z_{h}=y)=\pi^{X|Y=y}$.
 Then, this SDE would serve as a continuous-time interpolation of the second step.
@@
 SDE $\mathrm{d}Z_{t}=\mathrm{d}B_{t}$ $\mathrm{d}Z_{t}^{\leftarrow}=\nabla\log(\pi^{X}P_{h-t})(Z_{t}^{\leftarrow})\,\mathrm{d}t+\mathrm{d}B_{t}$
-Fokker-Planck $\partial_{t}\mu_{t}=\frac{1}{2}\Delta\mu_{t}$ $\partial_{t}\mu_{t}^{\leftarrow}=-\operatorname{div}\bigl{(}\mu_{t}^{\leftarrow}\nabla\log(\pi^{X}P_{h-t})\bigr{)}+\frac{1}{2}\Delta\mu_{t}^{\leftarrow}$
-Table 1: The Fokker-Planck equations for the forward and backward heat flow describe how the laws of $Z_{t}$ and $Z_{t}^{\leftarrow}$ in ($\mathsf{FH}$) and ($\mathsf{BH}$) evolve over time.
-See §3.2 for details.
+Fokker–Planck $\partial_{t}\mu_{t}=\frac{1}{2}\Delta\mu_{t}$ $\partial_{t}\mu_{t}^{\leftarrow}=-\operatorname{div}\bigl(\mu_{t}^{\leftarrow}\nabla\log(\pi^{X}P_{h-t})\bigr)+\frac{1}{2}\Delta\mu_{t}^{\leftarrow}$
+Table 1: The Fokker–Planck equations for the forward and backward heat flow describe how the laws of $Z_{t}$ and $Z_{t}^{\leftarrow}$ in ($\mathsf{FH}$) and ($\mathsf{BH}$) evolve over time.
+See §C for details.
 Such a time reversal SDE is indeed possible!
@@
 Thus, if we initialize ($\mathsf{BH}$) with $Z_{0}^{\leftarrow}\sim\mu_{i+1}^{Y}$, then the law of $Z_{h}^{\leftarrow}$ corresponds to $\int\pi^{X|Y=y}\,\mathrm{d}\mu_{i+1}^{Y}(y)=\mu_{i+1}^{X}$.
+We mention this reverse process due to its elegant symmetry but it will not be needed in our mixing analysis.
 Remark 1.
 We note that $\mathsf{In\text{-}and\text{-}Out}$ is exactly the proximal sampling scheme [LST21, CCSW22, FYC23] for uniform distributions.
-The proximal sampler with a target density proportional to $\exp(-V(x))$ considers an augmented distribution $\pi(x,y)\propto\exp(-V(x)-\frac{1}{2h}\lVert x-y\rVert^{2})$ and then repeats the following two steps: (1) $y_{i+1}\sim\pi^{Y|X=x_{i}}=\mathcal{N}(x_{i},hI_{d})$ and then (2) $x_{i+1}\sim\pi^{X|Y=y_{i+1}}$.
+The proximal sampler with a target density proportional to $\exp(-V)$ considers an augmented distribution $\pi(x,y)\propto\exp(-V(x)-\frac{1}{2h}\,\lvert x-y\rvert^{2})$ and then repeats the following two steps: (1) $y_{i+1}\sim\pi^{Y|X=x_{i}}=\mathcal{N}(x_{i},hI_{d})$ and then (2) $x_{i+1}\sim\pi^{X|Y=y_{i+1}}$.
 Naïvely, the proximal sampler is implemented by performing rejection sampling, with the Gaussian centered at the minimizer of $\log\pi^{\cdot|Y=y_{i+1}}$ as the proposal.
-Realizing this would require a projection oracle (to $\mathcal{K}$), which is only known to be implementable with $\mathcal{O}(d^{2})$ membership queries.
+Realizing this would require a projection oracle (to $\mathcal{K}$), which is only known to be implementable with $O(d^{2})$ membership queries.
 $\mathsf{In\text{-}and\text{-}Out}$ completely avoids the need for a projection oracle.
 1.2 Results
-Our computational model is the classical general model for convex bodies [GLS88].
+Our computational model is the classical general model for convex bodies [GLS93].
 We assume $\operatorname{vol}\mathcal{K}>0$ in this paper.
@@
 Definition 1 (Convex body oracle).
-A well-defined membership oracle for a convex body $\mathcal{K}\subset\mathbb{R}^{d}$ is given by a point $x_{0}\in\mathcal{K}$, a number $D>0$, with the guarantee that $B_{1}(x_{0})\subseteq\mathcal{K}\subseteq B_{D}(x_{0})$, and an oracle that correctly answers YES or NO to any query of the form “$x\in\mathcal{K}$?”
+A well-defined membership oracle for a convex body $\mathcal{K}\subset\mathbb{R}^{d}$ is given by a point $x_{0}\in\mathcal{K}$, a number $D>0$, with the guarantee that $B_{1}(x_{0})\subseteq\mathcal{K}\subseteq B_{D}(x_{0})$, and an oracle that correctly answers YES or NO to any query of the form $x\in\mathcal{K}$?”
 Definition 2 (Warmness).
-A distribution $\mu$ is $M$-warm with respect to another distribution $\pi$ if for every $x$ in the support of $\pi$, we have $\mathrm{d}\mu(x)\leq M\,\mathrm{d}\pi(x)$.
+A distribution $\mu$ is said to be $M$-warm with respect to another distribution $\pi$ if for every $x$ in the support of $\pi$, we have $\mathrm{d}\mu(x)\leq M\,\mathrm{d}\pi(x)$.
 We now summarize our main result, which is further elaborated in §3.4.
 Below, $\pi^{\mathcal{K}}$ is the uniform distribution over $\mathcal{K}$, and $\mathcal{R}_{q}$ is the Rényi-divergence of order $q$ (see Definition 6).
-Theorem 3 (Informal version of Theorem 27).
+Theorem 3 (Succinct version of Theorem 23).
 For any given $\eta,\varepsilon\in(0,1)$, $q\geq 1$, and any convex body $\mathcal{K}$ given by a well-defined membership oracle, there exist choices of parameters $h,N$ such that $\mathsf{In\text{-}and\text{-}Out}$, starting from an $M$-warm distribution, with probability at least $1-\eta$, returns $X\sim\mu$ such that $\mathcal{R}_{q}(\mu\mathbin{\|}\pi^{\mathcal{K}})\leq\varepsilon$.
-The number of proper steps is $\widetilde{\mathcal{O}}(qd^{2}\Lambda\log^{2}\nicefrac{{M}}{{\eta\varepsilon}})$, and the expected total number of membership queries is $\widetilde{\mathcal{O}}(qMd^{2}\Lambda\log^{6}\nicefrac{{1}}{{\eta\varepsilon}})$, where $\Lambda$ is the largest eigenvalue of the covariance of $\pi^{\mathcal{K}}$.
-Remark 2.
-Despite our guarantee being in the much stronger “metric” of $\mathcal{R}_{q}$ compared to the $\mathsf{TV}$ guarantees of $\mathsf{Ball\ walk}$, we do not have to incur any additional asymptotic complexity.
-To obtain this result, one should choose the following values for the parameters: $h^{-1}=\widetilde{\Theta}(d^{2}\log\frac{qM\Lambda}{\eta}\log\log\frac{1}{\varepsilon})$, $N=\widetilde{\Theta}(\frac{qMd^{2}\Lambda\log^{5}(1/\varepsilon)}{\eta})$.
+The number of proper steps is $\widetilde{O}(qd^{2}\Lambda\log^{2}\nicefrac{{M}}{{\eta\varepsilon}})$, and the expected total number of membership queries is $\widetilde{O}(qMd^{2}\Lambda\log^{6}\nicefrac{{1}}{{\eta\varepsilon}})$, where $\Lambda$ is the largest eigenvalue of the covariance of $\pi^{\mathcal{K}}$.
+Even though our guarantee is in the much stronger “metric” of $\mathcal{R}_{q}$ compared to the $\mathsf{TV}$ guarantees of the $\mathsf{Ball\ walk}$, we do not incur any additional asymptotic complexity.
+To obtain this result, we choose the following values for the parameters: $h^{-1}=\widetilde{\Theta}(d^{2}\log\frac{qM\Lambda}{\eta}\log\log\frac{1}{\varepsilon})$, $N=\widetilde{\Theta}(\eta^{-1}qMd^{2}\Lambda\log^{5}(1/\varepsilon))$.
 See Lemma 13 for more details.
-Finally, while the assumption of warmness for the initialization may seem strong at the outset, for well-rounded convex bodies ($\mathbb{E}_{\pi}[\lVert X\rVert^{2}]\leq C^{2}d$ for some constant $C$), it is possible to generate an $\mathcal{O}(1)$ warm-start with complexity $\widetilde{\mathcal{O}}(d^{3})$.
-See [CV18, KZ25] for details.
+Remark 2 (Warm-start generation).
+While the assumption of warmness for the initialization may seem strong at the outset, for well-rounded convex bodies ($\mathbb{E}_{\pi}[\lvert\cdot\rvert^{2}]\leq C^{2}d$ for some constant $C$), it is possible to generate an $O(1)$ warm-start with complexity $\widetilde{O}(d^{3})$.
+We refer readers to [CV18, KZ25, KV25a] for details.
 We note that for $X\sim\pi^{\mathcal{K}}$,
@@
   1. 1.
-    $\mathsf{KL}(\mu\mathbin{\|}\pi)=\lim_{q\downarrow 1}\mathcal{R}_{q}(\mu\mathbin{\|}\pi)\leq\mathcal{R}_{q}(\mu\mathbin{\|}\pi)\leq\mathcal{R}_{q^{\prime}}(\mu\mathbin{\|}\pi)\leq\mathcal{R}_{\infty}(\mu\mathbin{\|}\pi)=\log\sup\frac{\mathrm{d}\mu}{\mathrm{d}\pi}$, $1<q\leq q^{\prime}$.
+    $\mathsf{KL}(\mu\mathbin{\|}\pi)=\lim_{q\downarrow 1}\mathcal{R}_{q}(\mu\mathbin{\|}\pi)\leq\mathcal{R}_{q}(\mu\mathbin{\|}\pi)\leq\mathcal{R}_{q^{\prime}}(\mu\mathbin{\|}\pi)\leq\mathcal{R}_{\infty}(\mu\mathbin{\|}\pi)=\log\sup\frac{\mathrm{d}\mu}{\mathrm{d}\pi}$ for $1<q\leq q^{\prime}$.
   2. 2.
@@
   4. 4.
-    $\mathcal{W}_{2}^{2}(\mu,\pi)\leq 2C_{\mathsf{PI}}(\pi)\,\chi^{2}(\mu\mathbin{\|}\pi)$ [Liu20] and $C_{\mathsf{PI}}(\pi^{\mathcal{K}})\lesssim\lVert\operatorname{cov}(\pi^{\mathcal{K}})\rVert_{\mathsf{op}}\log d$.
+    $\mathcal{W}_{2}^{2}(\mu,\pi)\leq 2C_{\mathsf{PI}}(\pi)\,\chi^{2}(\mu\mathbin{\|}\pi)$ [Liu20] and $C_{\mathsf{PI}}(\pi^{\mathcal{K}})\lesssim\lVert\operatorname{cov}\pi^{\mathcal{K}}\rVert_{\mathsf{op}}\log d$.
 The query complexity is better if the convex body is (near-)isotropic, i.e., the uniform distribution over the body has (near-)identity covariance.
-This relies on recent estimates of the worst-case Poincaré constant for isotropic log-concave distributions [KLS95, Kla23].
-The condition that the convex body is isotropic can be achieved in practice through a rounding procedure [JLLV21, KZ24].
-See §A for more details.
+This relies on recent estimates of the worst-case Poincaré constant for isotropic logconcave distributions [KLS95, Kla23].
+The condition that the convex body is isotropic can be achieved in practice through a rounding procedure [JLLV21, JLLV24, KZ24].
+See §1.4 for more details.
 Corollary 4.
-Assume that $\pi^{\mathcal{K}}$ is near-isotropic, i.e., the operator norm of its covariance is $\mathcal{O}(1)$.
+Assume that $\pi^{\mathcal{K}}$ is near-isotropic, i.e., the operator norm of its covariance is $O(1)$.
 Under the same setting as above, $\mathsf{In\text{-}and\text{-}Out}$ succeeds with probability $1-\eta$, returning $X\sim\mu$ such that $\mathcal{R}_{q}(\mu\mathbin{\|}\pi^{\mathcal{K}})\leq\varepsilon$.
-The number of proper steps is $\widetilde{\mathcal{O}}(qd^{2}\log^{2}\frac{M}{\eta\varepsilon})$, and the expected total number of membership queries is $\widetilde{\mathcal{O}}(qMd^{2}\log^{6}\tfrac{1}{\eta\varepsilon})$.
+The number of proper steps is $\widetilde{O}(qd^{2}\log^{2}\frac{M}{\eta\varepsilon})$, and the expected total number of membership queries is $\widetilde{O}(qMd^{2}\log^{6}\tfrac{1}{\eta\varepsilon})$.
 Our analysis will in fact show that the bound on the number of proper steps holds for general non-convex bodies and any feasible start in $\mathcal{K}$.
-This is deduced under an $M$-warm start in Corollaries 28 and 29.
+This is deduced under an $M$-warm start in Corollaries 24 and 25.
 We remark that such a bound for non-convex uniform sampling is not known for the $\mathsf{Ball\ walk}$ or the $\mathsf{Speedy\ walk}$.
@@
 For any given $\varepsilon\in(0,1)$ and set $\mathcal{K}\subset B_{D}(0)$ with $\operatorname{vol}(\mathcal{K})>0$, $\mathsf{In\text{-}and\text{-}Out}$ with variance $h$ and $M$-warm initial distribution achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after the following number of iterations:
-$m=\min\begin{cases}\mathcal{O}\bigl{(}qh^{-1}C_{\mathsf{PI}}(\pi^{X})\log\frac{M}{\varepsilon}\bigr{)}&\text{for }q\geq 2\,,\\ \mathcal{O}\bigl{(}qh^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{\log M}{\varepsilon}\bigr{)}&\text{for }q\geq 1\,.\end{cases}$
+$m=\min\begin{cases}O\bigl(qh^{-1}C_{\mathsf{PI}}(\pi^{X})\log\frac{M}{\varepsilon}\bigr)&\text{for }q\geq 2\,,\\ O\bigl(qh^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{\log M}{\varepsilon}\bigr)&\text{for }q\geq 1\,.\end{cases}$
 We have two different convergence results above under (8) and (7).
@@
   • We show that the number of rejections in Line 3 over the entire execution of the algorithm is bounded with high probability.
-To do this, we apply a detailed argument involving local conductance and the convexity of $\mathcal{K}$, which relies on techniques from [BNN06].
+To do this, we apply a detailed argument involving local conductance and the convexity of $\mathcal{K}$, which relies on techniques from [BNN13].
 For this step, we show that with the appropriate choice of variance $h=\widetilde{\Theta}(d^{-2})$ and threshold $N=\widetilde{\Theta}(T\eta^{-1})$, the entire algorithm succeeds with probability $1-\eta$.
 The expected number of rejections is polylogarithmic.
-While each individual component resembles pre-existing work in the literature, in their synthesis we will demonstrate how to interleave past developments in theoretical computer science, optimal transport, and functional analysis.
-The combination of these in this domain yields elegant and surprisingly simple proofs, as well as stronger results.
+While each individual component resembles pre-existing work in the literature, in their synthesis we will demonstrate how to interleave past relevant developments in theoretical computer science, optimal transport, and functional analysis.
+This yields elegant and surprisingly simple proofs, as well as stronger results.
 1.3 Discussion
 No need to be lazy.
-Previous uniform samplers like $\mathsf{Ball\ walk}$ are made lazy (i.e., with probability $1/2$, it does nothing), to ensure convergence to the target stationary distribution.
+Previous uniform samplers like the $\mathsf{Ball\ walk}$ are made lazy (i.e., with probability $1/2$, it does nothing), to ensure convergence to the target stationary distribution.
 However, our algorithm does not need this, as our sampler is shown to directly contract towards the target.
 Unified framework.
-We remark that Theorem 5 places the previously known mixing guarantees for $\mathsf{Ball\ walk},\mathsf{Speedy\ walk}$ in a unified framework.
-Existing tight guarantees for $\mathsf{Speedy\ walk}$ are in TV distance and based on the log-Sobolev constant, assuming an oracle for implementing each step [LV17].
-The known convergence guarantees of $\mathsf{Ball\ walk}$ (see §B for details), namely the mixing time of $\widetilde{\mathcal{O}}(Md^{2}D^{2}\log\frac{1}{\varepsilon})$ for TV distance, are for the composite algorithm [$\mathsf{Speedy\ walk}+$rejection sampling].
-Here $\mathsf{Speedy\ walk}$ records only the accepted steps of $\mathsf{Ball\ walk}$, so its stationary distribution differs slightly from the uniform distribution (and can be corrected with a post-processing step).
+We remark that Theorem 5 places the previously known mixing guarantees for the $\mathsf{Ball\ walk},\mathsf{Speedy\ walk}$ in a unified framework.
+Existing tight guarantees for the $\mathsf{Speedy\ walk}$ are in TV distance and based on the log-Sobolev constant, assuming an oracle for implementing each step [LV24].
+The known convergence guarantees of the $\mathsf{Ball\ walk}$ (see §A for details), namely the mixing time of $\widetilde{O}(Md^{2}D^{2}\log\frac{1}{\varepsilon})$ for TV distance, are for the composite algorithm $\mathsf{Speedy\ walk}+$rejection sampling].
+Here, the $\mathsf{Speedy\ walk}$ records only the accepted steps of the $\mathsf{Ball\ walk}$, so its stationary distribution differs slightly from the uniform distribution (and can be corrected with a post-processing step).
 On the other hand, $\mathsf{In\text{-}and\text{-}Out}$ actually converges to $\pi^{\mathcal{K}}$ without any adjustments and achieves stronger Rényi divergence bounds in the same asymptotic complexity.
@@
 Effective step size.
-The $\mathsf{Ball\ walk}$’s largest possible step size is of order $1/\sqrt{d}$ (see §B) to keep the rejection probability bounded by a constant.
-This bound could also be viewed as an “effective” step size of $\mathsf{In\text{-}and\text{-}Out}$, since the $\ell_{2}$-norm of the Gaussian $\mathcal{N}(0,hI)$ is concentrated around $\sqrt{hd}$ and we will set the variance $h$ of $\mathsf{In\text{-}and\text{-}Out}$ to $\widetilde{\mathcal{O}}(1/d^{2})$, so we have $\sqrt{hd}\approx 1/\sqrt{d}$.
-What has really changed?
-$\mathsf{In\text{-}and\text{-}Out}$ has clear similarities to both $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$.
-What then are the changes that allow us to use continuous-time interpolation?
-One step of $\mathsf{Ball\ walk}$ is [random step ($y\in B_{\delta}(x)$) $+$ Metropolis-filter (accept if $y\in\mathcal{K}$)].
-This filtering is an abrupt discrete step, and it is unclear how to control contraction.
-It could be replaced by a step of $\mathsf{Speedy\ walk}$ ($x\sim\text{Unif}(B_{\delta}(y)\cap\mathcal{K})$).
-Then, each iteration of $\mathsf{In\text{-}and\text{-}Out}$ can be viewed as a Gaussian version of a $\mathsf{Ball\ walk^{\prime}s\ proposal}+\mathsf{Speedy\ walk}$ algorithm.
-How can we compare $\mathsf{In\text{-}and\text{-}Out}$ with $\mathsf{Speedy\ walk}?$ Iterating speedy steps leads to a biased distribution.
-As clarified in Remark 3, one step of (a Gaussian version of) $\mathsf{Speedy\ walk}$ can be understood as a step of backward heat flow.
-Therefore, if one can control the isoperimetric constants of the biased distribution along the trajectory of the backward flow, then contraction of $\mathsf{Speedy\ walk}$ toward the biased distribution will follow from the simultaneous backward analysis.
+The $\mathsf{Ball\ walk}$’s largest possible step size is of order $1/\sqrt{d}$ (see §A) to keep the rejection probability bounded by a constant.
+This bound could also be viewed as an “effective” step size of $\mathsf{In\text{-}and\text{-}Out}$, since the $\ell_{2}$-norm of the Gaussian $\mathcal{N}(0,hI)$ is concentrated around $\sqrt{hd}$ and we will set the variance $h$ of $\mathsf{In\text{-}and\text{-}Out}$ to $\widetilde{O}(d^{-2})$, so we have $\sqrt{hd}\approx 1/\sqrt{d}$.
+[…]
+How can we compare $\mathsf{In\text{-}and\text{-}Out}$ with the $\mathsf{Speedy\ walk}?$ Iterating speedy steps leads to a biased distribution.
+One step of (a Gaussian version of) the $\mathsf{Speedy\ walk}$ can be understood as a step of backward heat flow.
+Therefore, if one can control the isoperimetric constants of the biased distribution along the trajectory of the backward flow, then contraction of the $\mathsf{Speedy\ walk}$ toward the biased distribution will follow from the simultaneous backward analysis.
+Subsequent work.
+The ideas and methods of this paper have led to subsequent progress.
+First, [KZ25] shows that the output guarantee can be strengthened to $\mathcal{R}_{\infty}$ under LSI.
+Second, the diffusion-based approach is extended to the general logconcave sampling in [KV25b], beyond the uniform distribution.
+Lastly, the $\mathcal{R}_{\infty}$-warmness requirement has been relaxed to significantly weaker $\mathcal{R}_{\widetilde{O}(1)}$-warmness in [KV25a].
+1.4 Related work
+Sampling from constrained logconcave distributions is a fundamental task arising in many fields.
+Uniform sampling with convex constraints is its simplest manifestation, which was first studied as a core subroutine for a randomized volume-computation algorithm [DFK91].
+Since then, this fundamental problem has been studied for over three decades [LS90, LS93, KLS97, LV06, BEL18, BDMP17].
+We review these algorithms, grouping them under three categories — geometric random walks, structured samplers, and diffusion-type samplers.
+Below, $\mathcal{K}$ is convex.
+Geometric random walks.
+We discuss two geometric random walks – $\mathsf{Ball\ walk}$ [LS93, KLS97] and $\mathsf{Hit\text{-}and\text{-}Run}$ [Smi84, Lov99].
+The $\mathsf{Ball\ walk}$ is a simple metropolized random walk; it draws $y$ uniformly at random from a ball of radius $\delta$ centered at a current point $x$, and moves to $y$ if $y\in\mathcal{K}$ and stays at $x$ otherwise.
+In the literature, the $\mathsf{Ball\ walk}$ actually refers to a composite algorithm consisting of $\mathsf{Speedy\ walk}+$ rejection sampling], where the $\mathsf{Speedy\ walk}$ records only the accepted steps of the $\mathsf{Ball\ walk}$ (see §A for details).
+The step size $\delta$ should be set to $O(d^{-1/2})$ to avoid stepping outside of $\mathcal{K}$.
+[KLS97] showed that the $\mathsf{Ball\ walk}$ needs $\widetilde{O}(Md^{2}D^{2}\log\frac{1}{\varepsilon})$ membership queries to be $\varepsilon$-close to $\pi^{\mathcal{K}}$ in $\mathsf{TV}$, where $D$ is the diameter of $\mathcal{K}$, and the warmness parameter $M$ measures the closeness of the initial distribution to the target uniform distribution $\pi^{\mathcal{K}}$.
+$\mathsf{Hit\text{-}and\text{-}Run}$ is another zeroth-order algorithm that needs no step size; it picks a uniform random line $\ell$ passing a current point, and move to a uniform random point on $\ell\cap\mathcal{K}$.
+[LV06] shows that, if we define the second moment as $R^{2}:=\mathbb{E}_{\pi^{\mathcal{K}}}[\lvert X-\mathbb{E}X\rvert^{2}]$, then $\mathsf{Hit\text{-}and\text{-}Run}$ requires $O(d^{2}R^{2}\log\frac{M}{\varepsilon})$ queries.
+Notably, this algorithm has a poly-logarithmic dependence on $M$ as opposed to the $\mathsf{Ball\ walk}$.
+Both algorithm are affected by skewed shape of $\mathcal{K}$ (i.e., large $D$ or $R$), so these samplers are combined with pre-processing step called rounding.
+This procedure finds a linear transformation that makes the geometry of $\mathcal{K}$ less skewed and so more amenable to sampling.
+In literature, there exists a randomized algorithm [JLLV21] that rounds $\mathcal{K}$ and generates a good warm start (i.e., $M=O(1)$), with the $\mathsf{Ball\ walk}$ used as a core subroutine.
+We refer readers to [KZ24] for a streamlined proof, with $\mathsf{In\text{-}and\text{-}Out}$ used as a main sampler.
+This algorithm takes up $\widetilde{O}(d^{3.5})$ queries in total, and from a well-rounded position with a good warm start, the $\mathsf{Ball\ walk}$ only needs $\widetilde{O}(d^{2}\log\frac{1}{\varepsilon})$ queries to sample from $\pi^{\mathcal{K}}$.
+Structured samplers.
+The aforementioned samplers based on geometric random walks require only access to the membership oracle of the convex body without any additional structural assumptions.
+The alternate paradigm of geometry-aware sampling attempts to exploit the structure of convex constraints, with the aim of expediting the convergence of the resultant sampling schemes.
+One common assumption is to make available a self-concordant barrier function $\phi$ which has regularity on its high-order derivatives and blows up when approaching the boundary $\partial\mathcal{K}$.
+The Hessian of $\phi$ encodes the local geometry of the constraint, and the samplers often work directly with $\nabla^{2}\phi$.
+The first canonical example of such a zeroth-order sampler is $\mathsf{Dikin\ walk}$ used when $\mathcal{K}$ is given by $m$ linear constraints [KN12]; it draws a uniform sample from an ellipsoid (characterized by $\nabla^{2}\phi$) of fixed radius around a current point, and is often combined with a Metropolis adjustment.
+[KN12] shows that $\mathsf{Dikin\ walk}$ mixes in $O(md\log\frac{M}{\varepsilon})$ steps, although each iteration is slightly more expensive than one membership query.
+This algorithm requires no rounding, but still needs a good warm-start, which can be achieved by an annealing-type algorithm using $\widetilde{O}(md)$ iterations of $\mathsf{Dikin\ walk}$ [KV24].
+$\mathsf{Riemannian\ Hamiltonian\ Monte\ Carlo}$ is a structured sampler that exploits the first-order information of the potential (i.e., $\nabla\log(1/\pi)$) [GC11]; its proposal is given as the solution to the Hamilton’s ODE equation, followed by the Metropolis-filter.
+In the linear-constraint setting above, it requires $O(md^{2/3}\log\frac{M}{\varepsilon})$ many iterations to achieve $\varepsilon$-close distance to $\pi^{\mathcal{K}}$ [LV18].
+This sampler is further analyzed for practical ODE solvers [KLSV23] and for more sophisticated self-concordant barriers [GKV24].
+Similarly, $\mathsf{Mirror\ Langevin}$ [ZPFP20, Jia21, AC21, LTVW22] is a class of algorithms which converts the constrained problem into an unconstrained one obtained by considering the pushforward of the constrained space by $\nabla\phi$.
+The algorithm can also be metropolized [SWW24].
+The best known rate for this algorithm is $\widetilde{O}(d\log\frac{1}{\varepsilon})$ under some strong assumptions on $\phi$.
+Diffusion-based samplers.
+Samplers based on discretizations of Itô diffusions, stochastic processes which rapidly mix to $\pi$ in continuous time, have long been used for sampling without constraints [BGHM95, DT12, Dal17, Che25].
+While the underlying stochastic processes generalize easily to constrained settings, the discretization analysis relies crucially on the smoothness of the target distribution.
+This is clearly impossible to achieve in the constrained setting, so some techniques are required to circumvent this difficulty.
+These algorithms, however, generalize easily to the more general problem of sampling from distributions of the form $\tilde{\pi}^{X}\propto e^{-f}\mathds{1}_{\mathcal{K}}$, by incorporating first-order information from $f$.
+The first approach for adapting diffusion-based samplers [BEL18, Leh23] iterates a two-step procedure.
+First, a random step is taken, with $x_{k+1/2}\sim\mathcal{N}(x_{k},2hI_{d})$ for some appropriately chosen step hh, [^2] A gradient step can be added in the more general case, for sampling from $\tilde{\pi}^{X}$.
+and then project it to $\mathcal{K}$, i.e., $x_{k+1}=\mathsf{proj}_{\mathcal{K}}(x_{k+1/2})$.
+The complexity is given in terms of queries to a projection oracle, each call to which can be implemented with a polynomial number of membership oracle queries; a total of $\widetilde{O}(\nicefrac{{d^{2}D^{3}}}{{\varepsilon^{4}}})$ queries are needed to be $\varepsilon$-close in $\mathcal{W}_{2}$ to $\pi^{X}$.
+Another approach, which uses an algorithmically designed “soft” penalty instead of a projection, was proposed in [GHZ24], and achieves a rate estimate of $\widetilde{O}(\nicefrac{{d}}{{\varepsilon^{10}}})$.
+A second approach, suggested by [BDMP17], considers a different proximal scheme, which performs a “soft projection” onto $\mathcal{K}$, by taking steps like $\mathcal{N}((1-h\lambda^{-1})x_{k}+h\,\mathsf{proj}_{\mathcal{K}}(x_{k}),2hI_{d})$.
+It is called Moreau–Yosida regularized Langevin, named after an analogous regularization scheme for constrained optimization.
+This scheme also relies on access to a projection oracle for $\mathcal{K}$, and quantifies their query complexity accordingly.
+Their final rate estimate is $\widetilde{O}(\nicefrac{{d^{5}}}{{\varepsilon^{6}}})$ to be $\varepsilon$-close in $\mathsf{TV}$ distance to $\pi^{X}$.
+Observing the prior work integrating diffusion-based sampling with convex constraints, the dependence on the key parameters $d,\varepsilon$, while polynomial, are many orders worse than the rates for zeroth-order samplers such as $\mathsf{Ball\ walk}$ and $\mathsf{Hit\text{-}and\text{-}Run}$.
+In contrast, our analysis not only recovers but in some sense surpasses the known rates for $\mathsf{Ball\ walk}$ and $\mathsf{Hit\text{-}and\text{-}Run}$, while harmonizing well with the continuous-time perspective of diffusions.
+Proximal schemes for sampling.
+The Gibbs sampling scheme used in this paper was inspired by the restricted Gaussian oracle introduced in [LST21] (in turn inspired by Gaussian cooling [CV18]), which alternately iterates between a pure Gaussian step, and a “proximal” step (which we elaborate in our exposition).
+This scheme was given novel interpretations by [CCSW22], which showed that it interpolates the forward and backward heat flows, in the sense defined by [KP23].
+The backward heat flow itself is intimately related to stochastic localization schemes, invented and popularized in [Eld13, Che21].
+This formulation proved surprisingly powerful, allowing many existing rates in unconstrained sampling to be recovered from a relatively simple analysis.
+This was further extended by [FYC23] to achieve the current state-of-the-art rate in unconstrained sampling.
+Finally, [GLL+23] suggests that this could be applied to tackle some constrained problems.
+However, the assumptions in this final mentioned work are not compatible with the uniform sampling problem on general convex bodies.
 2 Preliminaries
-[…]
-To quantify the convergence rate, we introduce some common divergences between distributions.
+Unless otherwise specified, we will use $\lvert\cdot\rvert$ for the $2$-norm on $\mathbb{R}^{d}$.
+We write $a=O(b)$, $a\lesssim b$ to mean that $a\leq cb$ for some universal constant $c>0$.
+Similarly, we write $a\gtrsim b,a=\Omega(b)$ for $a\geq cb$, while $a=\Theta(b)$ means $a\lesssim b,b\lesssim a$ simultaneously.
+We will also use $a=\widetilde{O}(b)$ to denote $a=O(b\operatorname{polylog}b)$.
+Lastly, we will use measure and density interchangeably when there is no confusion.
+To quantify the convergence rate, we recall common notions of divergence between distributions.
 Definition 6 (Distance and divergence).
@@
 where $\mathsf{FI}(\mu\mathbin{\|}\nu):=\mathbb{E}_{\mu}[\lvert\nabla\log\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\rvert^{2}]$ is the Fisher information of $\mu$ with respect to $\nu$.
-We state two important lemmas which are needed for our proofs. The first is the data-processing inequality for Rényi divergence and $f$-divergence, given below.
+We state two important lemmas which are needed for our proofs. The first is the data-processing inequality (DPI) for Rényi divergence and $f$-divergence, given below.
 Lemma 9 (Data-processing inequality).
@@
 3 Analysis
-We begin this section by proving some introductory lemmas, which will streamline our later exposition.
-We first record two fundamental lemmas, which introduce the mathematical formalism for our analysis.
-The first is the existence of forward and backward heat flows (Lemma 22), which will interpolate each line in Algorithm 1.
-These flow equations describe how the laws of $Z_{t}$ and $Z_{t}^{\leftarrow}$ in ($\mathsf{FH}$) and ($\mathsf{BH}$) evolve respectively over time.
+We begin this section by proving the stationarity of the target $\pi^{X}$.
 Lemma 11.
-The forward heat flow equation with initial distribution $\mu_{0}$ is given by
-$\partial_{t}\mu_{t}=\frac{1}{2}\Delta\mu_{t}\,,$
-and its backward heat flow equation is given by
-$\partial_{t}\mu_{t}^{\leftarrow}=-\operatorname{div}\bigl{(}\mu_{t}^{\leftarrow}\nabla\log(\pi^{X}P_{h-t})\bigr{)}+\frac{1}{2}\Delta\mu_{t}^{\leftarrow}\quad\text{with }\mu_{0}^{\leftarrow}=\mu_{h}\,.$
-These admit (weak) solutions on $[0,h]$ for any initial distribution $\mu_{0}$ with $\frac{\mathrm{d}\mu_{0}}{\mathrm{d}\pi^{X}}\leq M<\infty$.
-One successful iteration of $\mathsf{In\text{-}and\text{-}Out}$ is exactly the same as the composition of running the forward heat flow and then backward heat flow, both for time $h$.
-Lemma 12.
-Let $\mu_{k}^{X}$ be the law of the $k$-th iterate $x_{k}$ of $\mathsf{In\text{-}and\text{-}Out}$.
-If ($\mathsf{FH}$) is initialized with $\mathsf{law}(Z_{0})=\mu_{k}^{X}$, then $\mathsf{law}(Z_{h})=\mu_{k+1}^{Y}$.
-If ($\mathsf{BH}$) is initialized with $\mathsf{law}(Z_{0}^{\leftarrow})=\mu_{k+1}^{Y}$, then $\mathsf{law}(Z_{h}^{\leftarrow})=\mu_{k+1}^{X}$.
+$\pi^{X}$ is stationary under iterations of $\mathsf{In\text{-}and\text{-}Out}$.
 Proof.
-We explicitly show that the forward and backward heat flows indeed interpolate the two discrete steps given in Algorithm 1.
-For the forward part, we have $Z_{h}=Z_{0}+\zeta$ for $\zeta\sim\mathcal{N}(0,hI_{d})$, so
-$\mathsf{law}(Z_{h})=\mathsf{law}(Z_{0})*\mathcal{N}(0,hI_{d})=\mu_{k}^{X}*\mathcal{N}(0,hI_{d})=\mu_{k+1}^{Y}\,.$
-Regarding the backward part, it is known from [CCSW22, Lemma 14] that the construction of the time-reversal SDE ensures that $(Z_{h}^{\leftarrow},Z_{0}^{\leftarrow})$ and $(Z_{0},Z_{h})$ have the same joint distribution, when $Z_{0}\sim\pi^{X}$ (and so $Z_{h}\sim\pi^{Y}$).
-Hence, $\mathsf{law}(Z_{h}^{\leftarrow}|Z_{0}^{\leftarrow}=y)=\mathsf{law}(Z_{0}|Z_{h}=y)=\pi^{X|Y=y}$, where the last equality follows from $(Z_{0},Z_{h})\sim\pi$.
-Since we initialize ($\mathsf{BH}$) with $Z_{0}^{\leftarrow}=y\sim\mu_{k+1}^{Y}$, we have
-$\mathsf{law}(Z_{h}^{\leftarrow})=\int\mathsf{law}(Z_{h}^{\leftarrow}|Z_{0}^{\leftarrow}=y)\,\mu_{k+1}^{Y}(\mathrm{d}y)=\int\pi^{X|Y}(\cdot|y)\,\mu_{k+1}^{Y}(\mathrm{d}y)=\mu_{k+1}^{X}\,,$
-where the last follows from the definition of Line 3. ∎
-Our analysis of these objects consists of two parts: (1) demonstrating the mixing of $\mathsf{In\text{-}and\text{-}Out}$, i.e., how many outer iterations are needed to be sufficiently close to the uniform distribution, and (2) quantifying the failure probability and wasted steps in Line 3.
-For (1), we collect in §3.1 some important implications of functional inequalities, e.g.
-the Poincaré and log-Sobolev inequalities, for the uniform distribution.
+Note that the transition kernel of the forward and backward step is $\pi^{Y|X=\cdot}$ and $\pi^{X|Y=\cdot}$, respectively. Hence, the forward step brings $\pi^{X}$ to $\pi^{Y}$ as seen in
+$\int\pi^{Y|X}(\cdot\,|\,x)\,\pi^{X}(x)\,\mathrm{d}x=\int\pi^{X,Y}(x,\cdot)\,\mathrm{d}x=\pi^{Y}\,.$
+Also, the backward step brings $\pi^{Y}$ to $\pi^{X}$:
+$\int\pi^{X|Y}(\cdot\,|\,y)\,\pi^{Y}(y)\,\mathrm{d}y=\int\pi^{X,Y}(\cdot,y)\,\mathrm{d}y=\pi^{X}\,.$
+Therefore, $\pi^{X}$ is stationary under each iteration of $\mathsf{In\text{-}and\text{-}Out}$. ∎
+Our analysis for $\mathsf{In\text{-}and\text{-}Out}$ consists of two parts: (1) demonstrating its mixing, i.e., how many outer iterations are needed to be sufficiently close to the uniform distribution, and (2) quantifying the failure probability and wasted steps in Line 3.
+For (1), we collect in §3.1 some important implications of functional inequalities, e.g., the Poincaré and log-Sobolev inequalities, for the uniform distribution.
 Then in §3.2, we exploit the flow perspective of the algorithm to obtain the mixing guarantees.
-To this end, we revisit the proofs for the contraction results of forward and backward heat flows, and then extend them to our constrained setting:
-Theorem 13 (Adapted for uniform distributions, [CCSW22]).
-Let $\mu^{X}_{k}$ be the law of the $k$-th output of $\mathsf{In\text{-}and\text{-}Out}$ with initial distribution $\mu_{0}^{X}$.
+To this end, we revisit the proofs for the contraction results of forward heat flows in [CCSW22, KO25].
+Theorem 12.
+Let $\mu^{X}_{k}$ be the law of the $k$-th output of $\mathsf{In\text{-}and\text{-}Out}$ with initial distribution $\mu_{0}^{X}$ and step size $h>0$.
 Let $C_{\mathsf{LSI}}$ be the (8) constant of the uniform distribution $\pi^{X}$ over $\mathcal{K}$.
 Then, for any $q\geq 1$,
-$\mathcal{R}_{q}(\mu_{k}^{X}\mathbin{\|}\pi^{X})\leq\frac{\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})}{(1+h/C_{\mathsf{LSI}})^{2k/q}}\,.$
-For $C_{\mathsf{PI}}$ the (7) constant of $\pi^{X}$ and any $q\geq 2$,
-$\mathcal{R}_{q}(\mu_{k}^{X}\mathbin{\|}\pi^{X})\leq\begin{cases}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-\frac{2k\log(1+h/C_{\mathsf{PI}})}{q}&\text{if }k\leq\frac{q}{2\log(1+h/C_{\mathsf{PI}})}\,\bigl{(}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-1\bigr{)}\,,\\ (1+h/C_{\mathsf{PI}})^{-2(k-k_{0})/q}&\text{if }k\geq k_{0}:=\lceil\frac{q}{2\log(1+h/C_{\mathsf{PI}})}\,\bigl{(}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-1\bigr{)}\rceil\,.\end{cases}$
+$\mathcal{R}_{q}(\mu_{k}^{X}\mathbin{\|}\pi^{X})\leq\frac{\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})}{(1+h/C_{\mathsf{LSI}})^{k/q}}\,.$
+For $C_{\mathsf{PI}}$ the (7) constant of $\pi^{X}$,
+$\chi^{2}(\mu^{X}_{k}\mathbin{\|}\pi^{X})\leq\frac{\chi^{2}(\mu^{X}_{0}\mathbin{\|}\pi^{X})}{(1+h/C_{\mathsf{PI}})^{k}}\,.$
+Furthermore, for any $q\geq 2$,
+$\mathcal{R}_{q}(\mu_{k}^{X}\mathbin{\|}\pi^{X})\leq\begin{cases}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-\frac{k\log(1+h/C_{\mathsf{PI}})}{q}&\text{if }k\leq\frac{q}{2\log(1+h/C_{\mathsf{PI}})}\,\bigl(\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-1\bigr)\,,\\ (1+h/C_{\mathsf{PI}})^{-(k-k_{0})/q}&\text{if }k\geq k_{0}:=\lceil\frac{q}{2\log(1+h/C_{\mathsf{PI}})}\,\bigl(\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-1\bigr)\rceil\,.\end{cases}$
 The result reduces the problem of obtaining a mixing guarantee to that of demonstrating a functional inequality on the target distribution.
 For this, it is not strictly necessary that $\mathcal{K}$ be convex.
-Convexity of $\mathcal{K}$ is crucial this time unlike (1).
+As for (2), convexity of $\mathcal{K}$ is crucial this time unlike (1).
 We show in §3.3 that the failure probability remains under control by taking a suitable variance $h$ and threshold $N$, and that the expected number of trials per iteration is of order $\log N$, not $N$.
@@
 Let $\mathcal{K}$ be any convex body in $\mathbb{R}^{d}$ presented by a well-defined membership oracle, $\pi^{X}$ the uniform distribution over $\mathcal{K}$, and $\mu$ an $M$-warm initial distribution with respect to $\pi^{X}$.
-For any given $m\in\mathbb{N}$ and $\eta\in(0,1)$, set $Z=\frac{9mM}{\eta}(\geq 9)$, $h=\frac{\log\log Z}{2d^{2}\log Z}$ and $N=Z\log^{4}Z=\widetilde{\mathcal{O}}(\nicefrac{{mM}}{{\eta}})$.
-Then, the failure probability of one iteration of $\mathsf{In\text{-}and\text{-}Out}$ is at most $\eta/m$, and the expected membership queries per iteration is $\mathcal{O}\bigl{(}M\log^{4}\nicefrac{{mM}}{{\eta}}\bigr{)}$.
+For any given $m\in\mathbb{N}$ and $\eta\in(0,1)$, set $Z=\frac{9mM}{\eta}(\geq 9)$, $h=d^{-2}\frac{\log\log Z}{2\log Z}$ and $N=Z\log^{4}Z=\widetilde{O}(\nicefrac{{mM}}{{\eta}})$.
+Then, the failure probability of one iteration of $\mathsf{In\text{-}and\text{-}Out}$ is at most $\eta/m$, and the expected membership queries per iteration is $O(M\log^{4}\nicefrac{{mM}}{{\eta}})$.
 3.1 Functional inequalities
@@
 The most natural ones to consider in this setting are the Poincaré inequality (7) and log-Sobolev inequality (8).
-In §C, we provide a more detailed discussion of how these are related to other important notions of isoperimetry, such as the Cheeger and log-Cheeger inequalities.
+In §B, we provide a more detailed discussion of how these are related to other important notions of isoperimetry, such as the Cheeger and log-Cheeger inequalities.
 Below, we use $\mu,\nu$ to denote two arbitrary probability measures over $\mathbb{R}^{d}$.
@@
 Assume that $\nu$ satisfies (7) with parameter $C_{\mathsf{PI}}(\nu)$. For any probability measure $\mu$ over $\mathbb{R}^{d}$ with $\mu\ll\nu$, it holds that
-$\chi^{2}(\mu\mathbin{\|}\nu)\leq\frac{C_{\mathsf{PI}}(\nu)}{2}\,\mathbb{E}_{\nu}\bigl{[}\bigl{\|}\nabla\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\bigr{\|}^{2}\bigr{]}\,.$
+$\chi^{2}(\mu\mathbin{\|}\nu)\leq\frac{C_{\mathsf{PI}}(\nu)}{2}\,\mathbb{E}_{\nu}\bigl[\big\lvert\nabla\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\big\rvert^{2}\bigr]\,.$
 The Poincaré inequality implies functional inequalities for the Rényi divergence.
-Lemma 16 ([VW19, Lemma 9]).
+Lemma 15 ([VW23, Lemma 9]).
 Assume that $\nu$ satisfies (7) with parameter $C_{\mathsf{PI}}(\nu)$. For any $q\geq 2$ and probability measure $\mu$ over $\mathbb{R}^{d}$, it holds that
-$1-\exp(-\mathcal{R}_{q}(\mu\mathbin{\|}\nu))\leq\frac{q\,C_{\mathsf{PI}}(\nu)}{4}\,\mathsf{RF}_{q}(\mu\mathbin{\|}\nu)\,,$
-where $\mathsf{RF}_{q}(\mu\mathbin{\|}\nu):=q\,\mathbb{E}_{\nu}\bigl{[}\bigl{(}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\bigr{)}^{q}\lVert\nabla\log\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\rVert^{2}\bigr{]}\,/\,\mathbb{E}_{\nu}\bigl{[}\bigl{(}\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\bigr{)}^{q}\bigr{]}$ is the Rényi Fisher information of order $q$ of $\mu$ with respect to $\nu$.
+$1-\exp\bigl(-\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\bigr)\leq\frac{q\,C_{\mathsf{PI}}(\nu)}{4}\,\mathsf{RF}_{q}(\mu\mathbin{\|}\nu)\,$,
+where $\mathsf{RF}_{q}(\mu\mathbin{\|}\nu):=q\,\mathbb{E}_{\nu}[(\frac{\mathrm{d}\mu}{\mathrm{d}\nu})^{q}\,\lvert\nabla\log\frac{\mathrm{d}\mu}{\mathrm{d}\nu}\rvert^{2}]\,/\,\mathbb{E}_{\nu}[(\frac{\mathrm{d}\mu}{\mathrm{d}\nu})^{q}]$ is the Rényi Fisher information of order $q$ of $\mu$ with respect to $\nu$.
 The log-Sobolev inequality paired with the KL-divergence (8) can be understood as a special case of the following inequality [^3] Such inequalities are often called Polyak-Łojasiewicz inequalities, which say for $f:\mathbb{R}^{d}\to\mathbb{R}$, and all $y\in\mathbb{R}^{d}$ that $f(y)\leq c\,\lvert\nabla f(y)\rvert^{2}$ for some constant $c$, if $\min f(x)=0$.
 paired with the $q$-Rényi divergence for $q\geq 1$.
-Lemma 17 ([VW19, Lemma 5]).
+Lemma 16 ([VW23, Lemma 5]).
 Assume that $\nu$ satisfies (8) with parameter $C_{\mathsf{LSI}}(\nu)$. For any $q\geq 1$ and probability measure $\mu$ over $\mathbb{R}^{d}$, it holds that
@@
 Note that $\lim_{q\to 1}\mathcal{R}_{q}=\mathsf{KL}$ and $\mathsf{RF}_{1}=\mathsf{FI}$.
-We have collected below the functional inequalities used to establish the mixing of our algorithm (see §C for a detailed presentation).
+We have collected below the functional inequalities used to establish the mixing of our algorithm (see §B for a detailed presentation).
 Lemma 17.
@@
 The contraction of Algorithm 1 comes from analyzing Lines 2 and 3 through the perspective of heat flows (see §1.1).
-To exploit this view, we first revisit the previous contraction analysis in [CCSW22], which is carried out for distributions with smooth densities.
-Although the uniform distribution is not even continuous, we prove a technical lemma (Lemma 22) that enables us to extend previously known results to the uniform distribution.
-Lastly, combining the previous results with our technical lemma, we obtain clean contraction results of Algorithm 1 toward the uniform distribution $\pi^{X}$ in Theorem 23.
-Part I: Contraction analysis for smooth distributions.
-In this part, we review the contraction results for heat flow and its time-reversal [CCSW22], which are intimately connected with our algorithm.
-We also provide key technical ingredients needed for its proof, such as the computations for measures evolving under simultaneous forward/backward heat flows.
-We refer interested readers to §D for additional details.
-Only in Part I, we assume that $\nu$ denotes a probability measure with smooth density.
-Forward heat flow.
-We begin by introducing the “heat flow” equation (or also known as the Fokker-Planck equation), which describes the evolution of the law of $Z_{t}$ under ($\mathsf{FH}$),
-$\partial_{t}\mu_{t}=\frac{1}{2}\,\Delta\mu_{t}=\frac{1}{2}\operatorname{div}(\mu_{t}\nabla\log\mu_{t})\,.$ ($\mathsf{FP\text{-}FH}$)
-It is well known that one can realize this equation in discrete time through a Gaussian transition density, in the sense that, for $\mu_{h}$ (the solution at time $h>0$ to ($\mathsf{FP\text{-}FH}$) with initial condition $\mu_{0}$), and for any smooth function $f:\mathbb{R}^{d}\to\mathbb{R}$,
-$\mathbb{E}_{\mu_{h}}[f(x)]=\mathbb{E}_{\mu_{0}}[P_{h}f(x)]\,,$
-where $P_{h}f(x)=\mathbb{E}_{\mathcal{N}(x,hI_{d})}[f]$.333$\{P_{h}\}_{h\geq 0}$ is often called the heat semigroup.
-By this we can formally identify $\mu_{h}=\mu_{0}P_{h}$, and also write $\mu_{h}$ for the law of $Z_{h}$, where $\{Z_{h}\}_{h\geq 0}$ solves ($\mathsf{FH}$).
-Backward heat flow.
-Although there are many ways to define a “reversal” of $P_{h}$, we will use the notion of adjoint introduced by [KP21], which is the most immediately useful.
-[…]
-$\partial_{t}\mu_{t}^{\leftarrow}=-\operatorname{div}\bigl{(}\mu_{t}^{\leftarrow}\nabla\log(\nu P_{h-t})\bigr{)}+\frac{1}{2}\Delta\mu_{t}^{\leftarrow}\quad\text{for }t\in[0,h]\,.$ ($\mathsf{FP\text{-}BH}$)
-Write $\mu_{t}^{\leftarrow}=\mu_{0}^{\leftarrow}Q_{t}^{\nu,h}$, where $\{Q_{t}^{\nu,h}\}_{t\in[0,h]}$ is a family of transition densities.
-Write $\mathbf{P}_{0,h}$ for the joint distribution of the $(Z_{0},Z_{h})$-marginals of ($\mathsf{FH}$), when $Z_{0}\sim\nu$, and $\mathbf{P}_{0|h}$ for the conditional.
-Note that $\mathbf{P}_{h|0}(\cdot|x)=\mathcal{N}(x,hI_{d})$.
-It is also known that ($\mathsf{FP\text{-}BH}$) gives a time-reversal of the heat equation at the SDE level, in the sense that we can interpret $\delta_{x}Q_{h}^{\nu,h}=\mathbf{P}_{0|h}(\cdot|Z_{h}=x)$.
-Thus $\mu_{0}^{\leftarrow}Q_{h}^{\nu,h}=\int\mathbf{P}_{0|h}(\cdot|Z_{h}=x)\,\mu_{0}^{\leftarrow}(\mathrm{d}x)$, and $\nu P_{h}Q_{t}^{\nu,h}=\nu P_{h-t}$ for all $t\in[0,h]$.
-The ultimate purpose of this machinery is to affirm our earlier description of the Gibbs sampling procedure as alternating forward and backward heat flows.
-Indeed, notice that, if $\mu_{i}^{X}$ is the law of the iterate at some iteration $i$, then $\mu_{i}^{X}P_{h}$ is precisely $\mu_{i+1}^{Y}$ under our scheme, while $(\mu_{i}^{X}P_{h})Q_{h}^{\pi^{X},h}$ is $\mu_{i+1}^{X}$, assuming $Q_{h}^{\pi^{X},h}$ is well defined for non-smooth measures $\pi^{X}$.
-Thus, while Algorithm 1 is implemented via discrete steps, it can be exactly analyzed through arguments in continuous time.
-We shall see the benefits of this shortly.
-Instead of considering the change in metrics along the evolution of $\mu P_{t}$ with respect to “fixed” $\nu$, it will be useful to consider the simultaneous evolution of $\mu P_{t},\nu P_{t}$ (and similarly $\mu Q_{t},(\nu P_{h})Q_{t}$).
-This type of computation was carried out for specific metrics in earlier work [VW19, CCSW22].
-The following is a more generalized form of one appearing in [YFL+23, Lemma 2].
-In the lemma below, we consider an arbitrary diffusion equation with corresponding Fokker-Planck equation:
-$\mathrm{d}X_{t}=b_{t}(X_{t})\,\mathrm{d}t+\,\mathrm{d}B_{t}\,\quad\text{and}\quad\partial_{t}\mu_{t}=-\nabla\cdot(b_{t}\mu_{t})+\frac{1}{2}\Delta\mu_{t}\,$ (3.1)
-where $b_{t}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is smooth, $X_{t}\in\mathbb{R}^{d}$, and $\mu_{t}=\mathsf{law}(X_{t})$ if $X_{0}\sim\mu_{0}$.
-Lemma 19 (Decay along forward/backward heat flows).
-Let $(\mu_{t})_{t\geq 0},(\nu_{t})_{t\geq 0}$ denote the laws of the solutions to (3.1) starting at $\mu_{0},\nu_{0}$ respectively. Then, for any differentiable function $g$,
-$\partial_{t}g\bigl{(}D_{f}(\mu_{t}\mathbin{\|}\nu_{t})\bigr{)}=-\frac{1}{2}\,g^{\prime}\bigl{(}D_{f}(\mu_{t}\mathbin{\|}\nu_{t})\bigr{)}\times\mathbb{E}_{\mu_{t}}\Bigl{\langle}\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)},\nabla\log\frac{\mu_{t}}{\nu_{t}}\Bigr{\rangle}\,.$
+Unlike prior works on uniform sampling, we do not bound the $s$-conductance of this chain.
+Instead, we use the “calculus” of the space of probability measures; i.e., for some probability divergence $D$, we compute $\partial_{t}D(\mu*\gamma_{t}\mathbin{\|}\pi^{X}*\gamma_{t})$ in terms of $t$ and isoperimetric constants of $\pi^{X}$.
+The classical data processing inequality (DPI; see Lemma 9) shows that $f$-divergence between two distributions cannot increase when both are convolved with a Gaussian.
+To prove contraction, we will use a strong data-processing inequality (SDPI) for $f$-divergences, which gives a quantitative bound on the contraction.
+Such an inequality was known for some cases including the KL-divergence [Sta59], and was proven for $f(x)=x^{q}-1$ by Klartag and Ordentlich [KO25] who also established the SDPI more generally for this type of Gaussian convolution [AG76, PW16, CPW18], showing that finite fourth moment (i.e., $\mathbb{E}_{\pi^{X}}[\lvert\cdot\rvert^{4}]<\infty$) is a sufficient condition.
+In §C, we show how the identity is derived under suitable regularity assumptions (see Lemma 28).
+Proposition 1 (General de Bruijn identity).
+Let $\mu$ and $\nu$ be probability measures on $\mathbb{R}^{d}$ such that $\mathbb{E}_{\nu}[\lvert\cdot\rvert^{4}]<\infty$ and $\chi^{q}(\mu\mathbin{\|}\nu)<\infty$. Then, for any $t>0$ and $q>1$,
+$\partial_{t}\chi^{q}(\mu_{t}\mathbin{\|}\nu_{t})=-\frac{q(q-1)}{2}\,\mathbb{E}_{\nu_{t}}\bigl[\bigl(\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}\bigr)^{q}\bigl|\nabla\log\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}\bigr|^{2}\bigr]=-\frac{q-1}{2}\,\Bigl\|\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}\Bigr\|^{q}_{L^{q}(\nu_{t})}\mathsf{RF}_{q}(\mu_{t}\mathbin{\|}\nu_{t})\,.$
+Using the general de Bruijn identity with $\mathcal{R}_{q}=\frac{1}{q-1}\log(1+\chi^{q})$, the chain rule results in
+$\partial_{t}\mathcal{R}_{q}(\mu_{t}\mathbin{\|}\nu_{t})=\frac{1}{q-1}\,\frac{\partial_{t}\chi^{q}(\mu_{t}\mathbin{\|}\nu_{t})}{1+\chi^{q}(\mu_{t}\mathbin{\|}\nu_{t})}=\frac{1}{q-1}\,\frac{\partial_{t}\chi^{q}(\mu_{t}\mathbin{\|}\nu_{t})}{\lVert\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}\rVert^{q}_{L^{q}(\nu_{t})}}=-\frac{1}{2}\,\mathsf{RF}_{q}(\mu_{t}\mathbin{\|}\nu_{t})\,.$ (3.1)
+We also obtain $\partial_{t}\chi^{2}(\mu_{t}\mathbin{\|}\nu_{t})=-\mathbb{E}_{\nu_{t}}[\lvert\nabla\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\nu_{t}}\rvert^{2}]$.
+Before proceeding, we need a property of the stability of functional inequalities under the heat flow.
+Lemma 18 (Functional inequalities under Gaussian convolutions, [Cha04, Corollary 13]).
+The following inequality holds for any $t>0$ and $\pi$ with finite log-Sobolev and Poincaré constants,
+$C_{\mathsf{PI}}(\pi_{t})\leq C_{\mathsf{PI}}(\pi)+t\,,\qquad\text{and}\qquad C_{\mathsf{LSI}}(\pi_{t})\leq C_{\mathsf{LSI}}(\pi)+t\,.$
+We can now deduce contraction under the heat flow.
+Lemma 19 (Contraction under functional inequalities).
+Let $\mu,\nu$ be probability measures on $\mathbb{R}^{d}$ such that $\mathcal{R}_{q}(\mu\mathbin{\|}\nu)<\infty$. Then, for any $t>0$ and $q>1$,
+$\mathcal{R}_{q}(\mu_{t}\mathbin{\|}\nu_{t})\leq\frac{\mathcal{R}_{q}(\mu\mathbin{\|}\nu)}{(1+t/C_{\mathsf{LSI}}(\nu))^{1/q}}\,.$
+Also, $\chi^{2}(\mu_{t}\mathbin{\|}\nu_{t})\leq\frac{\chi^{2}(\mu\mathbin{\|}\nu)}{(1+t/C_{\mathsf{PI}}(\nu))}$. Furthermore, for any $q\geq 2$,
+$\mathcal{R}_{q}(\mu_{t}\mathbin{\|}\nu_{t})\leq\begin{cases}\mathcal{R}_{q}(\mu\mathbin{\|}\nu)-\frac{\log(1+t/C_{\mathsf{PI}}(\nu))}{q}&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\geq 1\,,\\ \frac{\mathcal{R}_{q}(\mu\mathbin{\|}\nu)}{(1+t/C_{\mathsf{PI}}(\nu))^{1/q}}&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)<1\,.\end{cases}$
 Proof.
-The case where $g\neq\mathsf{id}$ is an application of the chain rule, so it suffices to take $g=\mathsf{id}$ and simply differentiate an $f$-divergence.
-For brevity, we drop the variable $x$ of functions involved, and proceed as follows:
-$\partial_{t}D_{f}(\mu_{t}\mathbin{\|}\nu_{t})$ =$=\int\Bigl{\{}\bigl{(}f\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\partial_{t}\nu_{t}+\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\bigl{(}\frac{\mu_{t}}{\nu_{t}}\bigr{)}^{\prime}\nu_{t}\Bigr{\}}\,\mathrm{d}x$
-=$=\int\Bigl{\{}\partial_{t}\nu_{t}\Bigl{(}\bigl{(}f\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}-\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\Bigr{)}+\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\partial_{t}\mu_{t}\Bigr{\}}\,\mathrm{d}x$
-=$\underset{(i)}{=}\int\bigl{[}-\nabla\cdot(b_{t}\nu_{t})+\frac{1}{2}\Delta\nu_{t}\bigr{]}\Bigl{(}\bigl{(}f\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}-\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\Bigr{)}\,\mathrm{d}x$
-$\qquad+\int\bigl{[}-\nabla\cdot(b_{t}\mu_{t})+\frac{1}{2}\Delta\mu_{t}\bigr{]}\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\,\mathrm{d}x\,,$
-where in $(i)$ we substitute the F-P equation from (3.1).
-Integrating by parts (i.e., $\int f\operatorname{div}(\mathbf{G})=-\int\langle\nabla f,\mathbf{G}\rangle$ for a real-valued function $f$ and vector-valued function $\mathbf{G}$), we have that
-$\int\bigl{[}-\nabla\cdot(b_{t}\nu_{t})\bigr{]}\bigl{(}f\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\,\mathrm{d}x=\int\Bigl{\langle}b_{t}\nu_{t},\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\nabla\frac{\mu_{t}}{\nu_{t}}\Bigr{\rangle}\,\mathrm{d}x\,.$ (3.2)
-On the other hand, we have that
-$-\int\bigl{[}-\nabla\cdot(b_{t}\nu_{t})\bigr{]}\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\,\mathrm{d}x=-\int\Bigl{\langle}b_{t}\nu_{t},\frac{\mu_{t}}{\nu_{t}}\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}+\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\nabla\frac{\mu_{t}}{\nu_{t}}\Bigr{\rangle}\,\mathrm{d}x\,.$
-The second term cancels with the RHS of (3.2). We have a similar cancellation for the $\frac{1}{2}\Delta\nu_{t}$ term:
-$\int\frac{1}{2}\Delta\nu_{t}\,\bigl{(}f\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\,\mathrm{d}x=-\int\frac{1}{2}\Bigl{\langle}\nabla\nu_{t},\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\nabla\frac{\mu_{t}}{\nu_{t}}\Bigr{\rangle}\,\mathrm{d}x\,,$
-and
-$-\int\frac{1}{2}\Delta\nu_{t}\,\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\,\mathrm{d}x=\int\frac{1}{2}\Bigl{\langle}\nabla\nu_{t},\frac{\mu_{t}}{\nu_{t}}\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}+\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\nabla\frac{\mu_{t}}{\nu_{t}}\Bigr{\rangle}\,\mathrm{d}x\,.$
-Combining these, we are left with
-$\int\bigl{[}-\nabla\cdot(b_{t}\nu_{t})+$ $\frac{1}{2}\Delta\nu_{t}\bigr{]}\Bigl{(}\bigl{(}f\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}-\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\Bigr{)}\,\mathrm{d}x$
-$=-\int\Bigl{\langle}b_{t}\nu_{t}-\frac{1}{2}\nabla\nu_{t},\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\Bigr{\rangle}\,\mathrm{d}x$
-$=-\int\Bigl{\langle}b_{t}\mu_{t}-\frac{1}{2}\mu_{t}\nabla\log\nu_{t},\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\Bigr{\rangle}\,\mathrm{d}x\,.$
-Finally, we note that
-$\int\bigl{[}-\nabla\cdot(b_{t}\mu_{t})+\frac{1}{2}\Delta\mu_{t}\bigr{]}\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\,\mathrm{d}x$ $=\int\Bigl{\langle}b_{t}\mu_{t}-\frac{1}{2}\nabla\mu_{t},\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\Bigr{\rangle}\,\mathrm{d}x$
-$=\int\Bigl{\langle}b_{t}\mu_{t}-\frac{1}{2}\mu_{t}\nabla\log\mu_{t},\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\Bigr{\rangle}\,\mathrm{d}x\,.$
-Putting it all together, noticing that the drift terms cancel, we are left with
-$\partial_{t}D_{f}(\mu_{t}\mathbin{\|}\nu_{t})=-\int\frac{1}{2}\Bigl{\langle}\mu_{t}\nabla\log\frac{\mu_{t}}{\nu_{t}},\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\Bigr{\rangle}\,\mathrm{d}x=-\frac{1}{2}\mathbb{E}_{\mu_{t}}\Bigl{\langle}\nabla\log\frac{\mu_{t}}{\nu_{t}},\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\Bigr{\rangle}\,,$
-which completes the proof. ∎
-To recover the decay result for the $q$-Rényi divergence, one can substitute $g(x)=\frac{1}{q-1}\log x$ and $f(x)=x^{q}-1$.
-For the $\chi^{2}$-divergence, instead substitute $g(x)=x$ and $f(x)=x^{2}-1$.
-From this, we can obtain a single step of decay for the Rényi and $\chi^{2}$-divergences under different functional inequalities.
-Before proceeding, we need a standard lemma on functional inequalities under ($\mathsf{FH}$).
-Lemma 20 (Functional inequalities under Gaussian convolutions, [Cha04, Corollary 13]).
-The following inequality holds for any $\pi$ with finite log-Sobolev and Poincaré constants,
-$C_{\mathsf{PI}}(\pi P_{t})\leq C_{\mathsf{PI}}(\pi)+t\,,\qquad\text{and}\qquad C_{\mathsf{LSI}}(\pi P_{t})\leq C_{\mathsf{LSI}}(\pi)+t\,.$
-Combining the previous two lemmas, we can establish contraction between $\mu P_{h}Q_{h}$ and $\nu$ after one forward/backward iteration.
-Theorem 21 ([CCSW22, Theorem 3 and 4]).
-Assume $\nu$, a measure with smooth density, satisfies (8) with constant $C_{\mathsf{LSI}}$.
-For any $q\geq 1$ and initial distribution $\mu$ with a smooth density, denoting again $Q_{h}:=Q_{h}^{\nu,h}$,
-$\mathcal{R}_{q}(\mu P_{h}Q_{h}\mathbin{\|}\nu)\leq\frac{\mathcal{R}_{q}(\mu\mathbin{\|}\nu)}{(1+h/C_{\mathsf{LSI}})^{2/q}}\,.$
-If $\nu$ satisfies (7) with constant $C_{\mathsf{PI}}$, then it follows that
-$\chi^{2}(\mu P_{h}Q_{h}\mathbin{\|}\nu)\leq\frac{\chi^{2}(\mu\mathbin{\|}\nu)}{(1+h/C_{\mathsf{PI}})^{2}}\,.$
-Moreover, for all $q\geq 2$,
-$\mathcal{R}_{q}(\mu P_{h}Q_{h}\mathbin{\|}\nu)\leq\begin{cases}\mathcal{R}_{q}(\mu\mathbin{\|}\nu)-\frac{2\log(1+h/C_{\mathsf{PI}})}{q}&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\geq 1\,,\\ \frac{\mathcal{R}_{q}(\mu\mathbin{\|}\nu)}{(1+h/C_{\mathsf{PI}})^{2/q}}&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq 1\,.\end{cases}$
-Proof.
-Since the SDE in (3.1) captures the forward heat flow ($\mathsf{FH}$), we set $\mu_{0}$ and $\nu_{0}$ in Lemma 19 to $\mu$ and $\nu$, respectively, obtaining contraction along the forward heat flow as follows: Substituting the $q$-Rényi into Lemma 19, we have, from the definition of the Rényi divergence as $\mathcal{R}_{q}(\mu\mathbin{\|}\nu):=\frac{1}{q-1}\log\mathopen{}\mathclose{{}\left(D_{f}(\mu\mathbin{\|}\nu)+1}\right),$ with $f(x)=x^{q}-1$ and $g(x)=\frac{1}{q-1}\log(x+1)$,
-$\partial_{t}\mathcal{R}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})$ $=-\frac{q}{2}\,\frac{\mathbb{E}_{\mu P_{t}}\Bigl{[}\Bigl{\langle}\nabla\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q-1},\nabla\log\frac{\mu P_{t}}{\nu P_{t}}\Bigr{\rangle}\Bigr{]}}{(q-1)\,\mathbb{E}_{\nu P_{t}}\Bigl{[}\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q}\Bigr{]}}$
-=$\underset{(i)}{=}-\frac{q}{2}\,\frac{\mathbb{E}_{\mu P_{t}}\Bigl{[}\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q-2}\langle\nabla\frac{\mu P_{t}}{\nu P_{t}},\nabla\log\frac{\mu P_{t}}{\nu P_{t}}\rangle\Bigr{]}}{\mathbb{E}_{\nu P_{t}}\Bigl{[}\Bigl{(}\frac{\mu P_{t}}{\nu P_{t}}\Bigr{)}^{q}\Bigr{]}}$
-=$\underset{(ii)}{=}-\frac{q}{2}\,\frac{\mathbb{E}_{\nu P_{t}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{\mu P_{t}}{\nu P_{t}}}\right)^{q}\lVert\nabla\log\frac{\mu P_{t}}{\nu P_{t}}\rVert^{2}}\right]}{\mathbb{E}_{\nu P_{t}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{\mu P_{t}}{\nu P_{t}}}\right)^{q}}\right]}=-\frac{1}{2}\,\mathsf{RF}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})\,,$
-where in $(i)$, we use again that $\nabla\bigl{[}f^{\prime}\bigl{(}\frac{\mu_{t}}{\nu_{t}}\bigr{)}\frac{\mu_{t}}{\nu_{t}}\bigr{]}=\nabla\bigl{(}f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr{)}\cdot\frac{\mu_{t}}{\nu_{t}}+f^{\prime}\bigl{(}\frac{\mu_{t}}{\nu_{t}}\bigr{)}\nabla\frac{\mu_{t}}{\nu_{t}}$, and $(ii)$ uses that $\nabla\frac{\mu P_{t}}{\nu P_{t}}=\frac{\mu P_{t}}{\nu P_{t}}\nabla\log\frac{\mu P_{t}}{\nu P_{t}}$, and the last equality recalls the definition of the Rényi Fisher information.
-This yields
-$\partial_{t}\mathcal{R}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})=-\frac{1}{2}\mathsf{RF}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})\underset{(i)}{\leq}-\frac{1}{q}\frac{\mathcal{R}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})}{C_{\mathsf{LSI}}(\nu P_{t})}\underset{(ii)}{\leq}-\frac{1}{q}\frac{\mathcal{R}_{q}(\mu P_{t}\mathbin{\|}\nu P_{t})}{C_{\mathsf{LSI}}+t}\,,$
-where we used Lemma 17 in $(i)$ and Lemma 20 in $(ii)$. Applying Grönwall’s inequality (Lemma 10),
-$\mathcal{R}_{q}(\mu P_{h}\mathbin{\|}\nu P_{h})\leq\exp\Bigl{(}-\frac{1}{q}\int_{0}^{h}\frac{1}{C_{\mathsf{LSI}}+t}\,\mathrm{d}t\Bigr{)}\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq\frac{\mathcal{R}_{q}(\mu\mathbin{\|}\nu)}{(1+h/C_{\mathsf{LSI}})^{1/q}}\,.$
-Since the SDE (3.1) also captures the backward equation ($\mathsf{BH}$), we set $\mu_{0}$ and $\nu_{0}$ in Lemma 19 to $\mu P_{h}$ and $\tilde{\nu}:=\nu P_{h}$ respectively, obtaining contraction along the backward heat flow:
-$\partial_{t}\mathcal{R}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})$ $=-\frac{1}{2}\,\mathsf{RF}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})$
-$\leq-\frac{1}{q}\,\frac{\mathcal{R}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})}{C_{\mathsf{LSI}}(\tilde{\nu}Q_{t})}\underset{(i)}{\leq}-\frac{1}{q}\,\frac{\mathcal{R}_{q}(\mu P_{h}Q_{t}\mathbin{\|}\tilde{\nu}Q_{t})}{C_{\mathsf{LSI}}+{h-t}}\,,$
-where $(i)$ follows from that $\tilde{\nu}Q_{t}=\nu P_{h}Q_{t}=\nu P_{h-t}$ and $C_{\mathsf{LSI}}(\tilde{\nu}Q_{t})\leq C_{\mathsf{LSI}}+h-t$ due to Lemma 20. Applying Lemma 10 again yields
-$\mathcal{R}_{q}(\mu P_{h}Q_{h}\mathbin{\|}\nu)\leq\frac{\mathcal{R}_{q}(\mu P_{h}\mathbin{\|}\tilde{\nu})}{(1+h/C_{\mathsf{LSI}})^{1/q}}\,.$
-Composing these two inequalities leads to the decay rate claimed in the theorem.
+It is well-known that (8) implies (7), and that (7) ensures finite moments of any order due to exponential integrability (see [BGL14, Proposition 4.4.2]).
+Also, both $\mu_{t}$ and $\nu_{t}$ are smooth due to the Gaussian convolution.
+Under $C_{\mathsf{LSI}}(\nu)<\infty$, (3.1) implies that
+$\partial_{t}\mathcal{R}_{q}(\mu_{t}\mathbin{\|}\nu_{t})=-\frac{1}{2}\,\mathsf{RF}_{q}(\mu_{t}\mathbin{\|}\nu_{t})\underset{(i)}{\leq}-\frac{\mathcal{R}_{q}(\mu_{t}\mathbin{\|}\nu_{t})}{q\,C_{\mathsf{LSI}}(\nu_{t})}\underset{(ii)}{\leq}-\frac{1}{q}\,\frac{\mathcal{R}_{q}(\mu_{t}\mathbin{\|}\nu_{t})}{C_{\mathsf{LSI}}(\nu)+t}\,$,
+where we used Lemma 16 in $(i)$ and Lemma 18 in $(ii)$. Applying Grönwall’s inequality (Lemma 10), we deduce that
+$\mathcal{R}_{q}(\mu_{t}\mathbin{\|}\nu_{t})\leq\exp\Bigl(-\frac{1}{q}\int_{0}^{t}\frac{1}{C_{\mathsf{LSI}}(\nu)+s}\,\mathrm{d}s\Bigr)\,\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq\frac{\mathcal{R}_{q}(\mu\mathbin{\|}\nu)}{(1+t/C_{\mathsf{LSI}}(\nu))^{1/q}}\,.$
 The result in the $\chi^{2}$-divergence can be derived entirely analogously. For instance, the decay from the forward part can be shown as follows:
-$\partial_{t}\chi^{2}(\mu P_{t}\mathbin{\|}\nu P_{t})=-\frac{1}{2}\,\mathbb{E}_{\nu P_{t}}\Bigl{[}\Bigl{\|}\nabla\frac{\mu P_{t}}{\nu P_{t}}\Bigr{\|}^{2}\Bigr{]}\underset{(i)}{\leq}-\frac{\chi^{2}(\mu P_{t}\mathbin{\|}\nu P_{t})}{C_{\mathsf{PI}}(\nu P_{t})}\leq-\frac{\chi^{2}(\mu P_{t}\mathbin{\|}\nu P_{t})}{C_{\mathsf{PI}}+t}\,,$
+$\partial_{t}\chi^{2}(\mu_{t}\mathbin{\|}\nu_{t})=-\frac{1}{2}\,\mathbb{E}_{\nu_{t}}\bigl[\big\lvert\nabla\frac{\mu_{t}}{\nu_{t}}\big\rvert^{2}\bigr]\underset{(i)}{\leq}-\frac{\chi^{2}(\mu_{t}\mathbin{\|}\nu_{t})}{C_{\mathsf{PI}}(\nu_{t})}\leq-\frac{\chi^{2}(\mu_{t}\mathbin{\|}\nu_{t})}{C_{\mathsf{PI}}(\nu)+t}\,$,
 where $(i)$ follows from Lemma 14. Applying Grönwall’s inequality then gives
-$\chi^{2}(\mu P_{h}\mathbin{\|}\nu P_{h})\leq\exp\Bigl{(}-\int_{0}^{h}\frac{1}{C_{\mathsf{PI}}+t}\,\mathrm{d}t\Bigr{)}\,\chi^{2}(\mu\mathbin{\|}\nu)\leq\frac{\chi^{2}(\mu\mathbin{\|}\nu)}{1+h/C_{\mathsf{PI}}}\,.$
-The decay along the backward heat flow in $\chi^{2}$ is entirely analogous to the Rényi case. Then we combine two contraction results from the forward and backward flows, completing the proof.
+$\chi^{2}(\mu_{t}\mathbin{\|}\nu_{t})\leq\exp\Bigl(-\int_{0}^{t}\frac{1}{C_{\mathsf{PI}}(\nu)+s}\,\mathrm{d}s\Bigr)\,\chi^{2}(\mu\mathbin{\|}\nu)\leq\frac{\chi^{2}(\mu\mathbin{\|}\nu)}{1+t/C_{\mathsf{PI}}(\nu)}\,.$
 The result in the $\mathcal{R}_{q}$ under (7) can be shown in a similar manner.
-The only difference is that in forward and backward computations, one should use the functional inequality in Lemma 16 and the following standard inequalities:
-$1-\exp\bigl{(}-\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\bigr{)}\geq\begin{cases}\frac{1}{2}&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\geq 1\,,\\ \frac{1}{2}\mathcal{R}_{q}(\mu\mathbin{\|}\nu)&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq 1\,.\end{cases}\qed$
-Part II: Extension to constrained distributions.
-We now prove a technical lemma that extends the contraction results to constrained distributions.
-This lemma guarantees the existence of weak solutions to two stochastic processes that describe the evolution of distributions involved in Line 2 and 3 in $\mathsf{In\text{-}and\text{-}Out}$, in addition to lower-semicontinuity of $f$-divergence.
-We shall prove it for any measure that is absolutely continuous with respect to $\pi^{X}$, since this imposes no additional technical hurdles.
-Lemma 22.
-Let $\nu$ be a measure, absolutely continuous with respect to the uniform measure $\pi^{X}$. The forward and backward heat flow equations given by
-$\partial_{t}\mu_{t}$ $=\frac{1}{2}\Delta\mu_{t}\,,$
-$\partial_{t}\mu_{t}^{\leftarrow}$ $=-\operatorname{div}\bigl{(}\mu_{t}^{\leftarrow}\nabla\log(\nu P_{h-t})\bigr{)}+\frac{1}{2}\Delta\mu_{t}^{\leftarrow}\quad\text{with }\mu_{0}^{\leftarrow}=\mu_{h}\,,$
-admit solutions on $(0,h]$, and the weak limit $\lim_{t\to h}\mu_{t}^{\leftarrow}=\mu_{h}^{\leftarrow}$ exists for any initial measure $\mu_{0}$ with bounded support.
-Moreover, for any $f$-divergence with $f$ lower semi-continuous,
-$D_{f}(\mu_{h}^{\leftarrow}\mathbin{\|}\nu)\leq\lim_{t\downarrow 0}D_{f}(\mu_{h-t}^{\leftarrow}\mathbin{\|}\nu_{t})\,.$
-Proof.
-The existence of weak solutions for the forward equation is well-known, since $\mu_{0}$ can be weakly approximated by measures with continuous density, for which the heat equation admits a unique solution for all time.
-In particular, the weak solution is $C^{\infty}$ for $t>0$.
-The reverse SDE is more subtle, since $\nabla\log\nu P_{t}$ will in general cease to be Lipschitz as $t\to 0$. On the other hand, for any $h>0$, we can write explicitly
-$\mu_{h}(x)=\frac{1}{(2\pi h)^{d/2}}\int\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2h}\bigr{)}\,\mathrm{d}\mu_{0}(y)\,.$
-If one considers the system started at $\tilde{\mu}_{0}=\mu_{\epsilon}=\nu P_{\epsilon}$ and solve the forward-backward Fokker-Planck equations on times $[0,h-\epsilon]$, then $\tilde{\mu}_{h-\epsilon}=\mu_{h}=\mu_{0}^{\leftarrow}=\tilde{\mu}_{0}^{\leftarrow}$ and
-$\mu_{h-\epsilon}^{\leftarrow}(x)=\tilde{\mu}_{h-\epsilon}^{\leftarrow}(x)=\int\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(x)}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)\,.$
-This follows from that if we consider system started at time $\epsilon>0$, with initial distribution $\mu_{\epsilon}$, then we obtain the above through the Bayesian perspective on the forward and reverse heat semigroups, elaborated in §D.
-We note that $\tilde{\mu}_{h}^{\leftarrow}$ defined below is obviously integrable with respect to $x$, since this is a probability density function of $\mu_{h}Q$, where $Q$ is the backward kernel:
-$\tilde{\mu}_{h}^{\leftarrow}(x):=\int\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2h}\bigr{)}\nu(x)}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)\,.$
-Hence, $\tilde{\mu}_{h}^{\leftarrow}$ is well-defined over the support of $\nu$ almost surely.
-We now show the pointwise convergence of $\tilde{\mu}_{h-\epsilon}^{\leftarrow}$ to $\tilde{\mu}_{h}^{\leftarrow}$ as $\epsilon\to 0$.
-Note that $\nu P_{\epsilon}\to\nu$, as $\nu$ has a bounded support.
-Also, the denominator is independent of $\epsilon$ due to
-$\frac{1}{(2\pi(h-\epsilon))^{d/2}}\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z=\mathcal{N}\bigl{(}0,(h-\epsilon)I\bigr{)}*\nu P_{\epsilon}=\nu*\mathcal{N}(0,hI)\,.$
-Hence, for $\epsilon\leq d^{-1}$,
-$\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z}$ $\leq\Bigl{(}\frac{h}{h-\epsilon}\Bigr{)}^{d/2}\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2h}\bigr{)}}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}$
-$\lesssim\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2h}\bigr{)}}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}\,.$
-Since the last bound is integrable with respect to $\mu_{h}$, the dominated convergence theorem implies
-$\lim_{\epsilon\to 0}\int\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2(h-\epsilon)}\bigr{)}\nu P_{\epsilon}(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)=\int\frac{\exp\bigl{(}-\frac{\lVert x-y\rVert^{2}}{2h}\bigr{)}}{\int\exp\bigl{(}-\frac{\lVert z-y\rVert^{2}}{2h}\bigr{)}\nu(z)\,\mathrm{d}z}\,\mathrm{d}\mu_{h}(y)\,,$
-Thus, the pointwise convergence follows.
-Note that if we take $\nu(x)=\pi^{X}(x)=\frac{\mathds{1}_{\mathcal{K}}(x)}{\operatorname{vol}(\mathcal{K})}$, then $\tilde{\mu}_{h}^{\leftarrow}$ is the distribution of the backward step of our algorithm.
-In particular, this corresponds to first sampling $x\sim\mu_{h}$, then $y\sim Q_{h}^{\nu,h}(\cdot|x)$, which is precisely the law of $\mu_{h}^{\leftarrow}$ given by ($\mathsf{FP\text{-}BH}$).
-As for the second statement, it follows from Scheffé’s lemma [Bil95, Theorem 16.12] that the pointwise convergence of $\mu_{h-\varepsilon}^{\leftarrow}\to\mu_{h}^{\leftarrow}$ leads to its $\mathsf{TV}$-convergence, which in turn implies the weak convergence.
-It follows from lower semicontinuity of $D_{f}$ [AFP00, Theorem 2.34] that the weak convergence ensures $D_{f}(\mu_{h}^{\leftarrow}\mathbin{\|}\nu)\leq\lim_{t\downarrow 0}D_{f}(\mu_{h-t}^{\leftarrow}\mathbin{\|}\nu^{\leftarrow}_{h-t})$.
-∎
-In the sequel, we will only consider $\nu=\pi^{X}$.
-Since the Rényi divergence is a continuous function of the $\chi^{q}$ divergence (see Definition 6), which itself is an $f$-divergence, it enjoys the same lower-semicontinuity properties.
-Using this lower-semicontinuity together with the decay results in Theorem 21, we can easily derive the contraction results of $\mathsf{In\text{-}and\text{-}Out}$ in $\mathcal{R}_{q}$ and $\chi^{q}$ for any $q\geq 1$.
-We remark that this result does not require convexity of $\mathcal{K}$.
-Theorem 23.
-Let $\mu^{X}_{k}$ be the law of the $k$-th output of $\mathsf{In\text{-}and\text{-}Out}$ with initial distribution $\mu_{0}^{X}$.
-Let $C_{\mathsf{LSI}}$ be the (8) constant of the uniform distribution $\pi^{X}$ over $\mathcal{K}$.
-Then, for any $q\geq 1$,
-$\mathcal{R}_{q}(\mu_{k}^{X}\mathbin{\|}\pi^{X})\leq\frac{\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})}{(1+h/C_{\mathsf{LSI}})^{2k/q}}\,.$
-For $C_{\mathsf{PI}}$ the (7) constant of $\pi^{X}$,
-$\chi^{2}(\mu^{X}_{k}\mathbin{\|}\pi^{X})\leq\frac{\chi^{2}(\mu^{X}_{0}\mathbin{\|}\pi^{X})}{(1+h/C_{\mathsf{PI}})^{2k}}\,.$
-Furthermore, for any $q\geq 2$,
-$\mathcal{R}_{q}(\mu_{k}^{X}\mathbin{\|}\pi^{X})\leq\begin{cases}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-\frac{2k\log(1+h/C_{\mathsf{PI}})}{q}&\text{if }k\leq\frac{q}{2\log(1+h/C_{\mathsf{PI}})}\,\bigl{(}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-1\bigr{)}\,,\\ (1+h/C_{\mathsf{PI}})^{-2(k-k_{0})/q}&\text{if }k\geq k_{0}:=\lceil\frac{q}{2\log(1+h/C_{\mathsf{PI}})}\,\bigl{(}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})-1\bigr{)}\rceil\,.\end{cases}$
-Proof.
-Let us set $\mu_{0}=\mu_{0}^{X}$ and $\pi_{0}=\pi^{X}$.
-Then, $\mu_{h}=\mu_{0}^{\leftarrow}=\mu_{1}^{Y}$, $\pi_{h}=\pi_{0}^{\leftarrow}=\pi^{Y}$, and $\mu_{h}^{\leftarrow}=\mu_{1}^{X}$, $\pi_{h}^{\leftarrow}=\pi^{X}$.
-For small $\epsilon>0$, as $\mu_{\epsilon}=(\mu_{0}^{X})_{\epsilon}=\mu_{0}^{X}*\mathcal{N}(0,\epsilon I_{d})$ is $C^{\infty}$-smooth, we can now invoke the decay results with step size $h-\epsilon$ in Theorem 21.
-Thus, for contraction constants $C_{\epsilon}=(1+\frac{h-\epsilon}{C_{\mathsf{LSI}}+\epsilon})^{-2/q}$ and $C_{\epsilon}=(1+\frac{h-\epsilon}{C_{\mathsf{PI}}+\epsilon})^{-2}$ respectively when $\Phi=\mathcal{R}_{q}$ and $\Phi=\chi^{2}$,
-$\Phi(\mu_{h-\epsilon}^{\leftarrow}\mathbin{\|}\pi_{\epsilon})\leq C_{\epsilon}\cdot\Phi(\mu_{\epsilon}\mathbin{\|}\pi_{\epsilon})\leq C_{\epsilon}\cdot\Phi(\mu_{0}\mathbin{\|}\pi_{0})\,,$
-where we used the data-processing inequality for the last inequality. By the second result of Lemma 22, sending $\epsilon\to 0$ leads to
-$\Phi(\mu_{1}^{X}\mathbin{\|}\pi^{X})=\Phi(\mu_{h}^{\leftarrow}\mathbin{\|}\pi_{0})\leq C\cdot\Phi(\mu_{0}\mathbin{\|}\pi_{0})=C\cdot\Phi(\mu_{0}^{X}\mathbin{\|}\pi^{X})\,.$
-Repeating this argument $k$ times completes the proof. ∎
+The only difference is that in forward computation, one should use the functional inequality in Lemma 15 and the following standard inequalities:
+$1-\exp\bigl(-\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\bigr)\geq\begin{cases}\frac{1}{2}&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\geq 1\,,\\ \frac{1}{2}\mathcal{R}_{q}(\mu\mathbin{\|}\nu)&\text{if }\mathcal{R}_{q}(\mu\mathbin{\|}\nu)\leq 1\,.\end{cases}$
+This completes the proof. ∎
+Using the contraction result above, we conclude the mixing time of $\mathsf{In\text{-}and\text{-}Out}$.
+Proof of Theorem 12.
+Note that one iteration of $\mathsf{In\text{-}and\text{-}Out}$ corresponds to the composition of two transition kernels $P_{h}$ and $Q_{h}$, where $P_{h}(x,\cdot)=\mathcal{N}(x,hI_{d})$.
+The forward step ($\mathsf{FH}$) (i.e., transition kernel $P_{h}$) convolves the law $\mu_{k}^{X}$ of the $k$-th iterate with Gaussian $\gamma_{h}$, so the forward step invokes the contraction with rate dependent on $C_{\mathsf{PI}}(\pi^{X})$ or $C_{\mathsf{LSI}}(\pi^{X})$ by Lemma 19.
+Using the DPI (Lemma 9) to address the effect of $Q_{h}$, we obtain that
+$\mathcal{R}_{q}(\mu_{k+1}^{X}\mathbin{\|}\pi^{X})=\mathcal{R}_{q}(\mu_{k+1}^{Y}Q_{h}\mathbin{\|}\pi^{Y}Q_{h})\underset{\text{DPI}}{\leq}\mathcal{R}_{q}(\mu_{k+1}^{Y}\mathbin{\|}\pi^{Y})\leq\frac{\mathcal{R}_{q}(\mu_{k}^{X}\mathbin{\|}\pi^{X})}{(1+h/C_{\mathsf{LSI}}(\pi^{X}))^{1/q}}\,.$
+Induction on $k$ proves the first claim. The other results can be proven in the similar fashion. ∎
 3.3 Failure probability and wasted steps
-We begin by defining a suitable version of local conductance [KLS97].
+We begin by defining a Gaussian version of local conductance [KLS97].
 Definition 20 (Local conductance).
@@
 Refined analysis under a failure condition.
-Going forward, we assume an $M$-warm start as in previous work for uniform sampling algorithms. By induction we have $\frac{\mathrm{d}\mu^{X}_{i}}{\mathrm{d}\pi^{X}}\leq M$ for all $i$.
+Going forward, we assume an $M$-warm start as in previous work for uniform sampling algorithms.
+By induction we have $\frac{\mathrm{d}\mu^{X}_{i}}{\mathrm{d}\pi^{X}}\leq M$ for all $i$.
 Lemma 21 (Propagation of warm-start).
@@
 $\int_{0}^{s}\operatorname{vol}(\partial\mathcal{K}_{z})\,\mathrm{d}z=\operatorname{vol}(\mathcal{K}_{s})-\operatorname{vol}(\mathcal{K})\leq\operatorname{vol}\bigl((1+s)\,\mathcal{K}\bigr)-\operatorname{vol}(\mathcal{K})=\bigl((1+s)^{d}-1\bigr)\operatorname{vol}(\mathcal{K})\,$,
-ensures that $\mathsf{F}$ vanishes at $s=\infty$. Hence, bounding the first term in (3.4) by 00 results in
+ensures that $\mathsf{F}$ vanishes at $s=\infty$. Hence, bounding the first term in (3.3) by 0 results in
 $\int_{\mathcal{K}_{\delta}^{c}}\int_{d(y,\mathcal{K})}^{\infty}\frac{\exp(-\frac{1}{2h}z^{2})}{\sqrt{2\pi h}}\,\mathrm{d}z\,\mathrm{d}y$ $\leq\frac{1}{\sqrt{2\pi h}}\int_{\delta}^{\infty}\exp\bigl(-\frac{s^{2}}{2h}\bigr)\bigl(\underbrace{(1+s)^{d}}_{\leq\exp(sd)}-1\bigr)\operatorname{vol}(\mathcal{K})\,\mathrm{d}s$
@@
 Now we choose a suitable threshold $N$ for bounding the failure probability.
-Following (3.3) in the proof, one can notice that for $y\in\mathcal{K}^{c}_{\delta}$, $\delta=\Omega(1/d)$, and $h=\Theta(d^{-2})$,
+Following (3.2) in the proof, one can notice that for $y\in\mathcal{K}^{c}_{\delta}$, $\delta=\Omega(t/d)$, and $h=\Theta(d^{-2})$,
 $\ell(y)\leq\int_{d(y,\mathcal{K})}^{\infty}\frac{\exp(-\frac{1}{2h}z^{2})}{\sqrt{2\pi h}}\,\mathrm{d}z=\mathbb{P}_{\mathcal{N}(0,h)}(Z\geq\delta)\leq\exp(-\Omega(t^{2}))\,.$
@@
 Intuitively, one can ignore whatever happens in $\mathcal{K}^{c}_{\delta}$, since $\mathcal{K}_{\delta}$ takes up most of measure of $\pi^{Y}$.
-As the number of trials from $\mathcal{K}_{\delta}^{c}$ is at least $\exp(\Omega(t^{2}))$ in expectation, the most straightforward way to ignore algorithmic behaviors from $\mathcal{K}_{\delta}^{c}$ is simply to set the threshold to $N=\widetilde{\mathcal{O}}(\exp(t^{2}))$.
+As the number of trials from $\mathcal{K}_{\delta}^{c}$ is at least $\exp(\Omega(t^{2}))$ in expectation, the most straightforward way to ignore algorithmic behaviors from $\mathcal{K}_{\delta}^{c}$ is simply to set the threshold to $N=\widetilde{O}(\exp(t^{2}))$.
 Even though the threshold is $N$, the expected number of trials is much lower.
@@
 For $\mu_{h}:=\mu*\mathcal{N}(0,hI_{d})$, the failure probability is $\mathbb{E}_{\mu_{h}}[(1-\ell)^{N}]$.
-Since $\mathrm{d}\mu/\mathrm{d}\pi^{X}\leq M$ implies $\mathrm{d}\mu_{h}/\mathrm{d}(\pi^{X})_{h}=\mathrm{d}\mu_{h}/\mathrm{d}\pi^{Y}\leq M$, it follows that
+Since $\mathrm{d}\mu/\mathrm{d}\pi^{X}\leq M$ implies $\mathrm{d}\mu_{h}/\mathrm{d}(\pi^{X})_{h}=\mathrm{d}\mu_{h}/\mathrm{d}\pi^{Y}\leq M$ (as easily checked by the definition of convolution), it follows that
 $\mathbb{E}_{\mu_{h}}[(1-\ell)^{N}]\leq M\,\mathbb{E}_{\pi^{Y}}[(1-\ell)^{N}]\,.$
@@
 $\leq\exp\bigl(-\frac{t^{2}}{2c}+t\bigr)+\frac{\eta}{3mM}+\frac{e^{t}}{N}\,\log\frac{3mM}{\eta}\,$,
-where we used $\operatorname{vol}(\mathcal{K}_{\delta})\subset\operatorname{vol}\bigl{(}(1+\delta)\mathcal{K}\bigr{)}=(1+\delta)^{d}\operatorname{vol}(\mathcal{K})\leq e^{t}\operatorname{vol}(\mathcal{K})$.
+where we used $\operatorname{vol}(\mathcal{K}_{\delta})\subset\operatorname{vol}((1+\delta)\,\mathcal{K})=(1+\delta)^{d}\operatorname{vol}(\mathcal{K})\leq e^{t}\operatorname{vol}(\mathcal{K})$.
 Taking $c=\frac{\log\log Z}{2\log Z}$, $t=\sqrt{8}\log\log Z$, and $N=Z\,(\log Z)^{4}$, we can bound the last line by $\frac{\eta}{mM}$.
 Therefore,
-$\mathbb{E}_{\mu_{h}}[(1-\ell(\cdot))^{N}]\leq M\,\mathbb{E}_{\pi^{Y}}[(1-\ell(\cdot))^{N}]\leq\frac{\eta}{m}\,.$
+$\mathbb{E}_{\mu_{h}}[(1-\ell)^{N}]\leq M\,\mathbb{E}_{\pi^{Y}}[(1-\ell)^{N}]\leq\frac{\eta}{m}\,.$
 We now bound the expected number of trials per iteration.
@@
 $\int_{\mathbb{R}^{d}}\bigl(\frac{1}{\ell}\wedge N\bigr)\,\mathrm{d}\pi^{Y}$ $\leq\int_{\mathcal{K}_{\delta}}\frac{1}{\ell}\,\mathrm{d}\pi^{Y}+N\pi^{Y}(\mathcal{K}_{\delta}^{c})=\frac{\operatorname{vol}(\mathcal{K}_{\delta})}{\operatorname{vol}(\mathcal{K})}+N\pi^{Y}(\mathcal{K}_{\delta}^{c})$
-$\leq e^{t}+N\exp\bigl{(}-\frac{t^{2}}{2c}+t\bigr{)}\leq(\log Z)^{3}+3(\log Z)^{4}=\mathcal{O}\bigl{(}\log^{4}\frac{mM}{\eta}\bigr{)}\,.$
-Therefore, the expected number of trials per step is $\mathcal{O}(M\log^{4}\frac{mM}{\eta})$, and the claim follows since each trial uses one query to the membership oracle of $\mathcal{K}$. ∎
+$\leq e^{t}+N\exp\bigl(-\frac{t^{2}}{2c}+t\bigr)\leq\log^{3}Z+3\log^{4}Z=O\bigl(\log^{4}\frac{mM}{\eta}\bigr)\,.$
+Therefore, the expected number of trials per step is $O(M\log^{4}\frac{mM}{\eta})$, and the claim follows since each trial uses one query to the membership oracle of $\mathcal{K}$. ∎
 3.4 Putting it together
-We can now show that $\mathsf{In\text{-}and\text{-}Out}$ subsumes previous results on uniform sampling from convex bodies (such as $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$), providing detailed versions of the main results in §1.2.
+We can now show that $\mathsf{In\text{-}and\text{-}Out}$ subsumes previous results on uniform sampling from convex bodies (such as the $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$), providing detailed versions of the main results in §1.2.
 We first establish that the query complexity of $\mathsf{In\text{-}and\text{-}Out}$ matches that of the $\mathsf{Ball\ walk}$ under stronger divergences.
@@
 Theorem 23.
-For any given $\eta,\varepsilon\in(0,1)$, $q\geq 1$, $m\in\mathbb{N}$ defined below and any convex body $\mathcal{K}$ given by a well-defined membership oracle, consider $\mathsf{In\text{-}and\text{-}Out}$ (Algorithm 1) with an $M$-warm initial distribution $\mu_{0}^{X}$, $h=(2d^{2}\log\frac{9mM}{\eta})^{-1}$, and $N=\widetilde{\mathcal{O}}(\frac{mM}{\eta})$.
+For any given $\eta,\varepsilon\in(0,1)$, $q>1$, $m\in\mathbb{N}$ defined below and any convex body $\mathcal{K}$ given by a well-defined membership oracle, consider $\mathsf{In\text{-}and\text{-}Out}$ (Algorithm 1) with an $M$-warm initial distribution $\mu_{0}^{X}$, step size $h=(2d^{2}\log\frac{9mM}{\eta})^{-1}$, and threshold $N=\widetilde{O}(\frac{mM}{\eta})$.
 For $\pi^{X}$ the uniform distribution over $\mathcal{K}$,
-  • It achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{\mathcal{O}}(qd^{2}\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{2}\frac{M}{\eta\varepsilon})$ iterations.
-With probability $1-\eta$, the algorithm iterates this many times without failure, using $\widetilde{\mathcal{O}}(qMd^{2}\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{6}\frac{1}{\eta\varepsilon})$ expected number of membership queries in total.
-  • For isotropic $\pi^{X}$, with probability $1-\eta$, the algorithm achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ with $m=\widetilde{\mathcal{O}}(qd^{2}\log^{2}\frac{M}{\eta\varepsilon})$ iterations, using $\widetilde{\mathcal{O}}(qMd^{2}\log^{6}\frac{1}{\eta\varepsilon})$ membership queries in expectation.
+  • It achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{O}(qd^{2}\,\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{2}\frac{M}{\eta\varepsilon})$ iterations.
+With probability $1-\eta$, the algorithm iterates this many times without failure, using $\widetilde{O}(qMd^{2}\,\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{6}\frac{1}{\eta\varepsilon})$ expected number of membership queries in total.
+  • For isotropic $\pi^{X}$, with probability $1-\eta$, the algorithm achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ with $m=\widetilde{O}(qd^{2}\log^{2}\frac{M}{\eta\varepsilon})$ iterations, using $\widetilde{O}(qMd^{2}\log^{6}\frac{1}{\eta\varepsilon})$ membership queries in expectation.
 Proof.
@@
 The $M$-warm start assumption guarantees $\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})\lesssim\log M$.
-Due to $C_{\mathsf{PI}}(\pi^{X})=\mathcal{O}(\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\,\log d)$ (Lemma 18), $\mathsf{In\text{-}and\text{-}Out}$ can achieve $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{\mathcal{O}}(qd^{2}\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{2}\nicefrac{{M}}{{\eta\varepsilon}})$ iterations.
+Due to $C_{\mathsf{PI}}(\pi^{X})=O(\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\,\log d)$ (Lemma 17), $\mathsf{In\text{-}and\text{-}Out}$ can achieve $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ if it iterates at least
+$m\gtrsim qd^{2}\,\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log d\log\frac{M}{\eta\varepsilon}\log\frac{mM}{\eta}\quad\text{times.}$
+Note that $x\geq A\log Bx$ for $A,B\geq 1$ is satisfied when $x\gtrsim A\log AB=\widetilde{O}(A\log B)$. Using this, we set
+$m=\widetilde{O}\bigl(qd^{2}\,\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log\frac{M}{{\eta\varepsilon}}\log\frac{M}{\eta}\bigr)=\widetilde{O}\bigl(qd^{2}\,\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{2}\frac{M}{{\eta\varepsilon}}\bigr)\,.$
 Since each iteration has $\eta/m$-failure probability by Lemma 13, the union bound ensures that the total failure probability is at most $\eta$ throughout $m$ iterations.
-Lastly, each iteration requires $\widetilde{\mathcal{O}}(M\log^{4}\nicefrac{{1}}{{\eta\varepsilon}})$ membership queries in expectation by Lemma 14.
-Therefore, $\mathsf{In\text{-}and\text{-}Out}$ uses $\widetilde{\mathcal{O}}(qMd^{2}\min(D^{2},\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}})\log^{6}\nicefrac{{1}}{{\eta\varepsilon}})$ expected number of membership queries over $m$ iterations.
+Lastly, each iteration requires $\widetilde{O}(M\log^{4}\nicefrac{{1}}{{\eta\varepsilon}})$ membership queries in expectation by Lemma 13.
+Therefore, $\mathsf{In\text{-}and\text{-}Out}$ uses $\widetilde{O}(qMd^{2}\,\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{6}\nicefrac{{1}}{{\eta\varepsilon}})$ expected number of membership queries over $m$ iterations.
 Since $\mathcal{R}_{q}$ is non-decreasing in $q$, we can obtain the desired bound on $\mathcal{R}_{q}$ for $q\in[1,2)$.
-For isotropic $\pi^{X}$, we have $\operatorname{cov}\pi^{X}=I_{d}$, so the claim immediately follows from $C_{\mathsf{PI}}(\pi^{X})=\mathcal{O}(\log d)$ (see Lemma 18). ∎
+For isotropic $\pi^{X}$, we have $\operatorname{cov}\pi^{X}=I_{d}$, so the claim immediately follows from $C_{\mathsf{PI}}(\pi^{X})=O(\log d)$. ∎
 We now show that the number of proper steps is bounded as claimed for general non-convex bodies and any feasible start in $\mathcal{K}$.
 We first establish this result under an $M$-warm start (Theorem 5).
-Proof of Theorem 5..
-By the Rényi-decay under (8) in Theorem 23, $\mathsf{In\text{-}and\text{-}Out}$ can achieve $\varepsilon$-distance to $\pi^{X}$ after $\mathcal{O}\bigl{(}qh^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{\mathcal{R}_{q}(\mu^{X}_{1}\mathbin{\|}\pi^{X})}{\varepsilon}\bigr{)}$ iterations for $q\geq 1$.
+Proof of Theorem 5.
+By the Rényi-decay under (8) in Theorem 12, $\mathsf{In\text{-}and\text{-}Out}$ can achieve $\varepsilon$-distance to $\pi^{X}$ after $O\bigl(qh^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{\mathcal{R}_{q}(\mu^{X}_{1}\mathbin{\|}\pi^{X})}{\varepsilon}\bigr)$ iterations for $q\geq 1$.
 For $q\geq 2$, we use the decay result under (7).
 In this case, $\mathsf{In\text{-}and\text{-}Out}$ decays under two different rates depending on the value of $\mathcal{R}_{q}(\cdot\mathbin{\|}\pi^{X})$.
-It first needs $\mathcal{O}(qh^{-1}C_{\mathsf{PI}}(\pi^{X})\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X}))$ iterations until $\mathcal{R}_{q}(\cdot\mathbin{\|}\pi^{X})$ reaches $1$.
-Then, $\mathsf{In\text{-}and\text{-}Out}$ additionally needs $\mathcal{O}(qh^{-1}C_{\mathsf{PI}}(\pi^{X})\log\frac{1}{\varepsilon})$ iterations, and thus it needs $\mathcal{O}(qh^{-1}C_{\mathsf{PI}}(\pi^{X})\bigl{(}\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})+\log\frac{1}{\varepsilon}\bigr{)})$ iterations in total.
+It first needs $O(qh^{-1}C_{\mathsf{PI}}(\pi^{X})\,\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X}))$ iterations until $\mathcal{R}_{q}(\cdot\mathbin{\|}\pi^{X})$ reaches $1$.
+Then, $\mathsf{In\text{-}and\text{-}Out}$ additionally needs $O(qh^{-1}C_{\mathsf{PI}}(\pi^{X})\log\frac{1}{\varepsilon})$ iterations, and thus it needs $O(qh^{-1}C_{\mathsf{PI}}(\pi^{X})\bigl(\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})+\log\frac{1}{\varepsilon}\bigr))$ iterations in total.
 By substituting $\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})\lesssim\log M$, we complete the proof.
@@
 Corollary 24.
-For any given $\varepsilon\in(0,1)$ and set $\mathcal{K}\subset B_{D}(0)$, $\mathsf{In\text{-}and\text{-}Out}$ with variance $h$ and any feasible start $x_{0}\in\mathcal{K}$ achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{\mathcal{O}}(qh^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{d+D^{2}/h}{\varepsilon})$ iterations.
+For any given $\varepsilon\in(0,1)$ and set $\mathcal{K}\subset B_{D}(0)$, $\mathsf{In\text{-}and\text{-}Out}$ with variance $h$ and any feasible start $x_{0}\in\mathcal{K}$ achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{O}(qh^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{d+D^{2}/h}{\varepsilon})$ iterations.
 Proof.
 We first bound the warmness of $\mu^{X}_{1}$ w.r.t. $\pi^{X}$ when $\mu^{X}_{0}=\delta_{x_{0}}$. One can readily check that
-$\mu_{1}^{X}(x)=\mathds{1}_{\mathcal{K}}(x)\cdot\int\frac{\exp(-\frac{1}{2h}\,\lVert y-x\rVert^{2})\exp(-\frac{1}{2h}\,\lVert y-x_{0}\rVert^{2})}{(2\pi h)^{d/2}\int_{\mathcal{K}}\exp(-\frac{1}{2h}\,\lVert y-x\rVert^{2})\,\mathrm{d}x}\,\mathrm{d}y\,.$
+$\mu_{1}^{X}(x)=\mathds{1}_{\mathcal{K}}(x)\int\frac{\exp(-\frac{1}{2h}\,\lvert y-x\rvert^{2})\exp(-\frac{1}{2h}\,\lvert y-x_{0}\rvert^{2})}{(2\pi h)^{d/2}\int_{\mathcal{K}}\exp(-\frac{1}{2h}\,\lvert y-x\rvert^{2})\,\mathrm{d}x}\,\mathrm{d}y\,.$
 By Young’s inequality, $\lvert y-x\rvert^{2}\leq(\lvert y\rvert+D)^{2}\leq\frac{3}{2}\,\lvert y\rvert^{2}+3D^{2}$ for $x\in\mathcal{K}$. Hence,
@@
 Therefore, $M=\operatorname{ess\,sup}\frac{\mu_{1}^{X}}{\pi^{X}}\leq 2^{d/2}\exp(5h^{-1}D^{2})$.
-By Theorem 5 under (8), $\mathsf{In\text{-}and\text{-}Out}$ needs $\widetilde{\mathcal{O}}(qh^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{d+D^{2}/h}{\varepsilon})$ iterations.
+By Theorem 5 under (8), $\mathsf{In\text{-}and\text{-}Out}$ needs $\widetilde{O}(qh^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{d+D^{2}/h}{\varepsilon})$ iterations.
 ∎
@@
 Corollary 25.
-For any given $\varepsilon\in(0,1)$ and convex body $\mathcal{K}\subset B_{D}(0)$, $\mathsf{In\text{-}and\text{-}Out}$ with variance $h$ and an $M$-warm initial distribution achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{\mathcal{O}}(qh^{-1}D^{2}\log\frac{1}{\varepsilon})$ iterations.
-If $\pi^{X}$ is isotropic, then $\mathsf{In\text{-}and\text{-}Out}$ only needs $\widetilde{\mathcal{O}}(qh^{-1}D\log\frac{d+d^{2}/h}{\varepsilon})$ iterations.
+For any given $\varepsilon\in(0,1)$ and convex body $\mathcal{K}\subset B_{D}(0)$, $\mathsf{In\text{-}and\text{-}Out}$ with variance $h$ and a feasible start $x_{0}\in\mathcal{K}$ achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{O}(qh^{-1}D^{2}\log\frac{1}{\varepsilon})$ iterations.
+If $\pi^{X}$ is isotropic, then $\mathsf{In\text{-}and\text{-}Out}$ only needs $\widetilde{O}(qh^{-1}D\log\frac{d+d^{2}/h}{\varepsilon})$ iterations.
 Proof.
-For convex $\mathcal{K}$, it follows from Lemma 18 that $C_{\mathsf{LSI}}(\pi^{X})=\mathcal{O}(D^{2})$ and $C_{\mathsf{LSI}}(\pi^{X})=\mathcal{O}(D)$ for isotropic $\mathcal{K}$.
+For convex $\mathcal{K}$, it follows from Lemma 17 that $C_{\mathsf{LSI}}(\pi^{X})=O(D^{2})$ and $C_{\mathsf{LSI}}(\pi^{X})=O(D)$ for isotropic $\mathcal{K}$.
 The rest of the proof can be completed in a similar way.
 ∎
-For $h=\tilde{\Theta}(d^{-2})$, $\mathsf{In\text{-}and\text{-}Out}$ requires $\widetilde{\mathcal{O}}(qd^{2}D^{2})$ iterations and in particular $\widetilde{\mathcal{O}}(qd^{2}D)$ iteration for isotropic uniform distributions.
-These results match those of $\mathsf{Speedy\ walk}$ [KLM06, LV17] (see Theorem 31).
+For $h=\tilde{\Theta}(d^{-2})$, $\mathsf{In\text{-}and\text{-}Out}$ requires $\widetilde{O}(qd^{2}D^{2})$ iterations and in particular $\widetilde{O}(qd^{2}D)$ iteration for isotropic uniform distributions.
+These bounds match those of the $\mathsf{Speedy\ walk}$ [KLM06, LV24] (see Theorem 27), albeit with stronger guarantees on the output distribution.
 Acknowledgements.
@@
 This work was supported in part by NSF award 2106444, NSERC through the CGS-D award, and a Simons Investigator award.
-Appendix A Related work
-Sampling from constrained log-concave distributions is a fundamental task arising in many fields.
-Uniform sampling with convex constraints is its simplest manifestation, which was first studied as a core subroutine for a randomized volume-computation algorithm [DFK91].
-Since then, this fundamental problem has been studied for over three decades [LS90, LS93, KLS97, LV03, LV06, BEL18, BDMP17].
-We review these algorithms, grouping them under three categories — geometric random walks, structured samplers, and diffusion-type samplers.
-Below, $\mathcal{K}$ is convex.
-Geometric random walk.
-We discuss two geometric random walks – $\mathsf{Ball\ walk}$ [LS93, KLS97] and $\mathsf{Hit\text{-}and\text{-}Run}$ [Smi84, Lov99].
-$\mathsf{Ball\ walk}$ is a simple metropolized random walk; it draws $y$ uniformly at random from a ball of radius $\delta$ centered at a current point $x$, and moves to $y$ if $y\in\mathcal{K}$ and stays at $x$ otherwise.
-In the literature, $\mathsf{Ball\ walk}$ actually refers to a composite algorithm consisting of [$\mathsf{Speedy\ walk}+$ rejection sampling], where $\mathsf{Speedy\ walk}$ records only the accepted steps of $\mathsf{Ball\ walk}$ (see §B for details).
-The step size $\delta$ should be set to $\mathcal{O}(d^{-1/2})$ to avoid stepping outside of $\mathcal{K}$.
-[KLS97] showed that $\mathsf{Ball\ walk}$ needs $\widetilde{\mathcal{O}}(Md^{2}D^{2}\log\frac{1}{\varepsilon})$ membership queries to be $\varepsilon$-close to $\pi^{\mathcal{K}}$ in $\mathsf{TV}$, where $D$ is the diameter of $\mathcal{K}$, and the warmness parameter $M$ measures the closeness of the initial distribution to the target uniform distribution $\pi^{\mathcal{K}}$.
-$\mathsf{Hit\text{-}and\text{-}Run}$ is another zeroth-order algorithm that needs no step size; it picks a uniform random line $\ell$ passing a current point, and move to a uniform random point on $\ell\cap\mathcal{K}$.
-[LV06] shows that, if we define the second moment as $R^{2}:=\mathbb{E}_{\pi^{\mathcal{K}}}[\lVert X-\mathbb{E}X\rVert^{2}]$, then $\mathsf{Hit\text{-}and\text{-}Run}$ requires $\mathcal{O}(d^{2}R^{2}\log\frac{M}{\varepsilon})$ queries.
-Notably, this algorithm has a poly-logarithmic dependence on $M$ as opposed to $\mathsf{Ball\ walk}$.
-Both algorithm are affected by skewed shape of $\mathcal{K}$ (i.e., large $D$ or $R$), so these samplers are combined with pre-processing step called rounding.
-This procedure finds a linear transformation that makes the geometry of $\mathcal{K}$ less skewed and so more amenable to sampling.
-In literature, there exists a randomized algorithm [JLLV21] that rounds $\mathcal{K}$ and generates a good warm start (i.e., $M=\mathcal{O}(1)$), with $\mathsf{Ball\ walk}$ used as a core subroutine.
-We refer readers to [KZ24] for a streamlined proof, with $\mathsf{In\text{-}and\text{-}Out}$ used as a main sampler.
-This algorithm takes up $\widetilde{\mathcal{O}}(d^{3.5})$ queries in total, and in such position with the good warm start, $\mathsf{Ball\ walk}$ only needs $\widetilde{\mathcal{O}}(d^{2}\log\frac{1}{\varepsilon})$ queries to sample from $\pi^{\mathcal{K}}$.
-Structured samplers.
-The aforementioned samplers based on geometric random walks require only access to the membership oracle of the convex body without any additional structural assumptions.
-The alternate paradigm of geometry-aware sampling attempts to exploit the structure of convex constraints, with the aim of expediting the convergence of the resultant sampling schemes.
-One common assumption is to make available a self-concordant barrier function $\phi$ which has regularity on its high-order derivatives and blows up when approaching the boundary $\partial\mathcal{K}$.
-The Hessian of $\phi$ encodes the local geometry of the constraint, and the samplers often work directly with $\nabla^{2}\phi$.
-The first canonical example of such a zeroth-order sampler is $\mathsf{Dikin\ walk}$ used when $\mathcal{K}$ is given by $m$ linear constraints [KN12]; it draws a uniform sample from an ellipsoid (characterized by $\nabla^{2}\phi$) of fixed radius around a current point, and is often combined with a Metropolis adjustment.
-[KN12] shows that $\mathsf{Dikin\ walk}$ mixes in $\mathcal{O}(md\log\frac{M}{\varepsilon})$ steps, although each iteration is slightly more expensive than one membership query.
-This algorithm requires no rounding, but still needs a good warm-start, which can be achieved by an annealing-type algorithm using $\widetilde{\mathcal{O}}(md)$ iterations of $\mathsf{Dikin\ walk}$ [KV24].
-$\mathsf{Riemannian\ Hamiltonian\ Monte\ Carlo}$ is a structured sampler that exploits the first-order information of the potential (i.e., $\nabla\log(1/\pi)$) [GC11]; its proposal is given as the solution to the Hamilton’s ODE equation, followed by the Metropolis-filter.
-In the linear-constraint setting above, it requires $\mathcal{O}(md^{2/3}\log\frac{M}{\varepsilon})$ many iterations to achieve $\varepsilon$-close distance to $\pi^{\mathcal{K}}$ [LV18].
-This sampler is further analyzed for practical ODE solvers [KLSV23] and for more sophisticated self-concordant barriers [GKV23].
-Similarly, $\mathsf{Mirror\ Langevin}$ [ZPFP20, Jia21, AC21, LTVW22] is a class of algorithms which converts the constrained problem into an unconstrained one obtained by considering the pushforward of the constrained space by $\nabla\phi$.
-The algorithm can also be metropolized [SWW24].
-The best known rate for this algorithm is $\widetilde{\mathcal{O}}(d\log\frac{1}{\varepsilon})$ under some strong assumptions on $\phi$.
-Diffusion-type samplers.
-Samplers based on discretizations of Itô diffusions, stochastic processes which rapidly mix to $\pi$ in continuous time, have long been used for sampling without constraints [BGHM95, DT12, Dal17, Che24].
-While the underlying stochastic processes generalize easily to constrained settings, the discretization analysis relies crucially on the smoothness of the target distribution.
-This is clearly impossible to achieve in the constrained setting, so some techniques are required to circumvent this difficulty.
-These algorithms, however, generalize easily to the more general problem of sampling from distributions of the form $\tilde{\pi}^{X}\propto e^{-f}\mathds{1}_{\mathcal{K}}$, by incorporating first-order information from $f$.
-The first approach for adapting diffusion-based samplers [BEL15, BEL18, Leh23] iterates a two-step procedure.
-First, a random step is taken, with $x_{k+1/2}\sim\mathcal{N}(x_{k},2hI_{d})$ for some appropriately chosen step $h$, [^4] A gradient step can be added in the more general case, for sampling from $\tilde{\pi}^{X}$.
-and then project it to $\mathcal{K}$, i.e., $x_{k+1}=\mathsf{proj}_{\mathcal{K}}(x_{k+1/2})$.
-The complexity is given in terms of queries to a projection oracle, each call to which can be implemented with a polynomial number of membership oracle queries; a total of $\tilde{\mathcal{O}}(\nicefrac{{d^{2}D^{3}}}{{\varepsilon^{4}}})$ queries are needed to be $\varepsilon$-close in $\mathcal{W}_{2}$ to $\pi^{X}$.
-Another approach, which uses an algorithmically designed “soft” penalty instead of a projection, was proposed in [GHZ22], and achieves a rate estimate of $\tilde{\mathcal{O}}(\nicefrac{{d}}{{\varepsilon^{10}}})$.
-A second approach, suggested by [BDMP17], considers a different proximal scheme, which performs a “soft projection” onto $\mathcal{K}$, by taking steps like $\mathcal{N}((1-h\lambda^{-1})x_{k}+h\mathsf{proj}_{\mathcal{K}}(x_{k}),2hI_{d})$.
-It is called Moreau-Yosida regularized Langevin, named after an analogous regularization scheme for constrained optimization.
-This scheme also relies on access to a projection oracle for $\mathcal{K}$, and quantifies their query complexity accordingly.
-Their final rate estimate is $\tilde{\mathcal{O}}(\nicefrac{{d^{5}}}{{\varepsilon^{6}}})$ to be $\varepsilon$-close in $\mathsf{TV}$ distance to $\pi^{X}$.
-Observing the prior work integrating diffusion-based sampling with convex constraints, the dependence on the key parameters $d,\varepsilon$, while polynomial, are many orders worse than the rates for zeroth-order samplers such as $\mathsf{Ball\ walk},\mathsf{Hit\text{-}and\text{-}Run}$.
-In contrast, our analysis not only recovers but in some sense surpasses the known rates for $\mathsf{Ball\ walk},\mathsf{Hit\text{-}and\text{-}Run}$, while harmonizing well with the continuous-time perspective of diffusions.
-Proximal schemes for sampling.
-The Gibbs sampling scheme used in this paper was inspired by the restricted Gaussian oracle introduced in [LST21] (in turn inspired by Gaussian Cooling [CV18]), which alternately iterates between a pure Gaussian step, and a “proximal” step (which we elaborate in our exposition).
-This scheme was given novel interpretations by [CCSW22], which showed that it interpolates the forward and backward heat flows, in the sense defined by [KP21].
-The backward heat flow itself is intimately related to stochastic localization schemes, invented and popularized in [Eld13, Che21].
-This formulation proved surprisingly powerful, allowing many existing rates in unconstrained sampling to be recovered from a relatively simple analysis.
-This was further extended by [FYC23] to achieve the current state-of-the-art rate in unconstrained sampling.
-Finally, [GLL+23] suggests that this could be applied to tackle some constrained problems.
-However, the assumptions in this final mentioned work are not compatible with the uniform sampling problem on general convex bodies.
-Appendix B $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$
-We restate the previously known guarantees for uniform sampling by $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$.
+Appendix A $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$
+We restate the previously known guarantees for uniform sampling by the $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$.
 Below, let $B_{r}(x)$ denote the $d$-dimensional ball of radius $r$ centered at $x$.
@@
 Algorithm 2 $\mathsf{Ball\ walk}$
-$\mathsf{Ball\ walk}$ is particularly simple; draw a uniform random point from $B_{\delta}$ around the current point, and go there if the drawn point is inside of $\mathcal{K}$ and stay at the current point otherwise.
+The $\mathsf{Ball\ walk}$ is particularly simple; draw a uniform random point from $B_{\delta}$ around the current point, and go there if the drawn point is inside of $\mathcal{K}$ and stay at the current point otherwise.
 Its stationary distribution can be easily seen to be $\pi\propto\mathds{1}_{\mathcal{K}}$, the uniform distribution over $\mathcal{K}$.
-In the literature, there are two approaches to analyzing the convergence rate of this sampler: (i) a direct analysis via the $s$-conductance of $\mathsf{Ball\ walk}$ and (ii) an indirect approach which first passes through $\mathsf{Speedy\ walk}$.
+In the literature, there are two approaches to analyzing the convergence rate of this sampler: (i) a direct analysis via the $s$-conductance of the $\mathsf{Ball\ walk}$ and (ii) an indirect approach which first passes through the $\mathsf{Speedy\ walk}$.
 Direct analysis.
-The following $\mathsf{TV}$-guarantee is obtained by lower bounding the $s$-conductance of $\mathsf{Ball\ walk}$, which requires a one-step coupling argument and the Cheeger inequality for $\pi$.
+The following $\mathsf{TV}$-guarantee is obtained by lower bounding the $s$-conductance of the $\mathsf{Ball\ walk}$, which requires a one-step coupling argument and the Cheeger inequality for $\pi$.
 We refer interested readers to [Vem05, §5].
 Theorem 26 (Convergence of $\mathsf{Ball\ walk}$).
-For any $\varepsilon\in(0,1)$ and convex body $\mathcal{K}\subset\mathbb{R}^{d}$ presented by a well-defined membership oracle, let $\pi_{t}$ be the distribution after $t$ steps of $\mathsf{Ball\ walk}$ with an $M$-warm initial distribution $\pi_{0}$.
-Then, $\mathsf{Ball\ walk}$ with step size $\delta=\Theta(\frac{\varepsilon}{M\sqrt{d}})$ achieves $\lVert\pi_{t}-\pi\rVert_{\mathsf{TV}}\leq\varepsilon$ for $t\gtrsim d^{2}D^{2}\frac{M^{2}}{\varepsilon^{2}}\log\frac{M}{\varepsilon}$.
-If $\pi$ is isotropic, then $\mathsf{Ball\ walk}$ needs $\mathcal{O}(d^{2}\frac{M^{2}}{\varepsilon^{2}}\log d\log\frac{M}{\varepsilon})$ iterations.
-The mixing time of $\mathsf{Ball\ walk}$ under this approach has a polynomial dependence on $1/\varepsilon$, rather than a polylogarithmic dependence.
+For any $\varepsilon\in(0,1)$ and convex body $\mathcal{K}\subset\mathbb{R}^{d}$ presented by a well-defined membership oracle, let $\pi_{t}$ be the distribution after $t$ steps of the $\mathsf{Ball\ walk}$ with an $M$-warm initial distribution $\pi_{0}$.
+Then, the $\mathsf{Ball\ walk}$ with step size $\delta=\Theta(\frac{\varepsilon}{M\sqrt{d}})$ achieves $\lVert\pi_{t}-\pi\rVert_{\mathsf{TV}}\leq\varepsilon$ for $t\gtrsim d^{2}D^{2}\frac{M^{2}}{\varepsilon^{2}}\log\frac{M}{\varepsilon}$.
+If $\pi$ is isotropic, then the $\mathsf{Ball\ walk}$ needs $O(d^{2}\frac{M^{2}}{\varepsilon^{2}}\log d\log\frac{M}{\varepsilon})$ iterations.
+The mixing time of the $\mathsf{Ball\ walk}$ under this approach has a polynomial dependence on $1/\varepsilon$, rather than a polylogarithmic dependence.
 Indirect analysis through $\mathsf{Speedy\ walk}$.
-[KLS97] introduced $\mathsf{Speedy\ walk}$, which could be viewed as a version of $\mathsf{Ball\ walk}$ and converges to a speedy distribution (see Proposition 1), which is slightly biased from $\pi$.
-Then, $\mathsf{Speedy\ walk}$ is used together with another algorithmic component (rejection sampling) [KLS97, Algorithm 4.15] that converts the speedy distribution to the uniform distribution.
-In the literature, $\mathsf{Ball\ walk}$ often refers to ‘$\mathsf{Speedy\ walk}$ combined with the conversion step’, rather than a direct implementation of Algorithm 2.
-Strictly speaking, a mixing guarantee of this combined algorithm should not be referred to as a provable guarantee of $\mathsf{Ball\ walk}$.
+[KLS97] introduced the $\mathsf{Speedy\ walk}$, which could be viewed as a version of the $\mathsf{Ball\ walk}$ and converges to a speedy distribution (see Proposition 2), which is slightly biased from $\pi$.
+Then, the $\mathsf{Speedy\ walk}$ is used together with another algorithmic component (rejection sampling) [KLS97, Algorithm 4.15] that converts the speedy distribution to the uniform distribution.
+In the literature, the $\mathsf{Ball\ walk}$ often refers to $\mathsf{Speedy\ walk}$ combined with the conversion step’, rather than a direct implementation of Algorithm 2.
+Strictly speaking, a mixing guarantee of this combined algorithm should not be referred to as a provable guarantee of the $\mathsf{Ball\ walk}$.
 Input: initial distribution $\pi_{0}$, convex body $\mathcal{K}\subset\mathbb{R}^{d}$, iterations $T$, step size $\delta>0$.
@@
 Algorithm 3 $\mathsf{Speedy\ walk}$
-As opposed to $\mathsf{Ball\ walk}$, $\mathsf{Speedy\ walk}$ always takes some step at each iteration.
+As opposed to the $\mathsf{Ball\ walk}$, the $\mathsf{Speedy\ walk}$ always takes some step at each iteration.
 However, the problem of sampling from $x_{i}\sim\textnormal{Unif}\mathopen{}\mathclose{{\left(\mathcal{K}\cap B_{\delta}(x_{i-1})}}\right)$ in Line 3 is not straightforward.
-This step admits the following implementation based on rejection sampling, via a procedure denoted by ()($\ast$)( ∗ ):
+This step admits the following implementation based on rejection sampling, via a procedure denoted by $(\ast)$:
   • Propose $y\sim\textnormal{Unif}(B_{\delta}(x_{i-1}))$.
   • Set $x_{i+1}\leftarrow y$ if $y\in\mathcal{K}$. Otherwise, repeat the proposal.
-Each actual step (indexed by $i$) in $\mathsf{Speedy\ walk}$ is called a proper step, and rejected steps during ()($\ast$)( ∗ ) are called improper steps.
-For example, if $x_{1},x_{1},x_{2},x_{3},x_{3},x_{3},x_{4},\dotsc$ are the positions produced by $\mathsf{Ball\ walk}$, then only proper steps $x_{1},x_{2},x_{3},x_{4},\dotsc$ are recorded by $\mathsf{Speedy\ walk}$.
-To describe the theoretical guarantees of $\mathsf{Speedy\ walk}$, we define the local conductance $\ell(x)$ at $x\in\mathcal{K}$, which measures the success probability of the rejection sampling scheme in ()($\ast$)( ∗ ):
+Each actual step (indexed by $i$) in the $\mathsf{Speedy\ walk}$ is called a proper step, and rejected steps during $(\ast)$ are called improper steps.
+For example, if $x_{1},x_{1},x_{2},x_{3},x_{3},x_{3},x_{4},\dotsc$ are the positions produced by the $\mathsf{Ball\ walk}$, then only proper steps $x_{1},x_{2},x_{3},x_{4},\dotsc$ are recorded by the $\mathsf{Speedy\ walk}$.
+To describe the theoretical guarantees of the $\mathsf{Speedy\ walk}$, we define the local conductance $\ell(x)$ at $x\in\mathcal{K}$, which measures the success probability of the rejection sampling scheme in $(\ast)$:
 $\ell(x):=\frac{\operatorname{vol}(\mathcal{K}\cap B_{\delta}(x))}{\operatorname{vol}(B_{\delta}(x))}\,$,
@@
 Proposition 2 ([KLS97]).
-The stationary distribution $\nu$ of $\mathsf{Speedy\ walk}$ has density
-[…]
+The stationary distribution $\nu$ of the $\mathsf{Speedy\ walk}$ has density
+$\nu(x)=\frac{\ell(x)\,\mathds{1}_{\mathcal{K}}(x)}{\int_{\mathcal{K}}\ell(x)\,\mathrm{d}x}\,.$
 The speedy distribution $\nu$ is indeed different from the uniform distribution $\pi$, and this discrepancy is quantified in terms of the average conductance.
@@
 The best known result for $\mathsf{Speedy\ walk}$’s mixing is due to [KLM06] devising the blocking conductance and using the log-Cheeger inequality.
-When $\nu$ is isotropic (i.e., it has covariance proportional to the identity matrix), [LV17] improves the mixing bound via the log-Cheeger constant.
+When $\nu$ is isotropic (i.e., it has covariance proportional to the identity matrix), [LV24] improves the mixing bound via the log-Cheeger constant.
 Theorem 27 (Mixing of $\mathsf{Speedy\ walk}$).
-For any $\varepsilon\in(0,1)$ and convex body $\mathcal{K}\subset\mathbb{R}^{d}$ presented by a well-defined membership oracle, let $\nu_{t}$ be the distribution after $t$ proper steps of $\mathsf{Speedy\ walk}$ started at any feasible point $x_{0}\in\mathcal{K}$.
-Then, $\mathsf{Speedy\ walk}$ with step size $\delta=\Theta(d^{-1/2})$ achieves $\lVert\nu_{t}-\nu\rVert_{\mathsf{TV}}\leq\varepsilon$ for $t\gtrsim(D^{2}+\log(D\sqrt{d}))\,d^{2}\log\frac{1}{\varepsilon}$.
-From an $M$-warm start, the expected number of improper steps during $t$ iterations is $\widetilde{\mathcal{O}}(tM)$.
-When $\nu$ is isotropic, $\mathsf{Speedy\ walk}$ needs $\mathcal{O}(d^{2}D\log\frac{1}{\varepsilon}\log\log D)$ proper steps to achieve $\varepsilon$-$\mathsf{TV}$ distance to $\nu$.
+For any $\varepsilon\in(0,1)$ and convex body $\mathcal{K}\subset\mathbb{R}^{d}$ presented by a well-defined membership oracle, let $\nu_{t}$ be the distribution after $t$ proper steps of the $\mathsf{Speedy\ walk}$ started at any feasible point $x_{0}\in\mathcal{K}$.
+Then, the $\mathsf{Speedy\ walk}$ with step size $\delta=\Theta(d^{-1/2})$ achieves $\lVert\nu_{t}-\nu\rVert_{\mathsf{TV}}\leq\varepsilon$ for $t\gtrsim(D^{2}+\log(D\sqrt{d}))\,d^{2}\log\frac{1}{\varepsilon}$.
+From an $M$-warm start, the expected number of improper steps during $t$ iterations is $\widetilde{O}(tM)$.
+When $\nu$ is isotropic, the $\mathsf{Speedy\ walk}$ needs $\mathcal{O}(d^{2}D\log\frac{1}{\varepsilon}\log\log D)$ proper steps to achieve $\varepsilon$-$\mathsf{TV}$ distance to $\nu$.
 Then, [KLS97] uses the following post-processing step to obtain an approximately uniform distribution on $\mathcal{K}$, with a provable guarantee.
-$\mathcal{A}$: Call $\mathsf{Speedy\ walk}$ to obtain a sample $X\sim\nu_{t}$ until $\frac{2d}{2d-1}\,X\in\mathcal{K}$. If so, return $\bar{X}=\frac{2d}{2d-1}\,X$.
+$\mathcal{A}$: Call the $\mathsf{Speedy\ walk}$ to obtain a sample $X\sim\nu_{t}$ until $\frac{2d}{2d-1}\,X\in\mathcal{K}$. If so, return $\bar{X}=\frac{2d}{2d-1}\,X$.
 Proposition 5 ([KLS97, Theorem 4.16]).
@@
 For $\bar{\nu}=\mathsf{law}(\bar{X})$ given by $\mathcal{A}$, it holds that $\lVert\bar{\nu}-\pi\rVert_{\mathsf{TV}}\leq\varepsilon$, and the expected number of calls on the conversion algorithm is at most $2$.
-Combining the previous two results, we conclude that the total expected number of membership queries to obtain a sample $\varepsilon$-close to $\pi$ in $\mathsf{TV}$ is $\widetilde{\mathcal{O}}(Md^{2}D^{2}\log\frac{1}{\varepsilon})$, which now has a poly-logarithmic dependence on $1/\varepsilon$.
-Remark 3 (Backward heat flow analysis of $\mathsf{Speedy\ walk}$).
-Consider a Gaussian version of $\mathsf{Speedy\ walk}$, whose one-step corresponds to $x_{i+1}\sim\mathcal{N}(x_{i},hI_{d})|_{\mathcal{K}}$, and this transition kernel exactly matches integrating ($\mathsf{BH}$) for time $h$.
-Thus, $\nu Q^{\pi^{X},h}_{h}=\nu$ due to the stationarity of $\nu$ under $\mathsf{Speedy\ walk}$, where $Q^{\pi^{X},h}_{h}$ is the transition kernel defined by the backward heat flow for time $h$ that reverses $\pi^{X}*\mathcal{N}(0,hI_{d})$ to $\pi^{X}$.
-Hence, if we can control the LSI/PI constants of $\nu$ along the backward heat-flow’s trajectory, then we could directly analyze $\mathsf{Speedy\ walk}$ by emulating computations in Lemma 21.
-Appendix C Functional inequalities
+Combining the previous two results, we conclude that the total expected number of membership queries to obtain a sample $\varepsilon$-close to $\pi$ in $\mathsf{TV}$ is $\widetilde{O}(Md^{2}D^{2}\log\frac{1}{\varepsilon})$, which now has a poly-logarithmic dependence on $1/\varepsilon$.
+Appendix B Functional inequalities
 We provide full details on functional inequalities omitted in §3.1.
-We use $\mu$ and $\mu_{\mathsf{LC}}$ to denote a probability measure and log-concave probability measure over $\mathbb{R}^{d}$, respectively.
+We use $\mu$ and $\mu_{\mathsf{LC}}$ to denote a probability measure and logconcave probability measure over $\mathbb{R}^{d}$, respectively.
 Cheeger and PI constants.
@@
 $\mu^{+}(S):=\liminf_{\varepsilon\to 0}\frac{\mu(S^{\varepsilon})-\mu(S)}{\varepsilon}\,.$
-[Che70] established $C_{\mathsf{PI}}(\mu)\lesssim C_{\mathsf{Ch}}^{-2}(\mu)$ [^5] The opposite direction $C_{\mathsf{PI}}(\mu_{\mathsf{LC}})\gtrsim C_{\mathsf{Ch}}^{-2}(\mu_{\mathsf{LC}})$ also holds for log-concave distributions due to [Mil09], while $C_{\mathsf{PI}}(\mu)\gtrsim C_{\mathsf{Ch}}^{-2}(\mu)/d$ for general distributions due to [Led94]., and then [KLS95] showed that for covariance matrix $\Sigma_{\mu}:=\mathbb{E}_{\mu}[(\cdot-\mathbb{E}_{\mu}X)(\cdot-\mathbb{E}_{\mu}X)^{\mathsf{T}}]$,
-$C_{\mathsf{Ch}}(\mu_{\mathsf{LC}})\gtrsim\frac{1}{(\mathbb{E}_{\mu_{\mathsf{LC}}}[\lVert X-\mathbb{E}_{\mu_{\mathsf{LC}}}X\rVert^{2}])^{1/2}}=\frac{1}{(\operatorname{tr}\Sigma_{\mu_{\mathsf{LC}}})^{1/2}}\,.$ (C.1)
+[Che70] established $C_{\mathsf{PI}}(\mu)\lesssim C_{\mathsf{Ch}}^{-2}(\mu)$ [^4] The opposite direction $C_{\mathsf{PI}}(\mu_{\mathsf{LC}})\gtrsim C_{\mathsf{Ch}}^{-2}(\mu_{\mathsf{LC}})$ also holds for logconcave distributions due to [Led94]., and then [KLS95] showed that for covariance matrix $\Sigma_{\mu}:=\mathbb{E}_{\mu}[(\cdot-\mathbb{E}_{\mu}X)(\cdot-\mathbb{E}_{\mu}X)^{\mathsf{T}}]$,
+$C_{\mathsf{Ch}}(\mu_{\mathsf{LC}})\gtrsim\frac{1}{(\mathbb{E}_{\mu_{\mathsf{LC}}}[\lvert X-\mathbb{E}_{\mu_{\mathsf{LC}}}X\rvert^{2}])^{1/2}}=\frac{1}{(\operatorname{tr}\Sigma_{\mu_{\mathsf{LC}}})^{1/2}}\,.$ (B.1)
 This immediately leads to $C_{\mathsf{PI}}(\pi)\lesssim(\mathbb{E}_{\pi}[\lvert X-\mathbb{E}_{\pi}X\rvert^{2}])^{1/2}\leq D^{2}$ for the uniform distribution $\pi$ over a convex body $\mathcal{K}$ with diameter $D>0$.
-Kannan et al. proposed the KLS conjecture in the same paper, which says that for the spectral norm $\lVert\cdot\rVert_{2}$,
-$C_{\mathsf{Ch}}(\mu_{\mathsf{LS}})\gtrsim\frac{1}{\lVert\Sigma_{\mu_{\mathsf{LS}}}\rVert_{2}^{1/2}}\,.$
-While the original result in [KLS95] ensures $C_{\mathsf{Ch}}\gtrsim d^{-1/2}$ for an isotropic log-concave distribution (due to $\Sigma=I_{d}$), this conjecture indeed claims $C_{\mathsf{Ch}}\gtrsim 1$ for such case.
-Following a line of work [LV17, Che21, KL22, Kla23], the current bound is
-$C_{\mathsf{Ch}}(\mu_{\mathsf{LS}})\gtrsim\frac{(\log d)^{-1/2}}{\lVert\Sigma_{\mu_{\mathsf{LS}}}\rVert_{2}^{1/2}}\,,$
+Kannan et al. proposed the KLS conjecture in the same paper, which says that
+$C_{\mathsf{Ch}}(\mu_{\mathsf{LC}})\gtrsim\frac{1}{\lVert\Sigma_{\mu_{\mathsf{LC}}}\rVert_{\mathsf{op}}^{1/2}}\,.$
+While the original result in [KLS95] ensures $C_{\mathsf{Ch}}\gtrsim d^{-1/2}$ for an isotropic logconcave distribution (due to $\Sigma=I_{d}$), this conjecture indeed claims $C_{\mathsf{Ch}}\gtrsim 1$ for such case.
+Following a line of work [LV24, Che21, KL22, Kla23], the current bound is
+$C_{\mathsf{Ch}}(\mu_{\mathsf{LC}})\gtrsim\frac{(\log d)^{-1/2}}{\lVert\Sigma_{\mu_{\mathsf{LC}}}\rVert_{\mathsf{op}}^{1/2}}\,$,
 which implies that $C_{\mathsf{PI}}(\pi)\lesssim\log d$ when $\pi$ is isotropic for convex $\mathcal{K}$.
@@
 $C_{\mathsf{logCh}}(\mu):=\inf_{S\subset\mathbb{R}^{d}:\mu(S)\leq\frac{1}{2}}\frac{\mu^{+}(S)}{\mu(S)\sqrt{\log\frac{1}{\mu(S)}}}\,.$
-[Led94] established that $C_{\mathsf{LSI}}(\mu)\lesssim C_{\mathsf{logCh}}^{-2}(\mu)$ and $C_{\mathsf{LSI}}(\mu_{\mathsf{LS}})\gtrsim C_{\mathsf{logCh}}^{-2}(\mu_{\mathsf{LS}})$.
-[KLM06] showed that any log-concave distributions with support of diameter $D>0$ satisfy $C_{\mathsf{logCh}}(\mu_{\mathsf{LS}})\gtrsim D^{-1}$.
-Later in 2016, [LV17] improved this to $C_{\mathsf{logCh}}(\mu_{\mathsf{LS}})\gtrsim D^{-1/2}$ under isotropy.
+[Led94] established that $C_{\mathsf{LSI}}(\mu)\lesssim C_{\mathsf{logCh}}^{-2}(\mu)$ and $C_{\mathsf{LSI}}(\mu_{\mathsf{LC}})\gtrsim C_{\mathsf{logCh}}^{-2}(\mu_{\mathsf{LC}})$.
+[FK99] showed that any logconcave distributions with support of diameter $D>0$ satisfy $C_{\mathsf{logCh}}(\mu_{\mathsf{LC}})\gtrsim D^{-1}$.
+Later in 2016, [LV24] improved this to $C_{\mathsf{logCh}}(\mu_{\mathsf{LC}})\gtrsim D^{-1/2}$ under isotropy.
 Therefore, for convex $\mathcal{K}$, it follows that $C_{\mathsf{LSI}}(\pi)\lesssim D^{2}$ and that $C_{\mathsf{LSI}}(\pi)\lesssim D$ if $\pi$ is isotropic.
-Appendix D The Wasserstein geometry
-We present additional technical background on the Wasserstein geometry and Markov semigroup theory.
-Interested readers can refer to [Vil09, AGS05, Che24] for standard references on Wasserstein spaces and applications to sampling.
-Wasserstein gradient.
-Let $\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$ be the space of probability measures admitting densities on $\mathbb{R}^{d}$ with finite second moment.
-Although there are many ways to metrize $\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$, the geometry induced by the Wasserstein-$2$ distance $\mathcal{W}_{2}$ is a particularly useful structure for analysis.
-Under the $\mathcal{W}_{2}$-geometry, one can define a “gradient” of a functional defined over $\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$.
-Specifically, for a functional $\mathcal{F}:\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})\to\mathbb{R}\cup\{\infty\}$, the Wasserstein gradient of $\mathcal{F}$ at $\mu\in\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$ is defined as $\nabla_{\mathcal{W}_{2}}\mathcal{F}(\mu)=\nabla(\delta\mathcal{F})(\mu)\in L^{2}(\mu)$, where $\nabla$ is the standard gradient and $\delta\mathcal{F}$ is the first variation of $\mathcal{F}$ [^6] The first variation can be defined, for any measures $\nu_{0},\nu_{1}\in\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$, as $\lim_{t\downarrow 0}\frac{\mathcal{F}((1-t)\nu_{0}+t\nu_{1})-\mathcal{F}(\nu_{0})}{t}=\langle\delta\mathcal{F}(\nu_{0}),\nu_{1}-\nu_{0}\rangle$.
-This definition is unique up to an additive constant, which is irrelevant as we are only concerned with its gradient..
-Equipped with this $\mathcal{W}_{2}$-gradient, one can define the Wasserstein gradient flow of $\mathcal{F}$ that describes the evolution of a measure $\{\mu_{t}\}_{t\geq 0}$, from some initial measure $\mu_{0}$, as follows:
-$\partial_{t}\mu_{t}=\operatorname{div}\bigl{(}\mu_{t}\nabla_{\mathcal{W}_{2}}\mathcal{F}(\mu_{t})\bigr{)}\,.$
-More generally, we can identify the Wasserstein “velocity” for some measure $\mu_{t}$ as $v_{t}$ if the time derivative of $\mu_{t}$ can be written in the form
-$\partial_{t}\mu_{t}=-\operatorname{div}(\mu_{t}v_{t})\,.$
-Under this identification, the time derivative of a functional $\mathcal{F}$ on $\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$ with smooth Wasserstein gradient under these dynamics can be written as
-$\partial_{t}\mathcal{F}(\mu_{t})=\mathbb{E}_{\mu_{t}}\langle\nabla_{\mathcal{W}_{2}}\mathcal{F}(\mu_{t}),v_{t}\rangle\,,$
-when $v_{t}\in\overline{\{\nabla\psi:\psi\in C_{c}^{\infty}(\mathbb{R}^{d})\}}^{L^{2}(\mu_{t})}$, where {}¯$\overline{\{\cdot\}}^{L^{2}(\mu_{t})}$ denotes the closure of a set with respect to $L^{2}(\mu_{t})$.
-This is the appropriate notion of tangent space in this geometry.
-For instance, when we take the functional to be the entropy of the measure, $\mathcal{H}(\mu)\coloneqq\frac{1}{2}\int\mu\log\mu$, one can verify $\nabla_{\mathcal{W}_{2}}\mathcal{H}(\mu)=\frac{1}{2}\nabla\log\mu$.
-The heat flow equation can be written as $\partial_{t}\mu_{t}=\frac{1}{2}\Delta\mu_{t}=\frac{1}{2}\operatorname{div}(\nabla\mu_{t})=\frac{1}{2}\operatorname{div}(\mu_{t}\nabla\log\mu_{t})$, which indicates that the velocity of measures $\mu_{t}$ under the heat flow is $v_{t}=-\frac{1}{2}\nabla\log\mu_{t}$.
-Hence, we can notice that $\nabla_{\mathcal{W}_{2}}\mathcal{H}(\mu_{t})=-v_{t}$, and thus recover the heat flow as the Wasserstein gradient flow of the entropy of the measure.
-Fokker-Planck equation and time-reversal of SDE.
-Consider a stochastic differential equation $(X_{t})$ given by
-$\mathrm{d}X_{t}=-a_{t}(X_{t})\,\mathrm{d}t+\,\mathrm{d}B_{t}\qquad\text{with }X_{0}\sim\mu_{0}\,.$ (D.1)
+Appendix C Semigroup calculus for contraction under the heat flow
+In this part, we first recall Markov semigroups, which are well-established mathematical tools that abstract an underlying Markov process.
+Interested readers can refer to [BGL14, Che25] for references.
+We then review the contraction results for heat flow and its time-reversal [CCSW22], which are intimately connected with our algorithm.
+We also provide key technical ingredients needed for its proof, such as the computations for measures evolving under simultaneous forward/backward heat flows.
+Forward heat flow.
+We begin by introducing the “heat flow” equation (or also known as the Fokker–Planck equation), which describes the evolution of the law of $Z_{t}$ under ($\mathsf{FH}$),
+$\partial_{t}\mu_{t}=\frac{1}{2}\,\Delta\mu_{t}=\frac{1}{2}\operatorname{div}(\mu_{t}\nabla\log\mu_{t})\,.$ ($\mathsf{FP\text{-}FH}$)
+It is well known that one can realize this equation in discrete time through a Gaussian transition density, in the sense that, for $\mu_{h}$ (the solution at time $h>0$ to ($\mathsf{FP\text{-}FH}$) with initial condition $\mu_{0}$), and for any smooth function $f:\mathbb{R}^{d}\to\mathbb{R}$,
+$\mathbb{E}_{\mu_{h}}[f(x)]=\mathbb{E}_{\mu_{0}}[P_{h}f(x)]\,$,
+[…]
+Backward heat flow.
+Although there are many ways to define a “reversal” of $P_{h}$, we will use the notion of adjoint introduced by [KP23], which is the most immediately useful.
+Given some initial measure $\nu$ and some time horizon $h$, the adjoint corresponds to reversing ($\mathsf{FH}$) for times in $[0,h]$ when the initial distribution under consideration is $Z_{0}\sim\nu$.
+For other measures, it must be interpreted more carefully, and is given by the following partial differential equation starting from some measure $\mu_{0}^{\leftarrow}$ (see (C.1) and its derivation):
+$\partial_{t}\mu_{t}^{\leftarrow}=-\operatorname{div}\bigl(\mu_{t}^{\leftarrow}\nabla\log(\nu P_{h-t})\bigr)+\frac{1}{2}\,\Delta\mu_{t}^{\leftarrow}\quad\text{for }t\in[0,h]\,.$ ($\mathsf{FP\text{-}BH}$)
+Write $\mu_{t}^{\leftarrow}=\mu_{0}^{\leftarrow}Q_{t}^{\nu,h}$, where $\{Q_{t}^{\nu,h}\}_{t\in[0,h]}$ is a family of transition densities.
+Write $\mathbf{P}_{0,h}$ for the joint distribution of the $(Z_{0},Z_{h})$-marginals of ($\mathsf{FH}$), when $Z_{0}\sim\nu$, and $\mathbf{P}_{0|h}$ for the conditional.
+Note that $\mathbf{P}_{h|0}(\cdot|x)=\mathcal{N}(x,hI_{d})$.
+It is also known that ($\mathsf{FP\text{-}BH}$) gives a time-reversal of the heat equation at the SDE level, in the sense that we can interpret $\delta_{x}Q_{h}^{\nu,h}=\mathbf{P}_{0|h}(\cdot|Z_{h}=x)$.
+Thus $\mu_{0}^{\leftarrow}Q_{h}^{\nu,h}=\int\mathbf{P}_{0|h}(\cdot|Z_{h}=x)\,\mu_{0}^{\leftarrow}(\mathrm{d}x)$, and $\nu P_{h}Q_{t}^{\nu,h}=\nu P_{h-t}$ for all $t\in[0,h]$.
+The ultimate purpose of this machinery is to affirm our earlier description of the Gibbs sampling procedure as alternating forward and backward heat flows.
+Indeed, notice that, if $\mu_{i}^{X}$ is the law of the iterate at some iteration $i$, then $\mu_{i}^{X}P_{h}$ is precisely $\mu_{i+1}^{Y}$ under our scheme, while $(\mu_{i}^{X}P_{h})Q_{h}^{\pi^{X},h}$ is $\mu_{i+1}^{X}$, assuming $Q_{h}^{\pi^{X},h}$ is well defined for non-smooth measures $\pi^{X}$.
+Thus, while Algorithm 1 is implemented via discrete steps, it can be exactly analyzed through arguments in continuous time.
+Fokker–Planck equation and time-reversal of SDE.  The description above can be further generalized further as follows. Consider a stochastic differential equation $(X_{t})$ given by
+$\mathrm{d}X_{t}=-a_{t}(X_{t})\,\mathrm{d}t+\,\mathrm{d}B_{t}\qquad\text{with }X_{0}\sim\mu_{0}\,.$ (C.1)
 It is well known that measures $\mu_{t}$ described by
-$\partial_{t}\mu_{t}=\mathsf{div}(\mu_{t}a_{t})+\frac{1}{2}\Delta\mu_{t}\,,$ (D.2)
-correspond to $\mathsf{law}(X_{t})$. In this context, (D.2) is referred to as the Fokker-Planck equation corresponding to (D.1).
-From this equation, one can deduce the Fokker-Planck equation of the time reversal $\mu_{t}^{\leftarrow}:=\mu_{T-t}$:
+$\partial_{t}\mu_{t}=\mathsf{div}(\mu_{t}a_{t})+\frac{1}{2}\Delta\mu_{t}\,$, (C.2)
+correspond to $\mathsf{law}(X_{t})$. In this context, (C.2) is referred to as the Fokker–Planck equation corresponding to (C.1).
+From this equation, one can deduce the Fokker–Planck equation of the time reversal $\mu_{t}^{\leftarrow}:=\mu_{T-t}$:
 $\partial_{t}\mu_{t}^{\leftarrow}=-\operatorname{div}(\mu_{t}^{\leftarrow}a_{T-t})-\frac{1}{2}\Delta\mu_{t}^{\leftarrow}=-\operatorname{div}\bigl(\mu_{t}^{\leftarrow}(a_{T-t}+\nabla\log\mu_{T-t})\bigr)+\frac{1}{2}\Delta\mu_{t}^{\leftarrow}$
 In particular, this describes the evolution of $\mathsf{law}(X_{t})$ of the stochastic differential equation:
-$\mathrm{d}X_{t}=\bigl{(}a_{T-t}(X_{t})+\nabla\log\mu_{T-t}(X_{t})\bigr{)}\,\mathrm{d}t+\,\mathrm{d}B_{t}\qquad\text{with }X_{0}\sim\mu_{0}^{\leftarrow}=\mu_{T}\,.$ (D.3)
-While the law of this process will give $\mu_{T}^{\leftarrow}=\mu_{0}$ at time $T$, it is also true that it will give $\mu_{0|T}(\cdot|z)$ if one starts (D.3) at $X_{0}=z$.
+$\mathrm{d}X_{t}=\bigl(a_{T-t}(X_{t})+\nabla\log\mu_{T-t}(X_{t})\bigr)\,\mathrm{d}t+\,\mathrm{d}B_{t}\qquad\text{with }X_{0}\sim\mu_{0}^{\leftarrow}=\mu_{T}\,.$ (C.3)
+While the law of this process will give $\mu_{T}^{\leftarrow}=\mu_{0}$ at time $T$, it is also true that it will give $\mu_{0|T}(\cdot\,|\,z)$ if one starts (C.3) at $X_{0}=z$.
 This is a subtle fact, whose justification requires the introduction of a tool called Doob’s $h$-transform.
-The presentation of this subject is beyond the scope of this paper, and we refer interested readers to [KP21] as a reference to its application in this context.
+The presentation of this subject is beyond the scope of this paper, and we refer interested readers to [KP23] as a reference to its application in this context.
+Contraction under simultaneous evolution.
+Instead of considering the change in metrics along the evolution of $\mu P_{t}$ with respect to “fixed” $\nu$, it will be useful to consider the simultaneous evolution of $\mu P_{t},\nu P_{t}$ (and similarly $\mu Q_{t}^{\pi^{X},h},(\nu P_{h})Q_{t}^{\pi^{X},h}$).
+This type of computation was carried out for specific metrics in earlier work [VW23, CCSW22].
+The following is a more generalized form of one appearing in [YFL+23, Lemma 2].
+In the lemma below, we consider an arbitrary diffusion equation with corresponding Fokker–Planck equation:
+$\mathrm{d}X_{t}=b_{t}(X_{t})\,\mathrm{d}t+\,\mathrm{d}B_{t}\,\quad\text{and}\quad\partial_{t}\mu_{t}=-\nabla\cdot(b_{t}\mu_{t})+\frac{1}{2}\,\Delta\mu_{t}\,$ (C.4)
+where $b_{t}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is smooth, $X_{t}\in\mathbb{R}^{d}$, and $\mu_{t}=\mathsf{law}(X_{t})$ if $X_{0}\sim\mu_{0}$.
+Below, we derive contractions of two measures along the same stochastic process, which proves Proposition 1 under suitable regularity assumptions.
+Lemma 28 (Decay along forward/backward heat flows).
+Let $(\mu_{t})_{t\geq 0},(\nu_{t})_{t\geq 0}$ denote the laws of the solutions to (C.4) starting at $\mu_{0},\nu_{0}$ respectively.
+Assume that $\mu_{t}$ and $\nu_{t}$ has sufficient regularity and fast tail decay so that we can discard boundary terms in integration by parts and differentiate under the integral sign.
+Then, for any differentiable function $g$,
+$\partial_{t}g\bigl(D_{f}(\mu_{t}\mathbin{\|}\nu_{t})\bigr)=-\frac{1}{2}\,g^{\prime}\bigl(D_{f}(\mu_{t}\mathbin{\|}\nu_{t})\bigr)\times\mathbb{E}_{\mu_{t}}\bigl\langle\nabla\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr),\nabla\log\frac{\mu_{t}}{\nu_{t}}\bigr\rangle\,.$
+Proof.
+The case where $g\neq\mathsf{id}$ is an application of the chain rule, so it suffices to take $g=\mathsf{id}$ and simply differentiate an $f$-divergence.
+For brevity, we drop the variable $x$ from functions involved and proceed by differentiating under the integral sign and discarding boundary terms in integration by parts:
+$\partial_{t}D_{f}(\mu_{t}\mathbin{\|}\nu_{t})$ $=\int\Bigl\{\bigl(f\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\,\partial_{t}\nu_{t}+\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\bigl(\frac{\mu_{t}}{\nu_{t}}\bigr)^{\prime}\nu_{t}\Bigr\}\,\mathrm{d}x$
+$=\int\Bigl\{\partial_{t}\nu_{t}\Bigl(\bigl(f\circ\frac{\mu_{t}}{\nu_{t}}\bigr)-\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\frac{\mu_{t}}{\nu_{t}}\Bigr)+\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\,\partial_{t}\mu_{t}\Bigr\}\,\mathrm{d}x$
+$\underset{(i)}{=}\int\bigl[-\nabla\cdot(b_{t}\nu_{t})+\frac{1}{2}\,\Delta\nu_{t}\bigr]\Bigl(\bigl(f\circ\frac{\mu_{t}}{\nu_{t}}\bigr)-\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\frac{\mu_{t}}{\nu_{t}}\Bigr)\,\mathrm{d}x$
+$\qquad+\int\bigl[-\nabla\cdot(b_{t}\mu_{t})+\frac{1}{2}\,\Delta\mu_{t}\bigr]\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\,\mathrm{d}x\,$,
+where in $(i)$ we substitute the F-P equation from (C.4).
+Integrating by parts (i.e., $\int f\operatorname{div}(\mathbf{G})=-\int\langle\nabla f,\mathbf{G}\rangle$ for a real-valued function $f$ and vector-valued function $\mathbf{G}$), we have that
+$\int\bigl[-\nabla\cdot(b_{t}\nu_{t})\bigr]\bigl(f\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\,\mathrm{d}x=\int\bigl\langle b_{t}\nu_{t},\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\nabla\frac{\mu_{t}}{\nu_{t}}\bigr\rangle\,\mathrm{d}x\,.$ (C.5)
+On the other hand, we have that
+$\int\bigl[-\nabla\cdot(b_{t}\nu_{t})\bigr]\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\frac{\mu_{t}}{\nu_{t}}\,\mathrm{d}x=-\int\bigl\langle b_{t}\nu_{t},\frac{\mu_{t}}{\nu_{t}}\,\nabla\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)+\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\nabla\frac{\mu_{t}}{\nu_{t}}\bigr\rangle\,\mathrm{d}x\,.$
+The second term cancels with the RHS of (C.5). We have a similar cancellation for the $\frac{1}{2}\,\Delta\nu_{t}$ term:
+$\int\frac{1}{2}\,\Delta\nu_{t}\,\bigl(f\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\,\mathrm{d}x=-\int\frac{1}{2}\,\bigl\langle\nabla\nu_{t},\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\nabla\frac{\mu_{t}}{\nu_{t}}\bigr\rangle\,\mathrm{d}x\,$,
+and
+$\int\frac{1}{2}\,\Delta\nu_{t}\,\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\frac{\mu_{t}}{\nu_{t}}\,\mathrm{d}x=\int\frac{1}{2}\,\bigl\langle\nabla\nu_{t},\frac{\mu_{t}}{\nu_{t}}\,\nabla\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)+\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\nabla\frac{\mu_{t}}{\nu_{t}}\bigr\rangle\,\mathrm{d}x\,.$
+Combining these, we are left with
+$\int\bigl[-\nabla\cdot(b_{t}\nu_{t})+\frac{1}{2}\,\Delta\nu_{t}\bigr]\Bigl(\bigl(f\circ\frac{\mu_{t}}{\nu_{t}}\bigr)-\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\frac{\mu_{t}}{\nu_{t}}\Bigr)\,\mathrm{d}x$ $=-\int\bigl\langle b_{t}\nu_{t}-\frac{1}{2}\,\nabla\nu_{t},\nabla\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\frac{\mu_{t}}{\nu_{t}}\bigr\rangle\,\mathrm{d}x$
+$=-\int\bigl\langle b_{t}\mu_{t}-\frac{1}{2}\,\mu_{t}\nabla\log\nu_{t},\nabla\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\bigr\rangle\,\mathrm{d}x\,.$
+Finally, we note that
+$\int\bigl[-\nabla\cdot(b_{t}\mu_{t})+\frac{1}{2}\Delta\mu_{t}\bigr]\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\,\mathrm{d}x$ $=\int\bigl\langle b_{t}\mu_{t}-\frac{1}{2}\,\nabla\mu_{t},\nabla\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\bigr\rangle\,\mathrm{d}x$
+$=\int\bigl\langle b_{t}\mu_{t}-\frac{1}{2}\,\mu_{t}\nabla\log\mu_{t},\nabla\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\bigr\rangle\,\mathrm{d}x\,.$
+Putting it all together, noticing that the drift terms cancel, we are left with
+$\partial_{t}D_{f}(\mu_{t}\mathbin{\|}\nu_{t})=-\int\frac{1}{2}\,\bigl\langle\mu_{t}\nabla\log\frac{\mu_{t}}{\nu_{t}},\nabla\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\bigr\rangle\,\mathrm{d}x=-\frac{1}{2}\,\mathbb{E}_{\mu_{t}}\bigl\langle\nabla\log\frac{\mu_{t}}{\nu_{t}},\nabla\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\bigr\rangle\,$,
+which completes the proof. ∎
+To recover the decay result for the $q$-Rényi divergence, one can substitute $g(x)=\frac{1}{q-1}\log x$ and $f(x)=x^{q}-1$.
+For the $\chi^{2}$-divergence, instead substitute $g(x)=x$ and $f(x)=x^{2}-1$.
+From this, we can obtain a single step of decay for the Rényi and $\chi^{2}$-divergences under different functional inequalities.

### v4

@@
 With the proliferation of data and the increasing importance of machine learning, sampling has also become an essential algorithmic tool, with applications needing samplers in very high dimension, e.g., scientific computing [CV16, HCT+17, KLSV22], systems biology [LNP12, TSF+13], differential privacy [MT07, Mir17] and machine learning [BCJ+19, Sta20].
-Samplers for convex bodies are based on Markov chains (see §1.4 for a summary).
+Samplers for convex bodies are based on Markov chains (see Sec.
+§1.4 for a summary).
 Their analysis is based on bounding the conductance of the associated Markov chain, which in turn bounds the mixing rate.
@@
 This idea was utilized in Gaussian cooling [CV18] and later as the restricted Gaussian Oracle (RGO) [LST21, CCSW22].
-Under this notation, Algorithm 1 corresponds to a Gibbs sampling scheme from the two marginals of $\pi(x,y)$. To be precise, Line 2 and Line 3 correspond to sampling
+Under this notation, Algorithm 1 without the failure condition corresponds to a Gibbs sampling scheme from the two marginals of $\pi(x,y)$. To be precise, Line 2 and Line 3 correspond to sampling
 $y_{i+1}\sim\pi^{Y|X}(\cdot\,|\,x_{i})=\mathcal{N}(x_{i},hI_{d})\qquad\mbox{ and }\qquad x_{i+1}\sim\pi^{X|Y}(\cdot\,|\,y_{i+1})=\mathcal{N}(y_{i+1},hI_{d})|_{\mathcal{K}}\,.$
+It is well known that such a Gibbs sampling procedure will ensure the desired stationarity of $\pi(x,y)$.
 We implement the latter step through rejection sampling; if the number of trials in Line 3 hits the threshold $N$, then we halt and declare failure of the algorithm.
-It is well known that such a Gibbs sampling procedure will ensure the desired stationarity of $\pi(x,y)$.
-Note that, conditioned on the event that the algorithm does not fail, the resulting iterate will be an unbiased sample from the correct distribution.
+Note that the accepted point in Line 3, conditional on success (i.e., number of attempts is at most the cap), is still exactly $\pi^{X|Y=y_{i+1}}$.
+However, conditioning on “no failure over all $T$ iterations” (or equivalently “re-run-until-success”) re-weights trajectories, so the final output law of the capped INO is not exactly the same as the INO without the cap; see §D for details.
+Nonetheless, it is $\eta$-close to the intended target $\pi^{X}$ pointwise when the overall failure probability is $\eta$.
+Practical implementation via restart.
+In practice, one would restart the algorithm if it declares failure before completing the target number of iterations.
+This yields an always-terminating implementation whose output law equals that of Algorithm 1 conditioned on success.
+The expected number of restarts is at most $(1-\eta)^{-1}$, so the expected total query complexity increases by a multiplicative factor of only $(1-\eta)^{-1}\leq 2$ for $\eta\leq 1/2$.
 Stochastic perspective: heat flows.
@@
 Theorem 3 (Succinct version of Theorem 26).
-For any given $\eta,\varepsilon\in(0,1)$, $q\geq 1$, and any convex body $\mathcal{K}$ given by a well-defined membership oracle, there exist choices of parameters $h,N$ such that $\mathsf{In\text{-}and\text{-}Out}$, starting from an $M$-warm distribution, with probability at least $1-\eta$, returns $X\sim\mu$ such that $\mathcal{R}_{q}(\mu\mathbin{\|}\pi^{\mathcal{K}})\leq\varepsilon$.
+For any given $\eta,\varepsilon\in(0,1/2)$, $q\geq 2$, and any convex body $\mathcal{K}$ given by a well-defined membership oracle, there exist choices of parameters $h,N$ such that $\mathsf{In\text{-}and\text{-}Out}$, starting from an $M$-warm distribution, with probability at least $1-\eta$, returns $X\sim\mu$ such that $\mathcal{R}_{q}(\mu\mathbin{\|}\pi^{\mathcal{K}})\leq\varepsilon+4\eta$.
 The number of proper steps is $\widetilde{O}(qd^{2}\Lambda\log^{2}\nicefrac{{M}}{{\eta\varepsilon}})$, and the expected total number of membership queries is $\widetilde{O}(qMd^{2}\Lambda\log^{6}\nicefrac{{1}}{{\eta\varepsilon}})$, where $\Lambda$ is the largest eigenvalue of the covariance of $\pi^{\mathcal{K}}$.
@@
 See Lemma 15 for more details.
-Remark 2 (Warm-start generation).
+The complexity of the restart-on-failure algorithm immediately follows from this result as a simple corollary.
+Corollary 4 (Restart-on-failure).
+For $\varepsilon\in(0,\frac{1}{10})$ and $q\geq 2$, consider the restart variant that repeats Algorithm 1 from scratch until success with the same parameter choices.
+Then, the output sample $X\sim\mu^{\prime}$ satisfies $\mathcal{R}_{q}(\mu^{\prime}\,\|\,\pi^{\mathcal{K}})\leq\varepsilon$, and the expected total number of membership queries is $\widetilde{O}(qMd^{2}\Lambda\log^{6}\nicefrac{{1}}{{\varepsilon}})$.
+Remark 2 ((Warm-start generation)).
 While the assumption of warmness for the initialization may seem strong at the outset, for well-rounded convex bodies ($\mathbb{E}_{\pi}[\lvert\cdot\rvert^{2}]\leq C^{2}d$ for some constant $C$), it is possible to generate an $O(1)$ warm-start with complexity $\widetilde{O}(d^{3})$.
@@
 Assume that $\pi^{\mathcal{K}}$ is near-isotropic, i.e., the operator norm of its covariance is $O(1)$.
-Under the same setting as above, $\mathsf{In\text{-}and\text{-}Out}$ succeeds with probability $1-\eta$, returning $X\sim\mu$ such that $\mathcal{R}_{q}(\mu\mathbin{\|}\pi^{\mathcal{K}})\leq\varepsilon$.
+Under the same setting as above, $\mathsf{In\text{-}and\text{-}Out}$ succeeds with probability $1-\eta$, returning $X\sim\mu$ such that $\mathcal{R}_{q}(\mu\mathbin{\|}\pi^{\mathcal{K}})\leq\varepsilon+4\eta$.
 The number of proper steps is $\widetilde{O}(qd^{2}\log^{2}\frac{M}{\eta\varepsilon})$, and the expected total number of membership queries is $\widetilde{O}(qMd^{2}\log^{6}\tfrac{1}{\eta\varepsilon})$.
-Our analysis will in fact show that the bound on the number of proper steps holds for general non-convex bodies and any feasible start in $\mathcal{K}$.
-This is deduced under an $M$-warm start in Corollaries 24 and 25.
-We remark that such a bound for non-convex uniform sampling is not known for the $\mathsf{Ball\ walk}$ or the $\mathsf{Speedy\ walk}$.
-Theorem 5.
-For any given $\varepsilon\in(0,1)$ and set $\mathcal{K}\subset B_{D}(0)$ with $\operatorname{vol}(\mathcal{K})>0$, $\mathsf{In\text{-}and\text{-}Out}$ with variance $h$ and $M$-warm initial distribution achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after the following number of iterations:
-$m=\min\begin{cases}O\bigl(qh^{-1}C_{\mathsf{PI}}(\pi^{X})\log\frac{M}{\varepsilon}\bigr)&\text{for }q\geq 2\,,\\ O\bigl(qh^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{\log M}{\varepsilon}\bigr)&\text{for }q\geq 1\,.\end{cases}$
-We have two different convergence results above under (8) and (7).
-Under (8) we have a doubly-logarithmic dependence on the warmness parameter $M$.
-On the other hand, using (7), which is weaker than (8) (in general, $C_{\mathsf{PI}}\leq C_{\mathsf{LSI}}$), the dependence on $M$ is logarithmic.
-We discuss implications of our results further in §1.3.
+Building on the contraction results in [CCSW22, KO25], we show that the bound on the number of proper steps holds for general non-convex bodies and from any feasible start in $\mathcal{K}$.
+We remark that this bound for non-convex uniform sampling is not known for the $\mathsf{Ball\ walk}$ or the $\mathsf{Speedy\ walk}$.
+Corollary 6.
+For any given $\varepsilon\in(0,1)$ and set $\mathcal{K}\subset B_{D}(0)$, $\mathsf{In\text{-}and\text{-}Out}$ with variance $h$ and any feasible start $x_{0}\in\mathcal{K}$ achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{O}(qh^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{d+D^{2}/h}{\varepsilon})$ iterations.
+Corollary 7.
+For any given $\varepsilon\in(0,1)$ and convex body $\mathcal{K}\subset B_{D}(0)$, $\mathsf{In\text{-}and\text{-}Out}$ with variance $h$ and a feasible start $x_{0}\in\mathcal{K}$ achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{O}(qh^{-1}D^{2}\log\frac{1}{\varepsilon})$ iterations.
+If $\pi^{X}$ is isotropic, then $\mathsf{In\text{-}and\text{-}Out}$ only needs $\widetilde{O}(qh^{-1}D\log\frac{d+d^{2}/h}{\varepsilon})$ iterations.
+We note that, with the step size $h\asymp d^{-2}$, Corollary 7 immediately recovers mixing guarantees matching the ones for the $\mathsf{Speedy\ walk}$ obtained via blocking conductance and the log-Cheeger inequality [KLM06] (see Theorem 28).
 Outline of analysis.
@@
 The ideas and methods of this paper have led to subsequent progress.
-First, [KZ25] shows that the output guarantee can be strengthened to $\mathcal{R}_{\infty}$ under LSI.
-Second, the diffusion-based approach is extended to the general logconcave sampling in [KV25b], beyond the uniform distribution.
-Lastly, the $\mathcal{R}_{\infty}$-warmness requirement has been relaxed to significantly weaker $\mathcal{R}_{\widetilde{O}(1)}$-warmness in [KV25a].
+First, [KZ25] shows that the output guarantee can be strengthened to $\mathcal{R}_{\infty}$ under LSI, using $\widetilde{O}(d^{2}R^{2}\operatorname{polylog}\frac{1}{\varepsilon\eta})$ queries.
+Second, the diffusion-based approach is extended to general logconcave sampling in [KV25b], with the query complexity of $\widetilde{O}(d^{2}\,(R^{2}\vee 1)\operatorname{polylog}\frac{1}{\varepsilon\eta})$ for a $\mathcal{R}_{\infty}$ guarantee, beyond the uniform distribution.
+Lastly, the $\mathcal{R}_{\infty}$-warmness requirement has been relaxed to significantly weaker $\mathcal{R}_{\widetilde{O}(1)}$-warmness in [KV25a], improving the query complexity of warm-start generation from $\widetilde{O}(d^{2}R^{2})$ to $\widetilde{O}(d^{2}R^{3/2}\lVert\operatorname{cov}\pi\rVert^{1/4})$.
 1.4 Related work
@@
 $\mathsf{Hit\text{-}and\text{-}Run}$ is another zeroth-order algorithm that needs no step size; it picks a uniform random line $\ell$ passing a current point, and move to a uniform random point on $\ell\cap\mathcal{K}$.
-[LV06] shows that, if we define the second moment as $R^{2}:=\mathbb{E}_{\pi^{\mathcal{K}}}[\lvert X-\mathbb{E}X\rvert^{2}]$, then $\mathsf{Hit\text{-}and\text{-}Run}$ requires $O(d^{2}R^{2}\log\frac{M}{\varepsilon})$ queries.
+[LV06] shows that, if we define the second moment as $R^{2}:=\mathbb{E}_{\pi^{\mathcal{K}}}[\lvert X-\mathbb{E}X\rvert^{2}]$, then $\mathsf{Hit\text{-}and\text{-}Run}$ requires $O(d^{2}R^{2}\log\frac{M}{\varepsilon})$ iterations.
 Notably, this algorithm has a poly-logarithmic dependence on $M$ as opposed to the $\mathsf{Ball\ walk}$.
+Chen and Eldan [CE25] further showed that $\mathsf{Hit\text{-}and\text{-}Run}$ mixes in $\widetilde{O}(M^{11}d^{2}/\varepsilon^{11})$ times for isotropic convex bodies.
 Both algorithm are affected by skewed shape of $\mathcal{K}$ (i.e., large $D$ or $R$), so these samplers are combined with pre-processing step called rounding.
@@
 Then, the failure probability of one iteration of $\mathsf{In\text{-}and\text{-}Out}$ is at most $\eta/m$, and the expected membership queries per iteration is $O(M\log^{4}\nicefrac{{mM}}{{\eta}})$.
+The bias of conditioning on success can be bounded as follows.
+Lemma 16 (Bias from no failure).
+Let $\mathsf{Succ}$ be the event that the failure is not declared over the required number $m$ iterations, and suppose $\mathbb{P}(\mathsf{Succ})\geq 1-\eta$.
+Let $\mu_{m}$ be the law of the uncapped $\mathsf{In\text{-}and\text{-}Out}$ output after $m$ iterations, and $\mu_{m}^{\textup{cap}}$ be the law of Algorithm 1 conditioned on $\mathsf{Succ}$.
+Then, we have:
+$\frac{\mathrm{d}\mu_{m}^{\textup{cap}}}{\mathrm{d}\mu_{m}}$ $\leq\frac{1}{1-\eta}\quad\textup{so}\quad\mathcal{R}_{\infty}(\mu_{m}^{\textup{cap}}\|\mu_{m})\leq\log\frac{1}{1-\eta}\,$,
+$\mathcal{R}_{q}(\mu_{m}^{\textup{cap}}\,\|\,\pi^{X})$ $\leq\mathcal{R}_{q}(\mu_{m}\|\pi)+\frac{q}{q-1}\,\log\frac{1}{1-\eta}\qquad\text{for }q>1\,.$
 3.1 Functional inequalities
@@
 In §C, we show how the identity is derived under suitable regularity assumptions (see Lemma 29).
-Proposition 1 (General de Bruijn identity).
+Proposition 1 (General de Bruijn identity, [KO25]).
 Let $\mu$ and $\nu$ be probability measures on $\mathbb{R}^{d}$ such that $\mathbb{E}_{\nu}[\lvert\cdot\rvert^{4}]<\infty$ and $\chi^{q}(\mu\mathbin{\|}\nu)<\infty$. Then, for any $t>0$ and $q>1$,
@@
 $C_{\mathsf{PI}}(\pi_{t})\leq C_{\mathsf{PI}}(\pi)+t\,,\qquad\text{and}\qquad C_{\mathsf{LSI}}(\pi_{t})\leq C_{\mathsf{LSI}}(\pi)+t\,.$
-We can now deduce contraction under the heat flow.
+We record contraction results under the heat flow and functional inequalities in $q$-Rényi divergence, combining [CCSW22] and Proposition 1.
 Lemma 22 (Contraction under functional inequalities).
-Let $\mu,\nu$ be probability measures on $\mathbb{R}^{d}$ such that $\mathcal{R}_{q}(\mu\mathbin{\|}\nu)<\infty$. Then, for any $t>0$ and $q>1$,
+Let $\mu,\nu$ be probability measures on $\mathbb{R}^{d}$ such that $\mathcal{R}_{q}(\mu\mathbin{\|}\nu)<\infty$ and $\mathbb{E}_{\nu}[\lvert\cdot\rvert^{4}]<\infty$. Then, for any $t>0$ and $q>1$,
 $\mathcal{R}_{q}(\mu_{t}\mathbin{\|}\nu_{t})\leq\frac{\mathcal{R}_{q}(\mu\mathbin{\|}\nu)}{(1+t/C_{\mathsf{LSI}}(\nu))^{1/q}}\,.$
@@
 Proof of Theorem 14.
-Note that one iteration of $\mathsf{In\text{-}and\text{-}Out}$ corresponds to the composition of two transition kernels $P_{h}$ and $Q_{h}$, where $P_{h}(x,\cdot)=\mathcal{N}(x,hI_{d})$.
-The forward step ($\mathsf{FH}$) (i.e., transition kernel $P_{h}$) convolves the law $\mu_{k}^{X}$ of the $k$-th iterate with Gaussian $\gamma_{h}$, so the forward step invokes the contraction with rate dependent on $C_{\mathsf{PI}}(\pi^{X})$ or $C_{\mathsf{LSI}}(\pi^{X})$ by Lemma 19.
-Using the DPI (Lemma 9) to address the effect of $Q_{h}$, we obtain that
-$\mathcal{R}_{q}(\mu_{k+1}^{X}\mathbin{\|}\pi^{X})=\mathcal{R}_{q}(\mu_{k+1}^{Y}Q_{h}\mathbin{\|}\pi^{Y}Q_{h})\underset{\text{DPI}}{\leq}\mathcal{R}_{q}(\mu_{k+1}^{Y}\mathbin{\|}\pi^{Y})\leq\frac{\mathcal{R}_{q}(\mu_{k}^{X}\mathbin{\|}\pi^{X})}{(1+h/C_{\mathsf{LSI}}(\pi^{X}))^{1/q}}\,.$
+Note that one iteration of $\mathsf{In\text{-}and\text{-}Out}$ corresponds to the forward and backward step.
+The forward step ($\mathsf{FH}$) (i.e., transition kernel $P_{h}=\mathcal{N}(x,hI_{d})$) convolves the law $\mu_{k}^{X}$ of the $k$-th iterate with Gaussian $\gamma_{h}$, so the forward step invokes the contraction with rate dependent on $C_{\mathsf{PI}}(\pi^{X})$ or $C_{\mathsf{LSI}}(\pi^{X})$ by Lemma 22.
+Using the DPI (Lemma 11) to address the effect of the backward step ($\mathsf{BH}$), we obtain that
+$\mathcal{R}_{q}(\mu_{k+1}^{X}\mathbin{\|}\pi^{X})\underset{\text{DPI}}{\leq}\mathcal{R}_{q}(\mu_{k+1}^{Y}\mathbin{\|}\pi^{Y})\leq\frac{\mathcal{R}_{q}(\mu_{k}^{X}\mathbin{\|}\pi^{X})}{(1+h/C_{\mathsf{LSI}}(\pi^{X}))^{1/q}}\,.$
 Induction on $k$ proves the first claim. The other results can be proven in the similar fashion. ∎
@@
 Therefore, the expected number of trials per step is $O(M\log^{4}\frac{mM}{\eta})$, and the claim follows since each trial uses one query to the membership oracle of $\mathcal{K}$. ∎
+Lastly, we prove a simple lemma showing that the bias due to conditioning on success is bounded in terms of the failure probability $\eta$.
+Proof of Lemma 16.
+Let $X$ be the uncapped $\mathsf{In\text{-}and\text{-}Out}$ output after the desired number $m$ of iterations, and $\mathsf{Succ}$ be the event that the failure is not declared until $m$ iterations.
+Let $\mu_{m}^{X}$ be the law of $X$ and $\mu_{m}^{\textup{cap}}$ be the law of $X$ conditioned on $\mathsf{Succ}$.
+Then, by Bayes’ rule,
+$\frac{\mathrm{d}\mu_{m}^{\textup{cap}}}{\mathrm{d}\mu_{m}}(x)=\frac{\mathbb{P}(X=x\,|\,\mathsf{Succ})}{\mathbb{P}(X=x)}=\frac{\mathbb{P}(\mathsf{Succ}\,|\,X=x)\,\mathbb{P}(X=x)}{\mathbb{P}(\mathsf{Succ})\,\mathbb{P}(X=x)}\leq\frac{1}{\mathbb{P}(\mathsf{Succ})}\leq\frac{1}{1-\eta}\,.$
+Lastly, using the weak triangle inequality for Rényi divergence,
+$\mathcal{R}_{q}(\mu_{m}^{\textup{cap}}\,\|\,\pi)\leq\frac{q}{q-1}\,\mathcal{R}_{\infty}(\mu_{m}^{\textup{cap}}\,\|\,\mu_{m})+\mathcal{R}_{q}(\mu_{m}\,\|\,\pi)\leq\mathcal{R}_{q}(\mu_{m}\,\|\,\pi)+\frac{q}{q-1}\log\frac{1}{1-\eta}\,$,
+which completes the proof. ∎
 3.4 Putting it together
@@
 Theorem 26.
-For any given $\eta,\varepsilon\in(0,1)$, $q>1$, $m\in\mathbb{N}$ defined below and any convex body $\mathcal{K}$ given by a well-defined membership oracle, consider $\mathsf{In\text{-}and\text{-}Out}$ (Algorithm 1) with an $M$-warm initial distribution $\mu_{0}^{X}$, step size $h=(2d^{2}\log\frac{9mM}{\eta})^{-1}$, and threshold $N=\widetilde{O}(\frac{mM}{\eta})$.
-For $\pi^{X}$ the uniform distribution over $\mathcal{K}$,
-  • It achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{O}(qd^{2}\,\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{2}\frac{M}{\eta\varepsilon})$ iterations.
-With probability $1-\eta$, the algorithm iterates this many times without failure, using $\widetilde{O}(qMd^{2}\,\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{6}\frac{1}{\eta\varepsilon})$ expected number of membership queries in total.
-  • For isotropic $\pi^{X}$, with probability $1-\eta$, the algorithm achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ with $m=\widetilde{O}(qd^{2}\log^{2}\frac{M}{\eta\varepsilon})$ iterations, using $\widetilde{O}(qMd^{2}\log^{6}\frac{1}{\eta\varepsilon})$ membership queries in expectation.
+For any given $\eta,\varepsilon\in(0,1/2)$, $q\geq 2$, $m\in\mathbb{N}$ defined below and any convex body $\mathcal{K}$ given by a well-defined membership oracle, consider $\mathsf{In\text{-}and\text{-}Out}$ (Algorithm 1) with an $M$-warm initial distribution $\mu_{0}^{X}$, step size $h=(2d^{2}\log\frac{9mM}{\eta})^{-1}$, and threshold $N=\widetilde{O}(\frac{mM}{\eta})$.
+For $\pi^{X}$ the uniform distribution over $\mathcal{K}$, let $\mu_{m}^{X}$ be the law of the $m$-th iterate of $\mathsf{In\text{-}and\text{-}Out}$ conditioned on the success.
+  • With probability $1-\eta$, the algorithm iterates this many times without failure, using $\widetilde{O}(qMd^{2}\,\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{6}\frac{1}{\eta\varepsilon})$ expected number of membership queries in total and achieving $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon+4\eta$ after $m=\widetilde{O}(qd^{2}\,\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log^{2}\frac{M}{\eta\varepsilon})$ iterations.
+  • For isotropic $\pi^{X}$, with probability $1-\eta$, the algorithm achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon+4\eta$ with $m=\widetilde{O}(qd^{2}\log^{2}\frac{M}{\eta\varepsilon})$ iterations, using $\widetilde{O}(qMd^{2}\log^{6}\frac{1}{\eta\varepsilon})$ membership queries in expectation.
 Proof.
@@
 The $M$-warm start assumption guarantees $\mathcal{R}_{q}(\mu_{0}^{X}\mathbin{\|}\pi^{X})\lesssim\log M$.
-Due to $C_{\mathsf{PI}}(\pi^{X})=O(\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\,\log d)$ (Lemma 17), $\mathsf{In\text{-}and\text{-}Out}$ can achieve $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ if it iterates at least
+Due to $C_{\mathsf{PI}}(\pi^{X})=O(\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\,\log d)$ (Lemma 20), $\mathsf{In\text{-}and\text{-}Out}$ “without the cap” can achieve $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ if it iterates at least
 $m\gtrsim qd^{2}\,\lVert\operatorname{cov}\pi^{X}\rVert_{\mathsf{op}}\log d\log\frac{M}{\eta\varepsilon}\log\frac{mM}{\eta}\quad\text{times.}$
@@
 Since $\mathcal{R}_{q}$ is non-decreasing in $q$, we can obtain the desired bound on $\mathcal{R}_{q}$ for $q\in[1,2)$.
-For isotropic $\pi^{X}$, we have $\operatorname{cov}\pi^{X}=I_{d}$, so the claim immediately follows from $C_{\mathsf{PI}}(\pi^{X})=O(\log d)$. ∎
-We now show that the number of proper steps is bounded as claimed for general non-convex bodies and any feasible start in $\mathcal{K}$.
-We first establish this result under an $M$-warm start (Theorem 5).
-Proof of Theorem 5.
+For isotropic $\pi^{X}$, we have $\operatorname{cov}\pi^{X}=I_{d}$, so the claim immediately follows from $C_{\mathsf{PI}}(\pi^{X})=O(\log d)$.
+Lastly, as noted in Lemma 16, the conditional-on-success output distribution of $\mathsf{In\text{-}and\text{-}Out}$ introduces a small bias, bounded by $2\eta$.
+∎
+We show that the restart variant increases the expected total query complexity by at most a factor of $(1-\eta)^{-1}$.
+Proof of Corollary 4.
+Since the success probability of $m$ iterations without failure is at least $1-\eta$, the expected number of restarts is $(1-\eta)^{-1}$.
+Setting $4\eta=\varepsilon(<1/10)$ and replacing $\varepsilon$ by $\varepsilon/2$ in Theorem 26, the conclusion follows.
+∎
+We now bound the number of proper steps for general non-convex bodies and any feasible start in convex $\mathcal{K}$.
+We first record the following under an $M$-warm start, as a corollary of the contraction results in [CCSW22, KO25].
+Proposition 2.
+For any given $\varepsilon\in(0,1)$ and set $\mathcal{K}\subset B_{D}(0)$ with $\operatorname{vol}(\mathcal{K})>0$, $\mathsf{In\text{-}and\text{-}Out}$ with variance $h$ and $M$-warm initial distribution achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after the following number of iterations:
+$m=\min\begin{cases}O\bigl(qh^{-1}C_{\mathsf{PI}}(\pi^{X})\log\frac{M}{\varepsilon}\bigr)&\text{for }q\geq 2\,,\\ O\bigl(qh^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{\log M}{\varepsilon}\bigr)&\text{for }q\geq 1\,.\end{cases}$
+Proof.
 By the Rényi-decay under (10) in Theorem 14, $\mathsf{In\text{-}and\text{-}Out}$ can achieve $\varepsilon$-distance to $\pi^{X}$ after $O\bigl(qh^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{\mathcal{R}_{q}(\mu^{X}_{1}\mathbin{\|}\pi^{X})}{\varepsilon}\bigr)$ iterations for $q\geq 1$.
@@
 Next, we show that $\mathsf{In\text{-}and\text{-}Out}$ mixes from any start.
-Corollary 24.
-For any given $\varepsilon\in(0,1)$ and set $\mathcal{K}\subset B_{D}(0)$, $\mathsf{In\text{-}and\text{-}Out}$ with variance $h$ and any feasible start $x_{0}\in\mathcal{K}$ achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{O}(qh^{-1}C_{\mathsf{LSI}}(\pi^{X})\log\frac{d+D^{2}/h}{\varepsilon})$ iterations.
-Proof.
+Proof of Corollary 6.
 We first bound the warmness of $\mu^{X}_{1}$ w.r.t. $\pi^{X}$ when $\mu^{X}_{0}=\delta_{x_{0}}$. One can readily check that
@@
 We then obtain the following corollary for a convex body $\mathcal{K}$.
-Corollary 25.
-For any given $\varepsilon\in(0,1)$ and convex body $\mathcal{K}\subset B_{D}(0)$, $\mathsf{In\text{-}and\text{-}Out}$ with variance $h$ and a feasible start $x_{0}\in\mathcal{K}$ achieves $\mathcal{R}_{q}(\mu^{X}_{m}\mathbin{\|}\pi^{X})\leq\varepsilon$ after $m=\widetilde{O}(qh^{-1}D^{2}\log\frac{1}{\varepsilon})$ iterations.
-If $\pi^{X}$ is isotropic, then $\mathsf{In\text{-}and\text{-}Out}$ only needs $\widetilde{O}(qh^{-1}D\log\frac{d+d^{2}/h}{\varepsilon})$ iterations.
-Proof.
+Proof of Corollary 7.
 For convex $\mathcal{K}$, it follows from Lemma 20 that $C_{\mathsf{LSI}}(\pi^{X})=O(D^{2})$ and $C_{\mathsf{LSI}}(\pi^{X})=O(D)$ for isotropic $\mathcal{K}$.
@@
 The mixing time of the $\mathsf{Ball\ walk}$ under this approach has a polynomial dependence on $1/\varepsilon$, rather than a polylogarithmic dependence.
-Indirect analysis through $\mathsf{Speedy\ walk}$.
-[KLS97] introduced the $\mathsf{Speedy\ walk}$, which could be viewed as a version of the $\mathsf{Ball\ walk}$ and converges to a speedy distribution (see Proposition 2), which is slightly biased from $\pi$.
+Indirect analysis through $\mathsf{Speedy\ walk}$.[KLS97] introduced the $\mathsf{Speedy\ walk}$, which could be viewed as a version of the $\mathsf{Ball\ walk}$ and converges to a speedy distribution (see Proposition 3), which is slightly biased from $\pi$.
 Then, the $\mathsf{Speedy\ walk}$ is used together with another algorithmic component (rejection sampling) [KLS97, Algorithm 4.15] that converts the speedy distribution to the uniform distribution.
@@
 We use $\mu$ and $\mu_{\mathsf{LC}}$ to denote a probability measure and logconcave probability measure over $\mathbb{R}^{d}$, respectively.
-Cheeger and PI constants.
-The Cheeger isoperimetric constant $C_{\mathsf{Ch}}(\mu)$ measures how large surface area a measurable subset with larger volume has, defined by
+Cheeger and PI constants.  The Cheeger isoperimetric constant $C_{\mathsf{Ch}}(\mu)$ measures how large surface area a measurable subset with larger volume has, defined by
 $C_{\mathsf{Ch}}(\mu):=\inf_{S\subset\mathbb{R}^{d}}\frac{\mu^{+}(S)}{\min(\mu(S),\mu(S^{c}))}\,$,
@@
 We also provide key technical ingredients needed for its proof, such as the computations for measures evolving under simultaneous forward/backward heat flows.
-Forward heat flow.
-We begin by introducing the “heat flow” equation (or also known as the Fokker–Planck equation), which describes the evolution of the law of $Z_{t}$ under ($\mathsf{FH}$),
+Forward heat flow.  We begin by introducing the “heat flow” equation (or also known as the Fokker–Planck equation), which describes the evolution of the law of $Z_{t}$ under ($\mathsf{FH}$),
 $\partial_{t}\mu_{t}=\frac{1}{2}\,\Delta\mu_{t}=\frac{1}{2}\operatorname{div}(\mu_{t}\nabla\log\mu_{t})\,.$ ($\mathsf{FP\text{-}FH}$)
@@
 […]
-Backward heat flow.
-Although there are many ways to define a “reversal” of $P_{h}$, we will use the notion of adjoint introduced by [KP23], which is the most immediately useful.
+Backward heat flow.  Although there are many ways to define a “reversal” of $P_{h}$, we will use the notion of adjoint introduced by [KP23], which is the most immediately useful.
 Given some initial measure $\nu$ and some time horizon $h$, the adjoint corresponds to reversing ($\mathsf{FH}$) for times in $[0,h]$ when the initial distribution under consideration is $Z_{0}\sim\nu$.
@@
 From this, we can obtain a single step of decay for the Rényi and $\chi^{2}$-divergences under different functional inequalities.
+Appendix D Bias from failure
+By computing the one-step distribution, we can note that $\mathsf{In\text{-}and\text{-}Out}$ conditioned on success has a bias.
+To see this, consider $\mathsf{In\text{-}and\text{-}Out}$ with threshold $N$ and the initial measure $X_{1}\sim\mu$.
+Then, for a measurable set $A$ and the success event $S$,
+$\mathbb{P}(X_{2}\in A\mid S)$ $=\frac{\mathbb{P}(X_{2}\in A,S)}{\mathbb{P}(S)}=\frac{1}{\mathbb{P}(S)}\int\mathbb{P}(X_{2}\in A,X_{1}=x,S)\,\mathrm{d}x$
+$=\frac{1}{\mathbb{P}(S)}\int\mathbb{P}(X_{2}\in A,S\mid X_{1}=x)\,\mu(\mathrm{d}x)$
+$=\frac{1}{\mathbb{P}(S)}\int\mathbb{P}(X_{2}\in A,S,Y_{2}=y\mid X_{1}=x)\,\mu(\mathrm{d}x)\,\mathrm{d}y$
+$=\frac{1}{\mathbb{P}(S)}\int\mathbb{P}(X_{2}\in A,S\mid X_{1}=x,Y_{2}=y)\,\gamma(y-x)\,\mu(x)\,\mathrm{d}x\mathrm{d}y$
+$=\frac{1}{\mathbb{P}(S)}\int\mathbb{P}(X_{2}\in A,S\mid Y_{2}=y)\,\gamma(y-x)\,\mu(x)\,\mathrm{d}x\mathrm{d}y$
+$=\frac{1}{\mathbb{P}(S)}\int\mathbb{P}(S\mid Y_{2}=y)\mathbb{P}(X_{2}\in A\mid S,Y_{2}=y)\,\gamma(y-x)\,\mu(x)\,\mathrm{d}x\mathrm{d}y$
+$=\frac{1}{\mathbb{P}(S)}\int\bigl(1-\bigl(1-\ell(y)\bigr)^{N}\bigr)\,\pi^{X|Y=y}(A)\,\gamma(y-x)\,\mu(x)\,\mathrm{d}x\mathrm{d}y$
+$=\frac{\int\bigl(1-\bigl(1-\ell(y)\bigr)^{N}\bigr)\,\pi^{X|Y=y}(A)\,\gamma(y-x)\,\mu(x)\,\mathrm{d}x\mathrm{d}y}{1-\mathbb{E}_{\mu*\gamma}[(1-\ell)^{N}]}\,.$
+On the other hand, $\mathsf{In\text{-}and\text{-}Out}$ without cap (i.e., $N=\infty$) satisfies
+$\mathbb{P}(X_{2}\in A)=\int\pi^{X|Y=y}(A)\,\gamma(y-x)\,\mu(x)\,\mathrm{d}x\mathrm{d}y\,.$
+Thus, trajectories are weighted differently in the two distributions.